A Complete Bibliography of ACM Transactions on Knowledge Discovery from Data (TKDD)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

28 August 2024 Version 1.62

Title word cross-reference

(k, P) [WLX^{+}23]. 3 [XLL^{+}21, ZELS24]. ^{2} [BXF^{+}20]. ^{3} [LGS^{+}23]. _{h} [PXW^{+}22]. A∗ [SYZ^{+}24b]. B [LK20]. Δ [TC18]. k [BBC^{+}19, GEG^{+}08, MKGV07, MM12, NADR21, SB24, SCS20, ZLY^{+}20a]. L [MKGV07]. N [ML15, CBRB09]. t [SHF24].

-anonymity [MKGV07]. -ary [CBRB09]. -center [GEG^{+}08]. -Core [SCS20, WLX^{+}23]. -Dimensional [XLL^{+}21]. -diversity [MKGV07]. -hop [ZELS24]. -hypergraphs [LK20]. -Means [MM12]. -SNE [SHF24]. -Subgraph [NADR21].

19 [BBB^{+}22, LLL^{+}23b, ZHSL23].

2006 [BBD^{+}07]. 2007 [MG09]. 2008 [MG09]. 2009 [Wan10]. 2011 [GSTC12]. 2012 [ACPW13]. 2014 [WL16]. 2016 [Agg17]. 2018 [XL20]. 2019 [LGZ^{+}22].

3E [LLY^{+}21]. 3E-LDA [LLY^{+}21].

8M [BT21].

ABLE [HFL^{+}22]. ABRA [RU18]. Absence [SB21b]. Academic [WLX^{+}23]. Accelerating [JYY^{+}21]. Accident [LCX^{+}23]. According [DKSK22]. Accumulated [SWH^{+}23]. Accuracy [CSM^{+}24, MM12, XYW^{+}20]. Accurate [JSE^{+}23, LNG18, LJK18, SHF18, SOK^{+}20, SLO^{+}21, YWDP16, Kor10]. Accurately [HKR^{+}23]. ACM [ACPW13, BBD^{+}07, GL15, MG09, WL16, ZPSYY10, Agg22]. Across [CRC^{+}23, GLZ^{+}21, JH19, KZW^{+}22, MAHT18, WYWY19, ZTL15a, ZGQ^{+}24, BGJV12, Jia24, QZB^{+}23]. Action [CZW^{+}24, HLC19, SKW^{+}21, YLHY20, VALF12]. Active [BIPR13, CWF^{+}13, Dor19, HHZ^{+}18, LCZ^{+}24, WY15, XXO^{+}24, BG09]. Activity [CYT^{+}17, LTN^{+}21, SYZB24, WLL^{+}21, ZCL^{+}22]. AD [LLY^{+}24]. Ada [RZY^{+}23]. Ada-MIP [RZY^{+}23]. Adaptation [CWL^{+}24b, HBD^{+}24, LYZ^{+}24, LCZ^{+}24, RET^{+}23, SKW^{+}21, WN22b, CSF^{+}12, TLZ^{+}08]. Adapting [SRVM24]. Adaptive [AKM18, BLP21, CCT^{+}23, DGB16, GSG^{+}20, GW21, HCH24a, KCR^{+}24, KBR^{+}16, LLW^{+}21, LTN^{+}08, LFL^{+}23, NWW^{+}20, RZY^{+}23, SZD^{+}23, SJHY24, TLG^{+}23, TKT^{+}24, WDF22, XXZ19, YZL^{+}22, ZLZ^{+}19, ZZY^{+}23, ZLDL23]. ADATIME [RET^{+}23]. Addendum [vLCV^{+}19]. Adding [WGC^{+}23]. Addition [Pap15]. Additive [LJW^{+}22, SM23]. Address [CG15, PCPN24]. Addressing [RCM^{+}13]. Adjusting [HSB22, GVdSL24, RKC19]. Adopter [ZWH^{+}16]. Adversarial [CDC23, DKSL18, FJZT24, HCH24a, HZM^{+}22, LLL23a, LCW24, LGZ^{+}21, LZY^{+}24b, LJG^{+}24, PNB24, WLH24, XWZ^{+}22, ZLD^{+}23, ZRJ^{+}22, ZJJ^{+}24, ZBAG20]. Adversary [LCW24]. Advertising [OCD^{+}24, ZBL^{+}20]. Advice [HHZ^{+}18]. Affect [ACE20]. Against [AKM17, YLL^{+}21a]. AGENDA [LCZ^{+}24]. Agent [RQM^{+}24]. Aggregated [HYYQ15, KCR^{+}24]. Aggregation [CG23, SJHY24, WMW^{+}22, YZW^{+}24, GMT07]. Agnostic [WJR^{+}10, ZYD24b]. AI [CXL^{+}24]. Aided [JSF^{+}24, LZLJ24]. Air [CXWH24, RGL^{+}23, WSDL19, XZSY19]. Aleatory [BAC23]. Algebraic [Bur21]. Algorithm [BBK19, BT21, Con20, FKKD17, JSP15, LWS22, LSZ^{+}24, LYL^{+}22, LWY16, LYGG22, MHS20, RGAÁC^{+}24, SYZ^{+}24b, WSZZ14, WLD^{+}23, WGYC21, WHC^{+}23, YLL^{+}21a, AF09, Vad10, WC12]. Algorithms [AK15, AH24, AKM20, BGSW13, BCC^{+}21, CCV19, CPC10, DA18, DDL24, GBGL20, GCB^{+}21, GJ16, GNMQ21, GSWJ20, JZS23, LCCM19, LYGG22, ME11, SLK^{+}24, SHF18, SGCH19, SRVM24, WZL^{+}16, YSY^{+}22, YW24, ZXL22, ZLW^{+}21a, BBCG10, GEG^{+}08, ZZW09]. Algorithms-Experiments [GNMQ21]. Alignment [BGSW13, HST^{+}23, MAHT18, RYM^{+}24, XXG^{+}24, ZTT^{+}20, ZWYW24, ZRJ^{+}22]. Alleviate [CC19]. Allocation [GL22, ZP15]. Alzheimer [LCG^{+}18]. among [LZZ^{+}22b, VDMC22]. AMS [JH19]. Analysis [ACE20, ABS19, AZBW21, BHW^{+}17, CLG^{+}19, CNZ^{+}17, CMS24, GNMQ21, GSG^{+}20, HLCR20, LSF18, LLY^{+}21, LW24, LAN^{+}18, LZF^{+}15, LWG^{+}16, MYL^{+}21, NWW^{+}20, PTL22, RBF^{+}21, SMS22, STD^{+}18, WQZ^{+}16, XL15, XZLL21, YWZ^{+}24, YHL15, ZZYY16, ZHT20, ZWB22, ZCZQ19, CRST09, DD09, DAR09, GEG^{+}08, LNR08, LTN^{+}08, STP^{+}08, ZLT09]. analytical [DD09]. Analytics [AT17, CKC^{+}18, LQW15, SGM^{+}23, SGCH19, WLP18, vLCV^{+}18, vLCV^{+}19]. Analyzing [KUU10, LCZ^{+}09, WDDB20, ASHK14, DD09]. Anchor [LF23, WZW^{+}22]. Anesthesiologists [MFRJ16]. Annotation [BFRL13, HLL^{+}20, PL10, WLL^{+}22a, WDF22, MXC^{+}07, YCJK08]. Annotations [GLMW22]. Anomalies [CNZ^{+}17, DWD^{+}20, LSL^{+}22, PA18]. Anomalous [CYT^{+}17, GLG^{+}22, NLA23, WWHW19]. Anomaly [Ang20, BLH^{+}22, EE24, GJDX14, HBD^{+}24, HQYY14, JZS23, KBR^{+}16, LNG18, LYTZ22, LLY^{+}24, LZV24, LTZ12, LCN14, LLL^{+}21, RA16, SFDW19, WJR^{+}10, YWW^{+}21, YHL15, ZLD^{+}23]. anonymity [MKGV07]. Anonymization [MFHL10]. Anonymized [WFW^{+}11, LNR08]. anonymizing [CGL18]. Answering [BWD10, GFM21, MSC^{+}19, ALB09]. Any [SGM^{+}24]. App [LWG^{+}16, LCF19, OGT^{+}21]. App2Vec [WLD^{+}21]. Appliances [JKP^{+}21]. Application [DPDG18, LYGG22, LFL^{+}23, MM12, SHL19, SDS18, WLD^{+}21, YWC24, CSF^{+}12]. Applications [AHGA14, GBGL20, GCB^{+}21, GHT^{+}24, Jia24, LYWL12, LSF18, LK15, LBT^{+}23, MH22, MZC^{+}24, NGB18, RJK^{+}20, WZL^{+}16, ZXL22, ZH16, CT14, GEG^{+}08, STP^{+}08]. Applied [QSS20, HAKU^{+}08]. Approach [AKM18, BPW^{+}18, CPP20, CWW^{+}24, CGZW16, CRGP14, CTZ16, DA19, HF24, HCZ^{+}14, JV20, JCB^{+}16, JYD19, JS21, KN18, KHTR18, KSB^{+}21, KL23, LSL^{+}22, LVH^{+}21, LGS^{+}23, LYL^{+}20, PE20, PNB24, PL10, RTM18, SMS22, TYZZ10, WYG^{+}17, WCZ^{+}22, XML18, XSZY20, XSZ^{+}22, YWS^{+}22, YZH^{+}18, YDS^{+}15, ZY14, ZL23, ZHSL23, BG09, CX10, Web10]. Approaches [LLW^{+}21, WCX24, ZLY^{+}20a]. Approximate [KN18, KRBK19, NADR21, PCVR22, SCS20]. Approximately [YSY^{+}22]. Approximating [ME11, RU18]. Approximation [CWR23, HP20, HSBH19, LJW^{+}22, MLM21, SÖ22]. Architecture [GBBC24]. Area [BBB^{+}22]. ArieL [FJZT24]. ARIS [CYY^{+}22]. Arrangement [ZLY^{+}20b]. Arrays [dVKCC11]. Arrival [ZLZ^{+}24]. Art [CPC10]. Articles [ALW^{+}24, SG12]. Artificial [HL22, SB21b]. ary [CBRB09]. ASCOS [CG15]. Aspect [WZL^{+}23b, XJW^{+}21, ZCZQ19]. Aspect-Aware [WZL^{+}23b]. aspects [CT14]. Assessing [AKM17, GMMT07, VDMC22, WWHW19]. Assessment [BBB^{+}22, LWLW18, MFRJ16, YHW^{+}24]. Asset [CZW^{+}24]. Assignment [ADK^{+}16, GSWJ20]. Assignments [CO18]. Associated [HBD^{+}24]. Association [RU14, YWR^{+}19, Bal13, HDC07, ZZW09]. Associations [WV14, Web10]. Assurance [CTZ16, WCX24]. Asymmetric [CG15, OGV22, YL24]. Asynchronous [QZZZ24, SYXZ23]. Atmospheric [LHS^{+}21]. ATR [MCS^{+}18]. ATR-Vis [MCS^{+}18]. Attacking [DLZ^{+}24, MSR^{+}24, WLG^{+}24]. Attacks [LCW24, ZBAG20]. Attention [ACE20, CHY^{+}24, CSZ^{+}21, GYX^{+}22, HHL^{+}24, JSF^{+}24, KCR^{+}24, KS24, LRK^{+}19, LXC^{+}22, LLY^{+}24, LWG^{+}22, OJW^{+}23, SMW^{+}24, WZ21, XLT^{+}20, YSLL24, ZZYY22, ZYC^{+}24, ZZY^{+}24b, PMC22]. Attention-Aided [JSF^{+}24]. Attention-based [CSZ^{+}21, XLT^{+}20]. Attentional [LZP20]. Attentive [PCPN24, XLQ^{+}24, XSZY20]. Attraction [CLYC23, TC09]. Attractiveness [GSI19]. Attractor [SYD^{+}16]. Attribute [CZY11, KS21, LCW24, LZY^{+}24b, ZYC^{+}24, GEG^{+}08]. Attribute-Guided [KS21]. Attributed [AAB^{+}24, CSX21, CZY11, KZM^{+}23, LHW^{+}21, LFXC24, LOM22, PA18, SHH^{+}20, WWZ^{+}24, XKW^{+}14, ZYC^{+}24]. Attributes [LZL^{+}24, MYB19, MA16, SZLP16, ZZY^{+}24b]. Attributes-driven [ZZY^{+}24b]. Attribution [JH22, KSV^{+}16]. Augmentation [CQH^{+}24, HLC19, KCP^{+}24, KL23, LWWW24, LZL^{+}22, LHZ^{+}24b, LILJ24, WDF22]. Augmented [WOGZ24]. Author [MHKG19, PMC22, TS09]. Authorship [CMS24, JH22]. Autism [ASJ^{+}23]. Auto [DWD^{+}24, ZJ24, ZLW21b, WZWC23, YWK^{+}22]. Auto-Encoder [DWD^{+}24, ZJ24]. Auto-Encoders [ZLW21b]. Auto-STGCN [WZWC23]. Autoencoder [LLL23a, LLY^{+}24, ZWC22]. Automate [Var22]. Automatic [BMTT18, GMS21, YWK^{+}22, ZGC18, ZLW^{+}21a]. Automatically [WCZ^{+}24, YWW^{+}21]. AutoML [PFC^{+}24, PDK24]. Automotive [HCY^{+}23]. Autonomous [WZWC23, YYC^{+}21]. Auxiliary [LLX^{+}24, YWC24]. Average [CWY^{+}23]. Averages [CWR23, PCVR22, RU18]. avoidance [KRPS12]. Aware [CRP^{+}24, CXWH24, FYN22, GXLC21, GYX^{+}22, HZZ^{+}15, HYYQ15, HGS^{+}24, KS24, LSS^{+}22, LHW^{+}20, LWH^{+}24, QCD^{+}19, SYG^{+}24, SMY^{+}24, SLWW24, TKT^{+}24, TZR24, WLD^{+}21, WLL^{+}22a, WZL^{+}23b, XW22, XSZ^{+}22, ZP15, ZYC^{+}24, ZWYW24, ZJ24, ZCZQ19, ZCL^{+}22, CSZ^{+}21, CWR23, FLL^{+}22, GLW^{+}23, KCP^{+}24, LWCZ24, LWW24, NXZ^{+}24, SYZ^{+}24b, SSL^{+}23, WZL^{+}23a, XLQ^{+}24]. Awareness [ZMW^{+}24]. axiomatic [JLL14].

B [PXW^{+}22]. Back [CYL^{+}24]. Back-In-Time [CYL^{+}24]. Backdoor [YCW^{+}24]. Background [ZH23, ZFH23, KUU10]. Badges [KHTR18]. Bags [SWY^{+}24]. Balance [CZY11, KZW^{+}22, XYW^{+}20]. Balance-Subsampled [KZW^{+}22]. Balancing [BTBg22, IKK19, KCL^{+}20]. Bandit [LSH^{+}22, ZXL22]. BapFL [YCW^{+}24]. Based [ASZ21, AF13, BHW^{+}17, BBK19, BJ21, BBB^{+}22, CPP20, CWF^{+}13, CLG^{+}19, CLT^{+}20, CTC^{+}22, CLYC23, CXL^{+}24, DWW22, DIB24, FCWQ17, FLG^{+}21, FYN22, GYX^{+}22, GNMQ21, GZXF16, GW20, HL22, HZZ^{+}15, HSS^{+}17, HQYY14, HYYQ15, HCC^{+}18, HXY^{+}19, HF12, JKP^{+}21, JK22, KN18, KCR23, KL23, LYG^{+}19, LWB^{+}22, LYTZ22, LJX23, LGG^{+}23, LLL^{+}23b, LOM22, LLS^{+}21, LWW23a, LTB18, LTZ12, LLLW22, LZY^{+}24b, LZD16, MDV11, MM12, MZC^{+}24, PTL22, QZL23, RWD23, RQM^{+}24, RLDL24, RYM^{+}24, SRBC22, SMS22, SI21, SGCH23, SYZ^{+}24a, WSZZ14, WBL21, WZW^{+}22, WCZ^{+}22, WLD^{+}23, WWZL23, WQZ^{+}16, WLS^{+}23, WZX^{+}23, WHC^{+}23, YLL^{+}21a, YCC^{+}15, YWG^{+}22, YWW^{+}21, YTW^{+}16, YDW^{+}23, ZLT^{+}15, ZZYY16, ZWG^{+}19, ZXZ^{+}23, ZLT^{+}23, ZFWC18, ZL15, ZH21, ZPC^{+}16, ZMZS23, AF09, APU09, CMS23, CCTW23, CCG^{+}24, CSHZ21, CSZ^{+}21, CWWL24, DA19, DD09, Dor21, GLG^{+}22, GFHL23, HGV^{+}08, Han24, II08, JCQ^{+}23, KMH^{+}24, KKZ09, LK20]. based [LWWW24, LTN^{+}08, LWG^{+}24, NK21, NK20, PKSM21, RJK^{+}20, SZWR22, SGM^{+}24, SYZ^{+}24b, SHF24, TWC22, TC09, WND^{+}09, WCL19, WWW^{+}20, WLX^{+}23, WZZ^{+}24, XLT^{+}20, YHLC23, YL24, YSLL24, ZCS10, ZZG^{+}21, ZL23, ZELS24]. Based-on [RLDL24]. Basket [SOL22]. Batch [CWF^{+}13, WY15, ZELS24]. Bavarian [CWR23]. Bayes [LLW16]. Bayesian [CSSP15, CZH18, Con20, CO18, GY15, GSWJ20, HCZ^{+}14, JYD19, KRBK19, LYL^{+}22, LGF10, LJW^{+}22, OGAB14, PL10, WC12, XKW^{+}14, YZZ22, YLL^{+}24, ZZY24a]. BCI [YSW^{+}21]. Be [GW21, LWWX23, WFW^{+}11]. Behavior [CSZ^{+}21, CYT^{+}17, DWW22, LLW16, SOL22, WLG^{+}24, XL16, XLT^{+}20, YWZ^{+}24, YBTL24, APU09, LWG^{+}24, ZFY14]. Behavior2Vec [Che18]. Behavioral [BGWSB19, CPC10, HHL^{+}24, ZZY^{+}24b, LSY^{+}09]. Behaviors [Che18, JCB^{+}16, LSZ^{+}19, LZZ^{+}22b, WWHW19]. Bellwether [CRST09]. Benchmark [AKM17, LFY^{+}23, TDLM19]. Benchmarking [RET^{+}23, SB21a]. BERT [LYGG22]. Best [Agg17, WL16, XL20, GSTC12]. Beta [IAB22]. Beta-Liouville [IAB22]. Better [LCWC20, QZB^{+}23, TSRK20]. between [CZY11, MYB19, QZB^{+}23, Web10, XYW^{+}20, ZHLZ24]. Betweenness [CWR23, LK20, PV24, RU18]. Beyond [BCK^{+}18, LLG^{+}23, LGZ^{+}22, RZF^{+}24, TLG^{+}23, YJT^{+}24, MKGV07]. Bi [LZP20, WMZL24]. Bi-Directional [LZP20]. Bi-graph [WMZL24]. Biaffine [ZHLZ24]. Bias [AZD^{+}21, BSK24, HHL^{+}24].

Biasing [LWWX23]. Bid [OCD^{+}24]. Bidding [LCWC20, OCD^{+}24]. Bidirectional [CSZ^{+}21, GSG^{+}20]. Big [AKM20, CTZ16, GSWJ20, RCM^{+}13, SGCH19, TWCO16, YWDP16]. BigChat [TWCO16]. Bike [WGC^{+}23]. Bike-sharing [WGC^{+}23]. BiLabel [ZFW22]. BiLabel-Specific [ZFW22]. Binary [DSTA22, DSL^{+}14, LV18, ZYZZ21]. bioinformatics [ZKYW08]. biological [JMR08]. Biomedical [CTZ16, SMA^{+}08]. Bipartite [CZY^{+}22, LFL^{+}23, WQZ^{+}16, ZLDL23]. Birth [MHKG19]. Birthday [JSP15]. BISC [CX10]. bitmap [CX10]. Blocking [KSM09, NZW23, OJLD22, YLW^{+}19, dVKCC11]. Blockmodel [LYC^{+}21]. Blog [YWC^{+}16]. Bloom [CXH^{+}20, PXW^{+}22, dVKCC11]. Boltzmann [JSG^{+}19]. Boolean [DSTA22, MV14]. boost [Bal13]. Boosting [FH14, NK20]. Bootstrap [MM12]. Bot [XLT^{+}20]. Both [AZD^{+}21, BAC23, LWC^{+}23]. Bottleneck [YLHY20]. Bound [LWW23a]. Boundary [AZD^{+}21]. Bounded [LLZ^{+}21]. Bounds [MSC^{+}19, PV24]. Bradykinesia [LLL^{+}20]. Brain [LLW^{+}21, STD^{+}18]. Branch [LLD^{+}24]. Breaches [WFW^{+}11]. Breadth [Bur21]. Breadth-First [Bur21]. Breaking [CCG^{+}24, LHN^{+}20]. Bregman [GG08, ZZL15]. Bribery [SRBC22]. Bridge [GLF^{+}22]. Bridge-Item [GLF^{+}22]. Browsing [LGF10]. bubble [GG08]. Budget [LCWC20, ZBL^{+}20]. Budget-Constrained [LCWC20]. Bug [WCZ^{+}24]. Building [RA16, ZHLZ24, CGL^{+}23]. Bundle [LFC^{+}17]. Bursty [CFP19, XZXW18]. Business [OLL20, XSZY20]. Butterfly [SÖ22]. Buying [LFC^{+}17]. Buzz [CYOL16].

Cache [TAJY17]. Cache-Conscious [TAJY17]. Calibration [TSRK20]. Call [YZH^{+}18]. Camouflage [HSS^{+}17]. Can [WFW^{+}11, YCW^{+}24]. Cancer [XKH^{+}16]. CANDECOMP [PFS15]. candidate [THD^{+}08]. Candidates [WTD24]. Cannot [YHCL12]. Capsule [ZLT^{+}23]. Capture [MNK18]. Car [HXY^{+}19]. Cardinality [AT17]. Career [Lap20]. Carlo [PCVR22]. Cascade [PTL22, XZXW18, AGHN13]. Cascade-Based [PTL22]. Cascades [XZLL21]. Case [GLMW22]. CASIN [CT14]. Catastrophic [MH22]. Catching [JCB^{+}16]. Categorical [IPM12]. Categorization [LZL^{+}22, YLHY20]. Category [SMY^{+}24, ZCZQ19]. Category-Aware [SMY^{+}24]. Causal [CNZ^{+}17, LWWX23, LXZ^{+}24, PZJ^{+}24, SGC^{+}23, SMW^{+}24, YCL^{+}21, YLL21b, YYD22, ZWC^{+}23]. Causality [AZBW21, LLH^{+}24, LHS^{+}21, SYZ^{+}24a]. Causality-Based [SYZ^{+}24a]. CausalSE [WGC^{+}23]. CBR [GSWJ20]. CenEEGs [DPB^{+}20]. Censored [YMD^{+}24]. center [GEG^{+}08]. Centralities [PV24]. Centrality [BBC^{+}19, CWR23, CDSV16, LK20, LLS^{+}21, RU18, SKSÇ17, SGCH23]. Centralized [MFHL10, SYXZ23]. Centric [RTM18]. CFOF [Ang20]. CGC [CGZW16]. Chain [CCLZ18]. Chains [WSM^{+}18]. Chameleon [BBK19]. Change [HCH^{+}24b, JYD19, CBLH12]. Change-Points [JYD19]. Changes [TC18]. Characterization [OLL20, WZL^{+}15]. Characterizing [LGZ^{+}23, MRTW19, APU09]. Chargers [WZR^{+}23]. ChatGPT [YJT^{+}24]. Chernoff [PSFV13]. Chief [Agg22]. Children [ASJ^{+}23]. Chinese [YWC^{+}16]. CHIRP [WAD12]. Choices [LWG^{+}16]. Chromatic [BGG^{+}15]. Churner [Row16]. ciForager [CBLH12]. Circles [BGC14, ML14]. Citation [JSF^{+}24]. Cities [DDL21]. City [BBB^{+}22, CXWH24, WZR^{+}23]. Citywide [CLT^{+}20, DWH^{+}23, FLL^{+}22, XLL^{+}21]. Class [AZD^{+}21, BSS^{+}24, NWA20, RWD23, SI21, SKM^{+}22]. Class-Boundary [AZD^{+}21]. Classes [WYWY19]. Classical [Var22]. Classification [AGHN13, AF13, BES15, BAMK18, CMS23, CRC^{+}23, CS24, CSM^{+}24, DPB^{+}20, DWL^{+}24, Dor21, DPDG18, FH14, GBBC24, GY15, HSB22, HLH^{+}24, IKK19, JV20, JLD^{+}19, JDE^{+}12, LV18, LWWW24, LTB18, LZX22, LHG^{+}24, LFL^{+}23, MA16, MES22, PSYJ24, PNB24, SZWR22, WSZZ14, WNH15, WBL21, WPW23, WDH^{+}24, WPD^{+}24, WN22a, WHC^{+}23, XLS24, XXO^{+}24, XXZ19, YGL^{+}22, YWS^{+}22, YWL^{+}24, YHZ^{+}24, YHCL12, YDS^{+}15, ZZG^{+}21, ZFW22, ZWYZ24, ZLHZ24, JTYY10, LTN^{+}08, MWF08]. Classifier [BC18, Dor19, DIB24, JV20, SGM^{+}24, TSRK20, WAD12]. Classifiers [AHGA14, JK22]. ClassiNet [BAMK18]. Click [DLZ^{+}24, XYGC21]. Click-Through [XYGC21, DLZ^{+}24]. Client [MYC^{+}24]. Clinical [TR22]. Clique [WTD24]. Clique-Querying [WTD24]. Cliques [SMDHT21, JP09]. Closed [CBRB09]. Closeness [BBC^{+}19, CDSV16]. Closure [HDT^{+}18, YWR^{+}23, LTH^{+}13]. Cluster [JV20, KBR^{+}16, LFXC24, SHF24, DAR09, GEG^{+}08, LQW^{+}18]. Cluster-Contractive [SHF24]. Clustered [CGL18]. Clusterers [AHGA14]. Clustering [AH24, DMGC21, BBK19, BGWSB19, BGJV12, BGG^{+}15, CMZS15, CCT^{+}23, CZY11, CGZW16, CWWL24, DA18, DA19, DWD^{+}24, GMT07, GF23, GFHL23, GFL^{+}24, GWZZ17, HYYQ15, IPM12, KZM^{+}23, KKZ09, LQW^{+}18, LHK^{+}18, LZSY24, LFXC24, LSW^{+}24, LF18, LL24, LF23, MDV11, MM12, MHS20, MG24, NC22, PGR18, PFTR16, PKH^{+}17, PDK24, SMK18, SYD^{+}16, SNH^{+}13, TLL^{+}19, VSV15, WZL^{+}11, WSR^{+}16, WCS^{+}18, WLD^{+}23, WJLY23, WZX^{+}23, XZYL12, XKW^{+}14, ZZL15, ZZYY16, ZGHM21, ZSC^{+}23, ZP23, ZYD24b, ZLD14, ZL15, ZZY^{+}21, ZLDL23, ZH16, ZZZ^{+}20b, ZLW^{+}21a, BFPP07, CSZ^{+}09, DG10, GG08, HGV^{+}08, TC09, WND^{+}09, YLHY20]. Clustering-based [DA19]. Clustering-Oriented [LZSY24]. Clusterings [MYPB20, CFD10]. Clusters [ZHSL23, GG08, KUU10]. Clusterwise [LBT^{+}23]. CNN [PSYJ24]. Co [CGZW16, DCF^{+}21, LOM22, LWH^{+}24, WTD24, WMZL24, YLHY20, ZFH23, ZH16, DG10]. Co-Clustering [ZH16, YLHY20, DG10]. Co-contrastive [WMZL24]. Co-Embedding [LOM22]. Co-evolving [DCF^{+}21]. Co-Location [WTD24]. Co-Occurrence [ZFH23]. Co-Regularized [CGZW16]. Co-Training-Teaching [LWH^{+}24]. Coalitions [LLW^{+}21]. CoBjeason [RQM^{+}24]. CoCoS [SLO^{+}21]. Code [LWW^{+}21, TZR24]. Codes [WLC^{+}17]. Coding [BLP21]. Coevolution [GBW^{+}24]. Coevolving [AK15]. Cold [CC19, LYTZ22, LGZ^{+}21, ML15, WZL^{+}23a]. Cold-Start [LYTZ22, LGZ^{+}21, ML15, WZL^{+}23a]. Collaboration [RQM^{+}24]. collaborations [AMIL13]. Collaborative [AFW^{+}21, CTX^{+}17, Dor19, HDQ^{+}18, LLC^{+}21, LSZ^{+}24, LHZ^{+}24b, LWG^{+}24, MYC^{+}24, NK20, SZD^{+}23, YBTL24, ZLC20, Kor10]. Collaborator [WXW^{+}21]. Collections [HLCR20, RBBV18]. Collective [BG07, DMI12, LTB18, NLG16]. Color [BLP21]. Combination [LL24, TYZZ10]. Combinatorial [BCC^{+}21, LGG^{+}23]. Combining [AHGA14, GWZZ17, HKR^{+}23, ZSM22]. comeNgo [ZCF^{+}17]. Comment [RYM^{+}24]. Commerce [XML18]. Commonality [GF23]. Communication [Agg22, VFA^{+}15]. Communities [BWD10, CCG^{+}24, CRGP14, CO18, LPK^{+}15, PA18, SH15, KUU10, LCZ^{+}09, MS09]. Community [BHW^{+}17, BB17, CSG^{+}16, CSX21, DA18, DHZ24, GFM21, GW20, HSBH19, LWS22, LHK^{+}18, LHW^{+}20, LYT^{+}23, LSS^{+}11, NLZH20, NYLZ24, NXZ^{+}24, PPDSBLP16, PTL22, QZL23, QZB^{+}23, RKR18, Row16, SWD^{+}21, SGCH23, SHH^{+}20, SLL^{+}22, TSS22, WLHH23, ZYC^{+}24, ASHK14, ALB09]. Community-Based [QZL23]. CommunityDiff [DA18]. Compact [YHW^{+}24]. Comparative [HLCR20, RBF^{+}21, WZLG12, WZLG13]. Comparing [PFTR16]. Comparison [LBT^{+}23, MAHT18, SWY^{+}24]. Comparisons [DYS20]. Competencies [SRVM24]. Competition [LGZ^{+}22]. Competitive [HL22, YTL18]. Competitiveness [WQZ^{+}16]. Competitor [YSLL24]. Complementation [ZLHZ24]. Complements [ZWG^{+}19]. Completion [BMZY21, LLZ^{+}23, LSW^{+}24, SGCH19, SJHY24, WWG^{+}23, WOGZ24]. Completion-Guided [LSW^{+}24]. Complex [CSG^{+}16, CSX21, HXY^{+}19, KCR23, LLH^{+}24, PAS24, DG10]. Complexity [LWZ14]. Component [CLG^{+}19]. composite [WAD12, ZFY14]. Compositional [JMR08]. Compound [LYC^{+}23]. Comprehensive [EE24, KT09]. Comprehensively [FCWQ17]. Compressed [FAV23]. Compressing [SZLP16]. Compression [GA22]. Computation [KRBK19, SKSÇ17]. Computational [HCY^{+}23, Var22, Bal13, CT14]. Computer [WZL^{+}22]. Computing [BBC^{+}19, BC24, GJ16, HAK^{+}23, LT10, YLL19, LSY^{+}09]. Con&Net [WZW^{+}22]. Concentration [Ang20]. Concept [HKR^{+}23, HSB22, LYZ^{+}24, PE20, SZWR22, WZLZ21, WWG^{+}23, WWZL23, ZGY^{+}24]. Concept-based [SZWR22]. Concept-Driven [WZLZ21]. Concept-Enhanced [WWZL23]. Conditional [PNB24, WSC^{+}17, ZXZ^{+}23]. Confidence [LHZG13, WZX^{+}23, Bal13]. Confounder [KCL^{+}20]. Confounders [YMX^{+}23]. Confounding [LWWX23]. Congestion [SYZ^{+}24b].

Congestion-aware [SYZ^{+}24b]. Connected [TWCO16, GEG^{+}08]. Connecting [SG12]. Connectivity [CCV19, CPYT21, CS24, Pap15, ZZZ^{+}20a]. Conscious [TAJY17]. Consensus [LF18, ZLDL23]. Consistent [LNG18]. Constant [SYK22]. Constrained [LYWL12, LSH^{+}22, LZSY24, LCWC20, LLLW22]. Constraint [GSG^{+}20, ZBL^{+}20]. Constraints [CCV19, CPP20, GGLP15, HYYT21, PFTR16, XXG^{+}24, GMSS13]. Constructing [KT09]. Construction [GMS21, HXY^{+}19]. Consumer [XL16]. Consumption [DWW22]. Contact [FSXL23]. contamination [KSM09]. Content [AS21, FCWQ17, GWZZ17, LWW24, TKT^{+}24]. Content-Aware [TKT^{+}24, LWW24]. Context [BSS^{+}24, CO18, FLL^{+}22, FYN22, GNMQ21, HZZ^{+}15, HCZ^{+}14, IPM12, JSF^{+}24, LWLW18, WLD^{+}21, WLL^{+}22a, XLD^{+}24]. Context-Aware [HZZ^{+}15, WLD^{+}21, WLL^{+}22a, FLL^{+}22]. Context-Based [GNMQ21]. Context-Dependent [CO18]. Contexts [ZFWC18]. Contextual [AFZ22, ACE20]. Contextualized [BSK24]. Continual [NWA20]. Continuous [KHTR18, LWC^{+}23, SBZR19, YMPH20]. Continuous-Time [KHTR18, LWC^{+}23, SBZR19]. Contractive [SHF24]. Contrast [CLJC24, Han24]. Contrastive [FJZT24, HTT^{+}24, HCH24a, JRC^{+}23, LWWW24, LLH^{+}24, LHZ^{+}24b, LJG^{+}24, XLS24, YWL^{+}24, YBTL24, ZMZS23, ZCL24, WMZL24]. Control [CWY^{+}23, YSW^{+}21]. Convergence [LLW16]. Conversation [LLD^{+}24]. Conversation-Branch-Tweet [LLD^{+}24]. Conversational [LZL^{+}24]. Convex [ACK^{+}22, CNY^{+}16]. Convolution [CTC^{+}22, Dor21, KZM^{+}23, WN22a, ZZY24a, ZMZS23]. Convolution-based [Dor21]. Convolutional [CLT^{+}20, CFX^{+}24, Jia24, JZX^{+}23, LXC^{+}22, LFY^{+}23, LXR^{+}23, OJW^{+}23, SCCM22, SYZ^{+}24b, WZWC23, XLL^{+}21, XLS24, ZHLZ24, ZLHZ24, ZLW21b]. Convolutions [LLC^{+}22]. Cooperative [GBW^{+}24]. Coordinate [BNY20]. Coordinated [WSM^{+}18]. Coordination [ABSP^{+}18]. Core [CCG^{+}24, GBGL20, LZD16, SI21, SCS20, WLX^{+}23, XYGC21, GCB^{+}21]. Core-based [CCG^{+}24]. Coreset [CPP20]. Coreset-Based [CPP20]. Coronavirus [JLZ^{+}22]. Corpus [MN20, SZWR22, II08]. corpus-based [II08]. Corpus-level [SZWR22]. correct [BG09]. Corrected [Cos21]. Correction [LJX23, LSH20]. Correlate [MFRJ16]. Correlated [CLG^{+}19, CBLH12]. Correlation [BGG^{+}15, DSL^{+}14, GFL^{+}24, KCP^{+}24, WLC^{+}17, WDL^{+}23, WMW^{+}22, ZZY^{+}23, ZSM22, KKZ09, WXT^{+}24]. Correlation-aware [KCP^{+}24]. Correlations [CNZ^{+}17, SFP10, ZZXY21, SJR08]. Corresponding [HST^{+}23]. Corruption [ZLZ^{+}19]. Cost [CWW^{+}24, HO14, SI21, CRST09, Vad10]. Cost-Effective [CWW^{+}24, CRST09]. Cost-Sensitive [HO14, SI21, Vad10]. COTE [LTB18]. Counterfactual [XLD^{+}24, YWK^{+}22]. Counting [AKM20, BCK^{+}18, BLP21, DERU17, LJK18, SOK^{+}20, SLO^{+}21, DTU21, YSY^{+}22, BBCG10, CX10]. Counts [CL18, JSP15]. Coupled [EC20, GCG22, MHS20, RABR23, RGL^{+}23, WCS^{+}18]. CoupledGT [RGL^{+}23]. Course [SMY^{+}24]. Covariance [GJ16]. Cover [XZ23, ZELS24]. Covered [RQM^{+}24]. covering [WAD12]. COVID [LLL^{+}23b, ZHSL23, BBB^{+}22]. COVID-19 [LLL^{+}23b, ZHSL23, BBB^{+}22]. CP [WJLL21]. CPU [YLL19]. Crawling [LWS22]. criteria [HGV^{+}08]. Criterion [LWG^{+}24]. Criterion-based [LWG^{+}24]. Criticality [LWB^{+}22]. Critique [ABDM21]. Cross [BGJV12, CTX^{+}17, GLF^{+}22, GFL^{+}24, GLW^{+}23, HTT^{+}24, HCH24a, HGS^{+}24, LYN^{+}24, LHG^{+}24, MRC^{+}22, ML15, OGT^{+}21, WNH15, WZW^{+}22, WCHH23, WXT^{+}24, YLHY20, YHZ^{+}24, ZWYW24, ZSM22, JP09]. Cross-aware [GLW^{+}23]. Cross-Correlation [ZSM22]. Cross-Dependency [CTX^{+}17]. Cross-Domain [HCH24a, LYN^{+}24, MRC^{+}22, ML15, OGT^{+}21, GLF^{+}22, WXT^{+}24]. Cross-Event [HGS^{+}24]. cross-graph [JP09]. Cross-Guided [BGJV12]. Cross-Language [WNH15]. Cross-Lingual [ZWYW24]. Cross-modal [LHG^{+}24]. Cross-Network [WZW^{+}22, YHZ^{+}24]. Cross-session [WCHH23]. Cross-View [GFL^{+}24, HTT^{+}24, YLHY20]. Crowd [CMQ19, ELH23, FLL^{+}22, WGL^{+}23, XLL^{+}21, YGL^{+}22, ZZXY21]. Crowd-Machine [CMQ19]. Crowd-powered [YGL^{+}22]. CrowdAtlas [ELH23]. Crowds [WYX^{+}16]. Crowdsourced [DYS20, RYM^{+}24]. Crowdsourcing [LJX23, MPG^{+}23, TYW^{+}21, WMW^{+}22, WZX^{+}23, WZB^{+}23, YZW^{+}24, ZYH19]. CrowdTC [YGL^{+}22]. CrowdWT [TYW^{+}21]. Cryptocurrency [CZW^{+}24]. CSNL [Vad10]. CT [LSZ^{+}19]. Cube [KN18]. Cuckoo [PKSM21]. Cumulative [WJLL21]. Cup [LGZ^{+}22]. Curiosity [FN24]. Curriculum [ZCL24]. Customer [BJ21, LWW18]. Customizable [SGM^{+}23]. Customization [WZL^{+}23a]. Cutting [MSR^{+}24]. Cyber [TYG^{+}15]. Cyber-Physical [TYG^{+}15]. Cyberbullying [YLL^{+}21a]. Cybersecurity [YYS^{+}24]. Cyclic [LZD16].

DACHA [CTC^{+}22]. Daily [ZZXY21].

Damage [ASZ21]. Data

[AEEBEE20, AOEM17, AH24, ABSP^{+}18, ABW20, ASPT21, AT17, AF13, DMGC21, BMTT18, BLW14, BC18, CYY^{+}22, CMZS15, CG23, CQH^{+}24, CKC^{+}18, CTZ16, DWW22, DSL^{+}14, DGB16, EGT14, EC20, FAV23, GBW^{+}24, GLMW22, GZSMM24, GZXF16, GSWJ20, HKR^{+}23, HZL20, HH19, HSB22, HQWW22, HXY^{+}19, HML^{+}24, IPM12, JCC23, KN18, KRBK19, KFI^{+}23, KZW^{+}22, KBR^{+}16, KCP^{+}24, KL23, LYG^{+}19, LYTZ22, LSW^{+}24, LK15, LGF10, LCN14, LCLL17, LZX22, LGZ^{+}22, LILJ24, MFHL10, VDMC22, PC20, QSST18, RET^{+}23, RCM^{+}13, RGL^{+}23, RKC19, RPT10, SATK20, SBRE14, SMS22, SGCH19, SB21a, SLTA11, SHF24, SM23, TL14, TYG^{+}15, TAJY17, TWC22, TWCO16, THB18, Var22, VSV15, WCS^{+}18, WLW^{+}19, WDDB20, WGC^{+}23, WFW^{+}11, WHMY17, WHC^{+}23, XML18, XKH^{+}16, XZSY19, XSZY20, YHLC23, YWC24, YSLL24, YHCL12, YWDP16, YTW^{+}16, YYD22, YZL^{+}22, ZZG^{+}21, ZJL^{+}14, ZP15, ZHT20, ZDY^{+}22, ZP23]. Data [ZYD^{+}24a, ZH16, ZWSD24, ZJJ^{+}24, vLCV^{+}18, vLCV^{+}19, BG07, BC24, CCC09, CKMS08, DG10, GEG^{+}08, GMMT07, KRPS12, KKZ09, MVT12, MWF08, TC09, VCKP08, WND^{+}09, ZLT09, Agg22]. Data-Aware [ZP15]. Data-Driven [XML18]. Database [GZXF16, MFRJ16]. Databases [Lap20, NLA23, CRST09]. Dataset [PKSM21, SDS18, WYG^{+}17]. Datasets [LLW16, OGAB14, PGR18, PNB24, TSRK20, WNC^{+}18, AF09, JMR08, SSK^{+}10]. De-anonymizing [CGL18]. De-Biasing [LWWX23]. Deadlines [HYYT21]. Dealing [NGMQ22]. Death [MHKG19]. Debiased [CWL^{+}24a]. Decentralized [MSC^{+}19]. Deciphering [GXZ^{+}22]. Decision [GXZ^{+}22, SI21, SYZ^{+}24a, ZZYW22, Vad10, VCKP08]. Decision-Making [GXZ^{+}22]. Decomposable [LZD16]. Decomposition [DSTA22, GBGL20, GCB^{+}21, JK23, LYTZ22, ME11, PFS15, SCS20, Tat19, WJLL21, XWZ^{+}22, YWK^{+}22, ZZ23, ZSC^{+}23]. Decompositions [SCS14]. Deconfounding [LYG^{+}24]. Deep [ACE20, CGL^{+}23, CLG^{+}19, CS22, CRM^{+}23, CE24, CSHZ21, DWD^{+}24, FFZ^{+}24, FSXL23, GA22, HZL20, HGW^{+}21, JKP^{+}21, JCC23, LL19, LXC^{+}22, LLL^{+}23b, LWW^{+}21, LSH20, LGS^{+}23, LZL^{+}24, LWW^{+}23b, MWL^{+}22, NC22, SCCM22, WLD^{+}23, WJLY23, XWZ^{+}22, YYC^{+}21, YHZ^{+}24, ZCC^{+}24, ZYD24b]. DeepCPR [LZL^{+}24]. DeepDepict [HGW^{+}21]. DeepMeshCity [ZCC^{+}24]. Defined [HQYY14, CRST09]. Definitions [ZTT^{+}20]. Degree [AEEBEE20, NK21, WQZ^{+}16]. Deletion [YLW^{+}19]. Deliveries [CZZ^{+}23]. Delivery [LYG^{+}19]. DeltaCon [KSV^{+}16]. DeltaShield [VLK^{+}23]. Demand [LYG^{+}19, LYC^{+}23, WZR^{+}23, YYC^{+}21]. Demands [WLW^{+}19]. Demarcating [KSB^{+}21]. Dense [FLC23, JK23, RTG17, SHF18, GG08]. Densification [LKF07]. Density [AH24, BBC^{+}24, BLW14, CMZS15, hDRNlD21, HYYQ15, Tat19, THB18, WND^{+}09, TC09]. Density-Aware [HYYQ15]. Density-based [WND^{+}09]. Density-Friendly [Tat19]. Density-peaks [AH24]. Dependence [GHT^{+}24, ZZ10]. Dependencies [FXG^{+}16, LZD16, ZZXY21]. Dependency [CTX^{+}17, FLL^{+}22, JSF^{+}24, LK20, LZY^{+}24a, WPW23]. Dependent [CO18]. Deploying [LH22]. Deployment [SYLC16]. Descent [BNY20, WYG^{+}17]. Description [ASPT21, HGW^{+}21, MV14]. Descriptors [HAK^{+}23]. Design [HCY^{+}23, KSB^{+}21]. Designing [LNG18]. Destination [WYG^{+}17, HYYT21, RLDL24]. Detect [CYT^{+}17, LLD^{+}24, LSGZ19, LLW^{+}21, IHS07]. Detecting [GLG^{+}22, NLA23, TC18, WWHW19, YMD^{+}24, YWC^{+}16, ZGY^{+}24]. Detection [ABSP^{+}18, ASZ21, Ang20, BHW^{+}17, BMTT18, BLH^{+}22, CMZS15, CSX21, CZW^{+}24, CFP19, DHZ24, DDL21, EE24, FLC23, GFL^{+}24, HSBH19, HBD^{+}24, HSS^{+}17, HQYY14, HGS^{+}24, HCH^{+}24b, JZS23, JYD19, JKP^{+}21, JLZ^{+}22, KG24, KBR^{+}16, LNG18, LWCZ24, LHK^{+}18, LSF18, LHW^{+}20, LSL^{+}22, LYTZ22, LLY^{+}24, LZV24, LTZ12, LCN14, LLL^{+}21, MCSZ20, MN20, NLZH20, NYLZ24, NXZ^{+}24, PE20, PLS^{+}21, PPDSBLP16, PTL22, QZB^{+}23, RKR18, Row16, SBRE14, SFDW19, SB21a, SGCH23, SHH^{+}20, TSS22, VLK^{+}23, WWW^{+}16, WCL^{+}23a, WZZ^{+}24, WJR^{+}10, XZXW18, XLT^{+}20, YZH^{+}18, YWW^{+}21, YHL15, ZLD^{+}23, ZYC^{+}24, ZZY24a, ZCL24, CSF^{+}12, KRPS12, ZP09]. Detector [WZZ^{+}24]. Determinantal [QXBT16]. Development [DDL24, Row16]. Developmental [YCJK08]. DEWP [FFZ^{+}24]. DexDeepFM [CS22]. DGCN [XLL^{+}21]. Diagnosis [CNZ^{+}17, THB18, WLC^{+}17, XKH^{+}16]. Diagram [WZL^{+}22]. Dialogue [WGL^{+}23]. Diameter [MW24]. diameters [LKF07]. Dictionary [SZF^{+}23, SKW^{+}21, ZZG^{+}21]. Dictionary-based [ZZG^{+}21]. Diff [CMS24]. Diff-Vectors [CMS24]. Different [ADK^{+}16, CMS24, WYWY19]. Different-Sized [ADK^{+}16]. Differentiable [CFX^{+}24]. Differential [HF24]. Differentially [EC20, KFI^{+}23]. Differentiated [KCL^{+}20]. Differentiating [CC19]. Difficulty [ZWSD24]. Diffuse [TLG^{+}23]. Diffusion [GRLK12, RKR18, RABR23, SGM^{+}23, THRR16, WCL19]. Digger [LSGZ19]. diGRASS [ZZF24]. Dilated [LLC^{+}22]. DiMBERT [LWG^{+}22]. Dimension [MG24]. Dimensional [GLZ^{+}21, MFHL10, PGR18, SHF24, WHMY17, XKH^{+}16, KKZ09, PMC22, XLL^{+}21]. Dimensionality [CCT^{+}23, CV24, GNMQ21, MDV11, ZWB22, PSFV13, ZZ10]. DipMeans [MYPB20]. Directed [BGC14, MRTW19, ZZF24]. Directional [LZP20, YBTL24]. Dirichlet [GL22, IAB22, ZB20]. Disaggregation [JKP^{+}21]. Disambiguation [PMC22, ZWB22, TS09]. disclosure [LNR08]. Discounts [XML18]. Discover [PZJ^{+}24]. Discovering [AASL23, BWD10, BGC14, DWD^{+}20, HGWY23, ML14, PA18, SMA^{+}08, SG12, SDS18, WSC^{+}17, WTD24, YWH^{+}15, ZJL^{+}14, CBLH12]. Discovery [Agg22, CYOL16, CZW^{+}24, GGLP15, GFL^{+}24, LSS^{+}11, LZD16, PLS^{+}21, RU14, SFP10, SLW^{+}18, WZW^{+}22, WV14, WSM^{+}18, ZWC^{+}23, ZPSYY10, RPT10, SSK^{+}10]. Discrepancy [GFL^{+}24]. Discrete [LLL^{+}23b, LWW24]. Discriminant [BXF^{+}20, GCL^{+}19b, GSG^{+}20, LLY^{+}21, LW24, NWW^{+}20, WLP18, ZWB22, LTN^{+}08]. discrimination [RPT10]. Discriminative [HF12, PE20, WZLG13, WY15, WZLG12]. Discussions [MCS^{+}18]. Disease [LLL^{+}20, LCG^{+}18, WLC^{+}17]. Disentangled [LHZ^{+}23, LWG^{+}22, QWH^{+}24, WZL^{+}23c]. Disentanglement [SGC^{+}23]. Disorder [ASJ^{+}23]. Dispatching [YYC^{+}21]. Display [OCD^{+}24]. Dissimilarity [IPM12]. Distance [ABS19, HZW^{+}15, HML^{+}20, IPM12, JK22, LFL^{+}23, ZLW^{+}21a, AF09]. distance-based [AF09]. Distance-Preserving [LFL^{+}23]. Distances [Cos22, BC24]. Distant [ZHLZ24]. Distillation [LXC^{+}22, MYC^{+}24, XWZ^{+}22, ZLZ^{+}24]. Distinct [VFA^{+}15]. Distributed [Che18, CZH18, DHZ24, GJ16, GBW^{+}24, GL22, KRBK19, LZZ^{+}22a, MM12, MFHL10, PSP^{+}18, SLO^{+}21, XW22, YSY^{+}22]. Distribution [CWF^{+}13, CRC^{+}23, ELH23, LWB^{+}24, RYM^{+}24, SLWW24, ZB20, ZWYZ24, SMW^{+}24, PAS24]. Distributional [XXG^{+}24]. Distributions [CZY^{+}22, JYD19, NK21]. Disturbance [WLS^{+}24]. DiVA [SGM^{+}23]. Diverse [HKR^{+}23, WGL^{+}23, ZWYW24]. Diversification [DA19]. Diversity [CPP20, CS22, HSB22, LF23, WDL^{+}21, ZS24, ZYC^{+}24, MKGV07]. Diversity-enhanced [WDL^{+}21]. Diversity-Induced [LF23]. Divisibility [CXL^{+}24]. DMGF [LJL^{+}23]. DMGF-Net [LJL^{+}23]. DNformer [JYH^{+}23]. Do [LWZ14, MFRJ16, PFC^{+}24]. Document [CKC^{+}18, HLCR20, JRC^{+}23, LGF10, LHG^{+}24, NTNP18, QCD^{+}19, SZWR22, WZL^{+}11, WZLG13, WSR^{+}16, WZLG12]. Documents [GCG22, XRGZ24, THD^{+}08]. Doing [MFRJ16]. DOLPHIN [AF09]. Domain [AFZ22, CWL^{+}24b, CGZW16, DSL^{+}24, GLW^{+}23, HBD^{+}24, HCH24a, JXGZ19, LYN^{+}24, LCZ^{+}24, MRC^{+}22, ML15, OGT^{+}21, RET^{+}23, RYM^{+}24, SLWW24, WLH24, WN22b, YMX^{+}23, CSF^{+}12, GLF^{+}22, WXT^{+}24]. Domain-Aware [SLWW24]. Domain-Specific [AFZ22]. Domains [CRC^{+}23, LV18, ZGQ^{+}24]. domination [ZLT09]. Dots [SG12]. DP [CS24]. DP-GCN [CS24]. DQN [LYGG22]. Drift [HKR^{+}23, HSB22, LYZ^{+}24, OLL20, PE20, YMD^{+}24]. Drifting [RKC19]. Driven [AT17, BMTT18, PPDSBLP16, WZLZ21, WJLY23, WZB^{+}23, XML18, YMX^{+}23, HGWY23, ZZY^{+}24b]. Drosophila [YCJK08]. Drug [AEEBEE20]. Dual [CTC^{+}22, CSHZ21, GLW^{+}23, HTT^{+}24, LSH^{+}22, LSL^{+}22, LZY^{+}24b, WNH15, WZL^{+}23a, WXT^{+}24, WLS^{+}23, WZL^{+}16, XWW^{+}24, ZLT^{+}23, ZL23]. Dual-aware [GLW^{+}23]. Dual-Embedding [CSHZ21]. Dual-Level [LSH^{+}22]. Dual-MGAN [LSL^{+}22]. Dual-Side [LZY^{+}24b]. DuCape [ZLT^{+}23]. Duration [YMD^{+}24, ZCS10]. During [JLZ^{+}22]. DWE [JXGZ19]. DWE-Med [JXGZ19]. Dynamic [AK15, AH24, CLJC24, DWH^{+}23, DERU17, EE24, HYYT21, HCH^{+}24b, JXGZ19, JYH^{+}23, JZX^{+}23, KG24, LNG18, LFY^{+}23, LJL^{+}23, LWC^{+}23, PZJ^{+}24, QZL23, RCM^{+}13, RBBV18, RU18, RTG17, SBZR19, SOK^{+}20, SCS20, THRR16, WJLL21, WC15, WWZ^{+}24, XJW^{+}21, XLL^{+}21, YZZ22, ZCF^{+}17, ZH23, ZZY^{+}24b, ZMW^{+}24, ZZYW22, ZCL^{+}22, LCZ^{+}09]. Dynamical [CNZ^{+}17]. Dynamically [HSB22, MPG^{+}23]. Dynamics [ACK^{+}22, ABS19, IYSU12, KSB^{+}21, VFA^{+}15, WXB^{+}21, XJW^{+}21].

E-Commerce [XML18]. E-News [VAFZ19]. Early [CZW^{+}24, PNB24, CSF^{+}12]. Edge [BLH^{+}22, CTP^{+}16, HAK^{+}23, HLH^{+}24, LHZ^{+}23, Pap15, WWW^{+}20, WLS^{+}24]. Edge-based [WWW^{+}20]. Edge-enhanced [LHZ^{+}23]. Edge-Type [WLS^{+}24]. Edge2vec [WWW^{+}20]. Edges [HSY^{+}21, MSR^{+}24]. Editor [Agg22]. Editor-in-Chief [Agg22]. Editorial [EK12, TWCO16, vLCV^{+}18]. Editors [vLCV^{+}19]. EEG [DPB^{+}20, YSW^{+}21]. EffCause [PZJ^{+}24]. Effect [HML^{+}24, KCL^{+}20]. Effective [CWW^{+}24, LWW23a, WLX^{+}23, WLT19, XRGZ24, YZW^{+}24, CRST09]. Effects [GXZ^{+}22, WGC^{+}23, ZBX^{+}24, VALF12]. Efficiency [QZB^{+}23]. Efficient [AH24, BHW^{+}17, BNY20, BBCG10, CGL^{+}23, CG23, CWY^{+}23, DDL24, DWD^{+}24, HWZ^{+}24, JSP15, LWS22, LYWL12, LSL^{+}22, LJL^{+}23, LJK18, LWY16, LLX^{+}24, MM12, MN20, OJLD22, RU14, SKM^{+}22, DTU21, TLZ^{+}20, WDDB20, WLX^{+}23, WV14, YLK13, YZW^{+}24, YZH^{+}18, YWW^{+}21, YLL^{+}24, ZZW09, ZHL^{+}21, ZYZZ21, ZLY^{+}20a, AF09, CX10, HAKU^{+}08, WC12, YZL^{+}22]. Efficiently [PZJ^{+}24, WLR^{+}14, ZSM22]. EFMVFL [HWZ^{+}24]. Ego [ZMC^{+}24, ML14]. Ego-graphs [ZMC^{+}24]. Eigen [CTP^{+}16]. Eigen-Optimization [CTP^{+}16]. Elastic [Dor21]. Electric [WZR^{+}23]. elements [SSK^{+}10]. Eliminating [LWB^{+}24].

Embedding [AFW^{+}21, CSHZ21, Dor21, FAV23, GFM21, HLZH18, HSY^{+}21, JYY^{+}21, JBGW24, KCR^{+}24, LHW^{+}21, LTN^{+}21, LOM22, LCZ^{+}24, LLL^{+}21, LFL^{+}23, PZW^{+}18, QZL23, QZB^{+}23, RBF^{+}21, SYXW22, SWD^{+}21, SHH^{+}20, TLZ^{+}20, WWW^{+}20, WZW^{+}22, WLX^{+}23, WN22b, WZB^{+}23, XSZ^{+}22, YHZ^{+}24, ZWG^{+}19, ZGHM21, ZYZZ21, ZWLS22, ZLT^{+}23, ZLSC24]. Embeddings [GLF^{+}22, JXGZ19, JHJK22, RJK^{+}20, WMZL24]. Emerging [KP18, NWA20, YDS^{+}15]. Emerging-Pattern [YDS^{+}15]. EML [HGS^{+}24]. Emotion [CRC^{+}23, HGS^{+}24]. Emotion-Aware [HGS^{+}24]. Employing [LWLW18, ZHL^{+}21]. Enabled [ZWB22]. Enabling [HGW^{+}21]. Encoder [DWD^{+}24, ZJ24]. Encoders [ZLW21b]. Encoding [GSI19]. End [NWA20, HDC07]. End-to-End [NWA20]. Endogenous [KSB^{+}21]. Energy [DWW22, JKP^{+}21]. Enhance [BLW14]. Enhanced [AEEBEE20, CS22, LXR^{+}23, LL24, SYZB24, SMY^{+}24, WZZ^{+}22, WWZL23, WWZ^{+}24, YBTL24, GXL24, LHZ^{+}23, WDL^{+}21, XLQ^{+}24]. Enhancement [YSW^{+}21]. Enhancements [LLY^{+}21]. Enhancing [MDV11, SLK^{+}24, SMW^{+}24, XML18]. Ensemble [BC18, CS22, GBBC24, JV20, LQW^{+}18, LLC^{+}21, LHL^{+}22, NGMQ22, TLL^{+}19, WCS^{+}18, WHC^{+}23, ZSC^{+}23, ZLD14]. Ensembles [AHGA14, ABWI23, HSB22, LTB18, NGB18, PC20, RA16, DAR09]. Entangled [LLH^{+}24]. Enterprise [YWZ^{+}24]. entire [YCJK08]. Entities [KCR23]. Entity [AKM18, BG21, BIPR13, HCC^{+}18, KCR23, WTZL22, ZWYW24, BG07]. Entity-Based [HCC^{+}18]. Entropy [AZBW21, WSZ12, WSM^{+}18]. Enumerating [PSP^{+}18]. Environment [YMD^{+}24, ZMW^{+}24]. Environmental [HQWW22]. Environments [YSW^{+}21]. Epidemic [FSXL23]. Epistemic [BAC23]. Equivalence [JZS23]. Era [TWCO16]. Error [AKM17, LCLL21, LSH20, MGL^{+}20]. Error-Tolerant [MGL^{+}20]. Errors [IKK19, WJLL21, LCZ07]. Establishing [DWW22]. Estate [FXG^{+}16]. Estimate [Cos22]. Estimates [CMZS15]. Estimating [ELH23, JSP15, LLS^{+}21, PV24, WLR^{+}14]. Estimation [BLW14, HML^{+}24, KCL^{+}20, LYG^{+}19, NK21, Var22, WSDL19, WCZ^{+}22, YWC24, ZLZ^{+}19, ZZY^{+}23]. Estimations [MM12]. Estimators [THB18]. Euclidean [AH24, CV24, Cos22]. Evaluating [JHJK22, MHKG19]. Evaluation [GCL^{+}19a, GNMQ21, LQW^{+}18, MCSZ20, WCX24, XRGZ24]. Evaluator [GLG^{+}22]. Evasion [LV18]. Evasion-Robust [LV18]. Event [ABSP^{+}18, CFP19, HGS^{+}24, LOW^{+}21, OLL20, PZW^{+}18, PLS^{+}21, WLP18, ZGY^{+}22, APU09, KT09, VFA^{+}15]. event-based [APU09]. Event-Oriented [PZW^{+}18]. Events [HW21, LHN^{+}20, IHS07]. Evidence [DKSK22, Lap20, LL24]. Evolution [LGZ^{+}23, PLS^{+}21, THB18, ZCF^{+}17, ZGY^{+}24, LKF07]. Evolutionary [JV20, WWHW19, XZYL12, APU09, CSZ^{+}09]. evolutions [LCZ^{+}09]. Evolving [BC18, GBBC24, NADR21, WWHW19, ZFH23, CBLH12, DCF^{+}21]. Exact [LGF10]. examples [MS09]. Exclusive [YWZ^{+}24]. exhibition [HDC07]. Existing [QHW22]. Exogenous [KSB^{+}21]. Expanding [MS09]. Expansion [FFZ^{+}24]. Experimental [KSB^{+}21]. Experiments [GNMQ21]. Expert [DWD^{+}20, DA19, GFM21, HHZ^{+}18, LYT^{+}23, MSC^{+}19, MFRJ16, WLX^{+}23]. Explainability [KL23]. Explainability-Based [KL23]. Explainable [GXZ^{+}22, HL22, LZV24, WZL^{+}23c, ZH23]. Explainer [SGM^{+}24]. Explanations [SZWR22, SFDW19]. Explanatory [SATK20]. Explicit [KCP^{+}24, LYG^{+}24, XYW^{+}20]. Exploiting [BLW14, FCWQ17, HLL^{+}20, LYZ^{+}24, LLD^{+}24, NTNP18, PSFV13, PMC22, WMW^{+}22, XJW^{+}21, ZPC^{+}16]. Exploration [KN18, PA18, RBBV18, SATK20, ZZY^{+}21, vLCV^{+}18, vLCV^{+}19]. Exploratory [LZF^{+}15]. Exploring [DCF^{+}21, HSY^{+}21, LWC^{+}23, SMK18, WCS^{+}18, XYW^{+}20, YYC^{+}21, YSW^{+}21, ZWSD24]. Exponent [PAS24]. Exponential [LCLL21]. Exposure [CWL^{+}24a, ZS24]. Express [CZZ^{+}23, LYG^{+}19]. expression [YCJK08]. Extension [ZZYY16]. Extension-Based [ZZYY16]. Extensions [CSSP15, LML^{+}16]. Exterior [HCY^{+}23]. External [DKSK22]. Extracting [LHN^{+}20, KNV07]. Extraction [AFZ22, AS21, FYN22, GFHL23, LWLW18, LWH^{+}23, NC22, YWR^{+}19, ZGC18]. Extrapolation [LCZ07]. Extrema [ZZY24a]. Extreme [CS22]. Extremely [LZZ^{+}22a]. Eye [ASJ^{+}23]. Eye-gaze [ASJ^{+}23].

Face [HSS^{+}17]. Facelift [HCY^{+}23]. Factor [AF16, CSHZ21, HL22, Kor10, LLZ^{+}21, RWD23]. Factor-Bounded [LLZ^{+}21]. Factorization [BNY20, CZH18, CCLZ18, CS22, DSTA22, EC20, GXZ^{+}22, HNHD14, LCLL21, LSS^{+}11, LCF19, LLZ^{+}21, MV14, PKH^{+}17, RABR23, SI24, SGCH23, WDL^{+}21, WLD^{+}23, WLP18, ZHT20, ZGHM21]. Factorizations [DKA11]. Factors [CRM^{+}23, FCWQ17, HDQ^{+}18]. Failure [PSYJ24]. Fair [GZSMM24, LXZ^{+}24, LZY^{+}24b, SYZ^{+}24a, WSW24]. FairGAT [KS24]. Fairness [BSS^{+}24, CRP^{+}24, JCS^{+}23, KS24, WZLZ23, WCX24, ZMW^{+}24]. Fairness-Aware [CRP^{+}24, KS24]. Fake [DLZ^{+}24]. False [HGS^{+}24]. Families [PCVR22]. Far [WCZ^{+}24]. farms [ZP09]. Fast [AKM20, CPYT21, FSK09, JS21, KN18, KRBK19, MLM21, ME11, QXBT16, SFP10, SKSÇ17, SHF18, SOK^{+}20, SLO^{+}21, WNH15, WCL^{+}23a, WJR^{+}10, ZZG^{+}21, ZTT^{+}20, ZELS24]. Fast-But-Approximate [KN18]. Faster [BBC^{+}19, BLP21, DWD^{+}24, ZYD^{+}24a]. Fastest [SYZ^{+}24b]. FastHAND [YWZ^{+}24]. FastHGNN [LN24]. fatigue [CSF^{+}12]. Fault [CNZ^{+}17]. Feature [BJ21, CNY^{+}16, CCC09, GQH22, GFHL23, GSWJ20, GXZ^{+}22, JLD^{+}19, LZSY24, LHL^{+}22, LXZ^{+}24, PKSM21, SZD^{+}23, SFDW19, SYG^{+}24, TL14, THYL24, WYWY19, WHMY17, XWW^{+}24, XXZ19, YYF^{+}16, YWDP16, YLL21b, YYD22, YLL^{+}24, ZLLW23, ZZY^{+}23, ZGY^{+}22, ZH21, ZZYW22, ZGY^{+}24, YLK13]. Feature-Based [BJ21]. Feature-preserved [CCC09]. Features [BAMK18, GCL^{+}19b, HKR^{+}23, LWZ14, RLDL24, YDS^{+}15, ZFW22]. Featuring [YWC^{+}16]. FedEgo [ZMC^{+}24]. Federated [CWY^{+}23, HWZ^{+}24, MYC^{+}24, THYL24, WSW24, YHLC23, YCW^{+}24, ZMC^{+}24, ZFL^{+}24]. Feedback [DWD^{+}20, HDQ^{+}18, LYG^{+}24, XYW^{+}20]. FETILDA [XRGZ24]. Few [DWL^{+}24, LSL^{+}22, RYM^{+}24, WDH^{+}24, ZFL^{+}24]. Few-Label [ZFL^{+}24]. Few-Shot [DWL^{+}24, RYM^{+}24, WDH^{+}24]. Fiber [HW21]. Field [JLH^{+}13, TLG^{+}23]. fields [WC12]. FiFrauD [KG24]. Fight [YLL^{+}21a]. Filter [PXW^{+}22]. Filtering [CTX^{+}17, HDQ^{+}18, LLC^{+}21, LSZ^{+}24, LZY^{+}24b, LHZ^{+}24b, LWG^{+}24, NK20, YBTL24, ZSM22, Kor10]. Filters [dVKCC11]. Finance [ZZY^{+}24b]. Financial [KG24, XRGZ24]. Finding [BBC^{+}24, CKMS08, DA19, hDRNlD21, GFM21, JPJ^{+}22, RTG17, SSL^{+}23, WLX^{+}23]. Fine [XZSY19]. Fine-Grained [XZSY19]. First [Bur21, CRGP14]. Fisher [SZF^{+}23, PSFV13]. Fit [YHCL12]. Fitness [LGS^{+}23]. FitRec [LGS^{+}23]. Fitting [CHZ23]. Five [BCK^{+}18]. Fixed [DERU17]. Flexible [CGZW16, HWZ^{+}24, SHF18, ZB20]. Flow [BHW^{+}17, CLT^{+}20, CZW^{+}24, FLL^{+}22, JCS^{+}23, JDE^{+}12, LLL^{+}23b, LOM22, LH22, MRC^{+}22, RLDL24, SZLP16, XLL^{+}21, YWG^{+}22, ZZXY21]. Flow-Based [BHW^{+}17, LOM22]. Flow-Level [JDE^{+}12]. Flows [SAS16]. fly [GBBC24]. Following [ABW20]. Forecasting [CRM^{+}23, CXWH24, CHZ23, DCF^{+}21, DSL^{+}24, DDY^{+}24, FFZ^{+}24, JSF^{+}24, LCLL21, LLC^{+}22, LLL^{+}23b, LYC^{+}23, LYGG22, LGZ^{+}23, ZGY^{+}22, VALF12]. foreground [KUU10]. Forest [BC24, MWL^{+}22, LHN^{+}20, MWL^{+}22]. Forest-distances [BC24]. Forgetting [MH22]. Form [JRC^{+}23]. Formal [Bal13, CMS23]. Formulation [GHT^{+}24, KRPS12]. Forums [BWD10]. Fourier [LWZ14]. Framework [AHGA14, ACE20, ABW20, BBB^{+}22, BLW14, GF23, HLZH18, HCY^{+}23, KZM^{+}23, LZSY24, LT10, LWH^{+}24, MM12, PFTR16, PSYJ24, RYM^{+}24, TYG^{+}15, TLZ^{+}20, WLW^{+}19, WTD24, WXT^{+}24, WZZ^{+}24, WJR^{+}10, XRGZ24, XKW^{+}14, XLT^{+}20, ZWG^{+}19, ZWYZ24, ZWWH20, ZLD14, ZYH19, ASHK14, APU09, DG10, GG08, HGV^{+}08, JTYY10]. Fraud [HSS^{+}17, KG24]. FrauDetector [YZH^{+}18]. Fraudulent [YZH^{+}18]. Free [Ang20, CZH18, MW24, XZ23, BFPP07]. Frequency [MN20, WCZ^{+}22, ZB20, ZFH23]. Frequent [NADR21, RU14, RGAÁC^{+}24, ZLT^{+}15, CX10, GMSS13, HDC07, JP09, MXC^{+}07, THD^{+}08, TVK10, ZCS10]. Friendly [Tat19]. Friendship [NK21, WLS^{+}23]. FSPMiner [RGAÁC^{+}24]. Fukunaga [PSFV13]. FulBM [ZELS24]. Fully [DERU17, SOK^{+}20, SCS20, ZELS24, ZWYZ24]. Function [JK22, KSV^{+}16, YTH18].

Function-on-Function [YTH18]. Functional [LLW^{+}21, ZHT20, ZP23]. Functions [ES15, HST^{+}23, SATK20, SYZ^{+}24a]. Fused [LCG^{+}18]. Fusion [ASJ^{+}23, CWWL24, LJL^{+}23, LHG^{+}24, LF23, WPW23, YSLL24, ZYS^{+}21]. Fusion-based [YSLL24]. Fuzzy [LLLW22, LL24, RWD23, ZZZ^{+}20b].

G [BMTT18, ZLY^{+}20a]. G-RoI [BMTT18]. G-Skyline [ZLY^{+}20a]. Game [LFXC24, SRVM24, XLT^{+}20, ZLT09]. Games [JSV^{+}15]. Gaming [JSV^{+}15]. Gamma [YMD^{+}24]. GAN [LHZ^{+}24a, ZLD^{+}23]. Gap [LYWL12, ZKCY07]. Gap-Constrained [LYWL12]. Gate [LYN^{+}24]. Gated [CLT^{+}20]. Gauge [VMR17]. Gaussian [LCLL17, SM23]. gaze [ASJ^{+}23]. GBAGC [XKW^{+}14]. GCN [CS24, XLS24]. Gender [BSK24]. Gene [KZZ^{+}20, YCJK08]. General [CPP20, HLZH18, KZM^{+}23, WLL^{+}21, XKW^{+}14, ZJL^{+}14]. Generalization [DSL^{+}24, LWB^{+}24, SMW^{+}24, SYXZ23, SLWW24, YMX^{+}23]. Generalized [Cos22, LFXC24, WLD^{+}23]. Generalizing [AF16, HZL20]. Generate [RLDL24, VFA^{+}15]. Generating [Che18, DLZ^{+}24, SB21b, WTD24, WNC^{+}18]. Generation [ALW^{+}24, Han24, HGW^{+}21, LHZ^{+}24a, QWH^{+}24, TZR24, ZLY^{+}20b, THD^{+}08]. Generation-based [Han24]. Generative [CWL^{+}24b, MHS20, PNB24, RYM^{+}24, WDL^{+}23, XZYL12, ZLD^{+}23, ZJJ^{+}24]. Generator [HGW^{+}21]. Generic [HZL20, WSZZ14, WDF22]. genome [ZZW09]. genome-wide [ZZW09]. Genuine [SB21b]. Geo [CLYC23, HCY^{+}23]. Geo-Tagged [CLYC23]. Geographic [FXG^{+}16]. Geolocation [BPW^{+}18]. Geometric [CGL^{+}23]. Geometrically [BC18]. Geometry [SRVM24]. Geospatial [RGL^{+}23]. Geospatial-temporal [RGL^{+}23]. Geotagged [BMTT18]. GGATB [CSZ^{+}21]. GGATB-LSTM [CSZ^{+}21]. GLAD [YHL15]. Global [AFZ22, CSZ^{+}21, DERU17, LHZ^{+}23, LZG^{+}24, LYF^{+}24, XXG^{+}24, ZLLW23, YLL^{+}24]. glycobiology [HAKU^{+}08]. GN [JS21]. GNN [CSM^{+}24]. Going [RZF^{+}24]. Good [LCW24]. GOOWE [BC18]. GPS [JCC23]. GPU [YLL19]. GRACE [KZM^{+}23]. Gradient [AF16, WYG^{+}17]. Gradients [XWZ^{+}22]. Grained [XZSY19, ZFWC18]. GrammarViz [SLW^{+}18]. Granger [AZBW21, LHS^{+}21]. Granularity [LHG^{+}24, LYF^{+}24]. Graph [BHW^{+}17, BMZY21, BBK19, CCV19, CWL^{+}24b, CCLZ18, CTC^{+}22, CRM^{+}23, CCTW23, CRP^{+}24, CXWH24, CE24, CFX^{+}24, CLJC24, CQH^{+}24, CGZW16, CGL18, DWH^{+}23, DWL^{+}24, DWD^{+}24, Dor21, DIB24, FH14, FJZT24, GYX^{+}22, GF23, GFHL23, GWZZ17, Han24, HAK^{+}23, HST^{+}23, HSS^{+}17, HF24, HYYT21, HLH^{+}24, JPJ^{+}22, JYY^{+}21, JCB^{+}16, Jia24, JZX^{+}23, JS21, KCR^{+}24, KZM^{+}23, KKB22, KMH^{+}24, KG24, KCR23, KS24, KSV^{+}16, KCP^{+}24, LKF07, LFY^{+}23, LJL^{+}23, LHZ^{+}23, LWC^{+}23, LYT^{+}23, LLZ^{+}23, LLL23a, LWWW24, LLL^{+}24, LZY^{+}24a, LFXC24, LSW^{+}24, LCZ^{+}24, LJK18, LLG^{+}23, LWW^{+}21, LGZ^{+}23, LXR^{+}23, LCX^{+}23, LHZ^{+}24b, LILJ24, LZLJ24, LFL^{+}23, LF23, LJG^{+}24, MLM21, MZC^{+}24, MNK18, NLG16, NLA23, OGT^{+}21, Pap15, PE20, PKH^{+}17, QHW22, RZY^{+}23, RZF^{+}24, RQM^{+}24, RBF^{+}21, SKSÇ17, SYK22, SOK^{+}20, SLO^{+}21, DTU21, SYZ^{+}24b, SSL^{+}23, SJHY24, SMY^{+}24, TLG^{+}23, Tat19, WZZ^{+}22, WDF22, WZWC23, WWG^{+}23, WCHH23, WPW23, WGL^{+}23, WDH^{+}24, WOGZ24]. Graph [WQZ^{+}16, WLS^{+}23, WPDZ21, WMW^{+}22, WZB^{+}23, WZL^{+}23c, XLL^{+}21, XLS24, XKW^{+}14, XSZ^{+}22, XXO^{+}24, YYS^{+}24, YWR^{+}19, YSY^{+}22, YZZ22, YWL^{+}24, YL24, YSLL24, YHW^{+}24, YZH^{+}18, YBTL24, ZZ23, ZGHM21, ZHL^{+}21, ZLT^{+}23, ZL23, ZYC^{+}24, ZZY24a, ZHLZ24, ZZY^{+}24b, ZZF24, ZMC^{+}24, ZLZ^{+}24, ZYD24b, ZJ24, ZZY^{+}21, ZLDL23, ZLSC24, ZZZ^{+}20a, ZCL^{+}22, ZHSL23, ZMZS23, ZCL24, ZBAG20, JP09, WMZL24, SLL^{+}22]. Graph-Based [BBK19, GYX^{+}22, HSS^{+}17, KCR23, CCTW23, GFHL23, LWWW24, ZL23]. Graph-Enhanced [WZZ^{+}22]. Graph-Level [RZF^{+}24, LJG^{+}24]. Graph-Mining [YZH^{+}18]. Graph-Regularization [ZZ23]. Graphical [LYN^{+}24]. Graphlet [CL18]. Graphlets [RAC^{+}21]. Graphs [ANK14, AAB^{+}24, AKM20, BBC^{+}19, Bur21, BGC14, CTP^{+}16, CZY11, DPDG18, EE24, EGT14, HCH^{+}24b, KTA^{+}11, LRK^{+}19, NADR21, NLA23, PCPN24, PA18, QZB^{+}23, RU18, SÖ22, SMW^{+}24, WZL^{+}15, WBL21, WLX^{+}23, WN22b, XZ23, ZFH23, ZYC^{+}24, ZWYZ24, ZWC^{+}23, ZZY^{+}21, APU09, CBLH12, KNV07, ZMC^{+}24]. GRASP [HST^{+}23]. Grassmann [HLC19, WHG^{+}18]. Greedily [CDSV16]. Greedy [WGYC21]. Grey [NGB18]. Grid [YWC24, ZCC^{+}24, TC09]. GrOD [XWZ^{+}22]. Grounded [LWG^{+}22]. Group [BSS^{+}24, CXWH24, LCG^{+}18, LLLW22, SFP10, SRBC22, SYLC16, SH15, TC18, WWLM23, XWW^{+}24, YHLC23, YHL15, ZCF^{+}17, TLZ^{+}08]. Group-Aware [CXWH24]. Group-based [YHLC23]. Group-Wise [XWW^{+}24]. Grouping [CSZ^{+}21]. Groups [LSGZ19, DMG^{+}21]. Growth [DMI12]. GT [TC18]. Guarantee [hDRNlD21]. Guarantees [BIPR13, RU14]. Guest [EK12, TWCO16, vLCV^{+}19]. Guided [BMZY21, BGJV12, KS21, LSW^{+}24, LF18, SNH^{+}13, LZG^{+}24, THD^{+}08].

HADI [KTA^{+}11]. Handwritten [MG24]. Harnessing [YJT^{+}24]. HARP [WZ21]. Hash [PSYJ24]. Hashing [CSSP15, LWW^{+}23b, MWL^{+}22, WSZ^{+}16, ZZLZ21]. Hashtag [CMBT22]. Hashtags [CMBT22]. HCBST [SKM^{+}22]. Head [WWW^{+}16, WPD^{+}24]. Head-Modifier [WWW^{+}16]. Headline [ALW^{+}24]. Health [LVH^{+}21, TWCO16]. Healthcare [KFI^{+}23, MFHL10]. Heat [HYYQ15]. heavy [CKMS08]. Heterogeneity [YYF^{+}16]. Heterogeneous [AS21, CQH^{+}24, DKSK22, DPDG18, GMS21, HLZH18, HSY^{+}21, HFL^{+}22, JYY^{+}21, JZX^{+}23, KCR^{+}24, KFI^{+}23, LLL^{+}24, LAN^{+}18, LSGZ19, LVH^{+}21, LZZ^{+}22b, LWG^{+}24, PC20, PLS^{+}21, QCD^{+}19, QZL23, RAC^{+}21, SBZR19, SNH^{+}13, TLZ^{+}20, WLX^{+}23, WLS^{+}24, WLH24, WSZ^{+}16, WWZ^{+}24, WMW^{+}22, WN22b, YTL18, YSLL24, ZGHM21, ZWLS22, ZXL22, ZYS^{+}21, ZBX^{+}24, ZL15, ZHS^{+}19, HNH^{+}13]. Heterogeneous-Length [QCD^{+}19]. Heterophily [HLH^{+}24]. Heuristic [ZLZ^{+}19]. HGV4Risk [LZG^{+}24]. Hi [RZF^{+}24]. Hi-PART [RZF^{+}24]. Hidden [FCWQ17, WLHH23, FSK09, ZCS10]. Hierarchical [CMZS15, CHY^{+}24, CRGP14, FLC23, GY15, IAB22, JPJ^{+}22, LZG^{+}24, LHZ^{+}24a, LYN^{+}24, LTB18, LWY16, LLB^{+}22, PZW^{+}18, RZF^{+}24, SLL^{+}23, SWD^{+}21, WZLZ21, WZ21, WDL^{+}21, WPW23, WGL^{+}23, WDH^{+}24, ZLHZ24, ZLD14, CKMS08]. Hierarchies [SCS14]. High [DGB16, GFL^{+}24, LDS^{+}22, MFHL10, OJLD22, PGR18, SB24, SHF24, WZL^{+}23b, WTD24, WHMY17, WLT19, XKH^{+}16, ZHL^{+}21, ZZY^{+}21, KKZ09]. High-Dimensional [MFHL10, PGR18, SHF24, XKH^{+}16, KKZ09]. High-Order [GFL^{+}24, WZL^{+}23b, ZZY^{+}21]. High-Quality [ZHL^{+}21]. High-Speed [DGB16]. High-Utility [LDS^{+}22, WLT19].

High-Value [OJLD22]. Higher

[LLL^{+}24, PMC22, ZHL^{+}21]. Higher-Order [LLL^{+}24, ZHL^{+}21]. Hilbert [CV24]. Historical [CTC^{+}22]. History [XLQ^{+}24]. History-enhanced [XLQ^{+}24]. HITS [KMH^{+}24]. HITS-based [KMH^{+}24]. hitters [CKMS08]. HIVE [LTB18]. HIVE-COTE [LTB18]. HMM [HCZ^{+}14]. Holistic [YWS^{+}22]. Homogeneity [HTT^{+}24]. Homophily [LLG^{+}23]. Homoskedastic [SM23]. hop [ZELS24]. Household [ASPT21, JKP^{+}21]. Housing [LYG^{+}19]. Human [AKM17, BTBg22, DDL24, HLL^{+}20, LTN^{+}21, SKW^{+}21, VLK^{+}23, YLHY20, YWC^{+}16, YDW^{+}23]. Human- [VLK^{+}23]. Human-in-the-loop [DDL24]. HUSP [ZYD^{+}24a]. HUSP-SP [ZYD^{+}24a]. HW [MWL^{+}22]. HW-Forest [MWL^{+}22]. Hybrid [BPW^{+}18, CMQ19, LYTZ22, LWC^{+}23, SKM^{+}22, ZWC22]. Hyperbolic [ZWLS22]. hypercubes [WAD12]. Hyperedges [WN22a, AMIL13]. Hypergraph [HTT^{+}24, JBGW24, LZL^{+}23, LN24, WN22a, LLD^{+}24]. Hypergraphs [AASL23, BBC^{+}24, CYL^{+}24, SCS20, LK20]. Hyperparameter [WSW24]. Hyperparameters [SYXZ23]. Hypothesis [LZF^{+}15].

ICU [LSS^{+}22, WLC^{+}17]. ID [CXL^{+}24]. ID-SR [CXL^{+}24]. Identification [ABSP^{+}18, HL22, KP18, NLG16, ZTL15a]. Identified [LSL^{+}22]. Identify [HKR^{+}23, SH15]. Identifying [ASPT21, ERSK14, ZS24, ZWG^{+}19]. If [GW21]. iGRM [NGB18]. iHypR [AMIL13]. IID [YHLC23]. Image [LWB^{+}24, LZL^{+}22, LWW23a, PSYJ24, RBBV18, RQM^{+}24, XCL22, YSG^{+}21, ZZG^{+}21, ZZLZ21, ZZZ^{+}20a]. Images [MG24, YCJK08]. Imbalance [BSS^{+}24, GVdSL24, RWD23, SI21, SKM^{+}22]. Imbalance-Robust [GVdSL24]. Imbalanced [CHZ23]. Impact [GSI19].

Implicit [HDQ^{+}18, KCP^{+}24, LYG^{+}24, LXR^{+}23, LWG^{+}24, SGC^{+}23]. Importance [JYY^{+}21, ZH21]. Imprecise [YWC24]. Impression [LSH^{+}22]. Improve [SYXZ23, XL16]. Improved [BBK19, BJ21, DKSK22, NGB18, YWW^{+}21]. Improvement [CGL^{+}23]. improvements [WAD12]. Improving [CSM^{+}24, CDSV16, LHZ^{+}24b, ML15, WMZL24, XJZ21, YBTL24, ZGC18, ZXL22, ZWH^{+}16]. Imputation [ABWI23, MGL^{+}20, PFC^{+}24, PNB24, YWC24]. In-Processing [WZLZ23]. Inaccurate [NGMQ22]. Incoherent [CLY12]. Income [ASPT21]. Incomplete [LSW^{+}24, LCN14, LL24, NGMQ22, ZTT^{+}20, ZJJ^{+}24]. Incorporated [LCF19]. Incorporating [DWD^{+}20, HDQ^{+}18, SJHY24, ZGC18]. Increased [ZS24]. Incremental [CSX21, GQH22, STP^{+}08, WJLL21, YZH^{+}18, ZZYY16, ZGY^{+}22]. Incrementally [CBLH12]. Independence [ZXZ^{+}23]. Indexing [MRJ11, RBBV18, XJZ21, ZHSL23]. Indirect [WSR^{+}16]. Individual [HML^{+}24, LVH^{+}21, ZBX^{+}24]. Individuality [GF23]. Induced [LF23]. Inductive [DIB24, QZB^{+}23, SJHY24]. Industrial [GHT^{+}24]. Industry [SMS22]. Infer [YTL18]. Inference [CTX^{+}17, CZH18, CMQ19, DWH^{+}23, GY15, LH22, LGF10, RABR23, SRVM24, WZX^{+}23, WZB^{+}23, XZLL21, XZSY19, YCL^{+}21, BG09]. Inferring [ABW20, GRLK12, LYL^{+}20, THRR16, SSK^{+}10]. Infinite [CXL^{+}24]. Influence [BAC23, CYY^{+}22, CDV21, GRLK12, GW20, GW21, HK18, HLL^{+}20, LGG^{+}23, LXY^{+}17, QZL23, TKT^{+}24, WLL^{+}21, WGYC21, YMPH20, YWR^{+}23, ZTL^{+}15b, ZL15]. Influenced [ZTL^{+}15b]. Influencer [LCCM19]. Influencers [SAS16]. Influences [LSZ^{+}19]. Influential [GW21, DMG^{+}21]. Infomax [SLL^{+}22].

Informatics [YYF^{+}16]. Information [ASJ^{+}23, BLW14, CKC^{+}18, DMI12, ERSK14, GWZZ17, HGW^{+}21, HML^{+}20, HLZH18, HFL^{+}22, HGS^{+}24, JCS^{+}23, LWCZ24, LWLW18, LWC^{+}23, LLL23a, LZSY24, LAN^{+}18, LK15, LLX^{+}24, MRC^{+}22, MCS^{+}18, MZC^{+}24, ML15, PLS^{+}21, QSS20, RKR18, RZY^{+}23, RLDL24, SBZR19, SGM^{+}23, SGC^{+}23, SZLP16, SH15, SAS16, SNH^{+}13, VLK^{+}23, WCL19, WCHH23, WPW23, WMW^{+}22, XW22, XL15, XKH^{+}16, YLHY20, YWH^{+}15, YDW^{+}23, ZGHM21, ZHL^{+}21, ZWLS22, ZXL22, ZWH^{+}16, ZYS^{+}21, ZL15, ZHS^{+}19, ALB09, CT14]. Information-aware [LWCZ24]. information-seeker [ALB09]. Information-Theoretic [XKH^{+}16]. Informative [WGL^{+}23, MVT12]. Informed [GLMW22, RQM^{+}24]. infrequency [GMSS13]. Initiator [ABSP^{+}18]. Inner [CO18]. Input [CSM^{+}24]. Inputs [ADK^{+}16]. Insensitive [CYY^{+}22]. Insights [LGZ^{+}22]. Inspecting [WCZ^{+}24]. Instance [BFRL13, LL24, LGT^{+}24, MYL^{+}21, SWY^{+}24]. Instrumental [YWK^{+}22, YMX^{+}23]. Insufficient [LZL^{+}22]. Integrated [GXLC21, GSWJ20]. Integrating [CGZW16, SNH^{+}13, WZL^{+}11, ZLLW23]. Integration [RYM^{+}24, ZH23]. Integrity [LOW^{+}21]. Intelligence [HL22, WGL^{+}23, ZPSYY10]. Intelligence-Based [HL22]. Intelligent [SMS22]. Intention [CZW^{+}24, YSW^{+}21]. Intentions [JCQ^{+}23]. Inter [VFA^{+}15]. Inter-event [VFA^{+}15]. Interaction [JYY^{+}21, SYG^{+}24, APU09]. Interactions [CO18, LLH^{+}24, LZZ^{+}22b]. Interactive [KN18, LL19, LHG^{+}24, MCS^{+}18, PA18, RBBV18, SGM^{+}23, SLW^{+}18, WSM^{+}18, vLCV^{+}18, vLCV^{+}19]. Interconversion [NZW23]. Interdependence [WCS^{+}18]. Interest [BMTT18, LYL^{+}20, QHW22, XYGC21, OJW^{+}23, WCHH23]. Interesting [AASL23, RV20, WV14, Web10]. Interference [HML^{+}24, ZBX^{+}24]. Internal [MCSZ20]. Internet [NGB18]. Interpretable [FAV23, JRC^{+}23, SZWR22]. Interpretation [HP20]. Intersections [HXY^{+}19]. Interval [LSS^{+}22]. Intervention [CSM^{+}24, FSXL23]. Intricate [LZY^{+}24a]. Introduction [ACPW13, Agg17, BBD^{+}07, CT14, GL15, Han07, LSY^{+}09, SLTA11, WL16, XL20, ZKYW08, vLCV^{+}19]. intrusive [SCCM22]. Invariant [GJDX14]. inverse [GMSS13]. Involving [NS23]. IoT [DMGC21]. IREOS [SLK^{+}24]. Irregularly [TR22]. Isolation [LTZ12]. Isolation-Based [LTZ12]. Iss [vLCV^{+}19]. Issue [ACPW13, Agg17, EK12, GL15, SLTA11, TWCO16, Wan10, WL16, XL20, ZPSYY10, vLCV^{+}18, vLCV^{+}19, BBD^{+}07, CT14, GSTC12, LSY^{+}09, MG09, ZKYW08]. Item [FCWQ17, GLF^{+}22, NK20, ZLC20, WCHH23, WWLM23]. Item-based [NK20]. items [Web10]. Itemset [WCZ^{+}22, WLT19, CX10, GMSS13]. Itemsets [RU14, SB24, LNR08, MVT12, Web10]. iterated [WAD12]. Iterative [JS21, SYD^{+}16]. IV [YWK^{+}22].

Joint [AFZ22, GEG^{+}08, GFHL23, GFL^{+}24, LH22, RABR23, SKW^{+}21, TYW^{+}21, ZGHM21, ZFWC18]. Jointly [LZZ^{+}22b, YYF^{+}16, ZZXY21]. Jumps [MRTW19].

Katz [LLS^{+}21]. KDD [Agg17, LGZ^{+}22, XL20]. KDD’10 [EK12]. Kernel [CSSP15, HYYQ15, LCF19, LLL^{+}21, SHL19, ZZL15, SZF^{+}23]. Kernel-Incorporated [LCF19]. Kernelized [NC22, XKH^{+}16]. kernels [MWF08]. Key [JPJ^{+}22]. Keyword [AFZ22]. Kmeans [MYPB20]. KNN [YWW^{+}21, GVdSL24, RKC19].

KNN-Based [YWW^{+}21]. Know [MFRJ16]. Knowledge [Agg22, BMZY21, BWD10, CTC^{+}22, LXC^{+}22, LZY^{+}24a, LGZ^{+}21, LZLJ24, LML^{+}16, MZC^{+}24, RQM^{+}24, RBF^{+}21, SWH^{+}23, SRVM24, SJHY24, WNH15, WSR^{+}16, WYWY19, WWG^{+}23, WGL^{+}23, WXT^{+}24, WOGZ24, WMZL24, XWZ^{+}22, YHW^{+}24, ZZYY22, ZLT^{+}23, ZLZ^{+}24, ZLHZ24, ZPSYY10, KUU10, LNR08]. Knowledge-correlation [WXT^{+}24]. Knowledge-Sharing [BWD10]. Koontz [PSFV13]. KRAN [ZZYY22]. Kronecker [MNK18]. Krylov [HSBH19].

L2MM [JCC23]. Label [AZD^{+}21, BFRL13, GQH22, GCL^{+}19b, LHW^{+}20, LHL^{+}22, LFL^{+}23, LGT^{+}24, MFS24, GVdSL24, RKC19, SZD^{+}23, WDL^{+}23, WPD^{+}24, WMW^{+}22, WZX^{+}23, WHC^{+}23, XCL22, YYF^{+}16, YZW^{+}24, ZFW22, ZWB22, ZLLW23, ZZY^{+}23, ZFL^{+}24, JTYY10, PKSM21, SCCM22, WDF22]. Label-Specific [GCL^{+}19b, LHL^{+}22]. Labeled [MA16, NLA23, YWW^{+}21, HAKU^{+}08]. Labeling [YZL^{+}22, ZELS24]. Labels [LZZ^{+}22a, QZZZ24, WPD^{+}24]. Lag [SFP10, AZBW21]. Landmark [MDV11, ZELS24]. Landmark-Based [MDV11, ZELS24]. Langevin [WXB^{+}21]. Language [BSK24, CWW^{+}24, LWG^{+}22, WNH15, WZLZ21, WCZ^{+}24]. Laplacian [HCH^{+}24b, LCG^{+}18, SHF24]. Laplacian-based [SHF24]. Large [AOEM17, BHW^{+}17, CTP^{+}16, CZH18, CPYT21, CSX21, CWW^{+}24, CZY11, CKC^{+}18, DHZ24, DKSL18, GJ16, GY15, HSBH19, HQWW22, HLZH18, JYY^{+}21, JCB^{+}16, JDE^{+}12, KTA^{+}11, KRBK19, Lap20, LK20, LLW16, LLS^{+}21, MW24, VDMC22, OGAB14, PGR18, RBBV18, SHL19, SATK20, SGM^{+}23, SDS18, SLTA11, WLR^{+}14, WNH15, WZL^{+}15, WCZ^{+}24, WHMY17, XXZ19, XJZ21, YHCL12, ZP15, ZZLZ21, AF09, BBCG10, CRST09, GMSS13, KT09]. Large-Scale [AOEM17, BHW^{+}17, CZH18, CPYT21, CKC^{+}18, DHZ24, DKSL18, GY15, HLZH18, JYY^{+}21, KRBK19, LLS^{+}21, VDMC22, SDS18, SLTA11, WNH15, WHMY17, XXZ19, ZZLZ21, SATK20, XJZ21, BBCG10, GMSS13]. Largest [SMDHT21]. Lasso [LWY16, LCG^{+}18]. Latent [CSHZ21, FCWQ17, GSST16, GL22, HDQ^{+}18, IAB22, LZL^{+}22, WSZ12, WMW^{+}22]. law [NK21, PAS24]. Layer [PSYJ24, THYL24]. Layer-Wise [THYL24]. Layered [CTX^{+}17]. Layout [WBL21]. Layout-Based [WBL21]. LDA [GXL24, LLY^{+}21, YCJK08]. Leakage [KRPS12]. Learnable [LHZ^{+}24b, ZZ23]. Learning [AHGA14, AT17, ASZ21, BJ21, BXF^{+}20, BFRL13, CMBT22, CGL^{+}23, CDC23, CNY^{+}16, CLY12, CCLZ18, CLG^{+}19, CTC^{+}22, CZY^{+}22, CZZ^{+}23, CRC^{+}23, CCTW23, CE24, CLJC24, CQH^{+}24, Con20, CFD10, DKSK22, DKSL18, DWL^{+}24, DPDG18, FFZ^{+}24, FSXL23, FJZT24, GMS21, GQH22, GF23, GZXF16, GCL^{+}19b, GHT^{+}24, GA22, HZL20, HTT^{+}24, Han24, HHL^{+}24, HHZ^{+}18, HCH24a, HZW^{+}15, HML^{+}20, HZM^{+}22, HLZH18, HGS^{+}24, HWZ^{+}24, HLH^{+}24, HF12, IPM12, IHS07, JH22, JRC^{+}23, JYY^{+}21, JKP^{+}21, JYH^{+}23, JCC23, JHC^{+}23, Jia24, JDE^{+}12, JZX^{+}23, JS21, KRBK19, LL19, LXC^{+}22, LYTZ22, LWC^{+}23, LLL23a, LWWW24, LZG^{+}24, LLL^{+}24, LZY^{+}24a, LZSY24, LSW^{+}24, LLH^{+}24, LWZ14, LHL^{+}22, LLG^{+}23, LYL^{+}22, LSZ^{+}19, LSH20, LWG^{+}22, LYGG22, LGS^{+}23, LCX^{+}23, LZY^{+}24b, LLX^{+}24, LHZ^{+}24b, LYF^{+}24, LTH^{+}13, LFL^{+}23, LN24, LGT^{+}24, LJG^{+}24, MH22, MYC^{+}24, MES22, NGMQ22, OGV22, QZZZ24, RZY^{+}23, RZF^{+}24, RLDL24, RYM^{+}24, SM21, SZD^{+}23, SWY^{+}24, SWH^{+}23]. Learning [SCCM22, SLL^{+}23, SJR08, SMW^{+}24, SKW^{+}21, SLWW24, SM23, TLG^{+}23, TWC22, TKT^{+}24, THYL24, Var22, WZLZ23, WC12, WSZZ14, WY15, WLC^{+}17, WCL19, WLD^{+}23, WDL^{+}23, WZL^{+}23b, WWG^{+}23, WJLY23, WSW24, WDH^{+}24, WZZ^{+}24, WMZL24, WWZ^{+}24, WYX^{+}16, WPDZ21, WZL^{+}23c, XW22, XWW^{+}24, XKH^{+}16, XWZ^{+}22, XL16, XXG^{+}24, XXO^{+}24, YTL18, YYC^{+}21, YGL^{+}22, YHLC23, YL24, YW24, YCW^{+}24, YWG^{+}22, YLL^{+}24, YBTL24, YSW^{+}21, ZY14, ZXL22, ZLLW23, ZL23, ZCC^{+}24, ZMC^{+}24, ZFL^{+}24, ZFWC18, ZWWH20, ZGY^{+}22, ZMW^{+}24, ZBX^{+}24, ZJ24, ZRJ^{+}22, ZLDL23, ZCZQ19, ZLC20, ZLW21b, ZWSD24, ZGQ^{+}24, ZMZS23, ZCL24, DG10, JTYY10, WZL^{+}23a, YLK13, ZY13, ZFY14, SZF^{+}23]. Learning-Based [JKP^{+}21, WSZZ14, TWC22, WZZ^{+}24]. Left [YMD^{+}24]. Left-Truncated [YMD^{+}24]. Legal [BSK24]. Length [ASPT21, MV14, QCD^{+}19, SLW^{+}18]. Less [PTL22, RA16, SG12]. Level [AS21, CLYC23, CQH^{+}24, JDE^{+}12, LSH^{+}22, RZF^{+}24, SJHY24, XJW^{+}21, LWWW24, LGS^{+}23, LJG^{+}24, SZWR22]. Leveraging [GCL^{+}19b, LCW24, LCF19, MA16, YTW^{+}16]. Lifecycle [CYOL16]. Lifelong [SWH^{+}23]. Lifetime [BJ21, LYL^{+}20, WXW^{+}21]. Lightweight [YZW^{+}24]. Likelihood [DHZ24, NK21, FSK09]. Limited [BBC^{+}24]. Linear [ASPT21, GSG^{+}20, LLY^{+}21, LW24, LBT^{+}23, NWW^{+}20, OGAB14, PBMID14, YHCL12, ZWB22, GMSS13, LCZ07, Vad10, WC12]. Linearization [LK20]. Lingual [ZWYW24]. Link [CGL^{+}23, CDV21, DKA11, JYH^{+}23, RBF^{+}21, WZW^{+}22, WLS^{+}24, WLH24, WWZ^{+}24, YLW^{+}19, YL24, ZP09]. Link-building [CGL^{+}23]. Linkage [OJLD22, dVKCC11]. Linkages [HZW^{+}15].

Linking [CHY^{+}24]. links [KSM09]. Liouville [IAB22]. LIS [LSZ^{+}19]. Listing [AKM20, CC12]. Listwise [LWW24, WYX^{+}16]. Literature [KP18]. LLMs [YJT^{+}24]. LMACL [LHZ^{+}24b]. Load [SCCM22]. Local [ACK^{+}22, AFZ22, BLW14, CWY^{+}23, CS24, CRGP14, DERU17, HSBH19, JYD19, LHK^{+}18, LJK18, LYF^{+}24, NLZH20, NYLZ24, NXZ^{+}24, NWW^{+}20, OGV22, SH15, WLC^{+}17, WCL^{+}23a, WLHH23, YLL^{+}24, ZLLW23, ZZY24a, ZZY^{+}21, BBCG10]. Local-First [CRGP14]. Local-to-global [YLL^{+}24]. Locality [CSSP15, YSG^{+}21, ZTL^{+}15b]. Location [BTBg22, FYN22, LYT^{+}23, WTD24, WLS^{+}23, XSZY20, YCC^{+}15, YTW^{+}16, ZFWC18, ZPC^{+}16]. Location-Aware [FYN22]. Location-Based [WLS^{+}23, YCC^{+}15, YTW^{+}16, ZFWC18, ZPC^{+}16]. Locations [VDMC22]. Log [YWW^{+}21]. Logs [WQZ^{+}16]. Long [FLL^{+}22, HZL20, HCC^{+}18, JRC^{+}23, JHC^{+}23, LLB^{+}22, LHG^{+}24, XRGZ24, ZZY24a, ZLHZ24]. Long- [JHC^{+}23]. Long-Form [JRC^{+}23]. Long-range [FLL^{+}22]. Long-Tail [HCC^{+}18, ZZY24a]. Long-Tailed [ZLHZ24]. Longitudinal [BTBg22, PNB24]. loop [DDL24]. LoSAC [CWY^{+}23]. Loss [ES15, HP20]. Lost [MES22]. Low [CLY12, CCT^{+}23, JCC23, LSF18, WN22b, YSG^{+}21, ZSC^{+}23]. Low-Quality [JCC23]. Low-Rank [CLY12, LSF18, WN22b, CCT^{+}23, ZSC^{+}23]. LSAB [HHL^{+}24]. LSTM [LYGG22, CSZ^{+}21].

Machine [CS22, CMQ19, JLD^{+}19, JDE^{+}12, JSG^{+}19, LCLL21, LYGG22, RWD23, SMS22, SLL^{+}23, TWC22, WZLZ23, XW22]. Machines [PBMID14, AGHN13]. Main [AS21]. Maintenance [ZELS24]. Make [LCWC20]. Making [GXZ^{+}22]. Malice [CZW^{+}24]. Managing [KFI^{+}23]. Manifold [CCT^{+}23, HQYY14, HNHD14, HF12, LWW23a]. Manifolds [HLC19, WHG^{+}18]. Manipulation [CTP^{+}16]. Manipulations [SKSÇ17]. Manufacturing [ZH16]. Many [LWZ14]. Map [AOEM17, JCC23, PSYJ24]. Mapping [CMBT22, GCL^{+}19b, ZHLZ24, ZCL19]. MapReduce [ADK^{+}16, LDS^{+}22]. Margin [GZXF16, TVK10]. Marginal [CWF^{+}13]. Marginalized [WDF22]. Marked [SYZB24]. Market [LYGG22, XL16]. Marketing [LCCM19, ZPC^{+}16]. Markov [FSK09, IHS07, LWB^{+}22, WC12, ZCS10]. Markov-modulated [IHS07]. Masked [ZJ24]. Mass [LH22]. Massive [KSV^{+}16, DTU21, WCZ^{+}22, ZZY^{+}21, CC12]. Massive-Graph [KSV^{+}16]. Matching [AOEM17, BG21, BIPR13, CWF^{+}13, CZY^{+}22, CGL18, CXH^{+}20, JRC^{+}23, JCC23, LYT^{+}23, LWW^{+}21, MAHT18, WSC^{+}17, ZHL^{+}21, TLZ^{+}20]. Math [QWH^{+}24]. Mathematical [XL15]. Matrices [DSTA22, GJ16, YWG^{+}22]. Matrix [CZH18, CXH^{+}20, DKA11, EC20, HNHD14, LCF19, LLZ^{+}21, LJW^{+}22, MV14, PKH^{+}17, RABR23, SI24, SGCH23, WNH15, WDL^{+}21, WLD^{+}23, WWLM23, ZZ23, ZGHM21, ZSC^{+}23]. Matroid [CPP20]. Max [GZXF16]. Maximal [SMDHT21, TVK10]. Maximally [CLG^{+}19]. Maximization [BAC23, CPP20, CDV21, GW20, GW21, HK18, LGG^{+}23, LWW18, QZL23, TKT^{+}24, WLL^{+}21, WWLM23, WGYC21, YMPH20, YWR^{+}23, ZLGY24, ZZ10]. Maximizing [ZS24]. Maximum [BBC^{+}24, NK21, WSZ12, WSM^{+}18]. MBN [SOL22]. McNemar’s [MAHT18]. MCRapper [PCVR22]. MCS [LWS22]. MDL4BMF [MV14]. Mean [LHZG13, LW24]. Mean-Optimal [LW24]. Means [MM12]. Measure [AEEBEE20, ABS19, AF16, Ang20, CG15, Cos22, DSL^{+}14, SRVM24, ZHS^{+}19, ZLW^{+}21a, ZWSD24]. Measurement [WXB^{+}21, ZH21]. Measures [BSS^{+}24, LK15, DD09]. Measuring [BSK24, KNV07, LWB^{+}22, WC15]. Mechanism [CC19]. Mechanisms [KS21, XCL22]. Med [JXGZ19]. Media [AEEBEE20, BMTT18, CYT^{+}17, LPK^{+}15, LLD^{+}24, THRR16, TL14, WLG^{+}24, YHL15, YTW^{+}16, ZTL15a, ZH23, ZFH23]. Medical [CSZ^{+}21, JXGZ19, LWB^{+}24, YYF^{+}16]. MEDLINE [TS09]. meet [CBRB09]. Meets [CCLZ18, GF23, ZBL^{+}20]. MEGA [WOGZ24]. Melody [ZLY^{+}20b]. Membership [CO18, PXW^{+}22]. Memo [CZH18]. Memo-Free [CZH18]. Memory [DERU17, HZL20, JSE^{+}23, LJK18, QWH^{+}24, RKC19, SYZB24, YHCL12]. Memory-Efficient [LJK18]. Memory-Enhanced [SYZB24]. Mental [LVH^{+}21]. MeSH [XJZ21]. MeSHProbeNet [XJZ21]. MeSHProbeNet-P [XJZ21]. Meso [LHW^{+}21]. Message [BGSW13]. Message-Passing [BGSW13]. Meta [DWL^{+}24, HKR^{+}23, HFL^{+}22, HGS^{+}24, LLL^{+}24, SNH^{+}13, SLWW24, THYL24, WPW23, WZL^{+}23a, WOGZ24, YHLC23, ZMW^{+}24, ZHS^{+}19]. Meta-Features [HKR^{+}23]. Meta-Graph [WOGZ24]. Meta-Information [WPW23]. Meta-Learning [DWL^{+}24, THYL24, ZMW^{+}24, WZL^{+}23a, YHLC23]. Meta-Path [HFL^{+}22, LLL^{+}24, SNH^{+}13]. Meta-Structure [ZHS^{+}19]. Metagraph [LSS^{+}11]. Method [DMGC21, CTC^{+}22, CG23, CWY^{+}23, DHZ24, FLG^{+}21, HLC19, LLC^{+}21, LZL^{+}22, LHS^{+}21, LZX22, LLX^{+}24, OJLD22, QZB^{+}23, DTU21, WSZZ14, WBL21, WZW^{+}22, WCL^{+}23a, YLL19, YWW^{+}21, YDW^{+}23, DD09]. methodologies [CT14]. methodology [SSK^{+}10]. Methods [CMS23, CSSP15, Jia24, LBT^{+}23, LWW^{+}23b, MZC^{+}24, MG24, PDK24, Var22, ZP23, ZH21, DAR09]. Metric [GJDX14, HZM^{+}22, XKH^{+}16].

Metrics [HML^{+}20]. MGAN [LSL^{+}22].

MICCF [LZSY24]. Micro

[LHW^{+}21, YWC^{+}16]. Micro-Blog [YWC^{+}16]. Micro-Meso [LHW^{+}21]. Microaggregation [GZSMM24]. microarray [LTN^{+}08]. microarray-based [LTN^{+}08]. Microblog [GCL^{+}19a, PZW^{+}18, WWZL23]. Microblogging [YWH^{+}15]. Microblogs [CMBT22]. microeconomic [ZLT09]. MIE [ABWI23]. MIE-TS [ABWI23]. miner [HDC07, WLL^{+}22b, WCL^{+}23b]. Minimization [CPYT21, LCLL21, TLL^{+}19, WHG^{+}18]. minimize [KSM09]. Minimum [ASPT21, MV14, XZ23]. Mining [AEEBEE20, AK15, BB17, CL18, CO18, GM18, GLFV^{+}19, GLZ^{+}21, GLMW22, GZXF16, GLW^{+}23, HH19, HYYT21, HGY24, JH19, JP09, JCB^{+}16, KTA^{+}11, KR16, Lap20, LYWL12, LDS^{+}22, LHS^{+}21, LLB^{+}22, LGZ^{+}23, LXR^{+}23, LWH^{+}23, LML^{+}16, MHKG19, NADR21, PCVR22, PZW^{+}18, PLL^{+}10, QSST18, RCM^{+}13, RV20, RGAÁC^{+}24, Row16, SMDHT21, SMS22, SHF18, SB24, SAS16, SLTA11, TYG^{+}15, Var22, WLC^{+}17, WDDB20, WTD24, WZL^{+}16, WLT19, WLL^{+}22b, WCL^{+}23b, YYS^{+}24, YZH^{+}18, YDS^{+}15, ZKCY07, ZLT^{+}15, ZDY^{+}22, ZYD^{+}24a, ZWH^{+}16, AF09, CX10, GMMT07, GG08, GMSS13, HAKU^{+}08, JMR08, KRPS12, RPT10, THD^{+}08, TVK10, ZCS10, ZLT09, MFRJ16]. MIP [RZY^{+}23]. MIRROR [LXR^{+}23]. misclassification [BG09]. Misconceptions [RJK^{+}20]. Misinformation [LLD^{+}24, NZW23, ZZY24a]. MiSoSouP [RV20]. Missing [BAMK18, ERSK14, MGL^{+}20, WGC^{+}23, YYD22, BC24]. Misuse [SBRE14]. Mitigating [AZD^{+}21, BSK24]. Mix [KL23]. Mix-Up [KL23]. Mixed [CZZ^{+}23, CO18, MRC^{+}22, YWL^{+}24]. Mixed-Membership [CO18]. Mixture [LLW^{+}21, YMD^{+}24, ZCL19, SJR08]. mixture-of-subsets [SJR08]. ML [LGG^{+}23]. ML-Based [LGG^{+}23]. MLP [LYF^{+}24]. MMORPG [JPJ^{+}22, WZZ^{+}24]. Mobile [AOEM17, HLL^{+}20, HCZ^{+}14, LHZ^{+}24a, LWG^{+}16, LZLJ24, OGT^{+}21, QSST18, SDS18, WSDL19, XYZ^{+}20]. Mobility [BGWSB19, HLL^{+}20, VDMC22, WYG^{+}17, WLL^{+}22a, YDW^{+}23]. Modal [WSZ^{+}16, ACE20, CWW^{+}24, LHG^{+}24, YHZ^{+}24]. Mode [CWF^{+}13, SCS14, WY15, YTH18]. Mode-Sparsity [YTH18]. Model [AZD^{+}21, ASZ21, BMZY21, CDC23, CCLZ18, CS22, CWW^{+}24, DMI12, DSTA22, DDL24, DWW22, DGB16, GXZ^{+}22, IAB22, JCQ^{+}23, JYD19, JLZ^{+}22, LZP20, LWC^{+}23, LLL^{+}23b, LOM22, LDS^{+}22, LGF10, LHZ^{+}24b, LML^{+}16, MHS20, NGB18, SLK^{+}24, SLL^{+}23, VAFZ19, WZLZ21, WZ21, WWZL23, WOGZ24, WCZ^{+}24, WJR^{+}10, YZW^{+}24, YWG^{+}22, YZL^{+}22, ZCF^{+}17, ZCC^{+}24, ZCL19, DD09, HAKU^{+}08, LCZ07, SJR08, THD^{+}08, ZCS10]. Model-Agnostic [WJR^{+}10]. Modeling [AFZ22, ALB09, CYOL16, CE24, CYT^{+}17, FLL^{+}22, FXG^{+}16, FN24, GCG22, HHL^{+}24, HLCR20, HML^{+}24, HF12, IYSU12, JSF^{+}24, JHC^{+}23, KHTR18, LLC^{+}22, LFC^{+}17, LCG^{+}18, LZZ^{+}22b, PFC^{+}24, PCPN24, PL10, QCD^{+}19, QHW22, RGL^{+}23, SYZB24, SOL22, TBW23, TYW^{+}21, WZLZ23, WLL^{+}21, WCHH23, XL15, XZLL21, YYF^{+}16, YCC^{+}15, ZB20, ZZXY21, ZZY^{+}24b, LSY^{+}09]. Modelling [GXL24]. Models [ASPT21, BJ21, BSK24, CSHZ21, CO18, GMS21, GA22, JCC23, LRK^{+}19, LLW^{+}21, MNK18, STD^{+}18, WZL^{+}16, WSM^{+}18, XZYL12, YMD^{+}24, DD09, FSK09, WC12]. Modifier [WWW^{+}16]. Modular [JDE^{+}12]. modulated [IHS07]. Molecular [XKH^{+}16]. moment [DD09, SYZB24]. Money [WZZ^{+}24]. Monitoring [SCCM22, TDLM19, XZXW18]. Monte [PCVR22]. Monte-Carlo [PCVR22]. Most [WV14, MVT12]. Motif [BLP21, LHW^{+}20, SSK^{+}10, WLR^{+}14, BCK^{+}18]. Motif-Aware [LHW^{+}20]. Motifs [AK15, GSST16, HTT^{+}24, SYXW22, DTU21]. Motisf [SSL^{+}23]. Motisf-aware [SSL^{+}23]. Motives [LFC^{+}17]. Movement [ABW20]. Moving [WYG^{+}17, ZBL^{+}20]. MP [BXF^{+}20]. MSIPA [LSS^{+}22]. MULFE [LHL^{+}22]. Multi [ACE20, ASPT21, AFW^{+}21, BXF^{+}20, BFRL13, CTX^{+}17, CLYC23, CFX^{+}24, CQH^{+}24, CWW^{+}24, CGZW16, CHZ23, CWWL24, DWH^{+}23, DDY^{+}24, GLZ^{+}21, GMS21, GF23, GFHL23, GFL^{+}24, GCL^{+}19b, GHT^{+}24, Han24, HYYT21, HSY^{+}21, HWZ^{+}24, HCH^{+}24b, JSF^{+}24, KKB22, LWCZ24, LSS^{+}22, LSF18, LJL^{+}23, LHZ^{+}24a, LYN^{+}24, LSW^{+}24, LHL^{+}22, LLG^{+}23, LF18, LZX22, LLLW22, LGZ^{+}23, LGS^{+}23, LCX^{+}23, LYF^{+}24, LFL^{+}23, LF23, LGT^{+}24, LWG^{+}24, MH22, MHS20, MFS24, NLA23, GVdSL24, OGV22, OGT^{+}21, PKSM21, PXW^{+}22, PZW^{+}18, PMC22, QCD^{+}19, RQM^{+}24, RKC19, SOL22, SZD^{+}23, SCCM22, SLL^{+}23, SYG^{+}24, SKW^{+}21, SJHY24, WLW^{+}19, WDF22, WCZ^{+}22, WDL^{+}23, WWG^{+}23, WCHH23, WCL^{+}23a, WSW24, WPD^{+}24, WSZ^{+}16, WN22b, WZB^{+}23, WHC^{+}23, XYZ^{+}20, XCL22, YLHY20, YHZ^{+}24, ZFW22, ZLLW23, ZZY^{+}23, ZSC^{+}23, ZFWC18, ZGY^{+}22, ZYH19, ZLY^{+}20b, ZZZ^{+}20b, JTYY10]. Multi-Agent [RQM^{+}24]. Multi-Behavior [SOL22, LWG^{+}24]. Multi-Concept [WWG^{+}23]. Multi-Context [JSF^{+}24]. Multi-destination [HYYT21]. Multi-Dimensional [GLZ^{+}21, PMC22]. Multi-Document [QCD^{+}19]. Multi-Domain [CGZW16]. Multi-Fuzzy-Constrained [LLLW22]. Multi-Grained [ZFWC18]. Multi-granularity [LYF^{+}24].

Multi-Graph [LJL^{+}23, NLA23, OGT^{+}21]. Multi-Instance [BFRL13, LGT^{+}24]. Multi-interest [WCHH23]. Multi-Label [BFRL13, GCL^{+}19b, LHL^{+}22, LFL^{+}23, MFS24, GVdSL24, RKC19, WDL^{+}23, WPD^{+}24, WHC^{+}23, XCL22, ZFW22, ZLLW23, ZZY^{+}23, PKSM21, SCCM22, WDF22, JTYY10]. Multi-Layered [CTX^{+}17]. Multi-Level [CLYC23, CQH^{+}24, SJHY24, LGS^{+}23]. Multi-Modal [WSZ^{+}16, ACE20, CWW^{+}24, YHZ^{+}24]. Multi-Network [GMS21]. Multi-Objective [WSW24, MH22, MHS20, PKSM21]. Multi-output [CHZ23]. Multi-Party [BXF^{+}20, HWZ^{+}24]. Multi-relation [KKB22]. Multi-Resolution [ASPT21]. Multi-Scale [LSS^{+}22, LHZ^{+}24a]. Multi-Scenario [SYG^{+}24]. Multi-set [PXW^{+}22]. Multi-Source [WN22b, YHZ^{+}24, ZGY^{+}22, WLW^{+}19]. Multi-Stage [HSY^{+}21, SLL^{+}23]. Multi-style [ZLY^{+}20b]. Multi-Task [GMS21, GHT^{+}24, LCX^{+}23, OGV22, SYG^{+}24, ACE20, YLHY20, ZYH19]. Multi-Transaction [WCZ^{+}22]. Multi-User [XYZ^{+}20]. Multi-View [AFW^{+}21, CFX^{+}24, DWH^{+}23, GF23, GFHL23, GFL^{+}24, LSF18, LYN^{+}24, LSW^{+}24, LF18, LZX22, LGZ^{+}23, LFL^{+}23, LF23, PZW^{+}18, SZD^{+}23, WCL^{+}23a, WZB^{+}23, CWWL24, Han24, HCH^{+}24b, LWCZ24, LLG^{+}23, MHS20, SKW^{+}21, ZSC^{+}23, ZZZ^{+}20b]. Multi-Way [DDY^{+}24]. Multi-weight [HYYT21]. Multiclass [LYL^{+}20]. Multidimensional [BB17, MRTW19, PC20, PLL^{+}10]. Multidocument [WZL^{+}11]. Multifaceted [WZL^{+}23a]. Multigraphs [LJK18]. Multilabel [WSZZ14, ZZ10, ZY13]. Multilayer [Cos22, GBGL20]. multilevel [PLL^{+}10]. Multimedia [GZXF16]. Multimodal [GZXF16, LWG^{+}22].

Multimodal-Attention [LWG^{+}22].

Multinomial [KRBK19]. Multiobjective [SMK18]. Multiplayer [JSV^{+}15]. Multiple [ABWI23, CLY12, CLT^{+}20, CXH^{+}20, hDRNlD21, HTT^{+}24, HGW^{+}21, IYSU12, JH19, JYD19, LCWC20, LZZ^{+}22b, LHG^{+}24, NYLZ24, SHL19, SWY^{+}24, SYZ^{+}24a, TBW23, WLC^{+}17, WGYC21, WN22b, CFD10, WND^{+}09]. Multiple-instance [SWY^{+}24]. Multiplex [LWS22]. multirelational [JMR08]. Multiresolution [SCS14]. Multisource [CSF^{+}12, YSLL24]. Multitask [LYL^{+}20, ZY14, ZZL15, YLK13]. Multivariate [ES15, HBD^{+}24, LLY^{+}24, LCLL17, TR22, ZLD^{+}23]. Multiview [SMK18]. Music [ZLY^{+}20b]. Mutual [LLL23a, LZSY24, RZY^{+}23, ZLC20].

Naive [LLW16]. Name [PMC22, TS09]. Named [WTZL22]. Nationwide [CXWH24]. Natural [ABDM21]. NBA [VALF12]. Near [CCV19, JK22]. Near-Optimal [CCV19]. Nearest [AF13, LSZ^{+}24, WHC^{+}23]. Nearly [NK20]. Need [LWZ14, PFC^{+}24]. Negative [SI24, WLD^{+}23, WCL^{+}23b, ZHT20]. Neighbor [AF13, JCQ^{+}23, KCR^{+}24, LSZ^{+}24, MA16, WWZ^{+}24]. Neighbor-Based [AF13]. Neighbor-Enhanced [WWZ^{+}24]. Neighborhood [EGT14, LJX23]. Neighbors [JK22, KCP^{+}24, WHC^{+}23, Kor10]. Neighbors-Based [JK22]. Neighbours [XLS24]. Neighbours-to-Neighbours [XLS24]. Nested [WTZL22]. Net [LJL^{+}23, XSZY20]. Network [ANK14, AFW^{+}21, DMGC21, BGSW13, CSX21, CS24, CXWH24, CDSV16, FLL^{+}22, GYX^{+}22, GMS21, GSWJ20, HZL20, HGW^{+}21, HZZ^{+}15, HF24, HXY^{+}19, HSY^{+}21, JLZ^{+}22, KCR^{+}24, KFI^{+}23, KS21, LSS^{+}22, LHW^{+}21, LWB^{+}22, LWWX23, LFY^{+}23, LJL^{+}23, LHZ^{+}23, LLL23a, LCZ^{+}24, LYL^{+}22, LVH^{+}21, LGZ^{+}21, LHG^{+}24, MW24, MYB19, MRJ11, MNK18, OJW^{+}23, OGT^{+}21, PNB24, QZL23, DMG^{+}21, SYXW22, SGC^{+}23, SWD^{+}21, SYZ^{+}24b, SHH^{+}20, TLZ^{+}20, WCL19, WWW^{+}20, WZW^{+}22, WZZ^{+}22, WZWC23, WLH24, WLS^{+}23, WGYC21, WN22a, XLL^{+}21, XLQ^{+}24, XLS24, XZXW18, XYGC21, XXG^{+}24, YZZ22, YWL^{+}24, YHZ^{+}24, YL24, YSLL24, YLL^{+}24, ZWG^{+}19, ZTT^{+}20, ZGHM21, ZYZZ21, ZWLS22, ZZYY22, ZLT^{+}23, ZL23, ZZY^{+}24b, ZLZ^{+}24, ZYS^{+}21, ZLHZ24, ZRJ^{+}22, ZCL^{+}22, ZJJ^{+}24, ZMZS23, VALF12, JLL14, KSM09, MS09, YWW^{+}14]. Network-Based [ZLT^{+}23, SYZ^{+}24b]. Networked [GWZZ17, HML^{+}24, ZJJ^{+}24]. Networking [GBTL14]. Networks [AK15, ADR21, ABS19, BG21, BB17, CSG^{+}16, CDC23, CG15, CTX^{+}17, CL18, CLT^{+}20, CPYT21, CRP^{+}24, CFX^{+}24, CHY^{+}24, CGL18, Con20, Cos21, Cos22, CO18, DWH^{+}23, DHZ24, DPDG18, ERSK14, FLG^{+}21, GBGL20, GCB^{+}21, GXLC21, GJDX14, GRLK12, HSBH19, HQWW22, HLZH18, HDT^{+}18, HFL^{+}22, JCS^{+}23, JCB^{+}16, JYH^{+}23, Jia24, JDE^{+}12, JZX^{+}23, KRBK19, KP18, KMH^{+}24, KS24, LNG18, LWS22, LTN^{+}21, LXC^{+}22, LZL^{+}23, LFXC24, LAN^{+}18, LOM22, LLS^{+}21, LT10, LSGZ19, LLW^{+}21, LXR^{+}23, LILJ24, LN24, MSC^{+}19, MLM21, MA16, MRTW19, NS23, NZW23, NYLZ24, PAS24, PLS^{+}21, QZL23, QSS20, RABR23, RJK^{+}20, SYZB24, SBZR19, SYK22, SGM^{+}23, STD^{+}18, SZLP16, SH15, SNH^{+}13, SHH^{+}20, SSL^{+}23, SMY^{+}24, THRR16, TSS22, WLR^{+}14, WWHW19, WLL^{+}21, WJLL21, WCHH23, WLHH23, WLS^{+}24, WLH24, WC15, WLS^{+}23, WWZ^{+}24, WZL^{+}16, WPDZ21, WMW^{+}22, WZL^{+}23c, YLW^{+}19, YTL18, YWZ^{+}24, ZS24, ZP15]. Networks [ZXL22, ZLD^{+}23, ZYC^{+}24, ZHLZ24, ZFWC18, ZBX^{+}24, ZL15, ZHS^{+}19, ZPC^{+}16, ZBAG20, AMIL13, AGHN13, CT14, CC12, HNH^{+}13, KNV07, LCZ^{+}09, LTH^{+}13, ML14, WC12, ZFY14]. Neural [BG21, CLT^{+}20, CRP^{+}24, CXWH24, DWH^{+}23, FLG^{+}21, FLL^{+}22, GBBC24, HF24, KMH^{+}24, LLC^{+}21, LHZ^{+}23, LZL^{+}23, LILJ24, LN24, MLM21, OGT^{+}21, QHW22, SYZB24, SYK22, SSL^{+}23, SMY^{+}24, WCHH23, WLS^{+}23, WPDZ21, WMW^{+}22, WZL^{+}23c, XLQ^{+}24, XSZY20, XYW^{+}20, YL24, ZWG^{+}19, ZH23, ZYC^{+}24, ZLZ^{+}24, ZLHZ24, ZCL^{+}22, ZBAG20]. NeuSE [LLC^{+}21]. News [ALW^{+}24, BBB^{+}22, LHN^{+}20, QHW22, SG12, VAFZ19]. News-Based [BBB^{+}22]. Next [JHC^{+}23, OJW^{+}23, SOL22, WZZ^{+}22, WCHH23, ZCL^{+}22]. Next-item [WCHH23]. NGUARD [XLT^{+}20]. NMC [LYGG22]. NMC-BERT-LSTM-DQN-X [LYGG22]. NMF [CWWL24]. NNC [XLS24]. NNC-GCN [XLS24]. Node [CGL^{+}23, CS24, CFX^{+}24, CSM^{+}24, DWL^{+}24, ERSK14, GWZZ17, GW21, JHJK22, MLM21, SZLP16, SGCH23, WZL^{+}15, WN22a, XXO^{+}24, YWL^{+}24, YHZ^{+}24, ZWYZ24, ZLSC24]. Nodes [BCK^{+}18, JS21, NS23, WWHW19, WPDZ21, WN22a, ZHLZ24]. Nodes-Hyperedges [WN22a]. Noise [CYY^{+}22, Cos21, ABDM21, LJX23, LGT^{+}24, SM23, BFPP07]. noise-robust [BFPP07]. Noisy [HF24, LZZ^{+}22a]. Non [ACK^{+}22, CZY^{+}22, JYD19, LCLL17, MYPB20, PV24, SI24, SCCM22, WLD^{+}23, YLL21b, ZHT20, Vad10, YHLC23]. Non-causal [YLL21b]. Non-Convex [ACK^{+}22]. Non-Gaussian [LCLL17]. Non-IID [YHLC23]. Non-intrusive [SCCM22]. non-linear [Vad10]. Non-Local [JYD19]. Non-Negative [SI24, WLD^{+}23, ZHT20]. Non-Redundant [MYPB20]. Non-stationary [CZY^{+}22]. Non-uniform [PV24]. Nonnegative [BNY20, HNHD14, LLZ^{+}21, PKH^{+}17, SGCH23, WNH15, ZGHM21]. Nonoverlapping [WLL^{+}22b]. nonredundant [CFD10]. Nontransductive [AHGA14]. NOODLE [GFL^{+}24]. Norm [DDY^{+}24]. Normalization [DDY^{+}24]. Normalizing [LOM22]. Novel [CMS23, CV24, DA19, GBBC24, NWA20, RGAÁC^{+}24, SB24, WZ21, WQZ^{+}16, ZZYY16, ZHSL23]. Novelty [XYW^{+}20]. Nr [MYPB20]. Nr-DipMeans [MYPB20]. Nr-Kmeans [MYPB20]. nSimplex [CV24]. NTP [WLL^{+}22b]. NTP-Miner [WLL^{+}22b]. Nyström [ZZYY16].

Object [RQM^{+}24, SNH^{+}13]. Objective [WSW24, HGV^{+}08, MH22, MHS20, PKSM21]. Observable [TSS22]. Observational [HML^{+}24]. Observed [RABR23]. Occupancy [NGB18, ZLT^{+}15]. Occupancy-Based [ZLT^{+}15]. Occurrence [ZFH23]. Off [JKP^{+}21, WCL^{+}23b]. OLAP [QSS20]. On-Demand [YYC^{+}21]. On-Shelf [ZDY^{+}22, SB24]. On-the-fly [GBBC24]. On/Off [JKP^{+}21]. One [HP20, LGT^{+}24, OGAB14, WCL^{+}23b]. One-off [WCL^{+}23b]. Ones [SB21b]. ONION [GFHL23]. Online [ADR21, ASZ21, BC18, CL18, CZY^{+}22, Cos21, GXLC21, HHZ^{+}18, HQWW22, JSV^{+}15, KP18, LCCM19, LZZ^{+}22a, LHZG13, LSH^{+}22, LWZ14, LT10, MYB19, Row16, SHL19, DMG^{+}21, SWH^{+}23, THRR16, TKT^{+}24, WJLL21, WHMY17, XZLL21, YLW^{+}19, YWG^{+}22, YWDP16, ZS24, ZWH^{+}16, ZWWH20, ZMW^{+}24, ZZYW22, YLK13]. Online-Weighted [BC18]. Only [CWW^{+}24]. ONP [WCL^{+}23b]. ONP-Miner [WCL^{+}23b]. Ontology [CTZ16, MAHT18]. Open [JSE^{+}23, XXO^{+}24, YWS^{+}22]. Open-Set [JSE^{+}23]. Open-World [XXO^{+}24]. Opinion [ACK^{+}22, ABS19, JLZ^{+}22, KSB^{+}21].

Optimal

[Bur21, CCV19, LW24, XZ23, ZBL^{+}20].

Optimization

[ACK^{+}22, AHGA14, CTP^{+}16, CWY^{+}23, CCT^{+}23, DYS20, IAB22, JK22, LGG^{+}23, LCWC20, MHS20, OCD^{+}24, RZY^{+}23, SMK18, SYLC16, WSW24, XYZ^{+}20, XWW^{+}24, XXZ19, ZL15, ZYH19]. Optimized [SMS22]. Optimizing [ES15, SBRE14, XWW^{+}24]. Optimum [BC18]. Order [GFL^{+}24, LLL^{+}24, PMC22, WZL^{+}23b, YLL19, ZZY^{+}21, ZCS10, ZHL^{+}21]. ordered [HAKU^{+}08, ZLGY24]. Ordinal [GXZ^{+}22, ZZLZ21]. Ordinal-Preserving [ZZLZ21]. Organized [LLW^{+}21]. Oriented [LZSY24, PZW^{+}18, TDLM19, XLD^{+}24, XSZ^{+}22]. Origin [RLDL24]. Origin-destination [RLDL24]. Orthogonal [XWZ^{+}22]. other [VALF12]. Our [CDSV16]. Out-of-Distribution [SLWW24, ZWYZ24, SMW^{+}24]. Outcome [TDLM19]. Outcome-Oriented [TDLM19]. Outcomes [DKSK22, MFRJ16]. Outdoor [ZBL^{+}20]. Outlier [AF16, CMZS15, DDL21, LWCZ24, LSF18, LSL^{+}22, MCSZ20, MN20, PC20, SLK^{+}24, SB21a, WCL^{+}23a]. Outlierness [WDDB20]. Outliers [AF16, HW21, SI24, SB21b, AF09]. Outlying [WDDB20]. Output [CSM^{+}24, LWW23a, CHZ23]. Over-Smoothing [PCPN24]. Overcome [MH22]. Overlap [BBC^{+}24]. Overlapping [CRGP14, CO18, LHK^{+}18, LFXC24, NLZH20, NXZ^{+}24]. Own [CDSV16].

P [LGS^{+}23, XJZ21]. P2P [AGHN13]. Paced [LLH^{+}24, ZLDL23]. Package [ZLZ^{+}24]. Page [AS21, WNH15, ZP09]. Page-Level [AS21]. PageRank [CGL^{+}23, LXY^{+}17, ZGC18]. Pairs [WZL^{+}15]. Pairwise [DYS20, PFTR16]. Pairwised [HZW^{+}15]. Pandemic [JLZ^{+}22]. Papers [Agg17, WL16, XL20]. Paradigm [KMH^{+}24]. Paradox [JSP15, NK21]. Paradox-based [NK21]. PARAFAC [PFS15]. Parallel [AKM20, GLFV^{+}19, JLH^{+}13, WGYC21]. Parallelizable [PFS15]. Parallelized [BXF^{+}20]. Parameter [ZYD24b, BFPP07]. Parameter-Agnostic [ZYD24b]. Parameter-free [BFPP07]. Parameterization [CLG^{+}19]. ParCube [PFS15]. Parkinson [LLL^{+}20]. Parliamentary [MCS^{+}18]. Parsing [WZL^{+}22]. PART [RZF^{+}24]. Partial [LYL^{+}22, WHG^{+}18, XL15, ZWB22]. Partially [RABR23, TSS22]. Participation [LK20]. Participation-based [LK20]. Particle [XXZ19]. Partition [RZF^{+}24]. Partitioned [TAJY17, MWF08, VCKP08]. Partitioning [SYD^{+}16, SZLP16]. Party [BXF^{+}20, HWZ^{+}24]. Pass [OGAB14]. Passenger [LH22, WLW^{+}19]. Passing [BGSW13]. Patent [LWH^{+}23]. Path [HFL^{+}22, LLL^{+}24, LH22, LZL^{+}24, MRJ11, SNH^{+}13, YCJK08]. Paths [ASJ^{+}23, Lap20, MSR^{+}24]. PathSelClus [SNH^{+}13]. Patients [WLC^{+}17]. Pattern [CYOL16, DCF^{+}21, FLC23, GLFV^{+}19, HHL^{+}24, HGY24, KR16, LSS^{+}22, LYT^{+}23, LLB^{+}22, PCVR22, PSYJ24, RGAÁC^{+}24, WLL^{+}22b, WCL^{+}23b, YDS^{+}15, ZLT^{+}15, KKZ09, YCJK08]. Pattern-Aware [LSS^{+}22]. pattern-based [KKZ09]. Patterns [AASL23, BGWSB19, CLY12, JH19, LLC^{+}22, LDS^{+}22, NADR21, SLW^{+}18, SDS18, WDDB20, WTD24, WC15, YWH^{+}15, ZBAG20, CBRB09, MXC^{+}07, PLL^{+}10, ZKCY07]. PCA [CNY^{+}16, FKKD17]. peaks [AH24]. Peer [HLL^{+}20]. Percentage [LCLL21]. Percolation [CGL18]. Perfection [AKM17]. Performance [MSC^{+}19, RU14]. periodic [ZKCY07]. periods [HDC07]. Permanence [CSG^{+}16]. Persian [MG24]. Person [ZCL^{+}22]. Personal [LCW24]. Personalised [ZGC18]. Personality [KP18]. Personalizable [XJZ21]. Personalized [ALW^{+}24, CLYC23, HGW^{+}21, HCZ^{+}14, LSH^{+}22, LFC^{+}17, LGS^{+}23, THYL24, YHLC23, YCW^{+}24, YCC^{+}15, ZMC^{+}24]. Perspective [BT21, CSSP15, CTX^{+}17, HZM^{+}22, HLH^{+}24, LXZ^{+}24, LGZ^{+}23, WXT^{+}24, XWW^{+}24, ZZYY16, ZZY^{+}21]. Persuasion [ACK^{+}22, XZLL21, ZS24]. Perturbation [YW24]. Perturbations [ZBAG20]. Petabyte [LGF10]. Petabyte-Scale [LGF10]. Phase [Dor19]. Phone [AOEM17, QSST18, YZH^{+}18]. Photos [CLYC23]. Phrase [LWH^{+}23]. Physical [HZW^{+}15, TYG^{+}15]. Physician [WDL^{+}21]. Physics [HQYY14]. Physics-Based [HQYY14]. physiological [SSK^{+}10]. Pickups [CZZ^{+}23]. Pipeline [YLL19]. Pipes [ZCL19]. Placement [ZBL^{+}20]. Planning [HYYT21, WZR^{+}23]. Platform [SGM^{+}23]. Play [DKSL18]. Player [XLT^{+}20]. PM2.5 [YWC24]. POI [JHC^{+}23, WZZ^{+}22]. Point [HCH^{+}24b, LYL^{+}20, OJW^{+}23, QXBT16, SYZB24, SYD^{+}16]. Point-of-Interest [LYL^{+}20, OJW^{+}23]. Pointer [HGW^{+}21]. Points [JYD19]. Poisoning [WLG^{+}24]. Poisson [IHS07]. Polar [ABS19]. Pooling [RZF^{+}24]. Pop [ZLY^{+}20b]. Popularity [GCL^{+}19a]. Population [ASPT21, WNC^{+}18]. Populations [MNK18]. Portable [FLG^{+}21]. Portfolio [LHZG13, YWG^{+}22]. Poset [PCVR22]. Positive [WPDZ21]. Posts [NTNP18]. Potential [LFXC24, QHW22, ZHSL23]. potentially [Web10]. Power [FFZ^{+}24, NK21, VDMC22, PAS24, YJT^{+}24]. Power-law [NK21, PAS24]. powered [YGL^{+}22]. Practical [CCV19, CZZ^{+}23]. Practice [YJT^{+}24]. Pre [CYY^{+}22, WWZL23, WOGZ24, WMZL24]. Pre-Processing [CYY^{+}22]. Pre-Training [WWZL23, WOGZ24, WMZL24]. Precedence [GGLP15]. precise [HDC07].

Predict [ASPT21, LVH^{+}21, SM23, LTH^{+}13]. Predictability [YDW^{+}23]. Predicting [BAMK18, ERSK14, LCZ^{+}24, DMG^{+}21, VMR17, WLW^{+}19, ZCL^{+}22, ZTL^{+}15b]. Prediction [BMZY21, BJ21, CLT^{+}20, CSZ^{+}21, DKA11, FLL^{+}22, GCL^{+}19a, GXLC21, HFL^{+}22, IKK19, JYH^{+}23, KZW^{+}22, LSS^{+}22, LFY^{+}23, LJL^{+}23, LZG^{+}24, LCLL17, LSH20, LZZ^{+}22b, RGL^{+}23, RBF^{+}21, SHL19, SBZR19, SLL^{+}23, TWC22, WYG^{+}17, WCL19, WLD^{+}21, WZ21, WXW^{+}21, WZR^{+}23, WLS^{+}24, WLH24, WLS^{+}23, WWZ^{+}24, XJW^{+}21, XLL^{+}21, XYGC21, YL24, YSLL24, YWK^{+}22, ZZXY21, ZZY^{+}24b, ZCC^{+}24, ZLZ^{+}24, ZJJ^{+}24, BG09, HNH^{+}13, LSY^{+}09]. Predictions [GSI19]. Predictive [AT17, PFC^{+}24, TDLM19, SSK^{+}10]. Predictors [DLZ^{+}24, LCWC20, CRST09]. Preference [HLL^{+}20, LLZ^{+}23, LYG^{+}24]. Preferences [JHC^{+}23]. Preferred [LZL^{+}24]. Preprocessing [GZSMM24, HQWW22]. Prerequisite [SMY^{+}24]. Prerequisite-Enhanced [SMY^{+}24]. Presence [KHTR18, BC24, LNR08]. Preservation [VSV15]. Preserve [LOW^{+}21]. preserved [CCC09]. Preserving [CXL^{+}24, LLL^{+}21, LGS^{+}23, LFL^{+}23, SI24, WCL^{+}23a, XCL22, YSG^{+}21, ZZLZ21, ZZF24, ZLSC24, MWF08, VCKP08, ZMC^{+}24]. Pretext [SM23]. Price [XML18]. Pricing [MPG^{+}23]. Principal [CLG^{+}19]. Principle [ASPT21, WSZ12]. Principled [KSV^{+}16]. Prior [JYD19, ZB20, LNR08]. Prioritized [LAN^{+}18]. Privacy [BTBg22, CXL^{+}24, GBW^{+}24, HF24, LCW24, LK15, LT10, LGS^{+}23, MWF08, RTM18, SI24, VCKP08, WFW^{+}11, XCL22, ZMC^{+}24, MKGV07]. Privacy-Preserving [CXL^{+}24, LGS^{+}23, SI24, XCL22, MWF08, VCKP08, ZMC^{+}24]. Private [EC20, GZSMM24, HXY^{+}19, KFI^{+}23, NS23, NYLZ24]. Probabilistic [CZH18, CRM^{+}23, HP20, HO14, HLCR20, HML^{+}20, JCQ^{+}23, JLD^{+}19, KHTR18, LLW^{+}21, ZB20, ZCL19, HAKU^{+}08]. Probability [CWF^{+}13, ZZLZ21]. Probes [XJZ21]. Problem [CG15, NZW23, QWH^{+}24, XL16, ZTT^{+}20, GEG^{+}08]. Problems [ADK^{+}16, DDL21, GXZ^{+}22, MYL^{+}21, RWD23, SKM^{+}22]. Process [GLMW22, GGLP15, IAB22, TDLM19, YMD^{+}24]. Processes [DMI12, LZD16, OLL20, QXBT16, SYZB24, TC18, IHS07]. Processing [CYY^{+}22, LK15, WZLZ23]. Product [HGW^{+}21, LFC^{+}17, MNK18, XJW^{+}21, XL15, XZLL21, XL16, ZWG^{+}19, ZWH^{+}16]. Products [Che18, ZWG^{+}19]. Professionals [Lap20]. Profile [LYTZ22]. Profiles [BBB^{+}22, YCC^{+}15]. Profiling [LCX^{+}23, LZLJ24, TYZZ10, XKH^{+}16, TLZ^{+}08]. Profit [LWW18]. programs [GMSS13]. ProgressER [AKM18]. Progression [LCG^{+}18]. Progressive [AKM18, JH19, PV24]. Projection [YSG^{+}21, WAD12]. Projections [PBMID14, WAD12]. Prominence [AMIL13]. Prominent [ZJL^{+}14]. Promotion [ZPC^{+}16]. Propagation [KMH^{+}24, LHW^{+}20, LXY^{+}17, YWH^{+}15]. Properties [BSS^{+}24, MW24, VFA^{+}15, Bal13]. Protect [LCW24]. Protected [BSS^{+}24]. Protecting [RTM18]. ProtoMGAE [ZJ24]. Prototype [ZJ24]. Prototype-Aware [ZJ24]. Provable [SOK^{+}20]. Proximity [RZY^{+}23, RJK^{+}20, SYLC16, KNV07]. Pruning [WLT19]. Pseudo [DHZ24]. Pseudo-Likelihood [DHZ24]. Pseudodimension [RV20]. PSL [LYL^{+}22]. PSP [JH19]. PSP-AMS [JH19]. PU [WZZ^{+}24]. PU-Detector [WZZ^{+}24]. Public [JLZ^{+}22]. Publishing [GBW^{+}24, VSV15]. Pulse [DCF^{+}21]. Punitive [RKC19]. Purposes [LCZ^{+}24].

Put [ALW^{+}24, PPDSBLP16]. Puzzle [SRVM24].

Quality [CXWH24, CTZ16, GLMW22, JCC23, LQW^{+}18, MDV11, QZB^{+}23, RGL^{+}23, WSDL19, XW22, XZSY19, ZHL^{+}21]. Quality-Informed [GLMW22]. Quantification [MFS24]. Quantifiers [ES15]. Quantifying [YDW^{+}23]. Quantitative [LYGG22]. Quantum [LSZ^{+}24]. Quasi [SMDHT21, JP09]. Quasi-Cliques [SMDHT21, JP09]. Quaternion [ZLT^{+}23]. Queries [HCC^{+}18, WWW^{+}16, ZLY^{+}20a]. Query [AT17, HCC^{+}18, LLLW22, MSC^{+}19, PXW^{+}22, WQZ^{+}16, CRST09]. query-defined [CRST09]. Query-Driven [AT17]. Querying [WY15, WTD24]. Question [BWD10, GFM21, WMZL24, ALB09]. Question-Answering [BWD10].

Rademacher [CWR23, PCVR22, PV24, RU18]. Radii [KTA^{+}11]. Rail [ELH23]. Rainfall [TWC22]. Raising [SB24]. Random [BC24, LWZ14, NS23, NK20, PBMID14, MWF08, WC12, WAD12]. randomization [GMMT07]. Randomized [FKKD17, WXB^{+}21]. range [FLL^{+}22]. Rank [CLY12, DSTA22, DKSL18, HH19, LSF18, TLL^{+}19, TKT^{+}24, WYX^{+}16, WN22b, YSG^{+}21, CCT^{+}23, ZSC^{+}23, ZGC18]. Ranking [BES15, CNZ^{+}17, DYS20, FXG^{+}16, GJDX14, JLH^{+}13, LSH^{+}22, LWW23a, MLM21, MN20, SRBC22, YDE^{+}22, AMIL13, JLL14]. Ranking-Based [LWW23a]. Rapid [SYLC16]. Rare [KR16, MN20, NWA20]. Rare-Class [NWA20]. Rate [DLZ^{+}24, XYGC21]. Rating [LWH^{+}24, XL15, XZLL21, YCC^{+}15].

Ratings [FCWQ17, WZL^{+}23b]. Rationality [LQW15]. Rationalizing [LILJ24]. Ratios [BSS^{+}24]. Reader [QCD^{+}19, VAFZ19]. Reader-Aware [QCD^{+}19]. Real [AOEM17, BLH^{+}22, CS24, CYL^{+}24, FXG^{+}16, LCWC20, LSH20, MGL^{+}20, NZW23, OCD^{+}24, WSDL19, WZZ^{+}24]. Real-Time [AOEM17, BLH^{+}22, LCWC20, MGL^{+}20, OCD^{+}24, WSDL19, LSH20]. Real-World [CS24, CYL^{+}24]. Realistic [DLZ^{+}24, SB21a, VFA^{+}15, WNC^{+}18]. Really [PFC^{+}24]. Realtime [LOW^{+}21]. Reasoning [LZY^{+}24a, LZL^{+}24, MZC^{+}24, RQM^{+}24]. Receptive [TLG^{+}23]. reciprocity [LTH^{+}13]. Recognition [HLC19, HCZ^{+}14, JSE^{+}23, LTN^{+}21, LLL^{+}20, NWA20, SKW^{+}21, WTZL22, XCL22, YSW^{+}21]. Recommend [CCLZ18]. Recommendation [CWL^{+}24a, CMBT22, CLYC23, CCTW23, CSHZ21, CKC^{+}18, CDV21, CXL^{+}24, FYN22, GLF^{+}22, GYX^{+}22, HHL^{+}24, HZZ^{+}15, HCH24a, HF24, HCC^{+}18, JCQ^{+}23, JBGW24, JHC^{+}23, JZX^{+}23, LL19, LWWX23, LHZ^{+}23, LLL^{+}24, LSZ^{+}24, LYG^{+}24, LWG^{+}16, LFC^{+}17, LCF19, LGZ^{+}21, LJW^{+}22, LZL^{+}24, LZY^{+}24b, LYF^{+}24, LWG^{+}24, LWW24, ML15, OJW^{+}23, OGT^{+}21, PCPN24, QHW22, SOL22, SYG^{+}24, SMY^{+}24, WSZZ14, WDL^{+}21, WZZ^{+}22, WZL^{+}23b, WCHH23, WZL^{+}23a, WXT^{+}24, XYZ^{+}20, XYW^{+}20, YCC^{+}15, YTW^{+}16, ZZYY22, ZWH^{+}16, ZLC20, ZLW^{+}21a, ZMZS23]. Recommendations [FCWQ17, HTT^{+}24, LYN^{+}24, MRC^{+}22, SGC^{+}23]. Recommender [Che18, CC19, FN24, ABDM21, LH22, LGS^{+}23, WCX24, ZWC22, ZH23, ZYS^{+}21]. Recommending [FLG^{+}21, LPK^{+}15]. Reconstructing [CWL^{+}24a, EGT14]. Reconstruction [GSG^{+}20, LSH20]. Record [OJLD22, dVKCC11]. Records [WLL^{+}22a]. Recovery [HW21, HQWW22, XLQ^{+}24]. Recurrent [GCG22, LZP20, LFY^{+}23, ZHS^{+}19]. Recurring [HKR^{+}23]. Recursive [GY15, THB18]. Redefined [LYC^{+}21]. Redescriptions [GM18]. Reduce [AZD^{+}21]. Reduced [DMGC21]. Reducing [WJLL21]. Reduction [CYY^{+}22, CCT^{+}23, CV24, GNMQ21, GBTL14, MDV11, MG24, ZWB22, PSFV13, ZZ10]. Redundant [MYPB20]. Refining [Pap15, ZZYY22]. Reflect [BG09]. Reframing [HO14]. Regime [TBW23]. Region [BMTT18, ZCL^{+}22]. Region-of-Interest [BMTT18]. Region-Relation-Aware [ZCL^{+}22]. Regional [RLDL24]. regions [CBLH12]. Regression [Dor21, GXZ^{+}22, HO14, KCL^{+}20, LBT^{+}23, LWH^{+}24, MYL^{+}21, OGAB14, SHL19, SATK20, STD^{+}18, YTH18, YDE^{+}22, ZLZ^{+}19]. Regressions [LZZ^{+}22a]. Regularization [CC19, CWWL24, GY15, MG24, YTH18, YZZ22, ZZ23, ZY14]. Regularized [CGZW16, PKH^{+}17, WCL19]. Regulation [LSH^{+}22]. Reinforcement [CZZ^{+}23, CCTW23, FSXL23, FYN22, LL19, MYC^{+}24, YYC^{+}21, ZRJ^{+}22]. Relatedness [ZGC18]. Relation [CTC^{+}22, WZ21, WZL^{+}23a, ZWYW24, ZCL^{+}22, KKB22]. Relation-aware [WZL^{+}23a]. Relational [AK15, AKM18, GYX^{+}22, JLZ^{+}22, NGB18, BG07]. Relations [DSTA22, VDMC22, WZL^{+}23c, CBRB09, SMA^{+}08]. Relationship [LAN^{+}18, LHS^{+}21, MYB19, SBZR19, WSM^{+}18, GEG^{+}08, ZY13]. Relationships [CDC23, JSV^{+}15, LYC^{+}23, LXR^{+}23, PZJ^{+}24, PMC22, YTL18, ZY14]. Relative [PFTR16]. Release [KFI^{+}23]. Relevance [LGF10]. Relevant [BGJV12]. Reliability [WZB^{+}23]. Reliability-Driven [WZB^{+}23]. Reliable [CWW^{+}24]. Rematching [LYT^{+}23]. REMIAN [MGL^{+}20]. Remote [XZSY19]. Repetitive [LWC^{+}23]. Representation [CTC^{+}22, CG23, CCT^{+}23, CCTW23, CLJC24, Dor19, DPDG18, Han24, HLC19, HLH^{+}24, JH22, JS21, LWC^{+}23, LLL23a, LZG^{+}24, LLG^{+}23, LCX^{+}23, LYF^{+}24, LJG^{+}24, RZY^{+}23, RZF^{+}24, SYD^{+}16, TLG^{+}23, WHG^{+}18, WCL19, WXW^{+}21, WZW^{+}22, WWG^{+}23, WJLY23, WCL^{+}23a, WWZ^{+}24, YSG^{+}21, YSW^{+}21, ZFH23, ZL23, ZWYW24, ZFWC18, ZJ24, ZLW21b, ZGQ^{+}24]. Representations [Che18, DKSK22, JZX^{+}23, LZSY24, LWG^{+}22, WXB^{+}21, WZL^{+}23b, XRGZ24]. Representative [CYL^{+}24, WY15, ZLY^{+}20a, MS09]. Reputation [SRBC22, XML18]. Reputation-Based [SRBC22]. requirement [ZKCY07]. Reranking [XLD^{+}24]. Residual [ZXZ^{+}23]. Resistance [SRBC22]. Resisting [WLS^{+}24]. Resolution [AKM18, ASPT21, KCR23, BG07]. resolutions [WND^{+}09]. Resource [YZL^{+}22]. Resource-efficient [YZL^{+}22]. Response [SYZ^{+}24a]. Responsive [ZMW^{+}24]. REST [CWL^{+}24a]. Restricted [JSG^{+}19]. results [GMMT07]. Resume [Lap20]. Retail [LCLL21]. Retraceable [PAS24]. Retrieval [CKC^{+}18, DKSL18, LWW23a, LWW^{+}21, MCS^{+}18, PSYJ24, QWH^{+}24, WWZL23, ZZLZ21]. Retweet [ZTL^{+}15b]. Revealing [JSV^{+}15]. Reverse [YLL^{+}21a]. Reversion [LHZG13]. Review [LWH^{+}24, PGR18, QSS20, TDLM19, ZP23]. Review-Aware [LWH^{+}24]. Reviews [WZL^{+}23b, XJW^{+}21, ZWH^{+}16]. Revisiting [HLH^{+}24]. RGB [LLL^{+}20]. RHPTree [LLB^{+}22]. RHPTree-Risk [LLB^{+}22]. RIC [BFPP07]. Rich [HGW^{+}21]. Ride [TLZ^{+}20, WLW^{+}19]. Ride-matching [TLZ^{+}20]. Ridesharing [TLZ^{+}20]. Riemannian [DYS20]. Right [DSL^{+}14, YMD^{+}24]. Right-Censored [YMD^{+}24]. Risk [LZG^{+}24, LLB^{+}22, YHW^{+}24, LNR08].

Road [HXY^{+}19, ZL23]. Robust [BAC23, CGZW16, DWL^{+}24, GFHL23, GSG^{+}20, HW21, HNHD14, IKK19, LZZ^{+}22a, LV18, LTN^{+}21, LW24, LWH^{+}24, GVdSL24, OLL20, PKH^{+}17, SRBC22, TLL^{+}19, YSG^{+}21, ZZG^{+}21, ZLZ^{+}19, ZHS^{+}19, ZCL24, dVKCC11, BFPP07, GG08]. Robustness [HK18, HZM^{+}22]. RoI [BMTT18]. Role [CO18, HZZ^{+}15, HLH^{+}24, RJK^{+}20, WCL19, JLL14]. Role-Based [HZZ^{+}15, RJK^{+}20, WCL19]. Rounding [FKKD17]. Route [HYYT21, SYZ^{+}24b, XYZ^{+}20]. Routes [LH22]. Routine [QSST18]. Rule [HGWY23]. Rules [DGB16, HGWY23, RU14, WSC^{+}17, ZLGY24, Bal13]. Rumor [VMR17, YLW^{+}19, ZCL24]. Rumors [VMR17].

S2OSC [YWS^{+}22]. SA2E [LLY^{+}24]. SA2E-AD [LLY^{+}24]. Saiyan [Con20]. SAKE [LLS^{+}21]. Sales [LCLL21]. Same [CMS24]. Sample [DIB24, LWZ14]. Sampled [HXY^{+}19, TR22]. Samples [DLZ^{+}24, LZL^{+}22, WY15, YWW^{+}21]. Sampling [ANK14, ADR21, BIPR13, BLP21, CWF^{+}13, CYL^{+}24, Cos21, JYY^{+}21, KS21, LK20, LJK18, LLS^{+}21, LN24, MM12, NS23, NK21, PV24, QXBT16, RU14, RV20, SKM^{+}22, DTU21, SJHY24, WXB^{+}21, YLL^{+}21a, ZZYY16, CCC09]. Sanitization [BCC^{+}21]. Satellite [XSZY20]. Satisfaction [AEEBEE20, ALB09]. Satisfactions [LWW18]. Saturating [BNY20]. Scalable [BHW^{+}17, DIB24, HST^{+}23, JLL14, LDS^{+}22, LWW23a, LYC^{+}21, LLB^{+}22, MXP20, SGM^{+}23, SYD^{+}16, TLG^{+}23, WSZ^{+}16, YWDP16, ZZYW22, Kor10]. Scale [AOEM17, BHW^{+}17, CZH18, CPYT21, CKC^{+}18, DHZ24, DKSL18, GY15, HLZH18, JYY^{+}21, KRBK19, LSS^{+}22, LHZ^{+}24a, LLS^{+}21, LGF10, MW24, VDMC22, SHL19, SDS18, SLTA11, WNH15, WHMY17, XZ23, XXZ19, ZZLZ21, BBCG10, GMSS13, SATK20, XJZ21]. Scale-Free [MW24, XZ23]. Scaled [ZB20]. Scaling [CPC10, XZ23]. Scan [ASJ^{+}23, CCV19, MXP20]. Scan-Paths [ASJ^{+}23]. Scarcity [GQH22]. Scenario [SYG^{+}24]. Scheduling [SYXZ23, YZL^{+}22]. Scheme [CYY^{+}22]. Scholar2vec [WXW^{+}21]. Scholars [WXW^{+}21]. Science [LGZ^{+}22, WZL^{+}22]. Scientific [KZZ^{+}20, Var22]. Scientists [Var22]. SCOAL [DG10]. Score [KRBK19]. Scores [LT10]. Screening [MWL^{+}22, Web10]. SDA [BXF^{+}20]. Search [ACK^{+}22, Bur21, MSC^{+}19, PKSM21, SYZ^{+}24b, TAJY17, VAFZ19, WWW^{+}16, WZWC23, WSZ^{+}16, ZYZZ21, FSK09]. Search-based [PKSM21]. Searching [LWW^{+}21, CRST09]. Seed [GW21]. Seeded [ZHL^{+}21]. Seeding [GW20]. seeker [ALB09]. Segment [YZZ22]. Segment-Wise [YZZ22]. Segmentation [GLW^{+}23, ZZZ^{+}20a]. Selecting [DSL^{+}14]. Selection [CFX^{+}24, DPB^{+}20, DSTA22, GFHL23, GSWJ20, JV20, JYD19, JLD^{+}19, LHZG13, LXZ^{+}24, OGAB14, PKSM21, SLK^{+}24, SZD^{+}23, SNH^{+}13, TL14, WZLG13, WYWY19, WHMY17, XL16, XSZY20, XXZ19, YWG^{+}22, YWDP16, YLL21b, YYD22, YLL^{+}24, ZLLW23, ZZY^{+}23, ZZYW22, DD09, WZLG12, YLK13]. Selective [LQW^{+}18, RA16]. Self [CLJC24, Han24, HHL^{+}24, JH22, JBGW24, JHC^{+}23, LLH^{+}24, LLW^{+}21, LJG^{+}24, GVdSL24, PCPN24, PMC22, RZY^{+}23, RKC19, SM23, TR22, WCL^{+}23a, WMZL24, Web10, XWW^{+}24, XXZ19, ZLD^{+}23, ZLDL23, ZZZ^{+}20b]. Self-Adaptive [LLW^{+}21, XXZ19]. Self-Adjusting [GVdSL24, RKC19]. Self-Attention [HHL^{+}24, PMC22]. Self-Attentive [PCPN24]. Self-Optimizing [XWW^{+}24]. Self-Organized [LLW^{+}21]. Self-Paced [LLH^{+}24, ZLDL23]. Self-Representation [WCL^{+}23a]. Self-sufficient [Web10]. Self-Supervised [CLJC24, JH22, JHC^{+}23, TR22, Han24, JBGW24, LJG^{+}24, RZY^{+}23, WMZL24]. Self-supervision [SM23]. Self-training [ZLD^{+}23]. Self-weighted [ZZZ^{+}20b]. Semantic [CMBT22, GLW^{+}23, II08, LWLW18, LWW^{+}21, MXC^{+}07, WLL^{+}22a, WZL^{+}23c, YWR^{+}19, ZGC18, ZZLZ21, SMA^{+}08, GXL24]. Semantic-LDA [GXL24]. Semantics [AFZ22, GXL24, WPW23]. Semantics-enhanced [GXL24]. Semi [CCLZ18, CRC^{+}23, CQH^{+}24, CWWL24, DD09, Dor21, DIB24, LSL^{+}22, LYT^{+}23, LWH^{+}24, NGMQ22, WN22a, XLS24, YWS^{+}22, YWL^{+}24, ZWYZ24]. Semi-analytical [DD09]. Semi-Supervised [CRC^{+}23, CQH^{+}24, LYT^{+}23, LWH^{+}24, NGMQ22, WN22a, XLS24, YWS^{+}22, CCLZ18, CWWL24, Dor21, DIB24, LSL^{+}22, YWL^{+}24, ZWYZ24]. Semiautomated [BES15]. Semiconductor [ZH16]. Semisupervised [AHGA14, HCZ^{+}14]. SemRe [ZGC18]. SemRe-Rank [ZGC18]. Sensing [NGB18, XZSY19]. Sensitive [CSSP15, HO14, LZY^{+}24b, SI21, YWR^{+}23, Vad10]. Sensor [HQWW22, WSDL19]. Sentence [CMBT22, JH22, WZLG13, WZLG12]. Sentence-to-Hashtags [CMBT22]. Sentiment [CRC^{+}23, IKK19, XJW^{+}21, YWC^{+}16, ZCZQ19]. separability [ASHK14]. Sequence [BJ21, LZG^{+}24, LZL^{+}24, SOL22, WDDB20, ZJL^{+}14, ZDY^{+}22, ZYD^{+}24a, ZCS10]. Sequence-To-Sequence [BJ21]. Sequences [BTBg22, GLZ^{+}21, PXW^{+}22, PCPN24, XLT^{+}20, KT09, ZKCY07]. Sequential [GLFV^{+}19, GCG22, GHT^{+}24, HHL^{+}24, HF24, HGWY23, HGY24, IYSU12, JH19, LHZ^{+}23, LYN^{+}24, LDS^{+}22, LYF^{+}24, MRC^{+}22, MM12, PCPN24, SM21, SFDW19, WLL^{+}22b, WCL^{+}23b, XYZ^{+}20, ZLGY24, ZFWC18, PLL^{+}10]. Serendipity [XYW^{+}20]. Series [ABWI23, ABSP^{+}18, ABW20, AZBW21, CHZ23, DSL^{+}24, DDY^{+}24, GSST16, HBD^{+}24, LLC^{+}22, LLY^{+}24, LTB18, MFRJ16, PZJ^{+}24, RET^{+}23, RCM^{+}13, SHL19, SLW^{+}18, SGM^{+}24, TBW23, TR22, ZLD^{+}23, ZSM22, CRM^{+}23, CE24, MH22, ZJJ^{+}24]. Server [MYC^{+}24]. Server-Client [MYC^{+}24]. Services [WLW^{+}19, YYC^{+}21, YWH^{+}15, YTW^{+}16]. Session [JCQ^{+}23, KN18, WCHH23]. Session-Based [KN18, JCQ^{+}23]. Sessions [JCQ^{+}23]. Set [CXH^{+}20, JSE^{+}23, YWS^{+}22, PXW^{+}22, WAD12]. set-covering [WAD12]. Setting [BBB^{+}22]. Settings [WLL^{+}21]. SGD [SYXZ23]. sGrapp [SÖ22]. Share [XL16]. Shared [JZX^{+}23, JTYY10]. shared-subspace [JTYY10]. Sharing [BWD10, EC20, WGC^{+}23]. Shelf [ZDY^{+}22, SB24]. Shift [LWB^{+}24]. Shifted [ZB20]. Shifted-Scaled [ZB20]. Shifts [TBW23]. Shop [YTW^{+}16]. Shop-Type [YTW^{+}16]. Short [BAMK18, HZL20, JLZ^{+}22, JHC^{+}23, TWC22]. Short-Term [HZL20, JHC^{+}23, TWC22]. Short-Text [BAMK18, JLZ^{+}22]. Shortcut [Pap15]. Shortcuts [ZHLZ24]. Shortest [MRJ11, MSR^{+}24]. Shortest-Path [MRJ11]. Shot [DWL^{+}24, RYM^{+}24, WDH^{+}24]. shrinking [LKF07]. Side [LZY^{+}24b, LGT^{+}24, ZYS^{+}21]. SigGAN [CDC23]. SIGKDD [ACPW13, BBD^{+}07, MG09, WL16, GSTC12, GL15, Wan10]. Signals [Row16]. Signed [CDC23]. Significance [GNMQ21]. SILVAN [PV24]. Similar [JCQ^{+}23, LSGZ19, RGAÁC^{+}24]. Similarities [CZY11]. Similarity [CG15, CLYC23, GXLC21, KSV^{+}16, LTN^{+}21, SI21, TAJY17, WCL^{+}23a, WSZ^{+}16, ZXZ^{+}23, ZHS^{+}19, II08, JLL14]. Simple [CPC10, CC19, LJK18, YSY^{+}22]. SimRank [CG15]. Simulation [LLLW22].

Simultaneous [WZR^{+}23, DG10]. Single [HCH^{+}24b, LLL^{+}20, MAHT18]. Singular [ME11, WHG^{+}18]. SIR [JS21]. SIR-GN [JS21]. Siren [GM18]. Site [GBTL14, LGZ^{+}21]. Size [DERU17, LNG18]. Sized [ADK^{+}16]. Skeleton [LLW^{+}21]. Sky [TWC22]. Skyline [ZLY^{+}20a]. Slack [RWD23]. Slack-Factor-Based [RWD23]. Small [TSRK20, XZ23]. Small-World [XZ23]. Smart [DDL21, YSW^{+}21, ZZL15]. Smartphone [DWW22]. SMONE [JCQ^{+}23]. Smooth [HP20, TLG^{+}23]. Smoothed [DYS20]. Smoothing [PCPN24]. Smoothness [XSZ^{+}22]. Snapshot [LLC^{+}21]. SNE [SHF24]. Social [AEEBEE20, ABS19, BMTT18, BGC14, CWL^{+}24a, CL18, CGL18, CDV21, Cos21, CYT^{+}17, CXL^{+}24, DPDG18, ERSK14, GXLC21, GBTL14, HTT^{+}24, HDQ^{+}18, HNH^{+}13, HDT^{+}18, JCS^{+}23, JSV^{+}15, KP18, LPK^{+}15, LWWX23, LCW24, LLD^{+}24, LLL^{+}24, LLS^{+}21, LT10, LSGZ19, LJW^{+}22, LLLW22, MYB19, NS23, NZW23, Pap15, PLS^{+}21, PL10, QZL23, RABR23, DMG^{+}21, SAS16, TL14, TSS22, WWHW19, WWW^{+}20, WLL^{+}21, WJLL21, WLHH23, WLH24, WC15, WLS^{+}23, WGYC21, WLG^{+}24, XZXW18, YLW^{+}19, YWZ^{+}24, YHL15, YTW^{+}16, ZTL15a, ZS24, ZTL^{+}15b, ZCF^{+}17, ZH23, ZFH23, ZFWC18, ZL15, ZPC^{+}16, ZLC20, ZMZS23, CT14, KSM09, LCZ^{+}09, LSY^{+}09, LTH^{+}13, ML14, YWW^{+}14, ZFY14]. Socializing [JSV^{+}15]. Soft [SZD^{+}23]. Solution [LFY^{+}23, WSZ^{+}16, YCJK08]. Solutions [DDL21, ZTT^{+}20]. Solving [GMSS13]. Source [WN22b, YHZ^{+}24, ZGY^{+}22, WLW^{+}19]. SP [ZYD^{+}24a]. Space [AH24, CYY^{+}22, DDL24, HQYY14, JSP15, LHL^{+}22, MYL^{+}21]. Space-Efficient [JSP15]. Spaces [CV24, GQH22, ZWLS22]. spam [ZP09]. Span [GCB^{+}21]. Span-core [GCB^{+}21]. SPAP [WZR^{+}23]. Sparse [BGSW13, BXF^{+}20, Bur21, CNY^{+}16, CLY12, FH14, FKKD17, LLW16, LWLW18, LCG^{+}18, PFS15, STD^{+}18, DTU21, TR22, WYG^{+}17, WXB^{+}21, WHMY17, XYW^{+}20, YWC24, ZHT20, ZSC^{+}23, ZWWH20, ZZZ^{+}20a, ZLW21b]. Sparsely [HXY^{+}19, MA16]. Sparsification [AAB^{+}24, ZZF24]. Sparsity [YTH18]. Spatial [FLL^{+}22, LLL^{+}23b, LCLL17, MXP20, NXZ^{+}24, OJW^{+}23, SYLC16, WZZ^{+}22, WZWC23, WGC^{+}23, WJR^{+}10]. Spatial-aware [NXZ^{+}24]. Spatial-Proximity [SYLC16]. Spatial-Temporal [FLL^{+}22, WZZ^{+}22, WZWC23, LLL^{+}23b, OJW^{+}23]. Spatio [CLT^{+}20, CHY^{+}24, DCF^{+}21, LYC^{+}23, LCX^{+}23, QSST18, SYZ^{+}24b, ZZXY21, ZGY^{+}22]. Spatio-Temporal [CHY^{+}24, LYC^{+}23, LCX^{+}23, QSST18, SYZ^{+}24b, ZZXY21, ZGY^{+}22, CLT^{+}20, DCF^{+}21]. Spatiotemporal [LZY^{+}24a]. Speaker [GSI19]. Special [Agg17, EK12, GSTC12, GL15, SLTA11, TWCO16, Wan10, WL16, XL20, ZPSYY10, vLCV^{+}18, vLCV^{+}19, BBD^{+}07, CT14, LSY^{+}09, MG09, ZKYW08, ACPW13]. Specific [AFZ22, GCL^{+}19b, HZW^{+}15, LL19, LHL^{+}22, ZFW22]. Spectral [HST^{+}23, LHK^{+}18, MM12, TLL^{+}19, YDE^{+}22, ZZYY16, ZZF24, CSZ^{+}09]. Spectrum [ASJ^{+}23, ZZF24]. Spectrum-Preserving [ZZF24]. Speed [DGB16]. Sports [DKSL18, VALF12]. spread [KSM09]. SpTV [YLL19]. SR [CXL^{+}24]. SsAG [AAB^{+}24]. STA [OJW^{+}23]. STA-TCN [OJW^{+}23]. Stability [HK18]. Stable [KZW^{+}22, SLWW24]. Stacked [LLY^{+}24, ZLW21b]. STAD [ZLD^{+}23]. STAD-GAN [ZLD^{+}23]. Stage [ALW^{+}24, HSY^{+}21, SLL^{+}23, RYM^{+}24, YCJK08]. Standardised [GLMW22]. Star [PCPN24, ZWC^{+}23]. Start [CC19, LYTZ22, LGZ^{+}21, ML15, WZL^{+}23a].

State [Agg22]. Static [ANK14, AH24, JK23, RU18, WCZ^{+}24]. stationary [CZY^{+}22]. Stations [LH22, WGC^{+}23]. Statistical [AF16, GNMQ21, LK15]. Statistics [CCV19, LNG18, MXP20, WLR^{+}14]. Status [CZW^{+}24, LYL^{+}20]. Statutes [FLG^{+}21]. STGCN [WZWC23]. STHAN [LYC^{+}23]. Stochastic [CWY^{+}23, IAB22, LYC^{+}21, TC18, VAFZ19, YLL^{+}21a]. Stock [GYX^{+}22, ZZY^{+}24b]. Store [LGZ^{+}21]. Stories [LHN^{+}20]. Story [LHN^{+}20]. Strategies [ABW20, CWL^{+}24a, SB24, Var22, WLT19, WCX24]. Strategy [LHZG13, ZZY24a]. Stratification [ASJ^{+}23]. Streaks [ZJL^{+}14]. Stream [PE20, PZW^{+}18, THB18, TC09, WCZ^{+}22, WHC^{+}23, YSY^{+}22]. Streaming [ANK14, HQWW22, JK23, JSP15, KBR^{+}16, PKSM21, PLS^{+}21, SÖ22, YDS^{+}15, ZSM22, ZZYW22, CCC09, CKMS08]. Streams [BLH^{+}22, BLW14, BC18, CFP19, DERU17, DGB16, GBBC24, GL22, HKR^{+}23, HAK^{+}23, HSB22, JH19, KG24, LJK18, OLL20, RKC19, SFP10, SOK^{+}20, SLO^{+}21, DTU21, SAS16, ZGY^{+}24, WND^{+}09]. Strengthen [HDT^{+}18]. String [BCC^{+}21, II08]. Strong [LLLW22]. StructCoder [TZR24]. Structural [ASZ21, CZY11, JHJK22, JS21, LWG^{+}16, MW24, RJK^{+}20, ZHL^{+}21, ZH16]. Structure [CSG^{+}16, CS24, GW20, JH22, LWS22, LHW^{+}21, LLD^{+}24, LYL^{+}22, LXR^{+}23, LF23, MRJ11, MZC^{+}24, RABR23, RBBV18, SG12, SWD^{+}21, TZR24, WPW23, WJLY23, WLHH23, YLL^{+}24, ZWYW24, ZHS^{+}19, ZZY^{+}21, ASHK14]. Structure-Aware [TZR24, ZWYW24]. Structure-Driven [WJLY23]. Structure-Enhanced [LXR^{+}23]. Structure-Information-Based [MZC^{+}24]. Structured [FH14, GZXF16, GWZZ17, WPW23]. Structures [BB17, JPJ^{+}22, WDH^{+}24].

Student [LZZ^{+}22b]. Study

[SLK^{+}24, ZZW09]. style [ZLY^{+}20b]. Sub [GLG^{+}22]. Sub-tracks [GLG^{+}22]. Subclasses [NWA20]. Subgraph [CLJC24, NADR21, SSL^{+}23, WLS^{+}23, TVK10]. Subgraph-Based [WLS^{+}23]. Subgraphs [BBC^{+}24, hDRNlD21, PSP^{+}18, PE20, RTG17]. Subgroups [RV20]. subjective [HGV^{+}08]. Subsampled [KZW^{+}22]. Subsequence [SGM^{+}24]. Subsequence-based [SGM^{+}24]. Subsequences [LYWL12, RCM^{+}13]. subsets [SJR08]. Subspace [AT17, GFHL23, GFL^{+}24, HSBH19, MYPB20, NC22, JTYY10, KKZ09]. Substantial [WAD12]. Substitutes [ZWG^{+}19]. Subtensor [SHF18]. Subtensors [hDRNlD21]. subtrees [THD^{+}08]. Success [KZZ^{+}20]. Succinct [BLP21]. succinctly [MVT12]. Sufficient [LWW18, Web10]. Suffix [dVKCC11]. Suggest [LML^{+}16]. Suite [RET^{+}23]. Sum [WHG^{+}18]. summaries [KT09]. Summarization [AAB^{+}24, JPJ^{+}22, KKB22, LOW^{+}21, NTNP18, PA18, QCD^{+}19, WZL^{+}11, WZLG13, ZLSC24, WZLG12]. Summarizing [MVT12, WWLM23]. Supervised [CRC^{+}23, CLJC24, CQH^{+}24, DWL^{+}24, JH22, JRC^{+}23, JHC^{+}23, LYT^{+}23, LWH^{+}24, MG24, NGMQ22, TR22, WN22a, XLS24, YWS^{+}22, CCLZ18, CWWL24, Dor21, DIB24, GLW^{+}23, Han24, JBGW24, LSL^{+}22, LJG^{+}24, RZY^{+}23, WMZL24, YWL^{+}24, ZWYZ24]. Supervision [BGJV12, GLW^{+}23, NGMQ22, WSR^{+}16, SM23]. Support [PBMID14, RWD23, SMS22, AGHN13, CX10]. Supporting [LZF^{+}15]. Surgical [MFRJ16]. Survey [BG21, CRP^{+}24, CE24, EE24, GLFV^{+}19, GA22, KP18, KR16, LRK^{+}19, LGG^{+}23, LZV24, LBT^{+}23, LWW^{+}23b, MZC^{+}24, OCD^{+}24, PDK24, SI21, SB21b, WZLZ23, WTZL22, YYS^{+}24, YJT^{+}24, YCL^{+}21, ZGQ^{+}24, KKZ09]. Susceptibilities [LSZ^{+}19]. Susceptibility [ACK^{+}22, ZS24]. swap [GMMT07]. Swarm [JBGW24, XXZ19]. Sybils [YWW^{+}14]. Symbolic [CG23]. Symmetrization [ZZF24]. Sync [YWR^{+}19]. Synchronization [LZD16]. Synchronization-Core-Based [LZD16]. Synchronized [JCB^{+}16]. Synopsis [WCZ^{+}22]. Synthesis [BTBg22]. Synthetic [SB21a, WNC^{+}18]. System [CNZ^{+}17, CKC^{+}18, ELH23, GSWJ20, JDE^{+}12, LSZ^{+}24, LH22, LGS^{+}23, SRBC22, SYG^{+}24, TYG^{+}15, WSDL19]. Systematic [GHT^{+}24, PGR18, PTL22]. Systems [Che18, CC19, CZZ^{+}23, FN24, ABDM21, LSH^{+}22, LYG^{+}24, MAHT18, PSP^{+}18, PDK24, WGL^{+}23, WCX24, XML18, XZLL21, ZWC22, ZYS^{+}21].

Tabular [SM23]. Tag [YWR^{+}19]. Tagged [CLYC23]. Tail [HCC^{+}18, ZZY24a]. Tailed [ZLHZ24]. TAP [LCX^{+}23]. target [ZP09]. Targeted [HGY24]. Targeting [CPC10]. Task [DKSK22, GMS21, GHT^{+}24, LCX^{+}23, MAHT18, OGV22, SYLC16, SYG^{+}24, WDH^{+}24, WMW^{+}22, XSZ^{+}22, YYC^{+}21, ZY14, ACE20, YLHY20, ZYH19]. Task-Oriented [XSZ^{+}22]. Tasks [BGJV12, CLY12, LZZ^{+}22b, MPG^{+}23, TYW^{+}21]. TaSPM [HGY24]. Taxonomy [GCL^{+}19a, YW24, TLZ^{+}08]. TCN [OJW^{+}23]. TDAN [WLH24]. Teaching [LWH^{+}24]. team [VALF12]. Technical [LWH^{+}23]. Technique [AEEBEE20, CMS23, GXL24, LN24, SKM^{+}22, YLL^{+}21a]. Techniques [RJK^{+}20, WZLZ23]. Technologies [QSS20]. TechPat [LWH^{+}23]. Telling [LHN^{+}20]. Temporal [ADR21, ASJ^{+}23, BMZY21, CYOL16, CTC^{+}22, CG23, CLJC24, CHY^{+}24, CNZ^{+}17, CYT^{+}17, DKA11, FLL^{+}22, GCB^{+}21, JYH^{+}23, LLC^{+}22, LWC^{+}23, LZY^{+}24a, LYC^{+}23, LSZ^{+}19, LCX^{+}23, QSST18, RLDL24, SYZ^{+}24b, TC18, WZZ^{+}22, WZWC23, WC15, XJW^{+}21, XSZ^{+}22, ZZXY21, ZLT^{+}23, ZGY^{+}22, ZH16, CLT^{+}20, DCF^{+}21, LLL^{+}23b, OJW^{+}23, RGL^{+}23]. Tendency [BMZY21, KBR^{+}16]. Tendency-Guided [BMZY21]. Tensor [ASZ21, BNY20, DDY^{+}24, DKA11, HW21, HQWW22, PFS15, SCS14, SGCH19, WLP18, ZSC^{+}23, STP^{+}08]. Tensor-Based [ASZ21]. Tensorizing [JSG^{+}19]. Tensors [FLC23, JK23, YLL19]. Tenuous [SSL^{+}23]. Term [HZL20, JHC^{+}23, ZGC18, TWC22]. Test [KZW^{+}22, MAHT18, ZXZ^{+}23, ZWYZ24]. Test-time [ZWYZ24]. Testing [LZF^{+}15]. Text [BES15, BAMK18, CWW^{+}24, ES15, GBBC24, GA22, JLZ^{+}22, KL23, LWWW24, MHKG19, MN20, MES22, QCD^{+}19, WPW23, WPD^{+}24, YGL^{+}22, II08]. Text-level [LWWW24]. Text-Only [CWW^{+}24]. Their [BGWSB19, LCZ^{+}09, ZBAG20]. Theoretic [BES15, XKH^{+}16]. Theory [CT14, GBGL20, Jia24, VLK^{+}23, YWC24, YDW^{+}23, ZWSD24, STP^{+}08]. Things [NGB18]. Third [HWZ^{+}24]. Thousands [WCZ^{+}24]. Three [LLY^{+}21, LLZ^{+}23, LDS^{+}22, PPDSBLP16, RYM^{+}24, WLL^{+}22b, YLL19, ZYC^{+}24]. Three-Order [YLL19]. Three-stage [RYM^{+}24]. Three-Tier [LDS^{+}22]. Three-View [ZYC^{+}24]. Three-Way [WLL^{+}22b, LLZ^{+}23]. Threshold [SB24]. Thresholded [GJ16]. Thresholding [ZLZ^{+}19]. Tie [SLL^{+}23]. Tied [MNK18]. Tier [LDS^{+}22]. Tiered [DTU21]. Ties [HDT^{+}18, MYB19]. Tight [RU14]. Time [ABWI23, AOEM17, ABSP^{+}18, ABW20, AZBW21, BLH^{+}22, CRM^{+}23, CE24, CSZ^{+}21, CHZ23, CYL^{+}24, DSL^{+}24, DDY^{+}24, GXLC21, GYX^{+}22, GSST16, HBD^{+}24, KHTR18, LLC^{+}22, LWC^{+}23, LLY^{+}24, LCWC20, LTB18, MGL^{+}20, MH22, MFRJ16, OCD^{+}24, PZJ^{+}24, QXBT16, RET^{+}23, RCM^{+}13, SHL19, SBZR19, SYK22, SLW^{+}18, SGM^{+}24, TBW23, TR22, WSDL19, WBL21, XSZ^{+}22, XZ23, YWR^{+}19, YZZ22, ZHT20, ZLD^{+}23, ZLZ^{+}24, ZSM22, ZJJ^{+}24, LSH20, ZWYZ24]. Time-Aware [GXLC21, GYX^{+}22, XSZ^{+}22, CSZ^{+}21]. Time-Series [GSST16, LLC^{+}22, MFRJ16, PZJ^{+}24, SHL19, TR22, CRM^{+}23, CE24, MH22, ZJJ^{+}24]. Time-Sync [YWR^{+}19]. Time-Varying [QXBT16, WBL21, YZZ22]. Time-Warped [ZHT20]. Times [VFA^{+}15]. Timescales [IYSU12]. TipTap [NADR21]. TKDD [CT14, vLCV^{+}19, MG09, TWCO16, Wan10, ZPSYY10, vLCV^{+}18]. TKDD-CASIN [CT14]. Today [BT21]. Together [PPDSBLP16]. Token [OJLD22]. Token-Blocking [OJLD22]. Tolerant [MGL^{+}20]. TOMGPT [CWW^{+}24]. Top [ML15, SB24, BBC^{+}19]. Top- [SB24, ML15, BBC^{+}19]. Topic [GXL24, GCG22, HDQ^{+}18, HLCR20, HF12, IAB22, IYSU12, JLZ^{+}22, LZP20, QCD^{+}19, TLZ^{+}08, WZL^{+}23c]. Topic-Disentangled [WZL^{+}23c]. Topics [LML^{+}16, PZW^{+}18]. Topology [CS24, DIB24, GWZZ17, WZL^{+}22]. Total [BBC^{+}24]. Totally [ZLGY24]. Totally-ordered [ZLGY24]. Tourism [SMS22]. Tourist [CLYC23]. TPmod [BMZY21]. TRACE [FYN22]. Traceable [XWW^{+}24]. Tracing [FSXL23, WMZL24]. Tracking [BBB^{+}22]. tracks [GLG^{+}22]. Tradeoff [QZB^{+}23]. Trading [LYGG22, WZZ^{+}24]. Traffic [DMGC21, CLT^{+}20, DWH^{+}23, HW21, JDE^{+}12, LFY^{+}23, LJL^{+}23, LLL^{+}23b, LHZ^{+}24a, LCX^{+}23, RLDL24]. Trafficking [VLK^{+}23]. Training [CWW^{+}24, LZL^{+}22, LWH^{+}24, WWZL23, WOGZ24, XWZ^{+}22, ZWYZ24, WMZL24, ZLD^{+}23]. Trajectories [LQW15, MXP20, RTM18, TYG^{+}15, ZBL^{+}20]. Trajectory [CHY^{+}24, DDL21, GLG^{+}22, HXY^{+}19, JCC23, XLQ^{+}24]. Trajectory-User [CHY^{+}24]. Transaction [WCZ^{+}22]. Transactions [Agg22]. Transduction [MES22]. Transductive [MES22]. Transfer [AHGA14, AZBW21, BGJV12, DKSK22, JYH^{+}23, Jia24, LSS^{+}22, LXC^{+}22, LYTZ22, LYN^{+}24, LGZ^{+}21, LLX^{+}24, MES22, SZF^{+}23, WNH15, XWZ^{+}22, ZFY14]. Transferable [RYM^{+}24, SKW^{+}21, WLH24]. Transference [OGV22]. Transferring [WYWY19]. Transform [SM21, SCCM22]. Transformation [HYYQ15, LLL^{+}23b, LTB18, THYL24, XWW^{+}24]. Transformation-Based [LTB18]. Transformer [LZL^{+}23, TR22, TZR24]. transforms [PSFV13]. Transit [ELH23, LH22]. Transitivity [JSP15]. Translated [WSZ^{+}16]. Translations [DA19]. Transmission [GSI19, ZHSL23]. Transparency [GBW^{+}24]. Transport [DCF^{+}21]. Transportation [LH22, LYC^{+}23, LSH20]. Travel [FYN22]. Traversal [TAJY17]. Treatment [CSZ^{+}21, HML^{+}24, KCL^{+}20, ZBX^{+}24]. Tree [LLB^{+}22, MW24, RZF^{+}24, THD^{+}08, ZH21, Vad10]. Tree-Based [ZH21]. Trees [MRJ11, SI21, HAKU^{+}08, LCZ07, VCKP08]. Trends [KP18]. Triadic [HDT^{+}18, YWR^{+}23, LTH^{+}13]. Triangle [CC12, JSP15, PPDSBLP16, SOK^{+}20, SLO^{+}21, YSY^{+}22, BBCG10]. Triangle-Driven [PPDSBLP16]. Triangles [AKM20, DERU17, LJK18]. TRIÈST [DERU17]. Trifactorization [WNH15]. Triggered [PSYJ24]. Trillion [PSP^{+}18]. Trillions [RCM^{+}13]. Trip [CCTW23, LCZ^{+}24]. Triplet [SWY^{+}24]. Truncated [TLG^{+}23, YMD^{+}24]. Truss [CCG^{+}24]. Truss-based [CCG^{+}24]. Trust [GXLC21, HZZ^{+}15, HNH^{+}13]. Trustworthy [CXL^{+}24, ZGQ^{+}24]. Truth [WZX^{+}23, WZB^{+}23]. TS [ABWI23]. TTS [DDY^{+}24]. TTS-Norm [DDY^{+}24]. Tucker [JK23]. Twain [HDC07]. Tweet [LLD^{+}24].

Twitter [BPW^{+}18, CFP19, MCS^{+}18, VMR17]. Two [Dor19, SG12, HDC07]. Two-end [HDC07]. Type [LYZ^{+}24, WLS^{+}24, YTW^{+}16].

Ubiquitous [XZSY19]. Ultra [MW24, WHMY17]. Ultra-High [WHMY17]. Ultra-Large [MW24]. Unbiased [WZL^{+}15, ZH21]. Uncertain [AF13, LZX22]. Uncertainties [CHZ23]. Uncertainty [AZD^{+}21, BAC23, XLQ^{+}24]. Uncertainty-aware [XLQ^{+}24]. Uncoupled [NK20]. Uncovering [BBB^{+}22, CRGP14, WLHH23, YWW^{+}14, ZHSL23]. Underground [ZCL19]. Understanding [JHJK22, SGM^{+}24, WZL^{+}22, WGC^{+}23, XZLL21]. Undirected [MRTW19]. Unification [AF16]. Unified [GF23, LSW^{+}24, MHS20, PFTR16, SI21, WLW^{+}19, WZL^{+}23c, YLL21b, ZWWH20]. uniform [PV24]. Uniqueness [DSTA22]. Univariate [PC20]. Universal [VFA^{+}15]. Unknown [KZW^{+}22]. Unlabeled [WPDZ21]. Unobserved [YMX^{+}23]. Unpaired [YWC24]. Unsupervised [CNY^{+}16, GFHL23, HBD^{+}24, KG24, KCR23, KR16, MCSZ20, NC22, PZW^{+}18, SLK^{+}24, SZD^{+}23, SB21a, WWW^{+}16, YLHY20, ZLD^{+}23, ZRJ^{+}22]. Untrustworthy [TYG^{+}15]. Unweighted [BBC^{+}19]. Unwilling [GW21]. Urban [DCF^{+}21, ELH23, LGZ^{+}23, LZLJ24, RLDL24, WSDL19, XZSY19, XSZY20, ZCC^{+}24]. US-Rule [HGWY23]. Usage [SDS18, WLD^{+}21]. Use [SH15]. Used [WFW^{+}11]. User [Che18, CHY^{+}24, DWW22, FN24, GFM21, GBTL14, HHL^{+}24, HLL^{+}20, JHC^{+}23, KHTR18, LL19, LHZ^{+}24a, LYG^{+}24, LWG^{+}16, LZL^{+}24, LZLJ24, MYB19, NTNP18, QHW22, RTM18, Row16, SYZB24, SNH^{+}13, TYZZ10, WCL19, WZL^{+}23a, WWLM23, XYZ^{+}20, YCC^{+}15, ZTL15a, ZFY14]. User-Centric [RTM18]. User-Guided [SNH^{+}13].

User-item [WWLM23]. User-Preferred [LZL^{+}24]. User-Specific [LL19]. Users [AEEBEE20, BGWSB19, HCZ^{+}14, LPK^{+}15, LT10, ML15, DMG^{+}21, ZS24]. Using [AFZ22, AOEM17, ASPT21, CYY^{+}22, CTC^{+}22, CLYC23, CTZ16, DKA11, HZZ^{+}15, JSP15, KP18, LZL^{+}24, MHS20, MN20, MG24, PE20, PXW^{+}22, PNB24, QZL23, RKR18, RABR23, SATK20, STD^{+}18, SB24, SAS16, THB18, WNH15, WYG^{+}17, XKH^{+}16, YWR^{+}19, YMD^{+}24, YWC24, ZB20, ZYC^{+}24, ZGY^{+}22, dVKCC11, CGL^{+}23, HNH^{+}13, II08, KUU10, LSH20, SMS22, SJR08, ZP09]. Utility [BTBg22, BES15, GLZ^{+}21, HGWY23, LDS^{+}22, SB24, WWLM23, WTD24, WFW^{+}11, WLT19, XLD^{+}24, ZDY^{+}22, ZYD^{+}24a, ZLGY24]. Utility-driven [HGWY23]. Utility-Oriented [XLD^{+}24]. Utility-Theoretic [BES15]. utilizing [SMA^{+}08].

Vaccine [ZP15]. Valence [SLL^{+}23]. Valid [DPB^{+}20]. validity [HGV^{+}08]. Value [BJ21, MGL^{+}20, ME11, OJLD22, WPD^{+}24]. Values [WHG^{+}18]. Vanishing [CNZ^{+}17]. Variable [AZBW21, LZL^{+}22, OGAB14, SLW^{+}18, YWK^{+}22, YMX^{+}23, ZCS10]. Variable-Driven [YMX^{+}23]. Variable-lag [AZBW21]. Variable-Length [SLW^{+}18]. Variance [AZD^{+}21, CWR23, MNK18]. Variance-aware [CWR23]. Variant [HLC19]. Variate [HBD^{+}24]. Variation [ZLZ^{+}19]. Variational [CZH18, IAB22, LLL23a, ZWC22]. Varied [WGC^{+}23]. Various [LYWL12]. Varying [ACK^{+}22, BSS^{+}24, QXBT16, WBL21, YZZ22]. Vector [CXH^{+}20, JLD^{+}19, PBMID14, RWD23, SMS22, WXW^{+}21, AGHN13]. Vectors [CMS24, ZB20]. Vehicle [WZR^{+}23]. Veracity [VMR17]. Version [DDL24]. Vertical [HWZ^{+}24, ZFL^{+}24]. vertically [MWF08, VCKP08]. Very [GJ16, LK20, AF09]. Via [LHZ^{+}24a, MAHT18, YSW^{+}21, ZXL22, ZZY^{+}23, ACK^{+}22, BNY20, BLP21, CWL^{+}24a, CGL^{+}23, CCT^{+}23, CLJC24, CHY^{+}24, CNZ^{+}17, DDY^{+}24, DWD^{+}24, DYS20, DLZ^{+}24, EC20, FLL^{+}22, FSXL23, GMMT07, GMSS13, HQWW22, JYY^{+}21, JHC^{+}23, KCL^{+}20, LLZ^{+}23, LFXC24, LLH^{+}24, LSS^{+}11, LLL^{+}20, LHL^{+}22, LXR^{+}23, LCX^{+}23, MWF08, ML15, MRTW19, NC22, NK21, NK20, Pap15, QWH^{+}24, RZY^{+}23, SYXW22, SMW^{+}24, TLZ^{+}20, TLL^{+}19, TKT^{+}24, TYW^{+}21, THYL24, WZLG12, WZLG13, WNH15, WYWY19, WDL^{+}21, WDF22, WLP18, WZB^{+}23, WZL^{+}23c, WLG^{+}24, XLQ^{+}24, XML18, XLT^{+}20, YHLC23, YCJK08, YSG^{+}21, YWK^{+}22, ZZ10, ZTL^{+}15b, ZLZ^{+}19, ZSC^{+}23, ZZF24, ZLZ^{+}24, ZWC^{+}23, ZYH19, ZZYW22, ZJJ^{+}24]. Vibrancy [LGZ^{+}23]. Video [LLL^{+}20, YWR^{+}19]. View [AFW^{+}21, CWL^{+}24b, CFX^{+}24, DWH^{+}23, GF23, GFHL23, GFL^{+}24, LSF18, LZG^{+}24, LYN^{+}24, LSW^{+}24, LXY^{+}17, LF18, LZX22, LGZ^{+}23, LFL^{+}23, LF23, PZW^{+}18, SZD^{+}23, SKW^{+}21, WCL^{+}23a, WZB^{+}23, YLL21b, ZYC^{+}24, CWWL24, HTT^{+}24, Han24, HCH^{+}24b, LWCZ24, LLG^{+}23, MHS20, YLHY20, ZSC^{+}23, ZZZ^{+}20b]. View-guided [LZG^{+}24]. Viewable [RBBV18]. Viral [ZPC^{+}16]. Virtual [NZW23]. Vis [MCS^{+}18]. Visibility [LHS^{+}21]. Vision [LWG^{+}22]. Vision-Language [LWG^{+}22]. VisIRR [CKC^{+}18]. Visual [CLYC23, CKC^{+}18, MCS^{+}18, PA18, SGM^{+}23, STD^{+}18]. Visualization [CMZS15, FAV23, KBR^{+}16, SHF24]. Visualizing [DA18, WBL21, YWC^{+}16]. VOGUE [ZCS10]. Voice [ALW^{+}24, GSI19]. Vol [vLCV^{+}19]. VOT [GLG^{+}22]. VOT-based [GLG^{+}22]. Vote [LTB18]. Voting [LJX23]. Voting-Based [LJX23]. Voxel [LWB^{+}24]. Voxel-Wise [LWB^{+}24].

Vulnerability

[GBTL14, LWB^{+}22, YHW^{+}24].

Wafer [PSYJ24]. Walk [NS23]. Walks [MRTW19, NK20]. Warmth [GSI19]. Warnings [WCZ^{+}24]. Warped [ZHT20]. Warping [RCM^{+}13]. Wavelet [LLL^{+}23b, THB18]. Way [DDY^{+}24, WLL^{+}22b, LLZ^{+}23]. Weak [LWY16]. Weakly [DWL^{+}24, GLW^{+}23]. Weakly-supervised [GLW^{+}23]. Wealth [YWG^{+}22]. Web [NTNP18, SMA^{+}08, TYZZ10, WNH15, ZPSYY10]. Webpages [AS21]. weight [HYYT21]. Weighted [BBC^{+}24, BC18, CG15, DAR09, LHZG13, LJX23, LGZ^{+}21, WHC^{+}23, ZZZ^{+}20b]. Weighting [DIB24]. Weights [CC19, ZFH23]. Who [LGZ^{+}22, ZTL^{+}15b]. wide [ZZW09]. wild [YWW^{+}14]. Will [HDT^{+}18, LGZ^{+}22]. Win [LGZ^{+}22]. Wind [FFZ^{+}24]. Window [MWL^{+}22]. Wisdom [DMI12]. Wise [LWB^{+}24, THYL24, XWW^{+}24, YZZ22]. within [ELH23]. without [HWZ^{+}24, WTD24]. Word [JXGZ19, QWH^{+}24, II08]. Worker [WMW^{+}22]. Worker/Task [WMW^{+}22]. Workers [TYW^{+}21]. Workload [JSE^{+}23]. World [CS24, WSR^{+}16, XZ23, XXO^{+}24, CYL^{+}24]. Wrapper [CMQ19].

X [LYGG22, PAS24, RGAÁC^{+}24].

X-distribution [PAS24]. X-FSPMiner [RGAÁC^{+}24]. Xeljanz [AEEBEE20]. XML [THD^{+}08].

Years [MHKG19].

Zen [CV24]. Zero [HP20, RYM^{+}24]. Zero- [RYM^{+}24].

References

Ali:2024:SSS

[AAB^{+}24] Sarwan Ali, Muhammad Ahmad, Maham Anwer Beg, Imdad Ullah Khan, Safiullah Faizullah, and Muhammad Asad Khan. SsAG: Summarization and sparsification of attributed graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):141:1–141:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3651619.

Alam:2023:DIP

[AASL23] Md. Tanvir Alam, Chowdhury Farhan Ahmed, Md. Samiullah, and Carson Kai-Sang Leung. Discovering interesting patterns from hypergraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):1– 34, October 2023. ISSN 1556-472X.

Jurdi:2021:CNN

[ABDM21] Wissam Al Jurdi, Jacques Bou Abdo, Jacques Demerjian, and Abdallah Makhoul. Critique on natural noise in recommender systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):75:1–75:30, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447780.

Amelkin:2019:DMA

[ABS19] Victor Amelkin, Petko Bogdanov, and Ambuj K. Singh. A distance measure for the analysis of polar opinion dynamics in social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(4):38:1–38:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3332168.

Amornbunchornvej:2018:CED

[ABSP^{+}18] Chainarong Amornbunchornvej, Ivan Brugere, Ariana Strandburg-Peshkin, Damien R. Farine, Margaret C. Crofoot, and Tanya Y. Berger-Wolf. Coordination event detection and initiator identification in time series data. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5):53:1–53:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Amornbunchornvej:2020:FIF

[ABW20] Chainarong Amornbunchornvej and Tanya Berger-Wolf. Framework for inferring following strategies from time series of movement data. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):35:1–35:22, May 2020.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3385730.

Aleryani:2023:MIE

[ABWI23] Aliya Aleryani, Aaron Bostrom, Wenjia Wang, and Beatriz Iglesia. Multiple imputation ensembles for time series (MIE-TS). ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):44:1–44:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551643.

Akhtar:2020:DMT

[ACE20] Md Shad Akhtar, Dushyant Singh Chauhan, and Asif Ekbal. A deep multi-task contextual attention framework for multimodal affect analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):32:1–32:27, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3380744.

Abebe:2022:ODO

[ACK^{+}22] Rediet Abebe, T.-H. Hubert Chan, Jon Kleinberg, Zhibin Liang, David Parkes, Mauro Sozio, and Charalampos E. Tsourakakis. Opinion dynamics optimization by varying susceptibility to persuasion via non-convex local search. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):33:1–33:34, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3466617.

Agarwal:2013:ISI

[ACPW13] Deepak Agarwal, Rich Caruana, Jian Pei, and Ke Wang. Introduction to the Special Issue ACM SIGKDD 2012. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3):9:1–9:??, September 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Afrati:2016:APD

[ADK^{+}16] Foto Afrati, Shlomi Dolev, Ephraim Korach, Shantanu Sharma, and Jeffrey D. Ullman. Assignment problems of different-sized inputs in MapReduce. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 18:1–18:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ahmed:2021:OST

[ADR21] Nesreen K. Ahmed, Nick Duffield, and Ryan A. Rossi. Online sampling of temporal networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):59:1–59:27, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442202.

Abd-Elaziz:2020:EDM

[AEEBEE20] M. M. Abd-Elaziz, Hazem M.

El-Bakry, Ahmed Abou Elfetouh, and Amira Elzeiny. Enhanced data mining technique to measure satisfaction degree of social media users of Xeljanz drug. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):33:1–33:13, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3389433.

Angiulli:2009:DEA

[AF09] Fabrizio Angiulli and Fabio Fassetti. DOLPHIN: an efficient algorithm for mining distance-based outliers in very large datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(1):4:1– 4:??, March 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Angiulli:2013:NNB

[AF13] Fabrizio Angiulli and Fabio Fassetti. Nearest neighborbased classification of uncertain data. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(1):1:1–1:??, March 2013. CODEN ????

ISSN 1556-4681 (print), 1556-472X (electronic).

Angiulli:2016:TGU

[AF16] Fabrizio Angiulli and Fabio Fassetti. Toward generalizing the unification with statistical outliers: The gradient outlier factor measure. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (3):27:1–27:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ata:2021:MVC

[AFW^{+}21] Sezin Kircali Ata, Yuan Fang, Min Wu, Jiaqi Shi, Chee Keong Kwoh, and Xiaoli Li. Multi-view collaborative network embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):39:1–39:18, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441450.

Abulaish:2022:DSK

[AFZ22] Muhammad Abulaish, Mohd Fazil, and Mohammed J. Zaki. Domain-specific keyword extraction using joint modeling of local and global contextual semantics. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):70:1–70:30, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494560.

Aggarwal:2017:ISI

[Agg17] Charu C. Aggarwal. Introduction to special issue on the best papers from KDD 2016. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):39:1–39:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Aggarwal:2022:CEC

[Agg22] Charu C. Aggarwal. Communication from the Editorin-Chief: State of the ACM Transactions on Knowledge Discovery from Data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):21e:1–21e:2, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3463950.

Ang:2013:CPN

[AGHN13] Hock Hee Ang, Vivekanand Gopalkrishnan, Steven C. H. Hoi, and Wee Keong Ng. Classification in P2P networks with cascade support vector machines. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(4): 20:1–20:??, November 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Amagata:2024:EDP

[AH24] Daichi Amagata and Takahiro Hara. Efficient density-peaks clustering algorithms on static and dynamic data in Euclidean space. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):2:1–2:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3607873.

Acharya:2014:OFC

[AHGA14] Ayan Acharya, Eduardo R. Hruschka, Joydeep Ghosh, and Sreangsu Acharyya. An optimization framework for combining ensembles of classifiers and clusterers with applications to nontransductive semisupervised learning and transfer learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1): 1:1–1:??, August 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ahmed:2015:AMC

[AK15] Rezwan Ahmed and George Karypis. Algorithms for mining the coevolving relational motifs in dynamic networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):4:1–4:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Anderson:2017:AHE

[AKM17] Ashton Anderson, Jon Kleinberg, and Sendhil Mullainathan. Assessing human error against a benchmark of perfection. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (4):45:1–45:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Altowim:2018:PAP

[AKM18] Yasser Altowim, Dmitri V.

Kalashnikov, and Sharad Mehrotra. ProgressER: Adaptive progressive approach to relational entity resolution. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):33:1–33:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Arifuzzaman:2020:FPA

[AKM20] Shaikh Arifuzzaman, Maleq Khan, and Madhav Marathe. Fast parallel algorithms for counting and listing triangles in big graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):5:1–5:34, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3365676.

Agichtein:2009:MIS

[ALB09] Eugene Agichtein, Yandong Liu, and Jiang Bian. Modeling information-seeker satisfaction in community question answering. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(2): 10:1–10:??, April 2009. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ao:2024:PYV

[ALW^{+}24] Xiang Ao, Ling Luo, Xiting Wang, Zhao Yang, Jiun-Hung Chen, Ying Qiao, Qing He, and Xing Xie. Put your voice on stage: Personalized headline generation for news articles. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):54:1–54:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3629168.

Adali:2013:IPR

[AMIL13] Sibel Adali, Malik Magdon-Ismail, and Xiaohui Lu. iHypR: Prominence ranking in networks of collaborations with hyperedges 1. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(4):16:1–16:??, November 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Angiulli:2020:CCF

[Ang20]  Fabrizio Angiulli.  CFOF:

a concentration free measure for anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(1):4:1–4:53, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3362158.

Ahmed:2014:NSS

[ANK14] Nesreen K. Ahmed, Jennifer Neville, and Ramana Kompella. Network sampling: From static to streaming graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(2):7:1–7:??, June 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Algizawy:2017:RTL

[AOEM17] Essam Algizawy, Tetsuji Ogawa, and Ahmed El-Mahdy. Real-time large-scale map matching using mobile phone data. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):52:1– 52:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Asur:2009:EBF

[APU09] Sitaram Asur, Srinivasan Parthasarathy, and Duygu Ucar. An event-based framework for characterizing the evolutionary behavior of interaction graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4): 16:1–16:??, November 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Alarte:2021:PLM

[AS21] Julián Alarte and Josep Silva. Page-level main content extraction from heterogeneous Webpages. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (6):105:1–105:105, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451168.

Abrahao:2014:SFA

[ASHK14] Bruno Abrahao, Sucheta Soundarajan, John Hopcroft, and Robert Kleinberg. A separability framework for analyzing community structure. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(1):5:1–5:??, February 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Atyabi:2023:SCA

[ASJ^{+}23] Adham Atyabi, Frederick Shic, Jiajun Jiang, Claire E. Foster, Erin Barney, Minah Kim, Beibin Li, Pamela Ventola, and Chung Hao Chen. Stratification of children with autism spectrum disorder through fusion of temporal information in eye-gaze scan-paths. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):25:1–25:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3539226.

Amornbunchornvej:2021:ILM

[ASPT21] Chainarong Amornbunchornvej, Navaporn Surasvadi, Anon Plangprasopchok, and Suttipong Thajchayapong. Identifying linear models in multi-resolution population data using minimum description length principle to predict household income. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):15:1–15:30, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424670.

Anaissi:2021:OTB

[ASZ21] Ali Anaissi, Basem Suleiman, and Seid Miad Zandavi. Online tensor-based learning model for structural damage detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (6):107:1–107:18, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451217.

Anagnostopoulos:2017:QDL

[AT17] Christos Anagnostopoulos and Peter Triantafillou. Querydriven learning for predictive analytics of data subspace cardinality. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (4):47:1–47:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Amornbunchornvej:2021:VLG

[AZBW21] Chainarong Amornbunchornvej, Elena Zheleva, and Tanya Berger-Wolf. Variable-lag Granger causality and transfer entropy for time series analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (4):67:1–67:30, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441452.

Almeida:2021:MCB

[AZD^{+}21] Matthew Almeida, Yong Zhuang, Wei Ding, Scott E. Crouter, and Ping Chen. Mitigating class-boundary label uncertainty to reduce both model bias and variance. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):27:1–27:18, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429447.

Biswas:2023:RIM

[BAC23] Tarun Kumer Biswas, Alireza Abbasi, and Ripon Kumar Chakrabortty. Robust influence maximization under both aleatory and epistemic uncertainty. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):100:1–100:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587100.

Balcazar:2013:FCP

[Bal13] José L. Balcázar. Formal and computational properties of the confidence boost of association rules. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(4): 19:1–19:??, November 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bollegala:2018:CPM

[BAMK18] Danushka Bollegala, Vincent Atanasov, Takanori Maehara, and Ken-Ichi Kawarabayashi. ClassiNet — predicting missing features for short-text classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5): 55:1–55:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Boutemine:2017:MCS

[BB17] Oualid Boutemine and Mohamed Bouguessa. Mining community structures in multidimensional networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):51:1–51:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bechini:2022:NBF

[BBB^{+}22] Alessio Bechini, Alessandro Bondielli, José Luis Corcuera Bárcena, Pietro Ducange, Francesco Marcelloni, and Alessandro Renda. A newsbased framework for uncovering and tracking city area profiles: Assessment in Covid-19 setting. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):125:1–125:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532186.

Bergamini:2019:CTK

[BBC^{+}19] Elisabetta Bergamini, Michele Borassi, Pierluigi Crescenzi, Andrea Marino, and Henning Meyerhenke. Computing top-k closeness centrality faster in unweighted graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):53:1–53:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3344719.

Balalau:2024:FSM

[BBC^{+}24] Oana Balalau, Francesco Bonchi, T-H. Hubert Chan, Francesco Gullo, Mauro Sozio, and Hao Xie. Finding subgraphs with maximum total density and limited overlap in weighted hypergraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):95:1–95:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639410.

Becchetti:2010:EAL

[BBCG10] Luca Becchetti, Paolo Boldi, Carlos Castillo, and Aristides Gionis. Efficient algorithms for large-scale local triangle counting. ACM Transactions on Knowledge Discovery from Data (TKDD), 4 (3):13:1–13:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bayardop:2007:ISI

[BBD^{+}07] Roberto Bayardop, Kristin P.

Bennett, Gautam Das, Dimitrios Gunopulos, and Johannes Gunopulos. Introduction to special issue ACM SIGKDD 2006. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(3):9:1– 9:??, December 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Barton:2019:CIG

[BBK19] Tomas Barton, Tomas Bruna, and Pavel Kordik. Chameleon 2: an improved graphbased clustering algorithm. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):10:1–10:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bonab:2018:GGO

[BC18] Hamed R. Bonab and Fazli Can. GOOWE: Geometrically optimum and online-weighted ensemble classifier for evolving data streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2): 25:1–25:??, March 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bicego:2024:CRF

[BC24] Manuele Bicego and Ferdinando Cicalese. Computing random forest-distances in the presence of missing data. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):180:1–180:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3656345.

Bernardini:2021:CAS

[BCC^{+}21] Giulia Bernardini, Huiping Chen, Alessio Conte, Roberto Grossi, Grigorios Loukides, Nadia Pisanti, Solon P. Pissis, Giovanna Rosone, and Michelle Sweering. Combinatorial algorithms for string sanitization. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (1):8:1–8:34, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418683.

Bressan:2018:MCB

[BCK^{+}18] Marco Bressan, Flavio Chierichetti, Ravi Kumar, Stefano Leucci, and Alessandro Panconesi. Motif counting beyond five nodes. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):48:1– 48:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Berardi:2015:UTR

[BES15] Giacomo Berardi, Andrea Esuli, and Fabrizio Sebastiani. Utility-theoretic ranking for semiautomated text classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):6:1–6:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bohm:2007:RPF

[BFPP07] Christian Böhm, Christos Faloutsos, Jia-Yu Pan, and Claudia Plant. RIC: Parameterfree noise-robust clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(3):10:1–10:??, December 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Briggs:2013:IAM

[BFRL13] Forrest Briggs, Xiaoli Z. Fern, Raviv Raich, and Qi Lou. Instance annotation for multiinstance multi-label learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3):14:1– 14:??, September 2013. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bhattacharya:2007:CER

[BG07] Indrajit Bhattacharya and Lise Getoor. Collective entity resolution in relational data. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(1):5:1–5:??, March 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bilgic:2009:RCM

[BG09] Mustafa Bilgic and Lise Getoor. Reflect and correct: a misclassification prediction approach to active inference. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4): 20:1–20:??, November 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Barlaug:2021:NNE

[BG21] Nils Barlaug and Jon Atle Gulla. Neural networks for entity matching: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):52:1–52:37, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442200.

Burton:2014:DSC

[BGC14] Scott H. Burton and Christophe G.

Giraud-Carrier. Discovering social circles in directed graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(4):21:1– 21:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bonchi:2015:CCC

[BGG^{+}15] Francesco Bonchi, Aristides Gionis, Francesco Gullo, Charalampos E. Tsourakakis, and Antti Ukkonen. Chromatic correlation clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4):34:1–34:??, June 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bhattacharya:2012:CGC

[BGJV12] Indrajit Bhattacharya, Shantanu Godbole, Sachindra Joshi, and Ashish Verma. Cross-guided clustering: Transfer of relevant supervision across tasks. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(2):9:1– 9:??, July 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bayati:2013:MPA

[BGSW13] Mohsen Bayati, David F. Gleich, Amin Saberi, and Ying Wang. Message-passing algorithms for sparse network alignment. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(1):3:1–3:??, March 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ben-Gal:2019:CUT

[BGWSB19] Irad Ben-Gal, Shahar Weinstock, Gonen Singer, and Nicholas Bambos. Clustering users by their mobility behavioral patterns. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):45:1–45:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3322126.

Bae:2017:SEF

[BHW^{+}17] Seung-Hee Bae, Daniel Halperin, Jevin D. West, Martin Rosvall, and Bill Howe. Scalable and efficient flow-based community detection for largescale graph analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):32:1–32:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bellare:2013:ASE

[BIPR13] Kedar Bellare, Suresh Iyengar, Aditya Parameswaran, and Vibhor Rastogi. Active sampling for entity matching with guarantees. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3): 12:1–12:??, September 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bauer:2021:ICL

[BJ21] Josef Bauer and Dietmar Jannach. Improved customer lifetime value prediction with sequence-to-sequence learning and feature-based models. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):80:1– 80:37, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3441444.

Bhatia:2022:RTA

[BLH^{+}22] Siddharth Bhatia, Rui Liu, Bryan Hooi, Minji Yoon, Kijung Shin, and Christos Faloutsos. Real-time anomaly detection in edge streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):75:1–75:22, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494564.

Bressan:2021:FMC

[BLP21] Marco Bressan, Stefano Leucci, and Alessandro Panconesi. Faster motif counting via succinct color coding and adaptive sampling. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):96:1–96:27, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447397.

Boedihardjo:2014:FEL

[BLW14] Arnold P. Boedihardjo, Chang-Tien Lu, and Bingsheng Wang. A framework for exploiting local information to enhance density estimation of data streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1):2:1– 2:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Belcastro:2018:GRA

[BMTT18] Loris Belcastro, Fabrizio Marozzo, Domenico Talia, and Paolo Trunfio. G-RoI: Automatic region-of-interest detection driven by geotagged social media data. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):27:1–27:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bai:2021:TTG

[BMZY21] Luyi Bai, Xiangnan Ma, Mingcheng Zhang, and Wenting Yu. TPmod: a tendencyguided prediction model for temporal knowledge graph completion. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):41:1–41:17, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3443687.

Balasubramaniam:2020:ENT

[BNY20] Thirunavukarasu Balasubramaniam, Richi Nayak, and Chau Yuen. Efficient nonnegative tensor factorization via saturating coordinate descent. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):40:1–40:28, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3385654.

Bakerman:2018:TGH

[BPW^{+}18] Jordan Bakerman, Karl Pazdernik, Alyson Wilson, Geoffrey Fairchild, and Rian Bahran. Twitter geolocation: a hybrid approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):34:1–34:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bozdag:2024:MMG

[BSK24] Mustafa Bozdag, Nurullah Sevim, and Aykut Koç. Measuring and mitigating gender bias in legal contextualized language models. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):79:1–79:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3628602.

Brzezinski:2024:PFM

[BSS^{+}24] Dariusz Brzezinski, Julia Stachowiak, Jerzy Stefanowski, Izabela Szczech, Robert Susmaga, Sofya Aksenyuk, Uladzimir Ivashka, and Oleksandr Yasinskyi. Properties of fairness measures in the context of varying class imbalance and protected group ratios. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):170:1–170:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3654659.

Belohlavek:2021:ATP

[BT21] Radim Belohlavek and Martin Trnecka. The 8M algorithm from today’s perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (2):22:1–22:22, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428078.

Benarous:2022:SLH

[BTBg22] Maya Benarous, Eran Toch, and Irad Ben-gal. Synthesis of longitudinal human location sequences: Balancing utility and privacy. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):118:1–118:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529260.

Burkhardt:2021:OAB

[Bur21] Paul Burkhardt. Optimal algebraic breadth-first search for sparse graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):77:1–77:19, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446216.

Bouguessa:2010:DKS

[BWD10] Mohamed Bouguessa, Shengrui Wang, and Benoit Dumoulin. Discovering knowledgesharing communities in questionanswering forums. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(1):3:1–3:??, December 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bian:2020:MSM

[BXF^{+}20] Jiang Bian, Haoyi Xiong, Yanjie Fu, Jun Huan, and Zhishan Guo. MP^{2}SDA: Multi-party parallelized sparse discriminant learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):26:1–26:22, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3374919.

Chan:2012:CID

[CBLH12] Jeffrey Chan, James Bailey, Christopher Leckie, and Michael Houle. ciForager: Incrementally discovering regions of correlated change in evolving graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 6 (3):11:1–11:??, October 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cerf:2009:CPM

[CBRB09] Loïc Cerf, Jérémy Besson, Céline Robardet, and Jean-François Boulicaut. Closed patterns meet n-ary relations. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(1):3:1–3:??, March 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chu:2012:TLM

[CC12] Shumo Chu and James Cheng.

Triangle listing in massive networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(4): 17:1–17:??, December 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2019:DRW

[CC19] Hung-Hsuan Chen and Pu Chen.

Differentiating regularization weights — a simple mechanism to alleviate cold start in recommender systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):8:1–8:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3285954.

Chuang:2009:FPS

[CCC09] Kun-Ta Chuang, Hung-Leng Chen, and Ming-Syan Chen. Feature-preserved sampling over streaming data. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(4):15:1–15:??, January 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2024:BTB

[CCG^{+}24] Huiping Chen, Alessio Conte, Roberto Grossi, Grigorios Loukides, Solon P. Pissis, and Michelle Sweering. On breaking truss-based and core-based communities. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):135:1–135:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3644077.

Chen:2018:SSL

[CCLZ18] Chaochao Chen, Kevin Chen-Chuan Chang, Qibing Li, and Xiaolin Zheng. Semisupervised learning meets factorization: Learning to recommend with chain graph model. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):73:1– 73:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2023:LRR

[CCT^{+}23] Haoran Chen, Xu Chen, Hongwei Tao, Zuhe Li, and Xiao Wang. Low-rank representation with adaptive dimensionality reduction via manifold optimization for clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):128:1– 128:??, November 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3589767.

Chen:2023:TRR

[CCTW23] Lei Chen, Jie Cao, Haicheng Tao, and Jia Wu. Trip reinforcement recommendation with graph-based representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):57:1–57:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564609.

Cadena:2019:NOP

[CCV19] Jose Cadena, Feng Chen, and Anil Vullikanti. Near-optimal and practical algorithms for graph scan statistics with connectivity constraints. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):20:1–20:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309712.

Chakraborty:2023:SAM

[CDC23] Roshni Chakraborty, Ritwika Das, and Joydeep Chandra.

SigGAN: Adversarial model for learning signed relationships in networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):8:1–8:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532610.

Crescenzi:2016:GIO

[CDSV16] Pierluigi Crescenzi, Gianlorenzo D’angelo, Lorenzo Severini, and Yllka Velaj. Greedily improving our own closeness centrality in a network. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):9:1–9:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Coro:2021:LRS

[CDV21] Federico Coró, Gianlorenzo D’angelo, and Yllka Velaj. Link recommendation for social influence maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):94:1–94:23, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3449023.

Chen:2024:GTS

[CE24] Hongjie Chen and Hoda Eldardiry. Graph time-series modeling in deep learning: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):119:1–119:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638534.

Cui:2010:LMN

[CFD10] Ying Cui, Xiaoli Z. Fern, and Jennifer G. Dy. Learning multiple nonredundant clusterings. ACM Transactions on Knowledge Discovery from Data (TKDD), 4 (3):15:1–15:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Comito:2019:BED

[CFP19] Carmela Comito, Agostino Forestiero, and Clara Pizzuti. Bursty event detection in Twitter streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):41:1–41:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3332185.

Chen:2024:MVG

[CFX^{+}24] Zhaoliang Chen, Lele Fu, Shunxin Xiao, Shiping Wang, Claudia Plant, and Wenzhong Guo. Multi-view graph convolutional networks with differentiable node selection.

ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):6:1–6:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3608954.

Chen:2015:AAS

[CG15] Hung-Hsuan Chen and C. Lee Giles. ASCOS++: an asymmetric similarity measure for weighted networks to address the problem of Sim-Rank. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):15:1– 15:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2023:EAM

[CG23] Xinye Chen and Stefan Güttel. An efficient aggregation method for the symbolic representation of temporal data. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (1):5:1–5:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532622.

Chiasserini:2018:ACS

[CGL18] Carla-Fabiana Chiasserini, Michel Garetto, and Emili Leonardi. De-anonymizing clustered social networks by percolation graph matching.

ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):21:1–21:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Carchiolo:2023:ENP

[CGL^{+}23] Vincenza Carchiolo, Marco Grassia, Alessandro Longheu, Michele Malgeri, and Giuseppe Mangioni. Efficient node PageRank improvement via link-building using geometric deep learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):38:1–38:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551642.

Cheng:2016:CFR

[CGZW16] Wei Cheng, Zhishan Guo, Xiang Zhang, and Wei Wang. CGC: a flexible and robust approach to integrating co-regularized multi-domain graph for clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):46:1–46:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2018:BGD

[Che18]  Hung-Hsuan Chen.  Be-

havior2Vec: Generating distributed representations of users’ behaviors on products for recommender systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):43:1– 43:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2024:TUL

[CHY^{+}24] Wei Chen, Chao Huang, Yanwei Yu, Yongguo Jiang, and Junyu Dong. Trajectory-user linking via hierarchical spatiotemporal attention networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):85:1–85:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635718.

Cheng:2023:FIU

[CHZ23] Jiezhu Cheng, Kaizhu Huang, and Zibin Zheng. Fitting imbalanced uncertainties in multi-output time series forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):102:1–102:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584704.

Choo:2018:VVA

[CKC^{+}18] Jaegul Choo, Hannah Kim, Edward Clarkson, Zhicheng Liu, Changhyun Lee, Fuxin Li, Hanseung Lee, Ramakrishnan Kannan, Charles D.

Stolper, John Stasko, and Haesun Park. VisIRR: a visual analytics system for information retrieval and recommendation for large-scale document data. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):8:1–8:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cormode:2008:FHH

[CKMS08] Graham Cormode, Flip Korn, S. Muthukrishnan, and Divesh Srivastava. Finding hierarchical heavy hitters in streaming data. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(4):2:1–2:??, January 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2018:MGC

[CL18] Xiaowei Chen and John C. S.

Lui. Mining graphlet counts in online social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):41:1–41:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2019:MCP

[CLG^{+}19] Haoran Chen, Jinghua Li, Junbin Gao, Yanfeng Sun, Yongli Hu, and Baocai Yin. Maximally correlated principal component analysis based on deep parameterization learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):39:1–39:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3332183.

Chen:2024:SSD

[CLJC24] Ke-Jia Chen, Linsong Liu, Linpu Jiang, and Jingqiang Chen. Self-supervised dynamic graph representation learning via temporal subgraph contrast. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):18:1–18:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3612931.

Chen:2020:CTF

[CLT^{+}20] Cen Chen, Kenli Li, Sin G. Teo, Xiaofeng Zou, Keqin Li, and Zeng Zeng. Citywide traffic flow prediction based on multiple gated spatiotemporal convolutional neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):42:1–42:23, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3385414.

Chen:2012:LIS

[CLY12] Jianhui Chen, Ji Liu, and Jieping Ye. Learning incoherent sparse and lowrank patterns from multiple tasks. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (4):22:1–22:??, February 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2023:MLV

[CLYC23] Ling Chen, Dandan Lyu, Shanshan Yu, and Gencai Chen. Multi-level visual similarity based personalized tourist attraction recommendation using geo-tagged photos. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):92:1– 92:??, August 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3582015.

Cantini:2022:LSH

[CMBT22] Riccardo Cantini, Fabrizio Marozzo, Giovanni Bruno, and Paolo Trunfio. Learning sentence-to-hashtags semantic mapping for hashtag recommendation on microblogs. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (2):32:1–32:26, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466876.

Crescenzi:2019:HCM

[CMQ19] Valter Crescenzi, Paolo Merialdo, and Disheng Qiu. Hybrid crowd-machine wrapper inference. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):51:1–51:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3344720.

Canfora:2023:NCT

[CMS23] Gerardo Canfora, Francesco Mercaldo, and Antonella Santone. A novel classification technique based on formal methods. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):113:1–113:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3592796.

Corbara:2024:SDD

[CMS24] Silvia Corbara, Alejandro Moreo, and Fabrizio Sebastiani. Same or different? Diff-vectors for authorship analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):12:1–12:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3609226.

Campello:2015:HDE

[CMZS15] Ricardo J. G. B. Campello, Davoud Moulavi, Arthur Zimek, and Jörg Sander. Hierarchical density estimates for data clustering, visualization, and outlier detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):5:1–5:??, July 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chang:2016:CSP

[CNY^{+}16] Xiaojun Chang, Feiping Nie, Yi Yang, Chengqi Zhang, and Heng Huang. Convex sparse PCA for unsupervised feature learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):3:1–3:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cheng:2017:RCA

[CNZ^{+}17] Wei Cheng, Jingchao Ni, Kai Zhang, Haifeng Chen, Guofei Jiang, Yu Shi, Xiang Zhang, and Wei Wang. Ranking causal anomalies for system fault diagnosis via temporal and dynamical analysis on vanishing correlations. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):40:1–40:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Costa:2018:MOC

[CO18] Gianni Costa and Riccardo Ortale. Mining overlapping communities and inner role assignments through Bayesian mixed-membership models of networks with contextdependent interactions. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):18:1–18:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Constantinou:2020:LBN

[Con20] Anthony C. Constantinou. Learning Bayesian networks with the saiyan algorithm. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):44:1–44:21, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3385655.

Coscia:2021:NCS

[Cos21] Michele Coscia. Noise corrected sampling of online social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):29:1–29:21, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3434749.

Coscia:2022:GEM

[Cos22]  Michele Coscia.  General-

ized Euclidean measure to estimate distances on multilayer networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):119:1–119:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529396.

Chen:2010:BTA

[CPC10] Ye Chen, Dmitry Pavlov, and John F. Canny. Behavioral targeting: The art of scaling up simple algorithms. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(4):17:1–17:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ceccarello:2020:GCB

[CPP20] Matteo Ceccarello, Andrea Pietracaprina, and Geppino Pucci. A general coresetbased approach to diversity maximization under matroid constraints. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):60:1–60:27, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3402448.

Chen:2021:FCM

[CPYT21] Chen Chen, Ruiyue Peng, Lei Ying, and Hanghang Tong.

Fast connectivity minimization on large-scale networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):53:1–53:25, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442342.

Chen:2024:SSH

[CQH^{+}24] Ying Chen, Siwei Qiang, Mingming Ha, Xiaolei Liu, Shaoshuai Li, Jiabi Tong, Lingfeng Yuan, Xiaobo Guo, and Zhenfeng Zhu. Semisupervised heterogeneous graph learning with multi-level data augmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):44:1–44:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3608953.

Chen:2023:SSS

[CRC^{+}23] Yufu Chen, Yanghui Rao, Shurui Chen, Zhiqi Lei, Haoran Xie, Raymond Y. K. Lau, and Jian Yin. Semi-supervised sentiment classification and emotion distribution learning across domains. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):74:1–74:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571736.

Coscia:2014:UHO

[CRGP14] Michele Coscia, Giulio Rossetti, Fosca Giannotti, and Dino Pedreschi. Uncovering hierarchical and overlapping communities with a local-first approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1):6:1–6:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2023:GDF

[CRM^{+}23] Hongjie Chen, Ryan A. Rossi, Kanak Mahadik, Sungchul Kim, and Hoda Eldardiry. Graph deep factors for probabilistic time-series forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):26:1– 26:??, February 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3543511.

Chen:2024:FAG

[CRP^{+}24] April Chen, Ryan A. Rossi, Namyong Park, Puja Trivedi, Yu Wang, Tong Yu, Sungchul Kim, Franck Dernoncourt, and Nesreen K. Ahmed. Fairnessaware graph neural networks: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):138:1–138:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3649142.

Chen:2009:BAS

[CRST09] Bee-Chung Chen, Raghu Ramakrishnan, Jude W. Shavlik, and Pradeep Tamma. Bellwether analysis: Searching for cost-effective query-defined predictors in large databases. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(1):5:1–5:??, March 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2022:DED

[CS22] Ling Chen and Hongyu Shi. DexDeepFM: Ensemble diversity enhanced extreme deep factorization machine model. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):90:1– 90:17, October 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3505272.

Chen:2024:DGN

[CS24] Zhe Chen and Aixin Sun. DP-GCN: Node classification by connectivity and local topology structure on real-world network. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):140:1–140:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649460.

Chattopadhyay:2012:MDA

[CSF^{+}12] Rita Chattopadhyay, Qian Sun, Wei Fan, Ian Davidson, Sethuraman Panchanathan, and Jieping Ye. Multisource domain adaptation and its application to early detection of fatigue. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(4): 18:1–18:??, December 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chakraborty:2016:PCS

[CSG^{+}16] Tanmoy Chakraborty, Sriram Srinivasan, Niloy Ganguly, Animesh Mukherjee, and Sanjukta Bhowmick. Permanence and community structure in complex networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2):14:1–14:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cheng:2021:DEB

[CSHZ21] Weiyu Cheng, Yanyan Shen, Linpeng Huang, and Yanmin Zhu. Dual-embedding based deep latent factor models for recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):85:1–85:24, June 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447395.

Chowdhury:2024:INC

[CSM^{+}24] Anjan Chowdhury, Sriram Srinivasan, Animesh Mukherjee, Sanjukta Bhowmick, and Kuntal Ghosh. Improving node classification accuracy of GNN through input and output intervention. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):17:1–17:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610535.

Chakrabarti:2015:BPL

[CSSP15] Aniket Chakrabarti, Venu Satuluri, Atreya Srivathsan, and Srinivasan Parthasarathy. A Bayesian perspective on locality sensitive hashing with extensions for kernel methods. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):19:1– 19:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2021:ICD

[CSX21] Zhe Chen, Aixin Sun, and Xiaokui Xiao. Incremental community detection on large complex attributed network. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):109:1–109:20, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451216.

Chi:2009:ESC

[CSZ^{+}09] Yun Chi, Xiaodan Song, Dengyong Zhou, Koji Hino, and Belle L. Tseng. On evolutionary spectral clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4):17:1– 17:??, November 2009. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cheng:2021:GLG

[CSZ^{+}21] Lin Cheng, Yuliang Shi, Kun Zhang, Xinjun Wang, and Zhiyong Chen. GGATB-LSTM: Grouping and global attention-based time-aware bidirectional LSTM medical treatment behavior prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):45:1–45:16, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441454.

Chen:2014:ISI

[CT14] Wei Chen and Jie Tang. Introduction to special issue on computational aspects of social and information networks: Theory, methodologies, and applications (TKDD-CASIN). ACM Transactions on Knowledge Discovery from Data (TKDD), 8(1):1:1–1:??, February 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2022:DDG

[CTC^{+}22] Ling Chen, Xing Tang, Weiqi Chen, Yuntao Qian, Yansheng Li, and Yongjun Zhang. DACHA: a dual graph convolution based temporal knowledge graph representation learning method using historical relation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):46:1–46:18, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3477051.

Chen:2016:EOL

[CTP^{+}16] Chen Chen, Hanghang Tong, B. Aditya Prakash, Tina Eliassi-Rad, Michalis Faloutsos, and Christos Faloutsos. Eigen-optimization on large graphs by edge manipulation. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):49:1– 49:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2017:CDI

[CTX^{+}17] Chen Chen, Hanghang Tong, Lei Xie, Lei Ying, and Qing He. Cross-dependency inference in multi-layered networks: a collaborative filtering perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (4):42:1–42:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cui:2016:BOQ

[CTZ16] Licong Cui, Shiqiang Tao, and Guo-Qiang Zhang. Biomedical ontology quality assurance using a big data approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):41:1–41:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Connor:2024:NZN

[CV24] Richard Connor and Lucia Vadicamo. nSimplex Zen: a novel dimensionality reduction for Euclidean and Hilbert spaces. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):143:1–143:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3647642.

Chattopadhyay:2013:BMA

[CWF^{+}13] Rita Chattopadhyay, Zheng Wang, Wei Fan, Ian Davidson, Sethuraman Panchanathan, and Jieping Ye. Batch mode active sampling based on marginal probability distribution matching. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3): 13:1–13:??, September 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cai:2024:RDS

[CWL^{+}24a] Ruichu Cai, Fengzhu Wu, Zijian Li, Jie Qiao, Wei Chen, Yuexing Hao, and Hao Gu. REST: Debiased social recommendation via reconstructing exposure strategies. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):38:1–38:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3624986.

Cai:2024:GDA

[CWL^{+}24b] Ruichu Cai, Fengzhu Wu, Zijian Li, Pengfei Wei, Lingling Yi, and Kun Zhang. Graph domain adaptation: a generative view. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):60:1–60:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3631712.

Cousins:2023:BBC

[CWR23] Cyrus Cousins, Chloe Wohlgemuth, and Matteo Riondato.

Bavarian: Betweenness centrality approximation with variance-aware Rademacher averages. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6): 78:1–78:??, July 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577021.

Chen:2024:TRT

[CWW^{+}24] Yunkai Chen, Qimeng Wang, Shiwei Wu, Yan Gao, Tong Xu, and Yao Hu. TOMGPT: Reliable text-only training approach for cost-effective multi-modal large language model. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):171:1– 171:??, August 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3654674.

Cui:2024:SSM

[CWWL24] Guosheng Cui, Ruxin Wang, Dan Wu, and Ye Li. Semisupervised multi-view clustering based on NMF with fusion regularization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):157:1–157:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653022.

Chen:2023:LEL

[CWY^{+}23] Huiming Chen, Huandong Wang, Quanming Yao, Yong Li, Depeng Jin, and Qiang Yang. LoSAC: an efficient local stochastic average control method for federated optimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):58:1–58:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3566128.

Chen:2010:BBI

[CX10] Jinlin Chen and Keli Xiao. BISC: a bitmap itemset support counting approach for efficient frequent itemset mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(3):12:1– 12:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Concas:2020:MSM

[CXH^{+}20] Francesco Concas, Pengfei Xu, Mohammad A. Hoque, Jiaheng Lu, and Sasu Tarkoma. Multiple set matching with Bloom matrix and Bloom vector. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):21:1– 21:21, March 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.

acm.org/doi/abs/10.1145/
3372409.

Cui:2024:ISP

[CXL^{+}24] Jingyi Cui, Guangquan Xu, Jian Liu, Shicheng Feng, Jianli Wang, Hao Peng, Shihui Fu, Zhaohua Zheng, Xi Zheng, and Shaoying Liu. ID-SR: Privacypreserving social recommendation based on infinite divisibility for trustworthy AI. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):161:1–161:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639412.

Chen:2024:GAG

[CXWH24] Ling Chen, Jiahui Xu, Binqing Wu, and Jianlong Huang. Group-aware graph neural network for nationwide city air quality forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):55:1–55:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631713.

Choe:2024:RBT

[CYL^{+}24] Minyoung Choe, Jaemin Yoo, Geon Lee, Woonsung Baek, U. Kang, and Kijung Shin. Representative and back-in-time sampling from real-world hypergraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):156:1–156:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653306.

Chang:2016:LMB

[CYOL16] Yi Chang, Makoto Yamada, Antonio Ortega, and Yan Liu. Lifecycle modeling for buzz temporal pattern discovery. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2):20:1– 20:??, December 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Costa:2017:MTA

[CYT^{+}17] Alceu Ferraz Costa, Yuto Yamaguchi, Agma Juci Machado Traina, Caetano Traina Jr., and Christos Faloutsos. Modeling temporal activity to detect anomalous behavior in social media. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (4):49:1–49:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cai:2022:ANI

[CYY^{+}22] Jianghui Cai, Yuqing Yang, Haifeng Yang, Xujun Zhao, and Jing Hao. ARIS: a noise insensitive data pre-processing scheme for data reduction using influence space. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):110:1–110:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522592.

Chen:2018:LSB

[CZH18] Guangyong Chen, Fengyuan Zhu, and Pheng Ann Heng. Large-scale Bayesian probabilistic matrix factorization with memo-free distributed variational inference. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):31:1–31:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Cheng:2024:AFS

[CZW^{+}24] Ling Cheng, Feida Zhu, Yong Wang, Ruicheng Liang, and Huiwen Liu. From asset flow to status, action, and intention discovery: Early malice detection in cryptocurrency. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):50:1–50:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3626102.

Cheng:2011:CLA

[CZY11] Hong Cheng, Yang Zhou, and Jeffrey Xu Yu. Clustering large attributed graphs: a balance between structural and attribute similarities. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(2):12:1– 12:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Chen:2022:OLB

[CZY^{+}22] Weirong Chen, Jiaqi Zheng, Haoyu Yu, Guihai Chen, Yixin Chen, and Dongsheng Li. Online learning bipartite matching with non-stationary distributions. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):83:1–83:22, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502734.

Chen:2023:RLP

[CZZ^{+}23] Jinwei Chen, Zefang Zong, Yunlin Zhuang, Huan Yan, Depeng Jin, and Yong Li. Reinforcement learning for practical express systems with mixed deliveries and pickups. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):33:1– 33:??, April 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3546952.

Datta:2018:CVC

[DA18] Srayan Datta and Eytan Adar.

CommunityDiff: Visualizing community clustering algorithms. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):11:1– 11:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Dehghan:2019:TDE

[DA19] Mahdi Dehghan and Ahmad Ali Abin. Translations diversification for expert finding: a novel clusteringbased approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):32:1–32:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3320489.

Domeniconi:2009:WCE

[DAR09] Carlotta Domeniconi and Muna Al-Razgan. Weighted cluster ensembles: Methods and analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 2 (4):17:1–17:??, January 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Deng:2021:PUT

[DCF^{+}21] Jinliang Deng, Xiusi Chen, Zipei Fan, Renhe Jiang, Xuan Song, and Ivor W. Tsang. The pulse of urban transport: Exploring the co-evolving pattern for spatio-temporal forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):103:1–103:25, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3450528.

Dhurandhar:2009:SAM

[DD09] Amit Dhurandhar and Alin Dobra. Semi-analytical method for analyzing models and model selection measures based on moment analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(1):2:1–2:??, March 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Djenouri:2021:TOD

[DDL21] Youcef Djenouri, Djamel Djenouri, and Jerry Chun-Wei Lin. Trajectory outlier detection: New problems and solutions for smart cities. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):20:1–20:28, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3425867.

DiPalma:2024:EVS

[DDL24] Luciano Di Palma, Yanlei Diao, and Anna Liu. Efficient version space algorithms for human-in-the-loop model development. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):69:1–69:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3637443.

Deng:2024:TNF

[DDY^{+}24] Jiewen Deng, Jinliang Deng, Du Yin, Renhe Jiang, and Xuan Song. TTS-Norm: Forecasting tensor time series via multi-way normalization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):3:1–3:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3605894.

DeStefani:2017:TCL

[DERU17] Lorenzo De Stefani, Alessandro Epasto, Matteo Riondato, and Eli Upfal. TRIÈST: Counting local and global triangles in fully dynamic streams with fixed memory size. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):43:1– 43:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Deodhar:2010:SFS

[DG10] Meghana Deodhar and Joydeep Ghosh. SCOAL: a framework for simultaneous coclustering and learning from complex data. ACM Transactions on Knowledge Discovery from Data (TKDD), 4 (3):11:1–11:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Duarte:2016:AMR

[DGB16] João Duarte, João Gama, and Albert Bifet. Adaptive model rules from high-speed data streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (3):30:1–30:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Deng:2024:DPL

[DHZ24] Jiayi Deng, Danyang Huang, and Bo Zhang. Distributed pseudo-likelihood method for community detection in largescale networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):181:1–181:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3657300.

Dornaika:2024:SIS

[DIB24] Fadi Dornaika, Zoulfikar Ibrahim, and Alirezah Bosaghzadeh. Scalable and inductive semisupervised classifier with sample weighting based on graph topology. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):133:1–133:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643645.

Dunlavy:2011:TLP

[DKA11] Daniel M. Dunlavy, Tamara G.

Kolda, and Evrim Acar. Temporal link prediction using matrix and tensor factorizations. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (2):10:1–10:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Davvetas:2022:ETL

[DKSK22] Athanasios Davvetas, Iraklis A. Klampanos, Spiros Skiadopoulos, and Vangelis Karkaletsis. Evidence transfer: Learning improved representations according to external heterogeneous task outcomes. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):85:1– 85:22, October 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502732.

Di:2018:LSA

[DKSL18] Mingyang Di, Diego Klabjan, Long Sha, and Patrick Lucey. Large-scale adversarial sports play retrieval with learning to rank. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (6):69:1–69:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Duan:2024:ACT

[DLZ^{+}24] Mingxing Duan, Kenli Li, Weinan Zhang, Jiarui Qin, and Bin Xiao. Attacking click-through rate predictors via generating realistic fake samples. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):110:1–110:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643685.

Salve:2021:PIU

[DMG^{+}21] Andrea De Salve, Paolo Mori, Barbara Guidi, Laura Ricci, and Roberto Di Pietro. Predicting influential users in online social network groups. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):35:1–35:50, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441447.

Azevedo:2021:RNT

[DMGC21] Ricardo De Azevedo, Gabriel Resende Machado, Ronaldo Ribeiro Goldschmidt, and Ricardo Choren. A reduced network traffic method for IoT data clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (1):13:1–13:23, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3423139.

Das:2012:MIG

[DMI12] Sanmay Das and Malik Magdon-Ismail. A model for information growth in collective wisdom processes. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(2):6:1–6:??, July 2012. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Dornaika:2019:ATP

[Dor19] Fadi Dornaika. Active two phase collaborative representation classifier. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):36:1–36:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326919.

Dornaika:2021:EET

[Dor21] F. Dornaika. Elastic embedding through graph convolutionbased regression for semisupervised classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):56:1–56:11, June 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441456.

Dai:2020:CVE

[DPB^{+}20] Chenglong Dai, Dechang Pi, Stefanie I. Becker, Jia Wu, Lin Cui, and Blake Johnson. CenEEGs: Valid EEG selection for classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (2):18:1–18:25, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3371153.

DosSantos:2018:RLC

[DPDG18] Ludovic Dos Santos, Benjamin Piwowarski, Ludovic Denoyer, and Patrick Gallinari. Representation learning for classification in heterogeneous graphs with application to social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5):62:1– 62:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Duan:2014:SRC

[DSL^{+}14] Lian Duan, W. Nick Street, Yanchi Liu, Songhua Xu, and Brook Wu. Selecting the right correlation measure for binary data. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2): 13:1–13:??, September 2014.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Deng:2024:DGT

[DSL^{+}24] Songgaojun Deng, Olivier Sprangers, Ming Li, Sebastian Schelter, and Maarten de Rijke. Domain generalization in time series forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):113:1–113:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643035.

Desantis:2022:FBM

[DSTA22] Derek Desantis, Erik Skau, Duc P. Truong, and Boian Alexandrov. Factorization of binary matrices: Rank relations, uniqueness and model selection of Boolean decomposition. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):112:1–112:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522594.

Stefani:2021:TSE

[DTU21] Lorenzo De Stefani, Erisa Terolli, and Eli Upfal. Tiered sampling: an efficient method for counting sparse motifs in massive graph streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):79:1–79:52, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441299.

deVries:2011:RRL

[dVKCC11] Timothy de Vries, Hui Ke, Sanjay Chawla, and Peter Christen. Robust record linkage blocking using suffix arrays and Bloom filters. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(2):9:1–9:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Das:2020:DAI

[DWD^{+}20] Shubhomoy Das, Weng-Keen Wong, Thomas Dietterich, Alan Fern, and Andrew Emmott. Discovering anomalies by incorporating feedback from an expert. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):49:1–49:32, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3396608.

Ding:2024:TFD

[DWD^{+}24] Shifei Ding, Benyu Wu, Ling Ding, Xiao Xu, Lili Guo, Hongmei Liao, and Xindong Wu. Towards faster deep graph clustering via efficient graph auto-encoder. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):202:1–202:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674983.

Dai:2023:DMV

[DWH^{+}23] Shaojie Dai, Jinshuai Wang, Chao Huang, Yanwei Yu, and Junyu Dong. Dynamic multiview graph neural networks for citywide traffic inference. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):53:1–53:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564754.

Ding:2024:RGM

[DWL^{+}24] Kaize Ding, Jianling Wang, Jundong Li, James Caverlee, and Huan Liu. Robust graph meta-learning for weakly supervised few-shot node classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4): 83:1–83:??, May 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3630260.

Ding:2022:ESU

[DWW22] Ming Ding, Tianyu Wang, and Xudong Wang. Establishing smartphone user behavior model based on energy consumption data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):25:1–25:40, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3461459.

Dong:2020:RCP

[DYS20] Jialin Dong, Kai Yang, and Yuanming Shi. Ranking from crowdsourced pairwise comparisons via smoothed Riemannian optimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):19:1–19:26, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3372407.

Ermis:2020:DSD

[EC20] Beyza Ermis and A. Taylan CemgIl. Data sharing via differentially private coupled matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):28:1–28:27, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3372408.

Ekle:2024:ADD

[EE24] Ocheme Anthony Ekle and William Eberle. Anomaly detection in dynamic graphs: a comprehensive survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):192:1–192:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3669906.

Erdo:2014:RGN

[EGT14] Dóra Erdős, Rainer Gemulla, and Evimaria Terzi. Reconstructing graphs from neighborhood data. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(4): 23:1–23:??, August 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Elkan:2012:GES

[EK12] Charles Elkan and Yehuda Koren. Guest editorial for special issue KDD’10. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (4):18:1–18:??, February 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

E:2023:CEC

[ELH23] Jinlong E., Mo Li, and Jianqiang Huang. CrowdAtlas: Estimating crowd distribution within the urban rail transit system. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):48:1–48:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3558521.

Eyal:2014:PIM

[ERSK14] Ron Eyal, Avi Rosenfeld, Sigal Sina, and Sarit Kraus. Predicting and identifying missing node information in social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(3): 14:1–14:??, June 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Esuli:2015:OTQ

[ES15] Andrea Esuli and Fabrizio Sebastiani. Optimizing text quantifiers for multivariate loss functions. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4): 27:1–27:??, June 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Freris:2023:IEV

[FAV23] Nikolaos M. Freris, Ahmad Ajalloeian, and Michalis Vlachos. Interpretable embedding and visualization of compressed data. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):22:1–22:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3537901.

Feng:2017:RBC

[FCWQ17] Shanshan Feng, Jian Cao, Jie Wang, and Shiyou Qian. Recommendations based on comprehensively exploiting the latent factors hidden in items’ ratings and content. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):35:1–35:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Fan:2024:DDE

[FFZ^{+}24] Wei Fan, Yanjie Fu, Shun Zheng, Jiang Bian, Yuanchun Zhou, and Hui Xiong. DEWP: Deep expansion learning for wind power forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):71:1–71:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3637552.

Fei:2014:SSB

[FH14] Hongliang Fei and Jun Huan.

Structured sparse boosting for graph classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1):4:1–4:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Feng:2024:AAG

[FJZT24] Shengyu Feng, Baoyu Jing, Yada Zhu, and Hanghang Tong. ArieL: Adversarial graph contrastive learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):82:1–82:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638054.

Fountoulakis:2017:RRA

[FKKD17] Kimon Fountoulakis, Abhisek Kundu, Eugenia-Maria Kontopoulou, and Petros Drineas. A randomized rounding algorithm for sparse PCA. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):38:1–38:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Feng:2023:HDP

[FLC23] Wenjie Feng, Shenghua Liu, and Xueqi Cheng. Hierarchical dense pattern detection in tensors. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):81:1–81:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3577022.

Feng:2021:RSP

[FLG^{+}21] Yi Feng, Chuanyi Li, Jidong Ge, Bin Luo, and Vincent Ng. Recommending statutes: a portable method based on neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):16:1–16:22, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424671.

Feng:2022:CAS

[FLL^{+}22] Jie Feng, Yong Li, Ziqian Lin, Can Rong, Funing Sun, Diansheng Guo, and Depeng Jin. Context-aware spatialtemporal neural network for citywide crowd flow prediction via modeling long-range spatial dependency. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):49:1–49:21, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477577.

Fu:2024:MUC

[FN24] Zhe Fu and Xi Niu. Modeling users’ curiosity in recommender systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):26:1–26:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617598.

Fujiwara:2009:FLS

[FSK09] Yasuhiro Fujiwara, Yasushi Sakurai, and Masaru Kitsuregawa. Fast likelihood search for hidden Markov models. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4): 18:1–18:??, November 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Feng:2023:CTE

[FSXL23] Tao Feng, Sirui Song, Tong Xia, and Yong Li. Contact tracing and epidemic intervention via deep reinforcement learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):34:1–34:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3546870.

Fu:2016:MGD

[FXG^{+}16] Yanjie Fu, Hui Xiong, Yong Ge, Yu Zheng, Zijun Yao, and Zhi-Hua Zhou. Modeling of geographic dependencies for real estate ranking. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (1):11:1–11:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Fu:2022:TTR

[FYN22] Zhe Fu, Li Yu, and Xi Niu. TRACE: Travel reinforcement recommendation based on location-aware context extraction. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):65:1–65:22, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487047.

Gupta:2022:CDL

[GA22] Manish Gupta and Puneet Agrawal. Compression of deep learning models for text: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):61:1–61:55, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487045.

Ghahramanian:2024:NNE

[GBBC24] Pouya Ghahramanian, Sepehr Bakhshi, Hamed Bonab, and Fazli Can. A novel neural ensemble architecture for onthe-fly classification of evolving text streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (4):101:1–101:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639054.

Galimberti:2020:CDM

[GBGL20] Edoardo Galimberti, Francesco Bonchi, Francesco Gullo, and Tommaso Lanciano. Core decomposition in multilayer networks: Theory, algorithms, and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):11:1–11:40, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3369872.

Gundecha:2014:UVR

[GBTL14] Pritam Gundecha, Geoffrey Barbier, Jiliang Tang, and Huan Liu. User vulnerability and its reduction on a social networking site. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2): 12:1–12:??, September 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ge:2024:DCC

[GBW^{+}24] Yong-Feng Ge, Elisa Bertino, Hua Wang, Jinli Cao, and Yanchun Zhang. Distributed cooperative coevolution of data publishing privacy and transparency. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):20:1–20:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3613962.

Galimberti:2021:SCD

[GCB^{+}21] Edoardo Galimberti, Martino Ciaperoni, Alain Barrat, Francesco Bonchi, Ciro Cattuto, and Francesco Gullo. Span-core decomposition for temporal networks: Algorithms and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):2:1–2:44, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418226.

Guo:2022:RCT

[GCG22] Jinjin Guo, Longbing Cao, and Zhiguo Gong. Recurrent coupled topic modeling over sequential documents. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):8:1–8:32, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451530.

Gao:2019:TEM

[GCL^{+}19a] Xiaofeng Gao, Zhenhao Cao, Sha Li, Bin Yao, Guihai Chen, and Shaojie Tang. Taxonomy and evaluation for microblog popularity prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):15:1–15:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301303.

Guo:2019:LLS

[GCL^{+}19b] Yumeng Guo, Fulai Chung, Guozheng Li, Jiancong Wang, and James C. Gee. Leveraging label-specific discriminant mapping features for multilabel learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):24:1–24:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3319911.

Ge:2008:JCA

[GEG^{+}08] Rong Ge, Martin Ester, Byron J. Gao, Zengjian Hu, Binay Bhattacharya, and Boaz Ben-Moshe. Joint cluster analysis of attribute data and relationship data: The connected k-center problem, algorithms and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(2): 7:1–7:??, July 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Gu:2023:IMC

[GF23] Zhibin Gu and Songhe Feng. Individuality meets commonality: a unified graph learning framework for multi-view clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (1):7:1–7:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532612.

Gu:2023:OJU

[GFHL23] Zhibin Gu, Songhe Feng, Ruiting Hu, and Gengyu Lyu. ONION: Joint unsupervised feature selection and robust subspace extraction for graph-based multiview clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):70:1–70:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568684.

Gu:2024:NJC

[GFL^{+}24] Zhibin Gu, Songhe Feng, Zhendong Li, Jiazheng Yuan, and Jun Liu. NOODLE: Joint cross-view discrepancy discovery and high-order correlation detection for multi-view subspace clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):151:1–151:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653305.

Ghasemi:2021:UEE

[GFM21] Negin Ghasemi, Ramin Fatourechi, and Saeedeh Momtazi. User embedding for expert finding in community question answering. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):70:1–70:16, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441302.

Gupta:2008:BBC

[GG08] Gunjan Gupta and Joydeep Ghosh. Bregman bubble clustering: a robust framework for mining dense clusters. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(2):8:1–8:??, July 2008. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Greco:2015:PDU

[GGLP15] Gianluigi Greco, Antonella Guzzo, Francesco Lupia, and Luigi Pontieri. Process discovery under precedence constraints. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4): 32:1–32:??, June 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Guo:2024:MTL

[GHT^{+}24] Xiaobo Guo, Mingming Ha, Xuewen Tao, Shaoshuai Li, Youru Li, Zhenfeng Zhu, Zhiyong Shen, and Li Ma. Multitask learning with sequential dependence toward industrial applications: a systematic formulation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):107:1–107:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3640468.

Gao:2016:DAC

[GJ16] Zekai J. Gao and Chris Jermaine. Distributed algorithms for computing very large thresholded covariance matrices. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 12:1–12:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ge:2014:RMA

[GJDX14] Yong Ge, Guofei Jiang, Min Ding, and Hui Xiong. Ranking metric anomaly in invariant networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(2):8:1– 8:??, June 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Gionis:2015:ISI

[GL15] Aristides Gionis and Hang Li. Introduction to the special issue ACM SIGKDD 2013. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3):15:1– 15:??, April 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Guo:2022:DLD

[GL22] Yunyan Guo and Jianzhong Li. Distributed latent Dirichlet allocation on streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):9:1–9:20, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451528.

Gao:2022:CDR

[GLF^{+}22] Chen Gao, Yong Li, Fuli Feng, Xiangning Chen, Kai Zhao, Xiangnan He, and Depeng Jin. Cross-domain recommendation with bridge-item embeddings. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (1):2:1–2:23, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447683.

Gan:2019:SPS

[GLFV^{+}19] Wensheng Gan, Jerry Chun-Wei Lin, Philippe Fournier-Viger, Han-Chieh Chao, and Philip S. Yu. A survey of parallel sequential pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):25:1–25:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314107.

Gao:2022:TES

[GLG^{+}22] Fei Gao, Jiada Li, Yisu Ge, Jianwen Shao, Shufang Lu, and Libo Weng. A trajectory evaluator by sub-tracks for detecting VOT-based anomalous trajectory. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):67:1–67:19, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490032.

Goel:2022:QIP

[GLMW22] Kanika Goel, Sander J. J.

Leemans, Niels Martin, and Moe T. Wynn. Qualityinformed process mining: a case for standardised data quality annotations. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):97:1–97:47, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3511707.

Guo:2023:DAD

[GLW^{+}23] Yuhui Guo, Xun Liang, Bo Wu, Xiangping Zheng, and Xuan Zhang. Dualaware domain mining and cross-aware supervision for weakly-supervised semantic segmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):101:1–101:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3589343.

Gan:2021:UMA

[GLZ^{+}21] Wensheng Gan, Jerry Chun-Wei Lin, Jiexiong Zhang, Hongzhi Yin, Philippe Fournier-Viger, Han-Chieh Chao, and Philip S. Yu. Utility mining across multi-dimensional sequences. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (5):82:1–82:24, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3446938.

Galbrun:2018:MRS

[GM18] Esther Galbrun and Pauli Miettinen. Mining redescriptions with siren. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):6:1–6:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Gionis:2007:ADM

[GMMT07] Aristides Gionis, Heikki Mannila, Taneli Mielikäinen, and Panayiotis Tsaparas. Assessing data mining results via swap randomization. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(3):14:1–14:??, December 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Garciarena:2021:TAC

[GMS21] Unai Garciarena, Alexander Mendiburu, and Roberto Santana. Towards automatic construction of multi-network models for heterogeneous multi-task learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):33:1–33:23, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434748.

Guzzo:2013:SIF

[GMSS13] Antonella Guzzo, Luigi Moccia, Domenico Saccà, and Edoardo Serra. Solving inverse frequent itemset mining with infrequency constraints via large-scale linear programs. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(4): 18:1–18:??, November 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Gionis:2007:CA

[GMT07] Aristides Gionis, Heikki Mannila, and Panayiotis Tsaparas. Clustering aggregation. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(1):4:1–4:??, March 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ghosh:2021:CBE

[GNMQ21] Aindrila Ghosh, Mona Nashaat, James Miller, and Shaikh Quader. Context-based evaluation of dimensionality reduction algorithms-experiments and statistical significance analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (2):24:1–24:40, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428077.

Gu:2022:IFS

[GQH22] Shilin Gu, Yuhua Qian, and Chenping Hou. Incremental feature spaces learning with label scarcity. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):106:1–106:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3516368.

Gomez-Rodriguez:2012:IND

[GRLK12] Manuel Gomez-Rodriguez, Jure Leskovec, and Andreas Krause. Inferring networks of diffusion and influence. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (4):21:1–21:??, February 2012.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Guo:2020:RAL

[GSG^{+}20] Jipeng Guo, Yanfeng Sun, Junbin Gao, Yongli Hu, and Baocai Yin. Robust adaptive linear discriminant analysis with bidirectional reconstruction constraint. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):75:1–75:20, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3409478.

Gallardo:2019:IVE

[GSI19] Laura Fernández Gallardo and Ramon Sanchez-Iborra. On the impact of voice encoding and transmission on the predictions of speaker warmth and attractiveness. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(4):40:1–40:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3332146.

Grabocka:2016:LTS

[GSST16] Josif Grabocka, Nicolas Schilling, and Lars Schmidt-Thieme. Latent time-series motifs. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):6:1–6:??, August 2016.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ghosh:2012:SIB

[GSTC12] Joydeep Ghosh, Padhraic Smyth, Andrew Tomkins, and Rich Caruana. Special issue on best of SIGKDD 2011. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(4):14:1–14:??, December 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Guo:2020:NAF

[GSWJ20] Yuan Guo, Yu Sun, Kai Wu, and Kerong Jiang. New algorithms of feature selection and big data assignment for CBR system integrated by Bayesian network. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):23:1–23:20, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3373086.

Nicola:2024:IRM

[GVdSL24] Victor Gomes De Oliveira Martins Nicola, Karina Valdivia Delgado, and Marcelo de Souza Lauretto. Imbalancerobust multi-label self-adjusting kNN. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):190:1– 190:??, September 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3663575.

Guo:2020:IMS

[GW20] Jianxiong Guo and Weili Wu. Influence maximization: Seeding based on community structure. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):66:1–66:22, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3399661.

Guo:2021:AIM

[GW21] Jianxiong Guo and Weili Wu.

Adaptive influence maximization: If influential node unwilling to be the seed. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):84:1–84:23, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447396.

Guo:2017:CSN

[GWZZ17] Ting Guo, Jia Wu, Xingquan Zhu, and Chengqi Zhang. Combining structured node content and topology information for networked graph clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3): 29:1–29:??, April 2017. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Geeganage:2024:SET

[GXL24] Dakshi Kapugama Geeganage, Yue Xu, and Yuefeng Li. A semantics-enhanced topic modelling technique: Semantic-LDA. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):93:1–93:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3639409.

Gao:2021:TPO

[GXLC21] Xiaofeng Gao, Wenyi Xu, Mingding Liao, and Guihai Chen. Trust prediction for online social networks with integrated time-aware similarity. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):95:1–95:30, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447682.

Guo:2022:DFE

[GXZ^{+}22] Mengzhuo Guo, Zhongzhi Xu, Qingpeng Zhang, Xiuwu Liao, and Jiapeng Liu. Deciphering feature effects on decisionmaking in ordinal regression problems: an explainable ordinal factorization model. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):59:1–59:26, June 2022.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487048.

Gopal:2015:HBI

[GY15] Siddharth Gopal and Yiming Yang. Hierarchical Bayesian inference and recursive regularization for large-scale classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3): 18:1–18:??, April 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Gao:2022:GBS

[GYX^{+}22] Jianliang Gao, Xiaoting Ying, Cong Xu, Jianxin Wang, Shichao Zhang, and Zhao Li. Graph-based stock recommendation by timeaware relational attention network. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (1):4:1–4:21, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451397.

Gonzalez-Zelaya:2024:FPD

[GZSMM24] Vladimiro González-Zelaya, Julián Salas, David Megías, and Paolo Missier. Fair and private data preprocessing through microaggregation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):49:1– 49:??, April 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3617377.

Guo:2016:MDM

[GZXF16] Zhen Guo, Zhongfei (Mark) Zhang, Eric P. Xing, and Christos Faloutsos. Multimodal data mining in a multimedia database based on structured max margin learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(3):23:1– 23:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hassan:2023:CGD

[HAK^{+}23] Zohair Raza Hassan, Sarwan Ali, Imdadullah Khan, Mudassir Shabbir, and Waseem Abbas. Computing graph descriptors on edge streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(8):112:1–112:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3591468.

Hashimoto:2008:NEP

[HAKU^{+}08] Kosuke Hashimoto, Kiyoko Flora Aoki-Kinoshita, Nobuhisa Ueda, Minoru Kanehisa, and Hiroshi Mamitsuka. A new efficient probabilistic model for mining labeled ordered trees applied to glycobiology. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(1):6:1–6:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Han:2007:I

[Han07] Jiawei Han. Introduction.

ACM Transactions on Knowledge Discovery from Data (TKDD), 1(1):1:1–1:??, March 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Han:2024:GBM

[Han24] Yuehui Han. Generation-based multi-view contrast for selfsupervised graph representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):130:1–130:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3645095.

He:2024:VAD

[HBD^{+}24] Yifan He, Yatao Bian, Xi Ding, Bingzhe Wu, Jihong Guan, Ji Zhang, and Shuigeng Zhou. Variate associated domain adaptation for unsupervised multivariate time series anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):187:1–187:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663573.

Huang:2018:EBQ

[HCC^{+}18] Zhipeng Huang, Bogdan Cautis, Reynold Cheng, Yudian Zheng, Nikos Mamoulis, and Jing Yan. Entity-based query recommendation for long-tail queries. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (6):64:1–64:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hsu:2024:AAC

[HCH24a] Chi-Wei Hsu, Chiao-Ting Chen, and Szu-Hao Huang. Adaptive adversarial contrastive learning for crossdomain recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):57:1–57:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630259.

Huang:2024:LCP

[HCH^{+}24b] Shenyang Huang, Samy Coulombe, Yasmeen Hitti, Reihaneh Rabbany, and Guillaume Rabusseau. Laplacian change point detection for single and multi-view dynamic graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):63:1–63:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL

https://dl.acm.org/doi/10.
1145/3631609.

Huang:2023:GCD

[HCY^{+}23] Jingmin Huang, Bowei Chen, Zhi Yan, Iadh Ounis, and Jun Wang. GEO: a computational design framework for automotive exterior facelift. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):82:1–82:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578521.

Huai:2014:TPC

[HCZ^{+}14] Baoxing Huai, Enhong Chen, Hengshu Zhu, Hui Xiong, Tengfei Bao, Qi Liu, and Jilei Tian. Toward personalized context recognition for mobile users: a semisupervised Bayesian HMM approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2): 10:1–10:??, September 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Huang:2007:TTE

[HDC07] Jen-Wei Huang, Bi-Ru Dai, and Ming-Syan Chen. Twain: Two-end association miner with precise frequent exhibition periods. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(2):8:1– 8:??, August 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hu:2018:CFT

[HDQ^{+}18] Guang-Neng Hu, Xin-Yu Dai, Feng-Yu Qiu, Rui Xia, Tao Li, Shu-Jian Huang, and Jia-Jun Chen. Collaborative filtering with topic and social latent factors incorporating implicit feedback. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2): 23:1–23:??, March 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Duong:2021:DGF

[hDRNlD21] Quang huy Duong, Heri Ramampiaro, Kjetil Nørvåg, and Thu lan Dam. Density guarantee on finding multiple subgraphs and subtensors. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):76:1–76:32, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3446668.

Huang:2018:WTC

[HDT^{+}18] Hong Huang, Yuxiao Dong, Jie Tang, Hongxia Yang, Nitesh V. Chawla, and Xiaoming Fu. Will triadic closure strengthen ties in social networks? ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):30:1–30:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Huh:2012:DTM

[HF12] Seungil Huh and Stephen E.

Fienberg. Discriminative topic modeling based on manifold learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (4):20:1–20:??, February 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hu:2024:TDP

[HF24] Wentao Hu and Hui Fang. Towards differential privacy in sequential recommendation: a noisy graph neural network approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):125:1–125:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643821.

Huang:2022:AMP

[HFL^{+}22] Chenji Huang, Yixiang Fang, Xuemin Lin, Xin Cao, and Wenjie Zhang. ABLE: Meta-path prediction in heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(4):73:1– 73:21, August 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3494558.

Huang:2024:EEA

[HGS^{+}24] Yinqiu Huang, Min Gao, Kai Shu, Chenghua Lin, Jia Wang, and Wei Zhou. EML: Emotion-aware meta learning for cross-event false information detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):185:1–185:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3661485.

Halkidi:2008:CFB

[HGV^{+}08] M. Halkidi, D. Gunopulos, M. Vazirgiannis, N. Kumar, and C. Domeniconi. A clustering framework based on subjective and objective validity criteria. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(4):4:1–4:??, January 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hao:2021:DEI

[HGW^{+}21] Shaoyang Hao, Bin Guo, Hao Wang, Yunji Liang, Lina Yao, Qianru Wang, and Zhiwen Yu. DeepDepict: Enabling information rich, personalized product description generation with the deep multiple pointer generator network. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):83:1–83:16, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446982.

Huang:2023:URD

[HGWY23] Gengsen Huang, Wensheng Gan, Jian Weng, and Philip S. Yu. US-Rule: Discovering utility-driven sequential rules. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):10:1– 10:??, January 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3532613.

Huang:2024:TTS

[HGY24] Gengsen Huang, Wensheng Gan, and Philip S. Yu. TaSPM: Targeted sequential pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):114:1–114:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639827.

Henzgen:2019:MRD

[HH19] Sascha Henzgen and Eyke Hüllermeier. Mining rank data. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):59:1– 59:??, December 2019. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3363572.

Han:2024:LUB

[HHL^{+}24] Di Han, Yifan Huang, Junmin Liu, Kai Liao, and Kunling Lin. LSAB: User behavioral pattern modeling in sequential recommendation by learning self-attention bias. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):62:1–62:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3632625.

Hao:2018:OAL

[HHZ^{+}18] Shuji Hao, Peiying Hu, Peilin Zhao, Steven C. H. Hoi, and Chunyan Miao. Online active learning with expert advice. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5): 58:1–58:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

He:2018:SRI

[HK18] Xinran He and David Kempe. Stability and robustness in influence maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):66:1–66:??, October 2018.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Halstead:2023:CDM

[HKR^{+}23] Ben Halstead, Yun Sing Koh, Patricia Riddle, Mykola Pechenizkiy, and Albert Bifet. Combining diverse metafeatures to accurately identify recurring concept drift in data streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):107:1–107:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587098.

Han:2022:EAI

[HL22] Juhee Han and Younghoon Lee. Explainable artificial intelligence-based competitive factor identification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):10:1–10:11, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451529.

Hong:2019:VGM

[HLC19] Junyuan Hong, Yang Li, and Huanhuan Chen. Variant Grassmann manifolds: a representation augmentation method for action recognition. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):23:1–23:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3314203.

Hua:2020:PTM

[HLCR20] Ting Hua, Chang-Tien Lu, Jaegul Choo, and Chandan K. Reddy. Probabilistic topic modeling for comparative analysis of document collections. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (2):24:1–24:27, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3369873.

Huang:2024:RRH

[HLH^{+}24] Jincheng Huang, Ping Li, Rui Huang, Na Chen, and Acong Zhang. Revisiting the role of heterophily in graph representation learning: an edge classification perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):13:1–13:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3603378.

Hu:2020:EUP

[HLL^{+}20] Renjun Hu, Yanchi Liu, Yanyan Li, Jingbo Zhou, Shuai Ma, and Hui Xiong. Exploiting user preference and mobile peer influence for human mobility annotation. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):67:1–67:18, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3406600.

Huang:2018:GEF

[HLZH18] Xiao Huang, Jundong Li, Na Zou, and Xia Hu. A general embedding framework for heterogeneous information learning in large-scale networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):70:1– 70:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Huai:2020:LDM

[HML^{+}20] Mengdi Huai, Chenglin Miao, Yaliang Li, Qiuling Suo, Lu Su, and Aidong Zhang. Learning distance metrics from probabilistic information. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(5):53:1– 53:33, August 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3364320.

Huang:2024:MII

[HML^{+}24] Qiang Huang, Jing Ma, Jundong Li, Ruocheng Guo, Huiyan Sun, and Yi Chang. Modeling interference for individual treatment effect estimation from networked observational data. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):48:1–48:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3628449.

Huang:2013:STP

[HNH^{+}13] Jin Huang, Feiping Nie, Heng Huang, Yi-Cheng Tu, and Yu Lei. Social trust prediction using heterogeneous networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(4): 17:1–17:??, November 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Huang:2014:RMN

[HNHD14] Jin Huang, Feiping Nie, Heng Huang, and Chris Ding. Robust manifold nonnegative matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(3):11:1–11:??, June 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hernandez-Orallo:2014:PRC

[HO14] José Hernández-Orallo. Probabilistic reframing for costsensitive regression. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(4):17:1–17:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hasan:2020:NSA

[HP20] Md Kamrul Hasan and Christopher Pal. A new smooth approximation to the zero one loss with a probabilistic interpretation. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(1):1:1–1:28, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3365672.

Hu:2022:SDP

[HQWW22] Yue Hu, Ao Qu, Yanbing Wang, and Daniel B. Work. Streaming data preprocessing via online tensor recovery for large environmental sensor networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):122:1–122:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532189.

Huang:2014:PBA

[HQYY14] Hao Huang, Hong Qin, Shinjae Yoo, and Dantong Yu. Physics-based anomaly detection defined on manifold space. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2): 14:1–14:??, September 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hidalgo:2022:DAD

[HSB22] Juan I. G. Hidalgo, Silas G. T. C. Santos, and Roberto S. M. Barros. Dynamically adjusting diversity in ensembles for the classification of data streams with concept drift. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):31:1–31:20, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3466616.

He:2019:KSA

[HSBH19] Kun He, Pan Shi, David Bindel, and John E. Hopcroft. Krylov subspace approximation for local community detection in large networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):52:1–52:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3340708.

Hooi:2017:GBF

[HSS^{+}17] Bryan Hooi, Kijung Shin, Hyun Ah Song, Alex Beutel, Neil Shah, and Christos Faloutsos. Graph-based fraud detection in the face of camouflage. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (4):44:1–44:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hermanns:2023:GSG

[HST^{+}23] Judith Hermanns, Konstantinos Skitsas, Anton Tsitsulin, Marina Munkhoeva, Alexander Kyster, Simon Nielsen, Alexander M. Bronstein, Davide Mottin, and Panagiotis Karras. GRASP: Scalable graph alignment by spectral corresponding functions. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):50:1–50:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561058.

Huang:2021:MSN

[HSY^{+}21] Hong Huang, Yu Song, Fanghua Ye, Xing Xie, Xuanhua Shi, and Hai Jin. Multistage network embedding for exploring heterogeneous edges. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):5:1–5:27, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3415157.

Han:2024:DHH

[HTT^{+}24] Jiadi Han, Yufei Tang, Qian Tao, Yuhan Xia, and Liming Zhang. Dual homogeneity hypergraph motifs with cross-view contrastive learning for multiple social recommendations. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):158:1–158:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653976.

Hu:2021:RTR

[HW21] Yue Hu and Daniel B. Work. Robust tensor recovery with fiber outliers for traffic events. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):6:1–6:27, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3417337.

Huang:2024:EEF

[HWZ^{+}24] Yimin Huang, Wanwan Wang, Xingying Zhao, Yukun Wang, Xinyu Feng, Hao He, and Ming Yao. EFMVFL: an efficient and flexible multi-party vertical federated learning without a third party. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):51:1–51:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3627993.

Huang:2019:RNC

[HXY^{+}19] Yourong Huang, Zhu Xiao, Xiaoyou Yu, Dong Wang, Vincent Havyarimana, and Jing Bai. Road network construction with complex intersections based on sparsely sampled private car trajectory data. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):35:1–35:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326060.

Huang:2015:DAC

[HYYQ15] Hao Huang, Shinjae Yoo, Dantong Yu, and Hong Qin. Density-aware clustering based on aggregated heat kernel and its transformation. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4): 29:1–29:??, June 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Huang:2021:DGM

[HYYT21] Yu Huang, Josh Jia-Ching Ying, Philip S. Yu, and Vincent S. Tseng. Dynamic graph mining for multi-weight multidestination route planning with deadlines constraints.

ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):3:1–3:32, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3412363.

Han:2020:GLS

[HZL20] Huimei Han, Xingquan Zhu, and Ying Li. Generalizing long short-term memory network for deep learning from generic data. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (2):13:1–13:28, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3366022.

Huai:2022:RML

[HZM^{+}22] Mengdi Huai, Tianhang Zheng, Chenglin Miao, Liuyi Yao, and Aidong Zhang. On the robustness of metric learning: an adversarial perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):95:1– 95:25, October 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502726.

Hu:2015:PSD

[HZW^{+}15] Juhua Hu, De-Chuan Zhan, Xintao Wu, Yuan Jiang, and Zhi-Hua Zhou. Pairwised specific distance learning from physical linkages. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3): 20:1–20:??, April 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Hong:2015:CAR

[HZZ^{+}15] Liang Hong, Lei Zou, Cheng Zeng, Luming Zhang, Jian Wang, and Jilei Tian. Contextaware recommendation using role-based trust network. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):13:1–13:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ihou:2022:SVO

[IAB22] Koffi Eddy Ihou, Manar Amayri, and Nizar Bouguila. Stochastic variational optimization of a hierarchical Dirichlet process latent beta-Liouville topic model. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):84:1–84:48, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502727.

Ihler:2007:LDE

[IHS07] Alexander Ihler, Jon Hutchins, and Padhraic Smyth. Learning to detect events with Markovmodulated Poisson processes.

ACM Transactions on Knowledge Discovery from Data (TKDD), 1(3):13:1–13:??, December 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Islam:2008:STS

[II08] Aminul Islam and Diana Inkpen. Semantic text similarity using corpus-based word similarity and string similarity. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(2):10:1– 10:??, July 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Iqbal:2019:BPE

[IKK19] Mohsin Iqbal, Asim Karim, and Faisal Kamiran. Balancing prediction errors for robust sentiment classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):33:1–33:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3328795.

Ienco:2012:CDL

[IPM12] Dino Ienco, Ruggero G. Pensa, and Rosa Meo. From context to distance: Learning dissimilarity for categorical data clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(1):1:1–1:??, March 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Iwata:2012:SMT

[IYSU12] Tomoharu Iwata, Takeshi Yamada, Yasushi Sakurai, and Naonori Ueda. Sequential modeling of topic dynamics with multiple timescales. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(4):19:1–19:??, February 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jian:2024:SSS

[JBGW24] Meng Jian, Yulong Bai, Jingjing Guo, and Lifang Wu. Swarm self-supervised hypergraph embedding for recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):73:1–73:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638058.

Jiang:2016:CSB

[JCB^{+}16] Meng Jiang, Peng Cui, Alex Beutel, Christos Faloutsos, and Shiqiang Yang. Catching synchronized behaviors in large networks: a graph mining approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):35:1–35:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jiang:2023:LLM

[JCC23] Linli Jiang, Chao-Xiong Chen, and Chao Chen. L2MM: Learning to map matching with deep models for low-quality GPS trajectory data. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):39:1– 39:??, April 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550486.

Jia:2023:SSB

[JCQ^{+}23] Bohan Jia, Jian Cao, Shiyou Qian, Nengjun Zhu, Xin Dong, Liang Zhang, Lei Cheng, and Linjian Mo. SMONE: a session-based recommendation model based on neighbor sessions with similar probabilistic intentions. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):111:1–111:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587099.

Jalali:2023:FIF

[JCS^{+}23] Zeinab S. Jalali, Qilan Chen, Shwetha M. Srikanta, Weixiang Wang, Myunghwan Kim, Hema Raghavan, and Sucheta Soundarajan. Fairness of information flow in social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):79:1–79:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578268.

Jin:2012:MML

[JDE^{+}12] Yu Jin, Nick Duffield, Jeffrey Erman, Patrick Haffner, Subhabrata Sen, and Zhi-Li Zhang. A modular machine learning system for flow-level traffic classification in large networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(1):4:1–4:??, March 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jaysawal:2019:PAP

[JH19] Bijay Prasad Jaysawal and Jen-Wei Huang. PSP-AMS: Progressive mining of sequential patterns across multiple streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (1):5:1–5:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3281632.

Jafariakinabad:2022:SSR

[JH22] Fereshteh Jafariakinabad and Kien A. Hua. A self-supervised representation learning of sentence structure for authorship attribution. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):68:1–68:16, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3491203.

Jiang:2023:MLS

[JHC^{+}23] Shaowei Jiang, Wei He, Lizhen Cui, Yonghui Xu, and Lei Liu. Modeling long- and shortterm user preferences via selfsupervised learning for next POI recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):125:1–125:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3597211.

Jin:2022:TUE

[JHJK22] Junchen Jin, Mark Heimann, Di Jin, and Danai Koutra. Toward understanding and evaluating structural node embeddings. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (3):58:1–58:32, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3481639.

Jiang:2024:TLA

[Jia24] Meng Jiang. Transfer learning across graph convolutional networks: Methods, theory, and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):24:1–24:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617376.

Jirina:2022:DFO

[JK22] Marcel Jirina and Said Krayem.

The distance function optimization for the near neighborsbased classifiers. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):101:1–101:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434769.

Jang:2023:SST

[JK23] Jun-Gi Jang and U. Kang.

Static and streaming Tucker decomposition for dense tensors. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):66:1– 66:??, June 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3568682.

Jiang:2021:DLB

[JKP^{+}21] Jie Jiang, Qiuqiang Kong, Mark D. Plumbley, Nigel Gilbert, Mark Hoogendoorn, and Diederik M. Roijers. Deep learning-based energy disaggregation and on/off detection of household appliances.

ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):50:1–50:21, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441300.

Jiang:2019:PFS

[JLD^{+}19] Bingbing Jiang, Chang Li, Maarten De Rijke, Xin Yao, and Huanhuan Chen. Probabilistic feature selection and classification vector machine. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):21:1–21:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3309541.

Ji:2013:PFR

[JLH^{+}13] Ming Ji, Binbin Lin, Xiaofei He, Deng Cai, and Jiawei Han. Parallel field ranking. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3):15:1–15:??, September 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jin:2014:SAR

[JLL14] Ruoming Jin, Victor E. Lee, and Longjie Li. Scalable and axiomatic ranking of network role similarity. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(1):3:1– 3:??, February 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jiang:2022:NPO

[JLZ^{+}22] Yuanchun Jiang, Ruicheng Liang, Ji Zhang, Jianshan Sun, Yezheng Liu, and Yang Qian. Network public opinion detection during the coronavirus pandemic: a short-text relational topic model. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):52:1–52:27, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480246.

Jin:2008:CMM

[JMR08] Ying Jin, T. M. Murali, and Naren Ramakrishnan. Compositional mining of multirelational biological datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(1):2:1–2:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jiang:2009:MFC

[JP09] Daxin Jiang and Jian Pei.

Mining frequent cross-graph quasi-cliques. ACM Transactions on Knowledge Discovery from Data (TKDD), 2 (4):16:1–16:??, January 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jang:2022:FKS

[JPJ^{+}22] Jun-Gi Jang, Chaeheum Park, Changwon Jang, Geonsoo Kim, and U. Kang. Finding key structures in MMORPG graph with hierarchical graph summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):115:1–115:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522691.

Jha:2023:SCL

[JRC^{+}23] Akshita Jha, Vineeth Rakesh, Jaideep Chandrashekar, Adithya Samavedhi, and Chandan K. Reddy. Supervised contrastive learning for interpretable long-form document matching. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (2):27:1–27:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542822.

Joaristi:2021:SGF

[JS21] Mikel Joaristi and Edoardo Serra. SIR-GN: a fast structural iterative representation learning approach for graph nodes. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):100:1– 100:39, July 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450315.

Jang:2023:AOS

[JSE^{+}23] Jun-Gi Jang, Sooyeon Shim, Vladimir Egay, Jeeyong Lee, Jongmin Park, Suhyun Chae, and U. Kang. Accurate openset recognition for memory workload. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (9):126:1–126:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597027.

Ji:2024:CFM

[JSF^{+}24] Taoran Ji, Nathan Self, Kaiqun Fu, Zhiqian Chen, Naren Ramakrishnan, and Chang-Tien Lu. Citation forecasting with multi-context attention-aided dependency modeling. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):144:1–144:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649140.

Ju:2019:TRB

[JSG^{+}19] Fujiao Ju, Yanfeng Sun, Junbin Gao, Michael Antolovich, Junliang Dong, and Baocai Yin. Tensorizing restricted Boltzmann machine. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):30:1–30:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3321517.

Jha:2015:SES

[JSP15] Madhav Jha, C. Seshadhri, and Ali Pinar. A spaceefficient streaming algorithm for estimating transitivity and triangle counts using the birthday paradox. ACM Transactions on Knowledge Discovery from Data (TKDD), 9 (3):15:1–15:??, February 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jia:2015:SGR

[JSV^{+}15] Adele Lu Jia, Siqi Shen, Ruud Van De Bovenkamp, Alexandru Iosup, Fernando Kuipers, and Dick H. J. Epema. Socializing by gaming: Revealing social relationships in multiplayer online games. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (2):11:1–11:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ji:2010:SSL

[JTYY10] Shuiwang Ji, Lei Tang, Shipeng Yu, and Jieping Ye. A shared-subspace learning framework for multi-label classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(2):8:1–8:??, May 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Jan:2020:ECC

[JV20] Zohaib Md. Jan and Brijesh Verma. Evolutionary classifier and cluster selection approach for ensemble classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):7:1–7:18, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3366633.

Jha:2019:DMD

[JXGZ19] Kishlay Jha, Guangxu Xun, Vishrawas Gopalakrishnan, and Aidong Zhang. DWE-Med: Dynamic word embeddings for medical domain. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):19:1– 19:??, June 2019. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3310254.

Jiang:2019:BMS

[JYD19] Fei Jiang, Guosheng Yin, and Francesca Dominici. Bayesian model selection approach to multiple change-points detection with non-local prior distributions. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):48:1–48:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3340804.

Jiang:2023:DTL

[JYH^{+}23] Xin Jiang, Zhengxin Yu, Chao Hai, Hongbo Liu, Xindong Wu, and Tomas Ward. DNformer: Temporal link prediction with transfer learning in dynamic networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):43:1–43:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551892.

Ji:2021:ALS

[JYY^{+}21] Yugang Ji, Mingyang Yin, Hongxia Yang, Jingren Zhou, Vincent W. Zheng, Chuan Shi, and Yuan Fang. Accelerating large-scale heterogeneous interaction graph embedding learning via importance sampling. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (1):10:1–10:23, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418684.

Jerez:2023:EAD

[JZS23] Carlos Ivan Jerez, Jun Zhang, and Marcia R. Silva. On equivalence of anomaly detection algorithms. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):21:1–21:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3536428.

Jing:2023:LSR

[JZX^{+}23] Mengyuan Jing, Yanmin Zhu, Yanan Xu, Haobing Liu, Tianzi Zang, Chunyang Wang, and Jiadi Yu. Learning shared representations for recommendation with dynamic heterogeneous graph convolutional networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):59:1–59:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565575.

Kumar:2016:ACT

[KBR^{+}16] Dheeraj Kumar, James C.

Bezdek, Sutharshan Rajasegarar, Marimuthu Palaniswami, Christopher Leckie, Jeffrey Chan, and Jayavardhana Gubbi. Adaptive cluster tendency visualization and anomaly detection for streaming data. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 24:1–24:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kuang:2020:TEE

[KCL^{+}20] Kun Kuang, Peng Cui, Bo Li, Meng Jiang, Yashen Wang, Fei Wu, and Shiqiang Yang. Treatment effect estimation via differentiated confounder balancing and regression. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):6:1–6:25, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3365677.

Kuo:2024:CAG

[KCP^{+}24] Chuan-Wei Kuo, Bo-Yu Chen, Wen-Chih Peng, Chih-Chieh Hung, and Hsin-Ning Su. Correlation-aware graph data augmentation with implicit and explicit neighbors. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):109:1–109:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638057.

Kirielle:2023:UGB

[KCR23] Nishadi Kirielle, Peter Christen, and Thilina Ranbaduge.

Unsupervised graph-based entity resolution for complex entities. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):12:1– 12:??, January 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3533016.

Kaibiao:2024:ANG

[KCR^{+}24] Lin Kaibiao, Jinpo Chen, Chen Ruicong, Yang Fan, Zhang Yang, Lin Min, and Lu Ping. Adaptive neighbor graph aggregated graph attention network for heterogeneous graph embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):29:1– 29:??, January 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3616377.

Khokhar:2023:DPR

[KFI^{+}23] Rashid Hussain Khokhar, Benjamin C. M. Fung, Farkhund Iqbal, Khalil Al-Hussaeni, and Mohammed Hussain. Differentially private release of heterogeneous network for managing healthcare data. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):90:1–90:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580367.

Khodabandehlou:2024:FUF

[KG24] Samira Khodabandehlou and Alireza Hashemi Golpayegani. FiFrauD: Unsupervised financial fraud detection in dynamic graph streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):111:1–111:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3641857.

Khodadadi:2018:CTU

[KHTR18] Ali Khodadadi, Seyed Abbas Hosseini, Erfan Tavakoli, and Hamid R. Rabiee. Continuoustime user modeling in presence of badges: a probabilistic approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3): 37:1–37:??, April 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ke:2022:MRG

[KKB22] Xiangyu Ke, Arijit Khan, and Francesco Bonchi. Multirelation graph summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):82:1– 82:30, October 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3494561.

Kriegel:2009:CHD

[KKZ09] Hans-Peter Kriegel, Peer Kröger, and Arthur Zimek. Clustering high-dimensional data: a survey on subspace clustering, pattern-based clustering, and correlation clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(1):1:1–1:??, March 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kwon:2023:EBM

[KL23] Soonki Kwon and Younghoon Lee. Explainability-based mixup approach for text data augmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (1):13:1–13:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3533048.

Khan:2024:HBP

[KMH^{+}24] Mehak Khan, Gustavo B. M.

Mello, Laurence Habib, Paal Engelstad, and Anis Yazidi. HITS-based propagation paradigm for graph neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):100:1–100:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638779.

Kamat:2018:SBA

[KN18] Niranjan Kamat and Arnab Nandi. A session-based approach to fast-but-approximate interactive data cube exploration. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):9:1–9:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Koren:2007:MEP

[KNV07] Yehuda Koren, Stephen C. North, and Chris Volinsky. Measuring and extracting proximity graphs in networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(3):12:1–12:??, December 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Koren:2010:FNS

[Kor10] Yehuda Koren. Factor in the neighbors: Scalable and accurate collaborative filtering. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(1):1:1–1:??, January 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kaushal:2018:ETP

[KP18] Vishal Kaushal and Manasi Patwardhan. Emerging trends in personality identification using online social networks — a literature survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):15:1–15:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Koh:2016:URP

[KR16] Yun Sing Koh and Sri Devi Ravana. Unsupervised rare pattern mining: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):45:1–45:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Katib:2019:FAS

[KRBK19] Anas Katib, Praveen Rao, Kobus Barnard, and Charles Kamhoua. Fast approximate score computation on large-scale distributed data for learning multinomial Bayesian networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (2):14:1–14:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301304.

Kaufman:2012:LDM

[KRPS12] Shachar Kaufman, Saharon Rosset, Claudia Perlich, and Ori Stitelman. Leakage in data mining: Formulation, detection, and avoidance. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(4):15:1–15:??, December 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kumar:2021:AGN

[KS21] Suhansanu Kumar and Hari Sundaram. Attribute-guided network sampling mechanisms. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):69:1– 69:24, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3441445.

Kose:2024:FFA

[KS24] O. Deniz Kose and Yanning Shen. FairGAT: Fairnessaware graph attention networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):164:1– 164:??, August 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3645096.

Koley:2021:DEE

[KSB^{+}21] Paramita Koley, Avirup Saha, Sourangshu Bhattacharya, Niloy Ganguly, and Abir De. Demarcating endogenous and exogenous opinion dynamics: an experimental design approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):99:1–99:25, July 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3449361.

Kimura:2009:BLM

[KSM09] Masahiro Kimura, Kazumi Saito, and Hiroshi Motoda. Blocking links to minimize contamination spread in a social network. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(2):9:1– 9:??, April 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Koutra:2016:DPM

[KSV^{+}16] Danai Koutra, Neil Shah, Joshua T. Vogelstein, Brian Gallagher, and Christos Faloutsos. DeltaCon: Principled massive-graph similarity function with attribution. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(3):28:1–28:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kiernan:2009:CCS

[KT09] Jerry Kiernan and Evimaria Terzi. Constructing comprehensive summaries of large event sequences. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4): 21:1–21:??, November 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kang:2011:HMR

[KTA^{+}11] U. Kang, Charalampos E. Tsourakakis, Ana Paula Appel, Christos Faloutsos, and Jure Leskovec. HADI: Mining radii of large graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(2):8:1–8:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kandylas:2010:AKC

[KUU10] Vasileios Kandylas, S. Phineas Upham, and Lyle H. Ungar. Analyzing knowledge communities using foreground and background clusters. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(2):7:1–7:??, May 2010. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Kamhoua:2023:GGG

[KZM^{+}23] Barakeel Fanseu Kamhoua, Lin Zhang, Kaili Ma, James Cheng, Bo Li, and Bo Han. GRACE: a general graph convolution framework for attributed graph clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):31:1– 31:??, April 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3544977.

Kuang:2022:BSS

[KZW^{+}22] Kun Kuang, Hengtao Zhang, Runze Wu, Fei Wu, Yueting Zhuang, and Aijun Zhang. Balance-subsampled stable prediction across unknown test data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (3):45:1–45:21, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477052.

Kong:2020:GSS

[KZZ^{+}20] Xiangjie Kong, Jun Zhang, Da Zhang, Yi Bu, Ying Ding, and Feng Xia. The gene of scientific success. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):41:1–41:19, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3385530.

Liang:2018:PRA

[LAN^{+}18] Jiongqian Liang, Deepak Ajwani, Patrick K. Nicholson, Alessandra Sala, and Srinivasan Parthasarathy. Prioritized relationship analysis in heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):29:1– 29:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lappas:2020:MCP

[Lap20] Theodoros Lappas. Mining career paths from large resume databases: Evidence from IT professionals. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):37:1–37:38, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3379984.

Long:2023:MAC

[LBT^{+}23] Qiang Long, Adil Bagirov, Sona Taheri, Nargiz Sultanova, and Xue Wu. Methods and applications of clusterwise linear regression: a survey and comparison. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):36:1–36:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3550074.

Lagree:2019:AOI

[LCCM19] Paul Lagrée, Olivier Cappé, Bogdan Cautis, and Silviu Maniu. Algorithms for online influencer marketing. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):3:1–3:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3274670.

Liu:2019:LKI

[LCF19] Chenyang Liu, Jian Cao, and Shanshan Feng. Leveraging kernel-incorporated matrix factorization for app recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):31:1–31:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3320482.

Liu:2018:MAD

[LCG^{+}18] Xiaoli Liu, Peng Cao, André R.

Gonçalves, Dazhe Zhao, and Arindam Banerjee. Modeling Alzheimer’s disease progression with fused Laplacian sparse group lasso. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):65:1–65:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2017:SPM

[LCLL17] Xutong Liu, Feng Chen, Yen-Cheng Lu, and Chang-Tien Lu. Spatial prediction for multivariate non-Gaussian data. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):36:1– 36:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2021:EFM

[LCLL21] Chongshou Li, Brenda Cheang, Zhixing Luo, and Andrew Lim. An exponential factorization machine with percentage error minimization to retail sales forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):19:1–19:32, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3426238.

Liu:2014:ADI

[LCN14] Siyuan Liu, Lei Chen, and Lionel M. Ni. Anomaly detection from incomplete data. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2):11:1–11:??, September 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2024:ASG

[LCW24] Xiaoting Li, Lingwei Chen, and Dinghao Wu. Adversary for social good: Leveraging adversarial attacks to protect personal attribute privacy. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):46:1– 46:??, February 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3614098.

Lin:2020:BCR

[LCWC20] Chi-Chun Lin, Kun-Ta Chuang, Wush Chi-Hsuan Wu, and Ming-Syan Chen. Budgetconstrained real-time bidding optimization: Multiple predictors make it better. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):14:1–14:27, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375393.

Liu:2023:TTA

[LCX^{+}23] Zhi Liu, Yang Chen, Feng Xia, Jixin Bian, Bing Zhu, Guojiang Shen, and Xiangjie Kong. TAP: Traffic accident profiling via multi-task spatiotemporal graph representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):56:1–56:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564594.

Loh:2007:EEL

[LCZ07] Wei-Yin Loh, Chien-Wei Chen, and Wei Zheng. Extrapolation errors in linear model trees. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(2):6:1– 6:??, August 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lin:2009:ACT

[LCZ^{+}09] Yu-Ru Lin, Yun Chi, Shenghuo Zhu, Hari Sundaram, and Belle L. Tseng. Analyzing communities and their evolutions in dynamic social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(2):8:1–8:??, April 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liao:2024:APT

[LCZ^{+}24] Chengwu Liao, Chao Chen, Wanyi Zhang, Suiming Guo, and Chao Liu. AGENDA: Predicting trip purposes with a new graph embedding network and active domain adaptation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):206:1–206:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3677020.

Lin:2022:SMH

[LDS^{+}22] Jerry Chun-Wei Lin, Youcef Djenouri, Gautam Srivastava, Yuanfa Li, and Philip S. Yu. Scalable mining of high-utility sequential patterns with threetier MapReduce model. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):60:1–60:26, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487046.

Liu:2018:CGM

[LF18] Hongfu Liu and Yun Fu.

Consensus guided multi-view clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4): 42:1–42:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lu:2023:SDI

[LF23] Xun Lu and Songhe Feng.

Structure diversity-induced anchor graph fusion for multiview clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):17:1–17:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534931.

Liu:2017:MBM

[LFC^{+}17] Guannan Liu, Yanjie Fu, Guoqing Chen, Hui Xiong, and Can Chen. Modeling buying motives for personalized product bundle recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):28:1– 28:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lu:2023:DPE

[LFL^{+}23] Xun Lu, Songhe Feng, Gengyu Lyu, Yi Jin, and Congyan Lang. Distance-preserving embedding adaptive bipartite graph multi-view learning with application to multi-label classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):19:1–19:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3537900.

Li:2024:OGC

[LFXC24] Hui-Jia Li, Yuhao Feng, Chengyi Xia, and Jie Cao. Overlapping graph clustering in attributed networks via generalized cluster potential game. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):27:1–27:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597436.

Li:2023:DGC

[LFY^{+}23] Fuxian Li, Jie Feng, Huan Yan, Guangyin Jin, Fan Yang, Funing Sun, Depeng Jin, and Yong Li. Dynamic graph convolutional recurrent network for traffic prediction: Benchmark and solution. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):9:1–9:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532611.

Liu:2010:BBM

[LGF10] Chao Liu, Fan Guo, and Christos Faloutsos. Bayesian browsing model: Exact inference of document relevance from petabyte-scale data. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(4):19:1–19:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2023:SIM

[LGG^{+}23] Yandi Li, Haobo Gao, Yunxuan Gao, Jianxiong Guo, and Weili Wu. A survey on influence maximization: From an ML-based combinatorial optimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):133:1–133:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604559.

Liu:2023:PPP

[LGS^{+}23] Xiao Liu, Bonan Gao, Basem Suleiman, Han You, Zisu Ma, Yu Liu, and Ali Anaissi. Privacy-preserving personalized fitness recommender system P^{3}FitRec: a multi-level deep learning approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):76:1–76:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3572899.

Luan:2024:MIL

[LGT^{+}24] Tianxiang Luan, Shilin Gu, Xijia Tang, Wenzhang Zhuge, and Chenping Hou. Multiinstance learning with one side label noise. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):122:1–122:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3644076.

Liu:2021:KTW

[LGZ^{+}21] Yan Liu, Bin Guo, Daqing Zhang, Djamal Zeghlache, Jingmin Chen, Ke Hu, Sizhe Zhang, Dan Zhou, and Zhiwen Yu. Knowledge transfer with weighted adversarial network for cold-start store site recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):47:1–47:27, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3442203.

Liu:2022:WWW

[LGZ^{+}22] Hao Liu, Qingyu Guo, Hengshu Zhu, Fuzhen Zhuang, Shenwen Yang, Dejing Dou, and Hui Xiong. Who will win the data science competition? insights from KDD Cup 2019 and beyond. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):98:1–98:24, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3511896.

Liu:2023:CFU

[LGZ^{+}23] Hao Liu, Qingyu Guo, Hengshu Zhu, Yanjie Fu, Fuzhen Zhuang, Xiaojuan Ma, and Hui Xiong. Characterizing and forecasting urban vibrancy evolution: a multi-view graph mining perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):68:1–68:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568683.

Lin:2022:JPF

[LH22] Fandel Lin and Hsun-Ping Hsieh. A joint passenger flow inference and path recommender system for deploying new routes and stations of mass transit transportation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (1):6:1–6:36, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451393.

Liu:2024:CMM

[LHG^{+}24] Tengfei Liu, Yongli Hu, Junbin Gao, Yanfeng Sun, and Baocai Yin. Cross-modal multiple granularity interactive fusion network for long document classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):78:1–78:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3631711.

Li:2018:LSC

[LHK^{+}18] Yixuan Li, Kun He, Kyle Kloster, David Bindel, and John Hopcroft. Local spectral clustering for overlapping community detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):17:1–17:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lin:2022:MML

[LHL^{+}22] Yaojin Lin, Qinghua Hu, Jinghua Liu, Xingquan Zhu, and Xindong Wu. MULFE: Multi-label learning via labelspecific feature space ensemble. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (1):5:1–5:24, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451392.

Liu:2020:SFE

[LHN^{+}20] Bang Liu, Fred X. Han, Di Niu, Linglong Kong, Kunfeng Lai, and Yu Xu. Story Forest: Extracting events and telling stories from breaking news. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3): 31:1–31:28, May 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3377939.

Liu:2021:MMG

[LHS^{+}21] Bo Liu, Xi He, Mingdong Song, Jiangqiang Li, Guangzhi Qu, Jianlei Lang, and Rentao Gu. A method for mining Granger causality relationship on atmospheric visibility. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):92:1– 92:16, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3447681.

Li:2020:CDM

[LHW^{+}20] Pei-Zhen Li, Ling Huang, Chang-Dong Wang, Jian-Huang Lai, and Dong Huang. Community detection by motif-aware label propagation. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):22:1– 22:19, March 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3378537.

Li:2021:ANE

[LHW^{+}21] Juan-Hui Li, Ling Huang, Chang-Dong Wang, Dong Huang, Jian-Huang Lai, and Pei Chen. Attributed network embedding with micromeso structure. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):72:1–72:26, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441486.

Li:2023:EEG

[LHZ^{+}23] Yunyi Li, Yongjing Hao, Pengpeng Zhao, Guanfeng Liu, Yanchi Liu, Victor S. Sheng, and Xiaofang Zhou. Edgeenhanced global disentangled graph neural network for sequential recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):80:1–80:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3577928.

Li:2024:MUT

[LHZ^{+}24a] Tong Li, Shuodi Hui, Shiyuan Zhang, Huandong Wang, Yuheng Zhang, Pan Hui, Depeng Jin, and Yong Li. Mobile user traffic generation via multi-scale hierarchical GAN. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):189:1–189:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664655.

Liu:2024:LIG

[LHZ^{+}24b] Xinru Liu, Yongjing Hao, Lei Zhao, Guanfeng Liu, Victor S. Sheng, and Pengpeng Zhao. LMACL: Improving graph collaborative filtering with learnable model augmentation contrastive learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):177:1–177:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3657302.

Li:2013:CWM

[LHZG13] Bin Li, Steven C. H. Hoi, Peilin Zhao, and Vivekanand Gopalkrishnan. Confidence weighted mean reversion strategy for online portfolio selection. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(1):4:1–4:??, March 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2024:RGN

[LILJ24] Gang Liu, Eric Inae, Tengfei Luo, and Meng Jiang. Rationalizing graph neural networks with data augmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):86:1–86:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638781.

Luo:2024:SSG

[LJG^{+}24] Xiao Luo, Wei Ju, Yiyang Gu, Zhengyang Mao, Luchen Liu, Yuhui Yuan, and Ming Zhang. Self-supervised graphlevel representation learning with adversarial contrastive learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (2):34:1–34:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3624018.

Lim:2018:MEA

[LJK18] Yongsub Lim, Minsoo Jung, and U. Kang. Memoryefficient and accurate sampling for counting local triangles in graph streams: From simple to multigraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (1):4:1–4:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2023:DNE

[LJL^{+}23] He Li, Duo Jin, Xuejiao Li, Jianbin Huang, Xiaoke Ma, Jiangtao Cui, Deshuang Huang, Shaojie Qiao, and Jaesoo Yoo. DMGF-Net: an efficient dynamic multigraph fusion network for traffic prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (7):97:1–97:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3586164.

Liu:2022:BAM

[LJW^{+}22] Huafeng Liu, Liping Jing, Jingxuan Wen, Pengyu Xu, Jian Yu, and Michael K. Ng. Bayesian additive matrix approximation for social recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (1):7:1–7:34, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451391.

Li:2023:NWV

[LJX23] Huiru Li, Liangxiao Jiang, and Siqing Xue. Neighborhood weighted voting-based noise correction for crowdsourcing. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (7):96:1–96:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3586998.

Lin:2015:IMS

[LK15] Bing-Rong Lin and Daniel Kifer. Information measures in statistical privacy and data processing applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4):28:1–28:??, June 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lee:2020:LDS

[LK20] Kwang Hee Lee and Myoung Ho Kim. Linearization of dependency and sampling for participation-based betweenness centrality in very large B-hypergraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):25:1–25:41, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3375399.

Leskovec:2007:GED

[LKF07] Jure Leskovec, Jon Kleinberg, and Christos Faloutsos. Graph evolution: Densification and shrinking diameters. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(1):2:1–2:??, March 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lei:2019:IRU

[LL19] Yu Lei and Wenjie Li. Interactive recommendation with user-specific deep reinforcement learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):61:1–61:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3359554.

Liu:2024:EFC

[LL24] Zhe Liu and Sukumar Letchmunan. Enhanced fuzzy clustering for incomplete instance with evidence combination. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (3):72:1–72:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638061.

Liu:2022:RRH

[LLB^{+}22] Danlu Liu, Yu Li, William Baskett, Dan Lin, and Chi-Ren Shyu. RHPTree-risk hierarchical pattern tree for scalable long pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(4):63:1– 63:33, August 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3488380.

Li:2021:NNS

[LLC^{+}21] Dongsheng Li, Haodong Liu, Chao Chen, Yingying Zhao, Stephen M. Chu, and Bo Yang. NeuSE: a neural snapshot ensemble method for collaborative filtering. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (6):102:1–102:20, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3450526.

Li:2022:MTP

[LLC^{+}22] Yangfan Li, Kenli Li, Cen Chen, Xu Zhou, Zeng Zeng, and Keqin Li. Modeling temporal patterns with dilated convolutions for timeseries forecasting. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):14:1–14:22, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3453724.

Li:2024:ECB

[LLD^{+}24] Fangfang Li, Zhi Liu, Junwen Duan, Xingliang Mao, Heyuan Shi, and Shichao Zhang. Exploiting conversation-branch-tweet HyperGraph structure to detect misinformation on social media. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):33:1–33:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610297.

Lin:2023:MVG

[LLG^{+}23] Bei Lin, You Li, Ning Gui, Zhuopeng Xu, and Zhiwu Yu. Multi-view graph representation learning beyond homophily. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):114:1–114:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3592858.

Liang:2024:LEI

[LLH^{+}24] Yunji Liang, Lei Liu, Luwen Huangfu, Sagar Samtani, Zhiwen Yu, and Daniel D. Zeng. Learning entangled interactions of complex causality via self-paced contrastive learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):56:1– 56:??, April 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3632406.

Lin:2020:BRP

[LLL^{+}20] Bo Lin, Wei Luo, Zhiling Luo, Bo Wang, Shuiguang Deng, Jianwei Yin, and Mengchu Zhou. Bradykinesia recognition in Parkinson’s disease via single RGB video. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):16:1–16:19, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3369438.

Liu:2021:ADK

[LLL^{+}21] Huawen Liu, Enhui Li, Xinwang Liu, Kaile Su, and Shichao Zhang. Anomaly detection with kernel preserving embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):91:1–91:18, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447684.

Li:2023:VGA

[LLL23a] Dongjie Li, Dong Li, and Guang Lian. Variational graph autoencoder with adversarial mutual information learning for network representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):45:1–45:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3555809.

Li:2023:TFF

[LLL^{+}23b] Haoran Li, Zhiqiang Lv, Jianbo Li, Zhihao Xu, Yue Wang, Haokai Sun, and Zhaoyu Sheng. Traffic flow forecasting in the COVID-19: a deep spatial-temporal model based on discrete wavelet transformation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):64:1–64:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564753.

Li:2024:HMP

[LLL^{+}24] Munan Li, Kai Liu, Hongbo Liu, Zheng Zhao, Tomas E. Ward, and Xindong Wu. Heterogeneous meta-path graph learning for higher-order social recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):197:1–197:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673658.

Liu:2022:SGQ

[LLLW22] Guliu Liu, Lei Li, Guanfeng Liu, and Xindong Wu. Social group query based on multi-fuzzy-constrained strong simulation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):54:1–54:27, June 2022.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3481640.

Lin:2021:SEK

[LLS^{+}21] Mingkai Lin, Wenzhong Li, Lynda J. Song, Cam-Tu Nguyen, Xiaoliang Wang, and Sanglu Lu. SAKE: Estimating Katz centrality based on sampling for large-scale social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):66:1–66:21, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441646.

Li:2016:CBN

[LLW16] Xiang Li, Charles X. Ling, and Huaimin Wang. The convergence behavior of naive Bayes on large sparse datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):10:1–10:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2021:SAS

[LLW^{+}21] Kai Liu, Hongbo Liu, Tomas E. Ward, Hua Wang, Yu Yang, Bo Zhang, and Xindong Wu. Self-adaptive skeleton approaches to detect self-organized coalitions from brain functional networks through probabilistic mixture models. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):87:1–87:26, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447570.

Liu:2024:ETL

[LLX^{+}24] Bo Liu, Liangjiao Li, Yanshan Xiao, Kai Wang, Jian Hu, Junrui Liu, Qihang Chen, and Ruiguang Huang. An efficient transfer learning method with auxiliary information. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):22:1–22:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3612930.

Li:2021:LTE

[LLY^{+}21] Yanni Li, Bing Liu, Yongbo Yu, Hui Li, Jiacan Sun, and Jiangtao Cui. 3E-LDA: Three enhancements to linear discriminant analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):57:1–57:20, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442347.

Li:2024:SAS

[LLY^{+}24] Mengyao Li, Zhiyong Li, Zhibang Yang, Xu Zhou, Yifan Li, Ziyan Wu, Lingzhao Kong, and Ke Nai. SA2E-AD: a stacked attention autoencoder for anomaly detection in multivariate time series. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):174:1–174:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653677.

Liu:2021:FBN

[LLZ^{+}21] Kai Liu, Xiangyu Li, Zhihui Zhu, Lodewijk Brand, and Hua Wang. Factorbounded nonnegative matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (6):111:1–111:18, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451395.

Li:2023:TWP

[LLZ^{+}23] Lei Li, Zhiyuan Liu, Zan Zhang, Huanhuan Chen, and Xindong Wu. Three-way preference completion via preference graph. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):89:1–89:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580368.

Lorenzetti:2016:MTS

[LML^{+}16] Carlos Lorenzetti, Ana Maguitman, David Leake, Filippo Menczer, and Thomas Reichherzer. Mining for topics to suggest knowledge model extensions. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 23:1–23:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lu:2024:FNS

[LN24] Fengcheng Lu and Michael Kwok-Po Ng. FastHGNN: a new sampling technique for learning with hypergraph neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):184:1–184:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663670.

LaFond:2018:DSC

[LNG18] Timothy La Fond, Jennifer Neville, and Brian Gallagher. Designing size consistent statistics for accurate anomaly detection in dynamic networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):46:1– 46:??, July 2018. CODEN ????

ISSN 1556-4681 (print), 1556-472X (electronic).

Lakshmanan:2008:DRA

[LNR08] Laks V. S. Lakshmanan, Raymond T. Ng, and Ganesh Ramesh. On disclosure risk analysis of anonymized itemsets in the presence of prior knowledge. ACM Transactions on Knowledge Discovery from Data (TKDD), 2 (3):13:1–13:??, October 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liang:2022:NFB

[LOM22] Shangsong Liang, Zhuo Ouyang, and Zaiqiao Meng. A normalizing flow-based co-embedding model for attributed networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):55:1– 55:31, June 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3477049.

Lin:2021:PIR

[LOW^{+}21] Chen Lin, Zhichao Ouyang, Xiaoli Wang, Hui Li, and Zhenhua Huang. Preserve integrity in realtime event summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):49:1–49:29, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442344.

Li:2015:RUC

[LPK^{+}15] Lei Li, Wei Peng, Saurabh Kataria, Tong Sun, and Tao Li. Recommending users and communities in social media. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):17:1– 17:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2015:RAT

[LQW15] Siyuan Liu, Qiang Qu, and Shuhui Wang. Rationality analytics from trajectories. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):10:1–10:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2018:CQE

[LQW^{+}18] Feijiang Li, Yuhua Qian, Jieting Wang, Chuangyin Dang, and Bing Liu. Cluster’s quality evaluation and selective clustering ensemble. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5): 60:1–60:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lee:2019:AMG

[LRK^{+}19] John Boaz Lee, Ryan A.

Rossi, Sungchul Kim, Nesreen K. Ahmed, and Eunyee Koh. Attention models in graphs: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):62:1–62:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3363574.

Li:2018:MVL

[LSF18] Sheng Li, Ming Shao, and Yun Fu. Multi-view low-rank analysis with applications to outlier detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3): 32:1–32:??, April 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2019:DDS

[LSGZ19] Xiaoming Liu, Chao Shen, Xiaohong Guan, and Yadong Zhou. Digger: Detect similar groups in heterogeneous social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):2:1–2:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3267106.

Liu:2020:RTT

[LSH20] Shuai Liu, Guojie Song, and Wenhao Huang. Real-time transportation prediction correction using reconstruction error in deep learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):17:1–17:20, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3369871.

Li:2022:CDL

[LSH^{+}22] Zhao Li, Junshuai Song, Zehong Hu, Zhen Wang, and Jun Gao. Constrained duallevel bandit for personalized impression regulation in online ranking systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):23:1–23:23, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3461340.

Li:2022:DME

[LSL^{+}22] Zhe Li, Chunhua Sun, Chunli Liu, Xiayu Chen, Meng Wang, and Yezheng Liu. Dual-MGAN: an efficient approach for semi-supervised outlier detection with few identified anomalies. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):107:1–107:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522690.

Lin:2011:CDM

[LSS^{+}11] Yu-Ru Lin, Jimeng Sun, Hari Sundaram, Aisling Kelliher, Paul Castro, and Ravi Konuru. Community discovery via metagraph factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(3):17:1–17:??, August 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lee:2022:MMS

[LSS^{+}22] Wu Lee, Yuliang Shi, Hongfeng Sun, Lin Cheng, Kun Zhang, Xinjun Wang, and Zhiyong Chen. MSIPA: Multi-scale interval pattern-aware network for ICU transfer prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):17:1– 17:17, February 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3458284.

Liang:2024:DCG

[LSW^{+}24] Tianhai Liang, Qiangqiang Shen, Shuqin Wang, Yongyong Chen, Guokai Zhang, and Junxin Chen. Data completion-guided unified graph learning for incomplete multiview clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):188:1–188:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664290.

Liu:2009:ISI

[LSY^{+}09] Huan Liu, John Salerno, Michael Young, Rakesh Agrawal, and Philip S. Yu. Introduction to special issue on social computing, behavioral modeling, and prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(2): 6:1–6:??, April 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2019:CLL

[LSZ^{+}19] Shenghua Liu, Huawei Shen, Houdong Zheng, Xueqi Cheng, and Xiangwen Liao. CT LIS: Learning influences and susceptibilities through temporal behaviors. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):57:1–57:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3363570.

Li:2024:QNN

[LSZ^{+}24] Jiaye Li, Jinjing Shi, Jian Zhang, Yuhu Lu, Qin Li, Chunlin Yu, and Shichao Zhang. Quantum nearest neighbor collaborative filtering algorithm for recommendation system. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):201:1–201:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3674982.

Liu:2010:FCP

[LT10] Kun Liu and Evimaria Terzi. A framework for computing the privacy scores of users in online social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(1):6:1–6:??, December 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lines:2018:TSC

[LTB18] Jason Lines, Sarah Taylor, and Anthony Bagnall. Time series classification with HIVE-COTE: The hierarchical vote collective of transformationbased ensembles. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5): 52:1–52:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lou:2013:LPR

[LTH^{+}13] Tiancheng Lou, Jie Tang, John Hopcroft, Zhanpeng Fang, and Xiaowen Ding. Learning to predict reciprocity and triadic closure in social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(2):5:1–5:??, July 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lu:2008:ADA

[LTN^{+}08] Yijuan Lu, Qi Tian, Jennifer Neary, Feng Liu, and Yufeng Wang. Adaptive discriminant analysis for microarray-based classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(1):5:1– 5:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2021:SEN

[LTN^{+}21] Chenglin Li, Carrie Lu Tong, Di Niu, Bei Jiang, Xiao Zuo, Lei Cheng, Jian Xiong, and Jianming Yang. Similarity embedding networks for robust human activity recognition. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):98:1–98:17, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3448021.

Liu:2012:IBA

[LTZ12] Fei Tony Liu, Kai Ming Ting, and Zhi-Hua Zhou. Isolation-based anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(1):3:1–3:??, March 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2018:ERC

[LV18] Bo Li and Yevgeniy Vorobeychik. Evasion-robust classification on binary domains. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):50:1–50:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2021:HNA

[LVH^{+}21] Shikang Liu, Fatemeh Vahedian, David Hachen, Omar Lizardo, Christian Poellabauer, Aaron Striegel, and Tijana Milenković. Heterogeneous network approach to predict individuals’ mental health. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (2):25:1–25:26, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429446.

Li:2024:MOR

[LW24] Xiangyu Li and Hua Wang. On mean-optimal robust linear discriminant analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):191:1–191:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665500.

Li:2022:MNV

[LWB^{+}22] Hui-Jia Li, Lin Wang, Zhan Bu, Jie Cao, and Yong Shi. Measuring the network vulnerability based on Markov criticality. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (2):28:1–28:24, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3464390.

Li:2024:VWM

[LWB^{+}24] Feifei Li, Yuanbin Wang, Oya Beyan, Mirjam Schöneck, and Liliana Lourenco Caldeira. Voxel-wise medical image generalization for eliminating distribution shift. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):167:1–167:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643034.

Li:2023:HCT

[LWC^{+}23] Xiaona Li, Zhu Wang, Xindong Chen, Bin Guo, and Zhiwen Yu. A hybrid continuoustime dynamic graph representation learning model by exploring both temporal and repetitive information. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):123:1–123:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3596447.

Lai:2024:IAM

[LWCZ24] Jinrong Lai, Tong Wang, Chuan Chen, and Zibin Zheng. Information-aware multi-view outlier detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):98:1–98:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638354.

Liu:2016:SAU

[LWG^{+}16] Bin Liu, Yao Wu, Neil Zhenqiang Gong, Junjie Wu, Hui Xiong, and Martin Ester. Structural analysis of user choices for mobile app recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 17:1–17:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2022:DLV

[LWG^{+}22] Fenglin Liu, Xian Wu, Shen Ge, Xuancheng Ren, Wei Fan, Xu Sun, and Yuexian Zou. DiMBERT: Learning vision-language grounded representations with disentangled multimodal-attention. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):1:1–1:19, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447685.

Luo:2024:CBH

[LWG^{+}24] Xiao Luo, Daqing Wu, Yiyang Gu, Chong Chen, Luchen Liu, Jinwen Ma, Ming Zhang, Minghua Deng, Jianqiang Huang, and Xian-Sheng Hua. Criterion-based heterogeneous collaborative filtering for multi-behavior implicit recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):14:1–14:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611310.

Liu:2023:TTP

[LWH^{+}23] Ye Liu, Han Wu, Zhenya Huang, Hao Wang, Yuting Ning, Jianhui Ma, Qi Liu, and Enhong Chen. TechPat: Technical phrase extraction for patent mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):129:1–129:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3596603.

Lu:2024:CTT

[LWH^{+}24] Xiangkui Lu, Jun Wu, Junheng Huang, Fangyuan Luo, and Jianbo Yuan. Cotraining-teaching: a robust semi-supervised framework for review-aware rating regression. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):41:1– 41:??, February 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3625391.

Li:2018:ESC

[LWLW18] Peipei Li, Haixun Wang, Hongsong Li, and Xindong Wu. Employing semantic context for sparse information extraction assessment. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5):54:1–54:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Laishram:2022:MEA

[LWS22] Ricky Laishram, Jeremy D. Wendt, and Sucheta Soundarajan. MCS+: an efficient algorithm for crawling the community structure in multiplex networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):11:1–11:32, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451527.

Long:2018:PMS

[LWW18] Cheng Long, Raymond Chi-Wing Wong, and Victor Junqiu Wei. Profit maximization with sufficient customer satisfactions. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2): 19:1–19:??, March 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ling:2021:DGM

[LWW^{+}21] Xiang Ling, Lingfei Wu, Saizhuo Wang, Gaoning Pan, Tengfei Ma, Fangli Xu, Alex X. Liu, Chunming Wu, and Shouling Ji. Deep graph matching and searching for semantic code retrieval. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):88:1–88:21, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447571.

Lin:2023:ESM

[LWW23a] Dandan Lin, Victor Junqiu Wei, and Raymond Chi-Wing Wong. Effective and scalable manifold ranking-based image retrieval with output bound. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):61:1–61:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565574.

Luo:2023:SDH

[LWW^{+}23b] Xiao Luo, Haixin Wang, Daqing Wu, Chong Chen, Minghua Deng, Jianqiang Huang, and Xian-Sheng Hua. A survey on deep hashing methods. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (1):15:1–15:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532624.

Luo:2024:DLC

[LWW24] Fangyuan Luo, Jun Wu, and Tao Wang. Discrete listwise content-aware recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):7:1–7:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3609334.

Li:2024:GBT

[LWWW24] Ximing Li, Bing Wang, Yang Wang, and Meng Wang. Graph-based text classification by contrastive learning with text-level graph augmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):77:1–77:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638353.

Li:2023:CBS

[LWWX23] Qian Li, Xiangmeng Wang, Zhichao Wang, and Guandong Xu. Be causal: De-biasing social network confounding in recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (1):14:1–14:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3533725.

Liu:2016:EAW

[LWY16] Yashu Liu, Jie Wang, and Jieping Ye. An efficient algorithm for weak hierarchical lasso. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (3):32:1–32:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Lin:2014:SCR

[LWZ14] Ming Lin, Shifeng Weng, and Changshui Zhang. On the sample complexity of random Fourier features for online learning: How many random Fourier features do we need? ACM Transactions on Knowledge Discovery from Data (TKDD), 8(3): 13:1–13:??, June 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2022:KDA

[LXC^{+}22] Xingjian Li, Haoyi Xiong, Zeyu Chen, Jun Huan, Ji Liu, Cheng-Zhong Xu, and Dejing Dou. Knowledge distillation with attention for deep transfer learning of convolutional networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):42:1–42:20, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473912.

Liu:2023:MMI

[LXR^{+}23] Jiaying Liu, Feng Xia, Jing Ren, Bo Xu, Guansong Pang, and Lianhua Chi. MIRROR: Mining implicit relationships via structure-enhanced graph convolutional networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):55:1–55:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564531.

Liu:2017:IPV

[LXY^{+}17] Qi Liu, Biao Xiang, Nicholas Jing Yuan, Enhong Chen, Hui Xiong, Yi Zheng, and Yu Yang. An influence propagation view of PageRank. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):30:1–30:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ling:2024:FFS

[LXZ^{+}24] Zhaolong Ling, Enqi Xu, Peng Zhou, Liang Du, Kui Yu, and Xindong Wu. Fair feature selection: a causal perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):163:1–163:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643890.

Liu:2021:SRS

[LYC^{+}21] Xueyan Liu, Bo Yang, Hechang Chen, Katarzyna Musial, Hongxu Chen, Yang Li, and Wanli Zuo. A scalable redefined stochastic blockmodel. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):46:1–46:28, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442589.

Ling:2023:STD

[LYC^{+}23] Shuai Ling, Zhe Yu, Shaosheng Cao, Haipeng Zhang, and Simon Hu. STHAN: Transportation demand forecasting with compound spatiotemporal relationships. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):54:1–54:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3565578.

Long:2024:LGM

[LYF^{+}24] Chao Long, Huanhuan Yuan, Junhua Fang, Xuefeng Xian, Guanfeng Liu, Victor S. Sheng, and Pengpeng Zhao. Learning global and multigranularity local representation with MLP for sequential recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):87:1–87:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638562.

Li:2019:HDE

[LYG^{+}19] Qingyang Li, Zhiwen Yu, Bin Guo, Huang Xu, and Xinjiang Lu. Housing demand estimation based on express delivery data. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):43:1–43:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3332522.

Liang:2024:DUP

[LYG^{+}24] Yuliang Liang, Enneng Yang, Guibing Guo, Wei Cai, Linying Jiang, and Xingwei Wang. Deconfounding user preference in recommendation systems through implicit and explicit feedback. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):198:1–198:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673762.

Liu:2022:FMM

[LYGG22] Chang Liu, Jie Yan, Feiyue Guo, and Min Guo. Forecasting the market with machine learning algorithms: an application of NMC-BERT-LSTM-DQN-X algorithm in quantitative trading. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):62:1–62:22, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3488378.

Lu:2020:ILS

[LYL^{+}20] Xinjiang Lu, Zhiwen Yu, Chuanren Liu, Yanchi Liu, Hui Xiong, and Bin Guo. Inferring lifetime status of pointof-interest: a multitask multiclass approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):10:1–10:27, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3369799.

Ling:2022:PAP

[LYL^{+}22] Zhaolong Ling, Kui Yu, Lin Liu, Jiuyong Li, Yiwen Zhang, and Xindong Wu. PSL: an algorithm for partial Bayesian network structure learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):93:1–93:25, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3508071.

Li:2024:SGC

[LYN^{+}24] Huiyuan Li, Li Yu, Xi Niu, Youfang Leng, and Qihan Du. Sequential and graphical cross-domain recommendations with a multi-view hierarchical transfer gate. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):8:1–8:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3604615.

Li:2023:SSG

[LYT^{+}23] Lei Li, Mengjiao Yan, Zhenchao Tao, Huanhuan Chen, and Xindong Wu. Semisupervised graph pattern matching and rematching for expert community location. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):6:1–6:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532623.

Li:2022:PDB

[LYTZ22] Ziyue Li, Hao Yan, Fugee Tsung, and Ke Zhang. Profile decomposition based hybrid transfer learning for cold-start data anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):121:1–121:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3530990.

Li:2012:EMG

[LYWL12] Chun Li, Qingyan Yang, Jianyong Wang, and Ming Li. Efficient mining of gapconstrained subsequences and its various applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(1):2:1–2:??, March 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2024:CDA

[LYZ^{+}24] Jinpeng Li, Hang Yu, Zhenyu Zhang, Xiangfeng Luo, and Shaorong Xie. Concept drift adaptation by exploiting drift type. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):96:1–96:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638777.

Lu:2016:SCB

[LZD16] Faming Lu, Qingtian Zeng, and Hua Duan. Synchronizationcore-based discovery of processes with decomposable cyclic dependencies. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(3):31:1–31:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Liu:2015:SEH

[LZF^{+}15] Guimei Liu, Haojun Zhang, Mengling Feng, Limsoon Wong, and See-Kiong Ng. Supporting exploratory hypothesis testing and analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4):31:1– 31:??, June 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Li:2024:HHG

[LZG^{+}24] Youru Li, Zhenfeng Zhu, Xiaobo Guo, Shaoshuai Li, Yuchen Yang, and Yao Zhao. HGV4Risk: Hierarchical global view-guided sequence representation learning for risk prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):1:1–1:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3605895.

Lin:2022:LVA

[LZL^{+}22] Luyue Lin, Xin Zheng, Bo Liu, Wei Chen, and Yanshan Xiao. A latent variable augmentation method for image categorization with insufficient training samples. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (1):3:1–3:35, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451165.

Li:2023:HTN

[LZL^{+}23] Mengran Li, Yong Zhang, Xiaoyong Li, Yuchen Zhang, and Baocai Yin. Hypergraph transformer neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):63:1– 63:??, June 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3565028.

Liu:2024:DDP

[LZL^{+}24] Huiting Liu, Yu Zhang, Peipei Li, Cheng Qian, Peng Zhao, and Xindong Wu. DeepCPR: Deep path reasoning using sequence of user-preferred attributes for conversational recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):15:1–15:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3610775.

Liu:2024:UKG

[LZLJ24] Yu Liu, Zhilun Zhou, Yong Li, and Depeng Jin. Urban knowledge graph aided mobile user profiling. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):28:1–28:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3596604.

Li:2020:BDR

[LZP20] Shuangyin Li, Yu Zhang, and Rong Pan. Bi-directional recurrent attentional topic model. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(6):74:1– 74:30, October 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3412371.

Li:2024:MMI

[LZSY24] Hongyu Li, Lefei Zhang, Kehua Su, and Wei Yu. MICCF: a mutual information constrained clustering framework for learning clustering-oriented feature representations. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):205:1–205:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672402.

Li:2024:SEA

[LZV24] Zhong Li, Yuxuan Zhu, and Matthijs Van Leeuwen. A survey on explainable anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):23:1–23:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3609333.

Liu:2022:NMV

[LZX22] Bo Liu, Haowen Zhong, and Yanshan Xiao. New multiview classification method with uncertain data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):19:1–19:23, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458282.

Li:2024:ISD

[LZY^{+}24a] Xuefei Li, Huiwei Zhou, Weihong Yao, Wenchu Li, Baojie Liu, and Yingyu Lin. Intricate spatiotemporal dependency learning for temporal knowledge graph reasoning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):136:1–136:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648366.

Liu:2024:DSA

[LZY^{+}24b] Shenghao Liu, Yu Zhang, Lingzhi Yi, Xianjun Deng, Laurence T. Yang, and Bang Wang. Dual-side adversarial learning based fair recommendation for sensitive attribute filtering. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):165:1–165:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648683.

Lei:2022:ODR

[LZZ^{+}22a] Shuo Lei, Xuchao Zhang, Liang Zhao, Arnold P. Boedihardjo, and Chang-Tien Lu. Online and distributed robust regressions with extremely noisy labels. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):41:1–41:24, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473038.

Liu:2022:JMH

[LZZ^{+}22b] Haobing Liu, Yanmin Zhu, Tianzi Zang, Yanan Xu, Jiadi Yu, and Feilong Tang. Jointly modeling heterogeneous student behaviors and interactions among multiple prediction tasks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):16:1–16:24, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458023.

McDowell:2016:LNA

[MA16] Luke K. McDowell and David W.

Aha. Leveraging neighbor attributes for classification in sparsely labeled networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):2:1–2:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Mohammadi:2018:COA

[MAHT18] Majid Mohammadi, Amir Ahooye Atashin, Wout Hofman, and Yaohua Tan. Comparison of ontology alignment systems across single matching task via the McNemar’s test. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):51:1–51:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Makki:2018:AVV

[MCS^{+}18] Raheleh Makki, Eder Carvalho, Axel J. Soto, Stephen Brooks, Maria Cristina Ferreira De Oliveira, Evangelos Milios, and Rosane Minghim. ATR-Vis: Visual and interactive information retrieval for parliamentary discussions in Twitter. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):3:1–3:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Marques:2020:IEU

[MCSZ20] Henrique O. Marques, Ricardo J. G. B. Campello, Jürg Sander, and Arthur Zimek. Internal evaluation of unsupervised outlier detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):47:1–47:42, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3394053.

Magdalinos:2011:ECQ

[MDV11] Panagis Magdalinos, Christos Doulkeridis, and Michalis Vazirgiannis. Enhancing clustering quality through landmark-based dimensionality reduction. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (2):11:1–11:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Menon:2011:FAA

[ME11] Aditya Krishna Menon and Charles Elkan. Fast algorithms for approximating the singular value decomposition. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (2):13:1–13:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Moreo:2022:LTT

[MES22] Alejandro Moreo, Andrea Esuli, and Fabrizio Sebastiani. Lost in transduction: Transductive transfer learning in text classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):13:1–13:21, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3453146.

Mohammed:2010:CDA

[MFHL10] Noman Mohammed, Benjamin C. M. Fung, Patrick C. K. Hung, and Cheuk-Kwong Lee. Centralized and distributed anonymization for high-dimensional healthcare data. ACM Transactions on Knowledge Discovery from Data (TKDD), 4 (4):18:1–18:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Myers:2016:DAK

[MFRJ16] Risa B. Myers, John C. Frenzel, Joseph R. Ruiz, and Christopher M. Jermaine. Do anesthesiologists know what they are doing? Mining a surgical time-series database to correlate expert assessment with outcomes. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (3):24:1–24:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Moreo:2024:MLQ

[MFS24] Alejandro Moreo, Manuel Francisco, and Fabrizio Sebastiani. Multi-label quantification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):4:1–4:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3606264.

Mannila:2009:ATS

[MG09] Heikki Mannila and Dimitrios Gunopulos. ACM TKDD special issue ACM SIGKDD 2007 and ACM SIGKDD 2008. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4): 15:1–15:??, November 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Moradnia:2024:SCP

[MG24] Sajedeh Moradnia and Mousa Golalizadeh. Supervised clustering of Persian handwritten images using regularization and dimension reduction methods. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):118:1–118:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638060.

Ma:2020:RRT

[MGL^{+}20] Qian Ma, Yu Gu, Wang-Chien Lee, Ge Yu, Hongbo Liu, and Xindong Wu. REMIAN: Real-time and error-tolerant missing value imputation. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(6):77:1–77:38, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3412364.

Mahmoud:2022:MOL

[MH22] Reem A. Mahmoud and Hazem Hajj. Multi-objective learning to overcome catastrophic forgetting in timeseries applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):103:1–103:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502728.

Moghaz:2019:TME

[MHKG19] Dror Moghaz, Yaakov Hacohen-Kerner, and Dov Gabbay. Text mining for evaluating authors’ birth and death years. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):7:1–7:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3281631.

Mitra:2020:UMV

[MHS20] Sayantan Mitra, Mohammed Hasanuzzaman, and Sriparna Saha. A unified multiview clustering algorithm using multi-objective optimization coupled with generative model. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):2:1–2:31, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3365673.

Machanavajjhala:2007:DPB

[MKGV07] Ashwin Machanavajjhala, Daniel Kifer, Johannes Gehrke, and Muthuramakrishnan Venkitasubramaniam. L-diversity: Privacy beyond k-anonymity. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(1):3:1–3:??, March 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Mcauley:2014:DSC

[ML14] Julian Mcauley and Jure Leskovec. Discovering social circles in ego networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(1):4:1–4:??, February 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Mirbakhsh:2015:ITR

[ML15] Nima Mirbakhsh and Charles X.

Ling. Improving top-N recommendation for cold-start users via cross-domain information. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4): 33:1–33:??, June 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Maurya:2021:GNN

[MLM21] Sunil Kumar Maurya, Xin Liu, and Tsuyoshi Murata. Graph neural networks for fast node ranking approximation. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):78:1–78:32, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446217.

Mavroeidis:2012:SSF

[MM12] Dimitrios Mavroeidis and Panagis Magdalinos. A sequential sampling framework for spectral k-means based on efficient bootstrap accuracy estimations: Application to distributed clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(2):5:1–5:??, July 2012. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Mohotti:2020:EOD

[MN20] Wathsala Anupama Mohotti and Richi Nayak. Efficient outlier detection in text corpus using rare frequency and ranking. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):71:1–71:30, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3399712.

Moreno:2018:TKP

[MNK18] Sebastian Moreno, Jennifer Neville, and Sergey Kirshner. Tied Kronecker product graph models to capture variance in network populations. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):35:1–35:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Miao:2023:DPC

[MPG^{+}23] Xiaoye Miao, Huanhuan Peng, Yunjun Gao, Zongfu Zhang, and Jianwei Yin. On dynamically pricing crowdsourcing tasks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):30:1–30:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3544018.

Ma:2022:MIF

[MRC^{+}22] Muyang Ma, Pengjie Ren, Zhumin Chen, Zhaochun Ren, Lifan Zhao, Peiyu Liu, Jun Ma, and Maarten de Rijke. Mixed information flow for cross-domain sequential recommendations. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):64:1–64:32, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487331.

Maier:2011:INS

[MRJ11] Marc Maier, Matthew Rattigan, and David Jensen. Indexing network structure with shortest-path trees. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(3):15:1–15:??, August 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Murai:2019:CDU

[MRTW19] Fabricio Murai, Bruno Ribeiro, Don Towlsey, and Pinghui Wang. Characterizing directed and undirected networks via multidimensional walks with jumps. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):11:1– 11:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Mehler:2009:ENC

[MS09] Andrew Mehler and Steven Skiena. Expanding network communities from representative examples. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(2):7:1– 7:??, April 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ma:2019:PBD

[MSC^{+}19] Liang Ma, Mudhakar Srivatsa, Derya Cansever, Xifeng Yan, Sue Kase, and Michelle Vanni. Performance bounds of decentralized search in expert networks for query answering. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (2):18:1–18:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3300230.

Miller:2024:ASP

[MSR^{+}24] Benjamin A. Miller, Zohair Shafi, Wheeler Ruml, Yevgeniy Vorobeychik, Tina Eliassi-Rad, and Scott Alfeld. Attacking shortest paths by cutting edges. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):35:1–35:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3622941.

Miettinen:2014:MMD

[MV14] Pauli Miettinen and Jilles Vreeken. MDL4BMF: Minimum description length for Boolean matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(4):18:1–18:??, October 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Mampaey:2012:SDS

[MVT12] Michael Mampaey, Jilles Vreeken, and Nikolaj Tatti. Summarizing data succinctly with the most informative itemsets. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(4): 16:1–16:??, December 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ma:2024:SPS

[MW24] Fei Ma and Ping Wang. Structural properties on scale-free tree network with an ultralarge diameter. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):200:1–200:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3674146.

Mangasarian:2008:PPC

[MWF08] Olvi L. Mangasarian, Edward W. Wild, and Glenn M. Fung. Privacy-preserving classification of vertically partitioned data via random kernels. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(3):12:1– 12:??, October 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ma:2022:HFD

[MWL^{+}22] Pengfei Ma, Youxi Wu, Yan Li, Lei Guo, He Jiang, Xingquan Zhu, and Xindong Wu. HW-Forest: Deep forest with hashing screening and window screening. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):123:1–123:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532193.

Mei:2007:SAF

[MXC^{+}07] Qiaozhu Mei, Dong Xin, Hong Cheng, Jiawei Han, and Chengxiang Zhai. Semantic annotation of frequent patterns. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(3): 11:1–11:??, December 2007.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Matheny:2020:SSS

[MXP20] Michael Matheny, Dong Xie, and Jeff M. Phillips. Scalable spatial scan statistics for trajectories. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):73:1–73:24, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3394046.

Mahmoudi:2019:RBO

[MYB19] Amin Mahmoudi, Mohd Ridzwan Yaakub, and Azuraliza Abu Bakar. The relationship between online social network ties and user attributes. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):26:1–26:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314204.

Mai:2024:SCC

[MYC^{+}24] Weiming Mai, Jiangchao Yao, Gong Chen, Ya Zhang, Yiu-Ming Cheung, and Bo Han. Server-client collaborative distillation for federated reinforcement learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):9:1–9:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3604939.

Munoz:2021:ISA

[MYL^{+}21] Mario Andrés Muñoz, Tao Yan, Matheus R. Leal, Kate Smith-Miles, Ana Carolina Lorena, Gisele L. Pappa, and Rômulo Madureira Rodrigues. An instance space analysis of regression problems. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):28:1–28:25, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3436893.

Mautz:2020:NRS

[MYPB20] Dominik Mautz, Wei Ye, Claudia Plant, and Christian Böhm. Non-redundant subspace clusterings with Nr-Kmeans and Nr-DipMeans. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(5):55:1–55:24, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3385652.

Meng:2024:SIB

[MZC^{+}24] Siyuan Meng, Jie Zhou, Xuxin Chen, Yufei Liu, Fengyuan Lu, and Xinli Huang. Structureinformation-based reasoning over the knowledge graph: a survey of methods and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):210:1–210:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671148.

Nasir:2021:TAM

[NADR21] Muhammad Anis Uddin Nasir, Cigdem Aslay, Gianmarco De Francisci Morales, and Matteo Riondato. TipTap: Approximate mining of frequent k-subgraph patterns in evolving graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):48:1–48:35, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442590.

Na:2022:USE

[NC22] Gyoung S. Na and Hyunju Chang. Unsupervised subspace extraction via deep kernelized clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):20:1–20:15, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3459082.

Nesa:2018:IIG

[NGB18] Nashreen Nesa, Tania Ghosh, and Indrajit Banerjee. iGRM: Improved grey relational model and its ensembles for occupancy sensing in Internet of things applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):47:1–47:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Nashaat:2022:SSE

[NGMQ22] Mona Nashaat, Aindrila Ghosh, James Miller, and Shaikh Quader. Semi-supervised ensemble learning for dealing with inaccurate and incomplete supervision. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):43:1–43:33, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473910.

Nikolakopoulos:2020:BIB

[NK20] Athanasios N. Nikolakopoulos and George Karypis. Boosting item-based collaborative filtering via nearly uncoupled random walks. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):64:1–64:26, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3406241.

Nettasinghe:2021:MLE

[NK21] Buddhika Nettasinghe and Vikram Krishnamurthy. Maximum likelihood estimation of power-law degree distributions via friendship paradoxbased sampling. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (6):106:1–106:28, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451166.

Nguyen:2023:DAG

[NLA23] Hung T. Nguyen, Pierre J.

Liang, and Leman Akoglu. Detecting anomalous graphs in labeled multi-graph databases. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):16:1–16:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3533770.

Namata:2016:CGI

[NLG16] Galileo Mark Namata, Ben London, and Lise Getoor. Collective graph identification. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(3):25:1– 25:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ni:2020:LOC

[NLZH20] Li Ni, Wenjian Luo, Wenjie Zhu, and Bei Hua. Local overlapping community detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):3:1–3:25, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3361739.

Nakajima:2023:RWS

[NS23] Kazuki Nakajima and Kazuyuki Shudo. Random walk sampling in social networks involving private nodes. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):51:1–51:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3561388.

Nguyen:2018:EUP

[NTNP18] Minh-Tien Nguyen, Duc-Vu Tran, Le-Minh Nguyen, and Xuan-Hieu Phan. Exploiting user posts for Web document summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4): 49:1–49:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Nguyen:2020:EEC

[NWA20] Hung Nguyen, Xuejian Wang, and Leman Akoglu. End-to-end continual rare-class recognition with emerging novel subclasses. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):61:1–61:28, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3399660.

Nie:2020:ALL

[NWW^{+}20] Feiping Nie, Zheng Wang, Rong Wang, Zhen Wang, and Xuelong Li. Adaptive local linear discriminant analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):9:1–9:19, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3369870.

Ni:2024:LOS

[NXZ^{+}24] Li Ni, Hefei Xu, Yiwen Zhang, Wenjian Luo, Yingying Huang, and Victor S. Sheng. Local overlapping spatial-aware community detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):59:1– 59:??, April 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3634707.

Ni:2024:LCD

[NYLZ24] Li Ni, Rui Ye, Wenjian Luo, and Yiwen Zhang. Local community detection in multiple private networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):126:1–126:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3644078.

Ni:2023:MBP

[NZW23] Peikun Ni, Jianming Zhu, and Guoqing Wang. Misinformation blocking problem in virtual and real interconversion social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):87:1–87:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3578936.

Ou:2024:SBO

[OCD^{+}24] Weitong Ou, Bo Chen, Xinyi Dai, Weinan Zhang, Weiwen Liu, Ruiming Tang, and Yong Yu. A survey on bid optimization in real-time bidding display advertising. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):58:1–58:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3628603.

Ordonez:2014:BVS

[OGAB14] Carlos Ordonez, Carlos Garcia-Alvarado, and Veerabhadaran Baladandayuthapani. Bayesian variable selection in linear regression in one pass for large datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1):3:1–3:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ouyang:2021:MAC

[OGT^{+}21] Yi Ouyang, Bin Guo, Xing Tang, Xiuqiang He, Jian Xiong, and Zhiwen Yu. Mobile app cross-domain recommendation with multi-graph neural network. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):55:1–55:21, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442201.

Oliveira:2022:AMT

[OGV22] Saullo H. G. Oliveira, André R.

Gonçalves, and Fernando J. Von Zuben. Asymmetric multi-task learning with local transference. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):99:1–99:30, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3514252.

Ohare:2022:HVT

[OJLD22] Kevin O’hare, Anna Jurek-Loughrey, and Cassio De Campos. High-value tokenblocking: Efficient blocking method for record linkage. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):24:1– 24:17, April 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450527.

Ou:2023:STS

[OJW^{+}23] Junjie Ou, Haiming Jin, Xiaocheng Wang, Hao Jiang, Xinbing Wang, and Chenghu Zhou. STA-TCN: Spatialtemporal attention over temporal convolutional network for next point-of-interest recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):124:1–124:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3596497.

Ostovar:2020:RDC

[OLL20] Alireza Ostovar, Sander J. J. Leemans, and Marcello La Rosa. Robust drift characterization from event streams of business processes. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):30:1–30:57, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375398.

Perozzi:2018:DCA

[PA18] Bryan Perozzi and Leman Akoglu. Discovering communities and anomalies in attributed graphs: Interactive visual exploration and summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2): 24:1–24:??, March 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Papagelis:2015:RSG

[Pap15]  Manos Papagelis.  Refin-

ing social graph connectivity via shortcut edge addition. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):12:1– 12:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Pandey:2024:XDR

[PAS24] Pradumn Kumar Pandey, Aikta Arya, and Akrati Saxena. X-distribution: Retraceable power-law exponent of complex networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):117:1–117:??, June 2024.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639413.

Paul:2014:RPL

[PBMID14] Saurabh Paul, Christos Boutsidis, Malik Magdon-Ismail, and Petros Drineas. Random projections for linear support vector machines. ACM Transactions on Knowledge Discovery from Data (TKDD), 8 (4):22:1–22:??, October 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Pang:2020:HUO

[PC20] Guansong Pang and Longbing Cao. Heterogeneous univariate outlier ensembles in multidimensional data. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):68:1–68:27, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3403934.

Peng:2024:MSS

[PCPN24] Bo Peng, Ziqi Chen, Srinivasan Parthasarathy, and Xia Ning. Modeling sequences as star graphs to address over-smoothing in selfattentive sequential recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):207:1–207:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3676560.

Pellegrina:2022:MMC

[PCVR22] Leonardo Pellegrina, Cyrus Cousins, Fabio Vandin, and Matteo Riondato. MCRapper: Monte-Carlo Rademacher averages for poset families and approximate pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):124:1–124:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532187.

Poulakis:2024:SAM

[PDK24] Yannis Poulakis, Christos Doulkeridis, and Dimosthenis Kyriazis. A survey on AutoML methods and systems for clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):120:1–120:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643564.

Paudel:2020:ACD

[PE20] Ramesh Paudel and William Eberle. An approach for concept drift detection in a graph stream using discriminative subgraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):70:1–70:25, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3406243.

Paterakis:2024:DWR

[PFC^{+}24] George Paterakis, Stefanos Fafalios, Paulos Charonyktakis, Vassilis Christophides, and Ioannis Tsamardinos. Do we really need imputation in AutoML predictive modeling? ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):147:1–147:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3643643.

Papalexakis:2015:PSP

[PFS15] Evangelos E. Papalexakis, Christos Faloutsos, and Nicholas D. Sidiropoulos. ParCube: Sparse parallelizable CANDECOMP– PARAFAC tensor decomposition. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):3:1–3:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Pei:2016:CCP

[PFTR16] Yuanli Pei, Xiaoli Z. Fern, Teresa Vania Tjahja, and Rómer Rosales. Comparing clustering with pairwise and relative constraints: a unified framework. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 22:1–22:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Pandove:2018:SRC

[PGR18] Divya Pandove, Shivan Goel, and Rinkl Rani. Systematic review of clustering high-dimensional and large datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):16:1– 16:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Peng:2017:RGR

[PKH^{+}17] Chong Peng, Zhao Kang, Yunhong Hu, Jie Cheng, and Qiang Cheng. Robust graph regularized nonnegative matrix factorization for clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):33:1– 33:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Paul:2021:MOC

[PKSM21] Dipanjyoti Paul, Rahul Kumar, Sriparna Saha, and Jimson Mathew. Multi-objective cuckoo search-based streaming feature selection for multilabel dataset. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):93:1–93:24, July 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447586.

Plangprasopchok:2010:MSA

[PL10] Anon Plangprasopchok and Kristina Lerman. Modeling social annotation: a Bayesian approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(1):4:1–4:??, December 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Plantevit:2010:MMM

[PLL^{+}10] Marc Plantevit, Anne Laurent, Dominique Laurent, Maguelonne Teisseire, and Yeow Wei Choong. Mining multidimensional and multilevel sequential patterns. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(1):4:1–4:??, January 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Peng:2021:SSE

[PLS^{+}21] Hao Peng, Jianxin Li, Yangqiu Song, Renyu Yang, Rajiv Ranjan, Philip S. Yu, and Lifang He. Streaming social event detection and evolution discovery in heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):89:1–89:33, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447585.

Pooja:2022:EHO

[PMC22] Km Pooja, Samrat Mondal, and Joydeep Chandra. Exploiting higher order multidimensional relationships with self-attention for author name disambiguation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):88:1–88:23, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502730.

Pingi:2024:CGA

[PNB24] Sharon Torao Pingi, Richi Nayak, and Md Abul Bashar. Conditional generative adversarial network for early classification of longitudinal datasets using an imputation approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):132:1–132:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3644821.

Prat-Perez:2016:PTT

[PPDSBLP16] Arnau Prat-Pérez, David Dominguez-Sal, Josep-M. Brunat, and Josep-Lluis Larriba-Pey. Put three and three together: Triangle-driven community detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (3):22:1–22:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Peng:2013:EFF

[PSFV13] Jing Peng, Guna Seetharaman, Wei Fan, and Aparna Varde. Exploiting Fisher and Fukunaga–Koontz transforms in Chernoff dimensionality reduction. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(2):8:1–8:??, July 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Park:2018:ETS

[PSP^{+}18] Ha-Myung Park, Francesco Silvestri, Rasmus Pagh, Chin-Wan Chung, Sung-Hyon Myaeng, and U. Kang. Enumerating trillion subgraphs on distributed systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):71:1–71:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Piao:2024:IHL

[PSYJ24] Minghao Piao, Yi Sheng, Jinda Yan, and Cheng Hao Jin. Image hash layer triggered CNN framework for wafer map failure pattern retrieval and classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):75:1–75:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638053.

Prokhorenkova:2022:WLM

[PTL22] Liudmila Prokhorenkova, Alexey Tikhonov, and Nelly Litvak. When less is more: Systematic analysis of cascadebased community detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(4):78:1–78:22, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3494563.

Pellegrina:2024:SEB

[PV24] Leonardo Pellegrina and Fabio Vandin. SILVAN: Estimating betweenness centralities with progressive sampling and non-uniform Rademacher bounds. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (3):52:1–52:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3628601.

Pei:2022:BHB

[PXW^{+}22] Shuyu Pei, Kun Xie, Xin Wang, Gaogang Xie, Kenli Li, Wei Li, Yanbiao Li, and Jigang Wen. B_{h}BF: a Bloom filter using B_{h} sequences for multi-set membership query. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):89:1–89:26, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502735.

Pan:2024:EDD

[PZJ^{+}24] Yicheng Pan, Yifan Zhang, Xinrui Jiang, Meng Ma, and Ping Wang. EffCause: Discover dynamic causal relationships efficiently from timeseries. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):105:1– 105:??, June 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3640818.

Peng:2018:MEO

[PZW^{+}18] Min Peng, Jiahui Zhu, Hua Wang, Xuhui Li, Yanchun Zhang, Xiuzhen Zhang, and Gang Tian. Mining eventoriented topics in microblog stream with unsupervised multi-view hierarchical embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):38:1– 38:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Qiang:2019:HLT

[QCD^{+}19] Jipeng Qiang, Ping Chen, Wei Ding, Tong Wang, Fei Xie, and Xindong Wu. Heterogeneouslength text topic modeling for reader-aware multi-document summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):42:1–42:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3333030.

Qiu:2022:GNN

[QHW22] Zhaopeng Qiu, Yunfan Hu, and Xian Wu. Graph neural news recommendation with user existing and potential interest modeling. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):96:1–96:17, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3511708.

Queiroz-Sousa:2020:ROT

[QSS20] Paulo Orlando Queiroz-Sousa and Ana Carolina Salgado. A review on OLAP technologies applied to information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (1):8:1–8:25, February 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3370912.

Qin:2018:STR

[QSST18] Tian Qin, Wufan Shangguan, Guojie Song, and Jie Tang. Spatio-temporal routine mining on mobile phone data. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5):56:1– 56:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Qin:2024:MWP

[QWH^{+}24] Wei Qin, Xiaowei Wang, Zhenzhen Hu, Lei Wang, Yunshi Lan, and Richang Hong. Math word problem generation via disentangled memory retrieval. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):123:1–123:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3639569.

Qiao:2016:FST

[QXBT16] Maoying Qiao, Richard Yi Da Xu, Wei Bian, and Dacheng Tao. Fast sampling for timevarying determinantal point processes. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):8:1–8:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Qin:2023:TBT

[QZB^{+}23] Meng Qin, Chaorui Zhang, Bo Bai, Gong Zhang, and Dit-Yan Yeung. Towards a better tradeoff between quality and efficiency of community detection: an inductive embedding method across graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):127:1–127:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3596605.

Qin:2023:CBI

[QZL23] Xi Qin, Cheng Zhong, and Hai Xiang Lin. Communitybased influence maximization using network embedding in dynamic heterogeneous social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):119:1–119:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3594544.

Qian:2024:LAL

[QZZZ24] Yu-Yang Qian, Zhen-Yu Zhang, Peng Zhao, and Zhi-Hua Zhou. Learning with asynchronous labels. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):186:1–186:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3662186.

Rayana:2016:LMB

[RA16] Shebuti Rayana and Leman Akoglu. Less is more: Building selective anomaly ensembles. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):42:1– 42:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ramezani:2023:JID

[RABR23] Maryam Ramezani, Aryan Ahadinia, Amirmohammad Ziaei Bideh, and Hamid R. Rabiee. Joint inference of diffusion and structure in partially observed social networks using coupled matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):132:1–132:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3599237.

Rossi:2021:HG

[RAC^{+}21] Ryan A. Rossi, Nesreen K. Ahmed, Aldo Carranza, David Arbour, Anup Rao, Sungchul Kim, and Eunyee Koh. Heterogeneous graphlets. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):9:1–9:43, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418773.

Rayar:2018:VIS

[RBBV18] Frédéric Rayar, Sabine Barrat, Fatma Bouali, and Gilles Venturini. A viewable indexing structure for the interactive exploration of dynamic and large image collections. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):2:1–2:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Rossi:2021:KGE

[RBF^{+}21] Andrea Rossi, Denilson Barbosa, Donatella Firmani, Antonio Matinata, and Paolo Merialdo. Knowledge graph embedding for link prediction: a comparative analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):14:1–14:49, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424672.

Rakthanmanon:2013:ABD

[RCM^{+}13] Thanawin Rakthanmanon, Bilson Campana, Abdullah Mueen, Gustavo Batista, Brandon Westover, Qiang Zhu, Jesin Zakaria, and Eamonn Keogh. Addressing big data time series: Mining trillions of time series subsequences under dynamic time warping. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3): 10:1–10:??, September 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ragab:2023:ABS

[RET^{+}23] Mohamed Ragab, Emadeldeen Eldele, Wee Ling Tan, Chuan-Sheng Foo, Zhenghua Chen, Min Wu, Chee-Keong Kwoh, and Xiaoli Li. ADATIME: a benchmarking suite for domain adaptation on time series data. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):106:1–106:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587937.

Rodriguez-Gonzalez:2024:XFN

[RGAÁC^{+}24] Ansel Y. Rodríguez-González, Ramón Aranda, Miguel Á. Álvarez-Carmona, Angel Díaz-Pacheco, and Rosa María Valdovinos Rosas. X-FSPMiner: a novel algorithm for frequent similar pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):121:1–121:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643820.

Ren:2023:CCG

[RGL^{+}23] Siyuan Ren, Bin Guo, Ke Li, Qianru Wang, Qinfen Wang, and Zhiwen Yu. CoupledGT: Coupled geospatial-temporal data modeling for air quality prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):135:1–135:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604616.

Rossi:2020:PSR

[RJK^{+}20] Ryan A. Rossi, Di Jin, Sungchul Kim, Nesreen K. Ahmed, Danai Koutra, and John Boaz Lee. On proximity and structural role-based embeddings in networks: Misconceptions, techniques, and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):63:1–63:37, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3397191.

Roseberry:2019:MLP

[RKC19] Martha Roseberry, Bartosz Krawczyk, and Alberto Cano. Multi-label punitive kNN with self-adjusting memory for drifting data streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):60:1–60:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3363573.

Ramezani:2018:CDU

[RKR18] Maryam Ramezani, Ali Khodadadi, and Hamid R. Rabiee. Community detection using diffusion information. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):20:1–20:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Rong:2024:LGT

[RLDL24] Can Rong, Zhicheng Liu, Jingtao Ding, and Yong Li. Learning to generate temporal origin-destination flow based-on urban regional features and traffic information. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):150:1–150:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3649141.

Rowe:2016:MUD

[Row16] Matthew Rowe. Mining user development signals for online community churner detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(3):21:1– 21:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ruggieri:2010:DMD

[RPT10] Salvatore Ruggieri, Dino Pedreschi, and Franco Turini. Data mining for discrimination discovery. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(2):9:1–9:??, May 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Rong:2024:CRC

[RQM^{+}24] Huan Rong, Minfeng Qian, Tinghuai Ma, Di Jin, and Victor S. Sheng. CoBjeason: Reasoning covered object in image by multi-agent collaboration based on informed knowledge graph. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):116:1–116:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643565.

Rozenshtein:2017:FDD

[RTG17] Polina Rozenshtein, Nikolaj Tatti, and Aristides Gionis. Finding dynamic dense subgraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):27:1– 27:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Romero-Tris:2018:PPT

[RTM18] Cristina Romero-Tris and David Megías. Protecting privacy in trajectories with a user-centric approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):67:1–67:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Riondato:2014:EDA

[RU14] Matteo Riondato and Eli Upfal. Efficient discovery of association rules and frequent itemsets through sampling with tight performance guarantees. ACM Transactions on Knowledge Discovery from Data (TKDD), 8 (4):20:1–20:??, October 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Riondato:2018:AAB

[RU18] Matteo Riondato and Eli Upfal. ABRA: Approximating betweenness centrality in static and dynamic graphs with Rademacher averages. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5):61:1–61:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Riondato:2020:MMI

[RV20] Matteo Riondato and Fabio Vandin. MiSoSouP: Mining interesting subgroups with sampling and pseudodimension. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(5):56:1– 56:31, August 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3385653.

Ren:2023:SFB

[RWD23] Jinjun Ren, Yuping Wang, and Xiyan Deng. Slackfactor-based fuzzy support vector machine for class imbalance problems. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):88:1–88:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579050.

Rong:2024:TST

[RYM^{+}24] Huan Rong, Xin Yu, Tinghuai Ma, Victor S. Sheng, Yang Zhou, and Al-Rodhaan Mznah. Three-stage transferable and generative crowdsourced comment integration framework based on zero- and fewshot learning with domain distribution alignment. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):68:1–68:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3636511.

Ren:2024:HPG

[RZF^{+}24] Yuyang Ren, Haonan Zhang, Luoyi Fu, Shiyu Liang, Lei Zhou, Xinbing Wang, Xinde Cao, Fei Long, and Chenghu Zhou. Hi-PART: Going beyond graph pooling with hierarchical partition tree for graph-level representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):94:1–94:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3636429.

Ren:2023:AMA

[RZY^{+}23] Yuyang Ren, Haonan Zhang, Peng Yu, Luoyi Fu, Xinde Cao, Xinbing Wang, Guihai Chen, Fei Long, and Chenghu Zhou. Ada-MIP: Adaptive self-supervised graph representation learning via mutual information and proximity optimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):69:1–69:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568165.

Subbian:2016:MIU

[SAS16] Karthik Subbian, Charu Aggarwal, and Jaideep Srivastava. Mining influencers using information flows in social streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (3):26:1–26:??, February 2016.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Savva:2020:LSD

[SATK20] Fotis Savva, Christos Anagnostopoulos, Peter Triantafillou, and Kostas Kolomvatsos. Large-scale data exploration using explanatory regression functions. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):76:1–76:33, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410448.

Steinbuss:2021:BUO

[SB21a] Georg Steinbuss and Klemens Böhm. Benchmarking unsupervised outlier detection with realistic synthetic data. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):65:1–65:20, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441453.

Steinbuss:2021:GAO

[SB21b] Georg Steinbuss and Klemens Böhm. Generating artificial outliers in the absence of genuine ones — a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):30:1–30:37, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447822.

Singh:2024:MTH

[SB24] Kuldeep Singh and Bhaskar Biswas. Mining top-k high on-shelf utility itemsets using novel threshold raising strategies. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):131:1–131:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3645115.

Shabtai:2014:ODM

[SBRE14] Asaf Shabtai, Maya Bercovitch, Lior Rokach, and Yuval Elovici. Optimizing data misuse detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(3): 16:1–16:??, June 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sajadmanesh:2019:CTR

[SBZR19] Sina Sajadmanesh, Sogol Bazargani, Jiawei Zhang, and Hamid R. Rabiee. Continuoustime relationship prediction in dynamic heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):44:1–44:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3333028.

Singh:2022:MLD

[SCCM22] Shikha Singh, Emilie Chouzenoux, Giovanni Chierchia, and Angshul Majumdar. Multi-label deep convolutional transform learning for non-intrusive load monitoring. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):91:1–91:6, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502729.

Schifanella:2014:MTD

[SCS14] Claudio Schifanella, K. Selçuk Candan, and Maria Luisa Sapino. Multiresolution tensor decompositions with mode hierarchies. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(2): 10:1–10:??, June 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sun:2020:FDA

[SCS20] Bintao Sun, T.-H. Hubert Chan, and Mauro Sozio. Fully dynamic approximate k-core decomposition in hypergraphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4): 39:1–39:21, July 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3385416.

Silva:2018:DMA

[SDS18] Fabrício A. Silva, Augusto C. S. A. Domingues, and Thais R. M. Braga Silva. Discovering mobile application usage patterns from a largescale dataset. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5): 59:1–59:??, July 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Siddiqui:2019:SFE

[SFDW19] Md Amran Siddiqui, Alan Fern, Thomas G. Dietterich, and Weng-Keen Wong. Sequential feature explanations for anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):1:1–1:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3230666.

Sakurai:2010:FDG

[SFP10] Yasushi Sakurai, Christos Faloutsos, and Spiros Papadimitriou. Fast discovery of group lag correlations in streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(1):5:1–5:??, December 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Shahaf:2012:CTL

[SG12] Dafna Shahaf and Carlos Guestrin. Connecting two (or less) dots: Discovering structure in news articles. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(4):24:1– 24:??, February 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sheth:2023:CDI

[SGC^{+}23] Paras Sheth, Ruocheng Guo, Lu Cheng, Huan Liu, and Kasim Selçuk Candan. Causal disentanglement for implicit recommendations with network information. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):94:1–94:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582435.

Song:2019:TCA

[SGCH19] Qingquan Song, Hancheng Ge, James Caverlee, and Xia Hu. Tensor completion algorithms in big data analytics. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):6:1–6:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3278607.

Su:2023:NMF

[SGCH23] Sixing Su, Jiewen Guan, Bilian Chen, and Xin Huang. Nonnegative matrix factorization based on node centrality for community detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):84:1–84:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578520.

Sehnan:2023:DSI

[SGM^{+}23] Dhruv Sehnan, Vasu Goel, Sarah Masud, Chhavi Jain, Vikram Goyal, and Tanmoy Chakraborty. DiVA: a scalable, interactive and customizable visual analytics platform for information diffusion on large networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):47:1–47:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3558771.

Spinnato:2024:UTS

[SGM^{+}24] Francesco Spinnato, Riccardo Guidotti, Anna Monreale, Mirco Nanni, Dino Pedreschi, and Fosca Giannotti. Understanding any time series classifier with a subsequencebased explainer. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):36:1–36:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624480.

Soundarajan:2015:ULG

[SH15] Sucheta Soundarajan and John E. Hopcroft. Use of local group information to identify communities in networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3): 21:1–21:??, April 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Shin:2018:FAF

[SHF18] Kijung Shin, Bryan Hooi, and Christos Faloutsos. Fast, accurate, and flexible algorithms for dense subtensor mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):28:1– 28:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sun:2024:LBC

[SHF24] Yan Sun, Yi Han, and Jicong Fan. Laplacian-based cluster-contractive t-SNE for high-dimensional data visualization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):19:1–19:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3612932.

Sun:2020:NEC

[SHH^{+}20] Heli Sun, Fang He, Jianbin Huang, Yizhou Sun, Yang Li, Chenyu Wang, Liang He, Zhongbin Sun, and Xiaolin Jia. Network embedding for community detection in attributed networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):36:1–36:25, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3385415.

Sahoo:2019:LSO

[SHL19] Doyen Sahoo, Steven C. H. Hoi, and Bin Li. Large scale online multiple kernel regression with application to timeseries prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1): 9:1–9:??, January 2019. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Siers:2021:CIC

[SI21] Michael J. Siers and Md Zahidul Islam. Class imbalance and cost-sensitive decision trees: a unified survey based on a core similarity. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):4:1–4:31, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3415156.

Saha:2024:PPN

[SI24] Swapnil Saha and Hafiz Imtiaz. Privacy-preserving nonnegative matrix factorization with outliers. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):64:1–64:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632961.

Sun:2024:IML

[SJHY24] Kai Sun, Huajie Jiang, Yongli Hu, and Baocai Yin. Incorporating multi-level sampling with adaptive aggregation for inductive knowledge graph completion. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):128:1–128:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3644822.

Somaiya:2008:LCU

[SJR08] Manas Somaiya, Christopher Jermaine, and Sanjay Ranka. Learning correlations using the mixture-of-subsets model. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(4):3:1–3:??, January 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sowah:2022:HEH

[SKM^{+}22] Robert A. Sowah, Bernard Kuditchar, Godfrey A. Mills, Amevi Acakpovi, Raphael A. Twum, Gifty Buah, and Robert Agboyi. HCBST: an efficient hybrid sampling technique for class imbalance problems. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):57:1–57:37, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3488280.

Sariyuce:2017:GMF

[SKSÇ17] Ahmet Erdem Sariyüce, Kamer Kaya, Erik Saule, and Ümit V. Çatalyürek. Graph manipulations for fast centrality computation. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):26:1– 26:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sun:2021:JTD

[SKW^{+}21] Bin Sun, Dehui Kong, Shaofan Wang, Lichun Wang, and Baocai Yin. Joint transferable dictionary learning and view adaptation for multi-view human action recognition. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):32:1–32:23, April 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434746.

Schlieper:2024:EUO

[SLK^{+}24] Philipp Schlieper, Hermann Luft, Kai Klede, Christoph Strohmeyer, Bjoern Eskofier, and Dario Zanca. Enhancing unsupervised outlier model selection: a study on IREOS algorithms. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):178:1–178:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653719.

Sun:2022:GCI

[SLL^{+}22] Heli Sun, Yang Li, Bing Lv, Wujie Yan, Liang He, Shaojie Qiao, and Jianbin Huang. Graph Community Infomax. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):53:1–53:21, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480244.

Singh:2023:MSM

[SLL^{+}23] Karandeep Singh, Seungeon Lee, Giuseppe (Joe) Labianca, Jesse Michael Fagan, and Meeyoung Cha. Multistage machine learning model for hierarchical tie valence prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6): 83:1–83:??, July 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3579096.

Shin:2021:CFA

[SLO^{+}21] Kijung Shin, Euiwoong Lee, Jinoh Oh, Mohammad Hammoud, and Christos Faloutsos. CoCoS: Fast and accurate distributed triangle counting in graph streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):38:1–38:30, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441487.

Sun:2011:ISI

[SLTA11] Jimeng Sun, Yan Liu, Jie Tang, and Chid Apte. Introduction to special issue on large-scale data mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(2):7:1–7:??, February 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Senin:2018:GID

[SLW^{+}18] Pavel Senin, Jessica Lin, Xing Wang, Tim Oates, Sunil Gandhi, Arnold P. Boedihardjo, Crystal Chen, and Susan Frankenstein. Grammar-Viz 3.0: Interactive discovery of variable-length time series patterns. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (1):10:1–10:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sun:2024:TDA

[SLWW24] Mingchen Sun, Yingji Li, Ying Wang, and Xin Wang. Towards domain-aware stable meta learning for outof-distribution generalization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):203:1–203:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3676558.

Sharma:2021:STL

[SM21] Shalini Sharma and Angshul Majumdar. Sequential transform learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):86:1–86:18, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447394.

Syed:2023:SST

[SM23] Tahir Syed and Behroz Mirza.

Self-supervision for tabular data by learning to predict additive homoskedastic Gaussian noise as pretext. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):122:1–122:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3594720.

Sahay:2008:DSB

[SMA^{+}08] Saurav Sahay, Sougata Mukherjea, Eugene Agichtein, Ernest V. Garcia, Shamkant B. Navathe, and Ashwin Ram. Discovering semantic biomedical relations utilizing the Web. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(1):3:1–3:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sanei-Mehri:2021:MLM

[SMDHT21] Seyed-Vahid Sanei-Mehri, Apurba Das, Hooman Hashemi, and Srikanta Tirthapura. Mining largest maximal quasicliques. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):81:1– 81:21, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3446637.

Saha:2018:EMO

[SMK18] Sriparna Saha, Sayantan Mitra, and Stefan Kramer. Exploring multiobjective optimization for multiview clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):44:1– 44:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sharma:2022:IDA

[SMS22] Ms Promila Sharma, Uma Meena, and Girish Kumar Sharma. Intelligent data analysis using optimized support vector machine based data mining approach for tourism industry. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):94:1–94:20, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494566.

Sui:2024:EDG

[SMW^{+}24] Yongduo Sui, Wenyu Mao, Shuyao Wang, Xiang Wang, Jiancan Wu, Xiangnan He, and Tat-Seng Chua. Enhancing out-of-distribution generalization on graphs via causal attention learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):127:1–127:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3644392.

Sun:2024:PEC

[SMY^{+}24] Jianshan Sun, Suyuan Mei, Kun Yuan, Yuanchun Jiang, and Jie Cao. Prerequisiteenhanced category-aware graph neural networks for course recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):112:1–112:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643644.

Sun:2013:PIM

[SNH^{+}13] Yizhou Sun, Brandon Norick, Jiawei Han, Xifeng Yan, Philip S. Yu, and Xiao Yu. PathSelClus: Integrating meta-path selection with user-guided object clustering in heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(3): 11:1–11:??, September 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sheshbolouki:2022:SBA

[SÖ22] Aida Sheshbolouki and M. Tamer Özsu. sGrapp: Butterfly approximation in streaming graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):76:1–76:43, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495011.

Shin:2020:FAP

[SOK^{+}20] Kijung Shin, Sejoon Oh, Jisu Kim, Bryan Hooi, and Christos Faloutsos. Fast, accurate and provable triangle counting in fully dynamic graph streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (2):12:1–12:39, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375392.

Shen:2022:MTM

[SOL22] Yanyan Shen, Baoyuan Ou, and Ranzhen Li. MBN: Towards multi-behavior sequence modeling for next basket recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):81:1–81:23, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3497748.

Saude:2022:RRB

[SRBC22] João Saúde, Guilherme Ramos, Ludovico Boratto, and Carlos Caleiro. A robust reputationbased group ranking system and its resistance to bribery. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (2):26:1–26:35, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3462210.

Strukova:2024:AKI

[SRVM24] Sofia Strukova, José A. Ruipérez-Valiente, and Félix Gómez Mármol. Adapting knowledge inference algorithms to measure geometry competencies through a puzzle game. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):21:1– 21:??, January 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3614436.

Syed:2010:MDP

[SSK^{+}10] Zeeshan Syed, Collin Stultz, Manolis Kellis, Piotr Indyk, and John Guttag. Motif discovery in physiological datasets: a methodology for inferring predictive elements. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(1):2:1–2:??, January 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sun:2023:GNN

[SSL^{+}23] Heli Sun, Miaomiao Sun, Xuechun Liu, Linlin Zhu, Liang He, Xiaolin Jia, and Yuan Chen. Graph neural networks with motisfaware for tenuous subgraph finding. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):109:1–109:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589643.

Shi:2018:VAB

[STD^{+}18] Lei Shi, Hanghang Tong, Madelaine Daianu, Feng Tian, and Paul M. Thompson. Visual analysis of brain networks using sparse regression models. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):5:1–5:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Sun:2008:ITA

[STP^{+}08] Jimeng Sun, Dacheng Tao, Spiros Papadimitriou, Philip S. Yu, and Christos Faloutsos. Incremental tensor analysis: Theory and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(3):11:1–11:??, October 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Song:2021:NEH

[SWD^{+}21] Guojie Song, Yun Wang, Lun Du, Yi Li, and Junshan Wang. Network embedding on hierarchical community structure network. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):62:1–62:23, June 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434747.

Shui:2023:LOL

[SWH^{+}23] Changjian Shui, William Wang, Ihsen Hedhli, Chi Man Wong, Feng Wan, Boyu Wang, and Christian Gagné. Lifelong online learning from accumulated knowledge. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):52:1–52:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563947.

Shu:2024:MIL

[SWY^{+}24] Senlin Shu, Deng-Bao Wang, Suqin Yuan, Hongxin Wei, Jiuchuan Jiang, Lei Feng, and Min-Ling Zhang. Multipleinstance learning from triplet comparison bags. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):90:1–90:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638776.

Shao:2016:SCI

[SYD^{+}16] Junming Shao, Qinli Yang, Hoang-Vu Dang, Bertil Schmidt, and Stefan Kramer. Scalable clustering by iterative partitioning and point attractor representation. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):5:1–5:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Song:2024:MSM

[SYG^{+}24] Derun Song, Enneng Yang, Guibing Guo, Li Shen, Linying Jiang, and Xingwei Wang. Multi-scenario and multitask aware feature interaction for recommendation system. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):142:1–142:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3651312.

Sato:2022:CTG

[SYK22] Ryoma Sato, Makoto Yamada, and Hisashi Kashima. Constant time graph neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):92:1–92:31, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502733.

Shen:2016:SPO

[SYLC16] Chih-Ya Shen, De-Nian Yang, Wang-Chien Lee, and Ming-Syan Chen. Spatial-proximity optimization for rapid task group deployment. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):47:1–47:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Shao:2022:NEM

[SYXW22] Ping Shao, Yang Yang, Shengyao Xu, and Chunping Wang. Network embedding via motifs. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):44:1–44:20, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473911.

Sun:2023:SHI

[SYXZ23] Jianhui Sun, Ying Yang, Guangxu Xun, and Aidong Zhang. Scheduling hyperparameters to improve generalization: From centralized SGD to asynchronous SGD. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):29:1–29:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3544782.

Su:2024:CBF

[SYZ^{+}24a] Cong Su, Guoxian Yu, Yongqing Zheng, Jun Wang, Zhengtian Wu, Xiangliang Zhang, and Carlotta Domeniconi. Causality-based fair multiple decision by response functions. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):61:1– 61:??, April 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3632529.

Sui:2024:CAS

[SYZ^{+}24b] Hongjie Sui, Huan Yan, Tianyi Zheng, Wenzhen Huang, Yunlin Zhuang, and Yong Li. Congestion-aware spatiotemporal graph convolutional network-based A∗ search algorithm for fastest route search. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):179:1– 179:??, August 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3657640.

Sahebi:2024:MMP

[SYZB24] Sherry Sahebi, Mengfan Yao, Siqian Zhao, and Reza Feyzi Behnagh. MoMENt: Marked point processes with memoryenhanced neural networks for user activity modeling. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):155:1–155:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649504.

Shi:2023:ACS

[SZD^{+}23] Dan Shi, Lei Zhu, Xiao Dong, Xuemeng Song, Jingjing Li, and Zhiyong Cheng. Adaptive collaborative soft label learning for unsupervised multiview feature selection. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(8):116:1–116:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3591467.

Shi:2023:KFD

[SZF^{+}23] Linrui Shi, Zheng Zhang, Zizhu Fan, Chao Xi, Zhengming Li, and Gaochang Wu. Kernel Fisher Dictionary Transfer Learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(8):108:1–108:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3588575.

Smith:2016:PNN

[SZLP16] Laura M. Smith, Linhong Zhu, Kristina Lerman, and Allon G. Percus. Partitioning networks with node attributes by compressing information flow. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 15:1–15:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Shi:2022:CLC

[SZWR22] Tian Shi, Xuchao Zhang, Ping Wang, and Chandan K. Reddy. Corpus-level and concept-based explanations for interpretable document classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (3):48:1–48:17, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477539.

Tang:2017:PSS

[TAJY17] Xun Tang, Maha Alabduljalil, Xin Jin, and Tao Yang. Partitioned similarity search with cache-conscious data traversal. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):34:1– 34:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tatti:2019:DFG

[Tat19]  Nikolaj Tatti.  Density-

friendly graph decomposition. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(5):54:1– 54:??, October 2019. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3344210.

Tajeuna:2023:MRS

[TBW23] Etienne Gael Tajeuna, Mohamed Bouguessa, and Shengrui Wang. Modeling regime shifts in multiple time series. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(8):115:1– 115:??, September 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592857.

Tu:2009:SDC

[TC09] Li Tu and Yixin Chen. Stream data clustering based on grid density and attraction. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(3):12:1–12:??, July 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Toth:2018:GDT

[TC18] Edward Toth and Sanjay Chawla. GTΔ: Detecting temporal changes in group stochastic processes. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):39:1–39:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Teinemaa:2019:OOP

[TDLM19] Irene Teinemaa, Marlon Dumas, Marcello La Rosa, and Fabrizio Maria Maggi. Outcome-oriented predictive process monitoring: Review and benchmark. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):17:1–17:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301300.

Trevino:2018:DSE

[THB18] Edgar S. García Treviño, Muhammad Zaid Hameed, and Javier A. Barria. Data stream evolution diagnosis using recursive wavelet density estimators. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (1):14:1–14:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tan:2008:TMG

[THD^{+}08] Henry Tan, Fedja Hadzic, Tharam S. Dillon, Elizabeth Chang, and Ling Feng. Tree model guided candidate generation for mining frequent subtrees from XML documents. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(2):9:1–9:??, July 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tahani:2016:IDD

[THRR16] Maryam Tahani, Ali M. A.

Hemmatyar, Hamid R. Rabiee, and Maryam Ramezani. Inferring dynamic diffusion networks in online media. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):44:1–44:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tu:2024:PFL

[THYL24] Jingke Tu, Jiaming Huang, Lei Yang, and Wanyu Lin. Personalized federated learning with layer-wise feature transformation via meta-learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):99:1–99:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638252.

Theocharidis:2024:ACA

[TKT^{+}24] Konstantinos Theocharidis, Panagiotis Karras, Manolis Terrovitis, Spiros Skiadopoulos, and Hady W. Lauw. Adaptive content-aware influence maximization via online learning to rank. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):146:1–146:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3651987.

Tang:2014:FSS

[TL14] Jiliang Tang and Huan Liu. Feature selection for social media data. ACM Transactions on Knowledge Discovery from Data (TKDD), 8 (4):19:1–19:??, October 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tang:2023:DSB

[TLG^{+}23] Hui Tang, Xun Liang, Yuhui Guo, Xiangping Zheng, Bo Wu, Sensen Zhang, and Zhiying Li. Diffuse and smooth: Beyond truncated receptive field for scalable and adaptive graph representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):75:1–75:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572781.

Tao:2019:RSE

[TLL^{+}19] Zhiqiang Tao, Hongfu Liu, Sheng Li, Zhengming Ding, and Yun Fu. Robust spectral ensemble clustering via rank minimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):4:1–4:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3278606.

Tang:2008:TTA

[TLZ^{+}08] Lei Tang, Huan Liu, Jianping Zhang, Nitin Agarwal, and John J. Salerno. Topic taxonomy adaptation for group profiling. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(4):1:1–1:??, January 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tang:2020:ERF

[TLZ^{+}20] Lei Tang, Zihang Liu, Yaling Zhao, Zongtao Duan, and Jingchi Jia. Efficient ridesharing framework for ride-matching via heterogeneous network embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):27:1–27:24, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3373839.

Tipirneni:2022:SST

[TR22] Sindhu Tipirneni and Chandan K. Reddy. Self-supervised transformer for sparse and irregularly sampled multivariate clinical time-series. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):105:1–105:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3516367.

Torvik:2009:AND

[TS09] Vetle I. Torvik and Neil R. Smalheiser. Author name disambiguation in MED-LINE. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(3):11:1– 11:??, July 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tuomo:2020:BCC

[TSRK20] Alasalmi Tuomo, Jaakko Suutala, Juha Röning, and Heli Koskimäki. Better classifier calibration for small datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):34:1–34:19, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3385656.

Tran:2022:CDP

[TSS22] Cong Tran, Won-Yong Shin, and Andreas Spitz. Community detection in partially observable social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):22:1–22:24, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3461339.

Thomas:2010:MMF

[TVK10] Lini T. Thomas, Satyanarayana R. Valluri, and Kamalakar Karlapalem. MAR-GIN: Maximal frequent subgraph mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 4 (3):10:1–10:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tey:2022:MLB

[TWC22] Fu Jie Tey, Tin-Yu Wu, and Jiann-Liang Chen. Machine learning-based shortterm rainfall prediction from sky data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):102:1–102:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3502731.

Tong:2016:GES

[TWCO16] Hanghang Tong, Fei Wang, Munmun De Choudhury, and Zoran Obradovic. Guest editorial: Special issue on connected health at big data era (BigChat): a TKDD special issue. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4): 37:1–37:??, July 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tang:2015:FMT

[TYG^{+}15] Lu-An Tang, Xiao Yu, Quanquan Gu, Jiawei Han, Guofei Jiang, Alice Leung, and Thomas La Porta. A framework of mining trajectories from untrustworthy data in cyber-physical system. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3):16:1– 16:??, February 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tu:2021:CCJ

[TYW^{+}21] Jinzheng Tu, Guoxian Yu, Jun Wang, Carlotta Domeniconi, Maozu Guo, and Xiangliang Zhang. CrowdWT: Crowdsourcing via joint modeling of workers and tasks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):12:1–12:24, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3421712.

Tang:2010:CAW

[TYZZ10] Jie Tang, Limin Yao, Duo Zhang, and Jing Zhang. A combination approach to Web user profiling. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(1):2:1– 2:??, December 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Tipirneni:2024:SSA

[TZR24] Sindhu Tipirneni, Ming Zhu, and Chandan K. Reddy. StructCoder: Structure-aware transformer for code generation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):70:1– 70:??, April 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3636430.

Vadera:2010:CCS

[Vad10] Sunil Vadera. CSNL: a costsensitive non-linear decision tree algorithm. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(2): 6:1–6:??, May 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Veloso:2019:SSM

[VAFZ19] Bráulio M. Veloso, Renato M. Assunção, Anderson A. Ferreira, and Nivio Ziviani. In search of a stochastic model for the e-news reader. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):65:1–65:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3362695.

deMelo:2012:FNO

[VALF12] Pedro O. S. Vaz de Melo, Virgilio A. F. Almeida, Antonio A. F. Loureiro, and Christos Faloutsos. Forecasting in the NBA and other team sports: Network effects in action. ACM Transactions on Knowledge Discovery from Data (TKDD), 6 (3):13:1–13:??, October 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Varde:2022:CES

[Var22] Aparna S. Varde. Computational estimation by scientific data mining with classical methods to automate learning strategies of scientists. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):86:1– 86:52, October 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502736.

Vaidya:2008:PPD

[VCKP08] Jaideep Vaidya, Chris Clifton, Murat Kantarcioglu, and A. Scott Patterson. Privacypreserving decision trees over vertically partitioned data. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(3):14:1–14:??, October 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Oliveira:2022:ALS

[VDMC22] Lucas Santos De Oliveira and Pedro O. S. Vaz-De-Melo and Aline Carneiro Viana. Assessing large-scale power relations among locations from mobility data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):38:1–38:31, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3470770.

VazDeMelo:2015:UDP

[VFA^{+}15] Pedro O. S. Vaz De Melo, Christos Faloutsos, Renato Assunção, Rodrigo Alves, and Antonio A. F. Loureiro. Universal and distinct properties of communication dynamics: How to generate realistic interevent times. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3):24:1–24:??, April 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

vanLeeuwen:2018:ETS

[vLCV^{+}18] Matthijs van Leeuwen, Polo Chau, Jilles Vreeken, Dafna Shahaf, and Christos Faloutsos. Editorial: TKDD special issue on interactive data exploration and analytics. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):1:1–1:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

vanLeeuwen:2019:ASI

[vLCV^{+}19] Matthijs van Leeuwen, Polo Chau, Jilles Vreeken, Dafna Shahaf, and Christos Faloutsos. Addendum to the special issue on Interactive Data Exploration and Analytics (TKDD, Vol. 12, Iss. 1): Introduction by the Guest Editors. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(1):13:1– 13:??, January 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Vajiac:2023:DIT

[VLK^{+}23] Catalina Vajiac, Meng-Chieh Lee, Aayushi Kulshrestha, Sacha Levy, Namyong Park, Andreas Olligschlaeger, Cara Jones, Reihaneh Rabbany, and Christos Faloutsos. DeltaShield: Information theory for humantrafficking detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):28:1–28:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563040.

Vosoughi:2017:RGP

[VMR17] Soroush Vosoughi, Mostafa ‘Neo’ Mohsenvand, and Deb Roy. Rumor gauge: Predicting the veracity of rumors on Twitter. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):50:1– 50:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Vlachos:2015:DPC

[VSV15] Michail Vlachos, Johannes Schneider, and Vassilios G. Vassiliadis. On data publishing with clustering preservation. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3):23:1– 23:??, April 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wilkinson:2012:SIS

[WAD12] Leland Wilkinson, Anushka Anand, and Tuan Nhon Dang. Substantial improvements in the set-covering projection classifier CHIRP (composite hypercubes on iterated random projections). ACM Transactions on Knowledge Discovery from Data (TKDD), 6(4):19:1–19:??, December 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2010:TSI

[Wan10] Wei Wang. TKDD special issue: SIGKDD 2009. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(4):16:1–16:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2021:LBC

[WBL21] Yunzhe Wang, George Baciu, and Chenhui Li. A layoutbased classification method for visualizing time-varying graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):54:1– 54:24, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3441301.

Wang:2012:LBN

[WC12] Zhenxing Wang and Laiwan Chan. Learning Bayesian networks from Markov random fields: an efficient algorithm for linear models. ACM Transactions on Knowledge Discovery from Data (TKDD), 6 (3):10:1–10:??, October 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wei:2015:MTP

[WC15] Wei Wei and Kathleen M. Carley. Measuring temporal patterns in dynamic social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):9:1–9:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2023:MCS

[WCHH23] Ting-Yun Wang, Chiao-Ting Chen, Ju-Chun Huang, and Szu-Hao Huang. Modeling cross-session information with multi-interest graph neural networks for the next-item recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):1:1–1:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532192.

Wang:2019:IDP

[WCL19] Zhitao Wang, Chengyao Chen, and Wenjie Li. Information diffusion prediction with network regularized role-based user representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):29:1–29:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314106.

Wang:2023:SRM

[WCL^{+}23a] Yu Wang, Chuan Chen, Jinrong Lai, Lele Fu, Yuren Zhou, and Zibin Zheng. A selfrepresentation method with local similarity preserving for fast multi-view outlier detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (1):2:1–2:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532191.

Wu:2023:OMO

[WCL^{+}23b] Youxi Wu, Mingjie Chen, Yan Li, Jing Liu, Zhao Li, Jinyan Li, and Xindong Wu. ONP-Miner: One-off negative sequential pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):37:1–37:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549940.

Wang:2018:CCE

[WCS^{+}18] Can Wang, Chi-Hung Chi, Zhong She, Longbing Cao, and Bela Stantic. Coupled clustering ensemble by exploring data interdependence. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (6):63:1–63:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wu:2024:FRS

[WCX24] Yao Wu, Jian Cao, and Guandong Xu. Fairness in recommender systems: Evaluation approaches and assurance strategies. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (1):10:1–10:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604558.

Wang:2022:SBA

[WCZ^{+}22] Guangtao Wang, Gao Cong, Ying Zhang, Zhen Hai, and Jieping Ye. A synopsis based approach for itemset frequency estimation over massive multitransaction stream. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):29:1–29:30, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3465238.

Wen:2024:AIT

[WCZ^{+}24] Cheng Wen, Yuandao Cai, Bin Zhang, Jie Su, Zhiwu Xu, Dugang Liu, Shengchao Qin, Zhong Ming, and Tian Cong. Automatically inspecting thousands of static bug warnings with large language model: How far are we? ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):168:1–168:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653718.

Wang:2020:EMO

[WDDB20] Tingting Wang, Lei Duan, Guozhu Dong, and Zhifeng Bao. Efficient mining of outlying sequence patterns for analyzing outlierness of sequence data. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):62:1–62:26, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3399671.

Wang:2022:GML

[WDF22] Lichen Wang, Zhengming Ding, and Yun Fu. Generic multi-label annotation via adaptive graph and marginalized augmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):12:1–12:20, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451884.

Wang:2024:LHT

[WDH^{+}24] Song Wang, Yushun Dong, Xiao Huang, Chen Chen, and Jundong Li. Learning hierarchical task structures for fewshot graph classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):67:1–67:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635473.

Wang:2021:HPR

[WDL^{+}21] Hao Wang, Shuai Ding, Yeqing Li, Xiaojian Li, and Youtao Zhang. Hierarchical physician recommendation via diversityenhanced matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):1:1–1:17, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418227.

Wang:2023:GML

[WDL^{+}23] Lichen Wang, Zhengming Ding, Kasey Lee, Seungju Han, Jae-Joon Han, Changkyu Choi, and Yun Fu. Generative multi-label correlation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):18:1–18:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3538708.

Webb:2010:SSI

[Web10]  Geoffrey I. Webb.  Self-

sufficient itemsets: an approach to screening potentially interesting associations between items. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(1):3:1– 3:??, January 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wong:2011:CUA

[WFW^{+}11] Raymond Chi-Wing Wong, Ada Wai-Chee Fu, Ke Wang, Philip S. Yu, and Jian Pei. Can the utility of anonymized data be used for privacy breaches? ACM Transactions on Knowledge Discovery from Data (TKDD), 5(3):16:1–16:??, August 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2023:CUV

[WGC^{+}23] Qianru Wang, Bin Guo, Lu Cheng, Zhiwen Yu, and Huan Liu. CausalSE: Understanding varied spatial effects with missing data toward adding new bike-sharing stations. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (2):20:1–20:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3536427.

Wang:2023:TID

[WGL^{+}23] Hao Wang, Bin Guo, Jiaqi Liu, Yasan Ding, and Zhiwen Yu. Towards informative and diverse dialogue systems over hierarchical crowd intelligence knowledge graph. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):105:1–105:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3583758.

Wu:2021:PGA

[WGYC21] Guanhao Wu, Xiaofeng Gao, Ge Yan, and Guihai Chen. Parallel greedy algorithm to multiple influence maximization in social network. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):43:1–43:21, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442341.

Wu:2023:WEC

[WHC^{+}23] Hongxin Wu, Meng Han, Zhiqiang Chen, Muhang Li, and Xilong Zhang. A weighted ensemble classification algorithm based on nearest neighbors for multi-label data stream. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):72:1–72:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3570960.

Wang:2018:PSM

[WHG^{+}18] Boyue Wang, Yongli Hu, Junbin Gao, Yanfeng Sun, and Baocai Yin. Partial sum minimization of singular values representation on Grassmann manifolds. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (1):13:1–13:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wu:2017:LSO

[WHMY17] Yue Wu, Steven C. H. Hoi, Tao Mei, and Nenghai Yu. Large-scale online feature selection for ultra-high dimensional sparse data. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):48:1–48:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2021:RCE

[WJLL21] Jingjing Wang, Wenjun Jiang, Kenli Li, and Keqin Li. Reducing cumulative errors of incremental CP decomposition in dynamic online social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):42:1–42:33, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441645.

Wang:2023:SDR

[WJLY23] Xiang Wang, Liping Jing, Huafeng Liu, and Jian Yu. Structure-driven representation learning for deep clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):1–25, October 2023. ISSN 1556-472X.

Wu:2010:MAF

[WJR^{+}10] Mingxi Wu, Chris Jermaine, Sanjay Ranka, Xiuyao Song, and John Gums. A modelagnostic framework for fast spatial anomaly detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(4):20:1–20:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2016:ISI

[WL16] Wei Wang and Jure Leskovec.

Introduction to the special issue of best papers in ACM SIGKDD 2014. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):

33:1–33:??, July 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2017:LMD

[WLC^{+}17] Sen Wang, Xue Li, Xiaojun Chang, Lina Yao, Quan Z. Sheng, and Guodong Long. Learning multiple diagnosis codes for ICU patients with local disease correlation mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):31:1– 31:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2021:ACA

[WLD^{+}21] Huandong Wang, Yong Li, Mu Du, Zhenhui Li, and Depeng Jin. App2Vec: Contextaware application usage prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):112:1– 112:21, July 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3451396.

Wang:2023:GDL

[WLD^{+}23] Dexian Wang, Tianrui Li, Ping Deng, Fan Zhang, Wei Huang, Pengfei Zhang, and Jia Liu. A generalized deep learning clustering algorithm based on non-negative matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (7):99:1–99:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584862.

Wu:2024:ASM

[WLG^{+}24] Chenwang Wu, Defu Lian, Yong Ge, Min Zhou, and Enhong Chen. Attacking social media via behavior poisoning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):169:1–169:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3654673.

Wang:2024:TTD

[WLH24] Huan Wang, Guoquan Liu, and Po Hu. TDAN: Transferable domain adversarial network for link prediction in heterogeneous social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):11:1– 11:??, January 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3610229.

Wang:2023:ULH

[WLHH23] Meng Wang, Boyu Li, Kun He, and John Hopcroft. Uncovering the local hidden community structure in social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):67:1–67:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3567597.

Wang:2021:MIM

[WLL^{+}21] Rui Wang, Yongkun Li, Shuai Lin, Hong Xie, Yinlong Xu, and John C. S. Lui. On modeling influence maximization in social activity networks under general settings. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):108:1–108:28, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451218.

Wang:2022:CAS

[WLL^{+}22a] Huandong Wang, Yong Li, Junjie Lin, Hancheng Cao, and Depeng Jin. Context-aware semantic annotation of mobility records. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):47:1–47:20, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3477048.

Wu:2022:NMN

[WLL^{+}22b] Youxi Wu, Lanfang Luo, Yan Li, Lei Guo, Philippe Fournier-Viger, Xingquan Zhu, and Xindong Wu. NTP-Miner: Nonoverlapping three-way sequential pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):51:1–51:21, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480245.

Wen:2018:EAD

[WLP18] Xidao Wen, Yu-Ru Lin, and Konstantinos Pelechrinis. Event analytics via discriminant tensor factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(6):72:1– 72:??, October 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2014:EEM

[WLR^{+}14] Pinghui Wang, John C. S. Lui, Bruno Ribeiro, Don Towsley, Junzhou Zhao, and Xiaohong Guan. Efficiently estimating motif statistics of large networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2):8:1–8:??, September 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wei:2023:DSB

[WLS^{+}23] Xuemei Wei, Yezheng Liu, Jianshan Sun, Yuanchun Jiang, Qifeng Tang, and Kun Yuan. Dual subgraphbased graph neural network for friendship prediction in location-based social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (3):42:1–42:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3554981.

Wang:2024:RET

[WLS^{+}24] Huan Wang, Ruigang Liu, Chuanqi Shi, Junyang Chen, Lei Fang, Shun Liu, and Zhiguo Gong. Resisting the edge-type disturbance for link prediction in heterogeneous networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):45:1–45:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3614099.

Wu:2019:HUI

[WLT19] Jimmy Ming-Tai Wu, Jerry Chun-Wei Lin, and Ashish Tamrakar. High-utility itemset mining with effective pruning strategies. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):58:1–58:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3363571.

Wang:2019:UFM

[WLW^{+}19] Yuandong Wang, Xuelian Lin, Hua Wei, Tianyu Wo, Zhou Huang, Yong Zhang, and Jie Xu. A unified framework with multi-source data for predicting passenger demands of ride services. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):56:1–56:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3355563.

Wang:2023:EEA

[WLX^{+}23] Yuxiang Wang, Jun Liu, Xiaoliang Xu, Xiangyu Ke, Tianxing Wu, and Xiaoxuan Gou. Efficient and effective academic expert finding on heterogeneous graphs through (k, P)-core based embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):85:1–85:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578365.

Wu:2022:EHG

[WMW^{+}22] Hanlu Wu, Tengfei Ma, Lingfei Wu, Fangli Xu, and Shouling Ji. Exploiting heterogeneous graph neural networks with latent worker/task correlation information for label aggregation in crowdsourcing. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):27:1–27:18, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460865.

Wang:2024:PTQ

[WMZL24] Wentao Wang, Huifang Ma, Yan Zhao, and Zhixin Li. Pretraining question embeddings for improving knowledge tracing with self-supervised bigraph co-contrastive learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):74:1–74:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638055.

Wu:2022:HCN

[WN22a] Hanrui Wu and Michael K. Ng. Hypergraph convolution on nodes-hyperedges network for semi-supervised node classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):80:1–80:19, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3494567.

Wu:2022:MGL

[WN22b] Hanrui Wu and Michael K. Ng. Multiple graphs and low-rank embedding for multisource heterogeneous domain adaptation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):77:1–77:25, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3492804.

Wu:2018:GRS

[WNC^{+}18] Hao Wu, Yue Ning, Prithwish Chakraborty, Jilles Vreeken, Nikolaj Tatti, and Naren Ramakrishnan. Generating realistic synthetic population datasets. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):45:1– 45:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wan:2009:DBC

[WND^{+}09] Li Wan, Wee Keong Ng, Xuan Hong Dang, Philip S. Yu, and Kuan Zhang. Densitybased clustering of data streams at multiple resolutions. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(3):14:1– 14:??, July 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2015:LSC

[WNH15] Hua Wang, Feiping Nie, and Heng Huang. Large-scale cross-language Web page classification via dual knowledge transfer using fast nonnegative matrix trifactorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):1:1–1:??, July 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2024:MMG

[WOGZ24] Yashen Wang, Xiaoye Ouyang, Dayu Guo, and Xiaoling Zhu. MEGA: Meta-graph augmented pre-training model for knowledge graph completion. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):30:1– 30:??, January 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3617379.

Wang:2024:VHL

[WPD^{+}24] Haobo Wang, Cheng Peng, Hede Dong, Lei Feng, Weiwei Liu, Tianlei Hu, Ke Chen, and Gang Chen. On the value of head labels in multi-label text classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):124:1–124:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643853.

Wu:2021:LGN

[WPDZ21] Man Wu, Shirui Pan, Lan Du, and Xingquan Zhu. Learning graph neural networks with positive and unlabeled nodes.

ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):101:1–101:25, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450316.

Wang:2023:MIF

[WPW23] Shaokang Wang, Li Pan, and Yu Wu. Meta-information fusion of hierarchical semantics dependency and graph structure for structured text classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):23:1–23:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3537971.

Wei:2016:NBG

[WQZ^{+}16] Qiang Wei, Dandan Qiao, Jin Zhang, Guoqing Chen, and Xunhua Guo. A novel bipartite graph based competitiveness degree analysis from query logs. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2): 21:1–21:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2017:DCM

[WSC^{+}17] Yihan Wang, Shaoxu Song, Lei Chen, Jeffrey Xu Yu, and Hong Cheng. Discovering conditional matching rules. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(4):46:1–46:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2019:RTE

[WSDL19] Yun Wang, Guojie Song, Lun Du, and Zhicong Lu. Realtime estimation of the urban air quality with mobile sensor system. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):49:1–49:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3356584.

Wu:2018:IDC

[WSM^{+}18] Hao Wu, Maoyuan Sun, Peng Mi, Nikolaj Tatti, Chris North, and Naren Ramakrishnan. Interactive discovery of coordinated relationship chains with maximum entropy models. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):7:1–7:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2016:WKI

[WSR^{+}16] Chenguang Wang, Yangqiu Song, Dan Roth, Ming Zhang, and Jiawei Han. World knowledge as indirect supervision for document clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2):13:1–13:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2024:FFL

[WSW24] Chunnan Wang, Xiangyu Shi, and Hongzhi Wang. Fair federated learning with multiobjective hyperparameter optimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):208:1–208:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3676968.

Wang:2012:LME

[WSZ12] Shaojun Wang, Dale Schuurmans, and Yunxin Zhao. The latent maximum entropy principle. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(2):8:1–8:??, July 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wei:2016:HTH

[WSZ^{+}16] Ying Wei, Yangqiu Song, Yi Zhen, Bo Liu, and Qiang Yang. Heterogeneous translated hashing: a scalable solution towards multi-modal similarity search. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):36:1–36:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2014:GML

[WSZZ14] Guangtao Wang, Qinbao Song, Xueying Zhang, and Kaiyuan Zhang. A generic multilabel learning-based classification algorithm recommendation method. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1):7:1–7:??, October 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2024:CQM

[WTD24] Lizhen Wang, Vanha Tran, and Thanhcong Do. A clique-querying mining framework for discovering high utility co-location patterns without generating candidates. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):25:1– 25:??, January 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3617378.

Wang:2022:NNE

[WTZL22] Yu Wang, Hanghang Tong, Ziye Zhu, and Yun Li. Nested named entity recognition: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):108:1–108:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522593.

Webb:2014:EDM

[WV14] Geoffrey I. Webb and Jilles Vreeken. Efficient discovery of the most interesting associations. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(3): 15:1–15:??, June 2014. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2023:MCR

[WWG^{+}23] Jiapu Wang, Boyue Wang, Junbin Gao, Yongli Hu, and Baocai Yin. Multiconcept representation learning for knowledge graph completion. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):11:1– 11:??, January 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3533017.

Wang:2019:DAA

[WWHW19] Huan Wang, Jia Wu, Wenbin Hu, and Xindong Wu. Detecting and assessing anomalous evolutionary behaviors of nodes in evolving social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (1):12:1–12:??, January 2019.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2023:SUI

[WWLM23] Yongjie Wang, Ke Wang, Cheng Long, and Chunyan Miao. Summarizing useritem matrix by group utility maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):86:1–86:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3578586.

Wang:2016:UHM

[WWW^{+}16] Zhongyuan Wang, Fang Wang, Haixun Wang, Zhirui Hu, Jun Yan, Fangtao Li, Ji-Rong Wen, and Zhoujun Li. Unsupervised head-modifier detection in search queries. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2):19:1–19:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2020:EEB

[WWW^{+}20] Changping Wang, Chaokun Wang, Zheng Wang, Xiaojun Ye, and Philip S. Yu. Edge2vec: Edge-based social network embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):45:1–45:24, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3391298.

Wei:2024:NER

[WWZ^{+}24] Xiangyu Wei, Wei Wang, Chongsheng Zhang, Weiping Ding, Bin Wang, Yaguan Qian, Zhen Han, and Chunhua Su. Neighbor-enhanced representation learning for link prediction in dynamic heterogeneous attributed networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):204:1– 204:??, September 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3676559.

Wang:2023:MRB

[WWZL23] Yashen Wang, Zhaoyu Wang, Huanhuan Zhang, and Zhirun Liu. Microblog retrieval based on concept-enhanced pre-training model. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):41:1–41:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3552311.

Wang:2021:SSR

[WXB^{+}21] Kafeng Wang, Haoyi Xiong, Jiang Bian, Zhanxing Zhu, Qian Gao, Zhishan Guo, Cheng-Zhong Xu, Jun Huan, and Dejing Dou. Sampling sparse representations with randomized measurement Langevin dynamics. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):21:1–21:21, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3427585.

Wang:2024:DPF

[WXT^{+}24] Yuhan Wang, Qing Xie, Mengzi Tang, Lin Li, Jingling Yuan, and Yongjian Liu. A dual perspective framework of knowledge-correlation for cross-domain recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):152:1– 152:??, July 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3652520.

Wang:2021:SVR

[WXW^{+}21] Wei Wang, Feng Xia, Jian Wu, Zhiguo Gong, Hanghang Tong, and Brian D. Davison. Scholar2vec: Vector representation of scholars for lifetime collaborator prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):40:1–40:19, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442199.

Wang:2015:QDR

[WY15] Zheng Wang and Jieping Ye. Querying discriminative and representative samples for batch mode active learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 9 (3):17:1–17:??, February 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2017:MDP

[WYG^{+}17] Liang Wang, Zhiwen Yu, Bin Guo, Tao Ku, and Fei Yi. Moving destination prediction using sparse dataset: a mobility gradient descent approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(3):37:1– 37:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2019:FST

[WYWY19] Zheng Wang, Xiaojun Ye, Chaokun Wang, and Philip S. Yu. Feature selection via transferring knowledge across different classes. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):22:1–22:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314202.

Wu:2016:LLR

[WYX^{+}16] Ou Wu, Qiang You, Fen Xia, Lei Ma, and Weiming Hu. Listwise learning to rank from crowds. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):4:1–4:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2021:HNH

[WZ21] Yashen Wang and Huanhuan Zhang. HARP: a novel hierarchical attention model for relation prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):17:1–17:22, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3424673.

Wu:2023:CTIb

[WZB^{+}23] Gongqing Wu, Xingrui Zhuo, Xianyu Bao, Xuegang Hu, Richang Hong, and Xindong Wu. Crowdsourcing truth inference via reliabilitydriven multi-view graph embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5): 65:1–65:??, June 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3565576.

Wang:2011:IDC

[WZL^{+}11] Dingding Wang, Shenghuo Zhu, Tao Li, Yun Chi, and Yihong Gong. Integrating document clustering and multidocument summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(3):14:1–14:??, August 2011. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2015:UCN

[WZL^{+}15] Pinghui Wang, Junzhou Zhao, John C. S. Lui, Don Towsley, and Xiaohong Guan. Unbiased characterization of node pairs over large graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3): 22:1–22:??, April 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wu:2016:MDN

[WZL^{+}16] Yubao Wu, Xiaofeng Zhu, Li Li, Wei Fan, Ruoming Jin, and Xiang Zhang. Mining dual networks: Models, algorithms, and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4): 40:1–40:??, July 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2022:CSD

[WZL^{+}22] Shaowei Wang, Lingling Zhang, Xuan Luo, Yi Yang, Xin Hu, Tao Qin, and Jun Liu. Computer science diagram understanding with topology parsing. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):114:1–114:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522689.

Wang:2023:MRA

[WZL^{+}23a] Chunyang Wang, Yanmin Zhu, Haobing Liu, Tianzi Zang, Ke Wang, and Jiadi Yu. Multifaceted relation-aware metalearning with dual customization for user cold-start recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):130:1–130:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597458.

Wang:2023:LAA

[WZL^{+}23b] Ke Wang, Yanmin Zhu, Haobing Liu, Tianzi Zang, and Chunyang Wang. Learning aspect-aware high-order representations from ratings and reviews for recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):3:1–3:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532188.

Wu:2023:LES

[WZL^{+}23c] Likang Wu, Hongke Zhao, Zhi Li, Zhenya Huang, Qi Liu, and Enhong Chen. Learning the explainable semantic relations via unified graph topic-disentangled neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(8):110:1– 110:??, September 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3589964.

Wang:2012:CDS

[WZLG12] Dingding Wang, Shenghuo Zhu, Tao Li, and Yihong Gong. Comparative document summarization via discriminative sentence selection. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(3):12:1–12:??, October 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2013:CDS

[WZLG13] Dingding Wang, Shenghuo Zhu, Tao Li, and Yihong Gong. Comparative document summarization via discriminative sentence selection. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(1):2:1–2:??, March 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Wang:2021:HCD

[WZLZ21] Yashen Wang, Huanhuan Zhang, Zhirun Liu, and Qiang Zhou. Hierarchical conceptdriven language model. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):104:1–104:22, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451167.

Wan:2023:PMT

[WZLZ23] Mingyang Wan, Daochen Zha, Ninghao Liu, and Na Zou. In-processing modeling techniques for machine learning fairness: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):35:1–35:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551390.

Wang:2023:SSD

[WZR^{+}23] Yizong Wang, Dong Zhao, Yajie Ren, Desheng Zhang, and Huadong Ma. SPAP: Simultaneous demand prediction and planning for electric vehicle chargers in a new city. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):60:1–60:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565577.

Wang:2022:CCN

[WZW^{+}22] Xueyuan Wang, Hongpo Zhang, Zongmin Wang, Yaqiong Qiao, Jiangtao Ma, and Honghua Dai. Con&Net: a cross-network anchor link discovery method based on embedding representation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):36:1–36:18, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469083.

Wang:2023:ASA

[WZWC23] Chunnan Wang, Kaixin Zhang, Hongzhi Wang, and Bozhou Chen. Auto-STGCN: Autonomous spatial-temporal graph convolutional network search. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):73:1– 73:??, June 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3571285.

Wu:2023:CTIa

[WZX^{+}23] Gongqing Wu, Liangzhu Zhou, Jiazhu Xia, Lei Li, Xianyu Bao, and Xindong Wu. Crowdsourcing truth inference based on label confidence clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):46:1–46:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3556545.

Wang:2022:GES

[WZZ^{+}22] Zhaobo Wang, Yanmin Zhu, Qiaomei Zhang, Haobing Liu, Chunyang Wang, and Tong Liu. Graph-enhanced spatialtemporal network for next POI recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):104:1–104:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3513092.

Wang:2024:PDP

[WZZ^{+}24] Yilin Wang, Sha Zhao, Shiwei Zhao, Runze Wu, Yuhong Xu, Jianrong Tao, Tangjie Lv, Shijian Li, Zhipeng Hu, and Gang Pan. PU-Detector: a PU learning-based framework for real money trading detection in MMORPG. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):91:1–91:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638561.

Xu:2022:PPM

[XCL22] Honghui Xu, Zhipeng Cai, and Wei Li. Privacy-preserving mechanisms for multi-label image recognition. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):69:1–69:21, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491231.

Xia:2021:ETD

[XJW^{+}21] Peike Xia, Wenjun Jiang, Jie Wu, Surong Xiao, and Guojun Wang. Exploiting temporal dynamics in product reviews for dynamic sentiment prediction at the aspect level. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):68:1–68:29, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441451.

Xun:2021:MPI

[XJZ21] Guangxu Xun, Kishlay Jha, and Aidong Zhang. MeSHProbeNet-P: Improving large-scale MeSH indexing with personalizable MeSH probes. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (1):11:1–11:14, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3421713.

Xiong:2016:KIT

[XKH^{+}16] Feiyu Xiong, Moshe Kam, Leonid Hrebien, Beilun Wang, and Yanjun Qi. Kernelized information-theoretic metric learning for cancer diagnosis using high-dimensional molecular profiling data. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):38:1–38:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Xu:2014:GGB

[XKW^{+}14] Zhiqiang Xu, Yiping Ke, Yi Wang, Hong Cheng, and James Cheng. GBAGC: a general Bayesian framework for attributed graph clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(1):5:1–5:??, August 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Xie:2015:MMA

[XL15] Hong Xie and John C. S. Lui. Mathematical modeling and analysis of product rating with partial information. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4): 26:1–26:??, June 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Xu:2016:PSP

[XL16] Silei Xu and John C. S.

Lui. Product selection problem: Improve market share by learning consumer behavior. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):34:1– 34:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Xiong:2020:ISI

[XL20] Hui Xiong and Chih-Jen Lin.

Introduction to the special issue on the best papers from KDD 2018. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):51e:1–51e:2, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3407901.

Xi:2024:UOR

[XLD^{+}24] Yunjia Xi, Weiwen Liu, Xinyi Dai, Ruiming Tang, Qing Liu, Weinan Zhang, and Yong Yu. Utility-oriented reranking with counterfactual context. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):193:1–193:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671004.

Xia:2021:GDD

[XLL^{+}21] Tong Xia, Junjie Lin, Yong Li, Jie Feng, Pan Hui, Funing Sun, Diansheng Guo, and Depeng Jin. 3DGCN: 3-Dimensional dynamic graph convolutional network for citywide crowd flow prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):110:1–110:21, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451394.

Xia:2024:HEU

[XLQ^{+}24] Tong Xia, Yong Li, Yunhan Qi, Jie Feng, Fengli Xu, Funing Sun, Diansheng Guo, and Depeng Jin. History-enhanced and uncertainty-aware trajectory recovery via attentive neural network. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):53:1–53:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3615660.

Xiao:2024:NGN

[XLS24] Feng Xiao, Youfa Liu, and Jia Shao. NNC-GCN: Neighboursto-neighbours contrastive graph convolutional network for semi-supervised classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):97:1–97:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638780.

Xu:2020:NAB

[XLT^{+}20] Jiarong Xu, Yifan Luo, Jianrong Tao, Changjie Fan, Zhou Zhao, and Jiangang Lu. NGUARD+: an attentionbased game bot detection framework via player behavior sequences. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):65:1–65:24, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3399711.

Xie:2018:ERP

[XML18] Hong Xie, Richard T. B. Ma, and John C. S. Lui. Enhancing reputation via price discounts in e-commerce systems: a data-driven approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3):26:1–26:??, April 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Xia:2024:FEF

[XRGZ24] Bolun (Namir) Xia, Vipula Rawte, Aparna Gupta, and Mohammed Zaki. FETILDA: Evaluation framework for effective representations of long financial documents. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):182:1–182:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3657299.

Xu:2022:TAG

[XSZ^{+}22] Yonghui Xu, Shengjie Sun, Huiguo Zhang, Chang’an Yi, Yuan Miao, Dong Yang, Xiaonan Meng, Yi Hu, Ke Wang, Huaqing Min, Hengjie Song, and Chuanyan Miao. Timeaware graph embedding: a temporal smoothness and task-oriented approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):56:1–56:23, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480243.

Xu:2020:ANA

[XSZY20] Yanan Xu, Yanyan Shen, Yanmin Zhu, and Jiadi Yu. AR 2 Net: an attentive neural approach for business location selection with satellite data and urban data. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (2):20:1–20:28, March 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3372406.

Xiao:2022:TQI

[XW22] Houping Xiao and Shiyu Wang. Toward quality of information aware distributed machine learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):109:1–109:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522591.

Xiao:2024:TGW

[XWW^{+}24] Meng Xiao, Dongjie Wang, Min Wu, Kunpeng Liu, Hui Xiong, Yuanchun Zhou, and Yanjie Fu. Traceable groupwise self-optimizing feature transformation learning: a dual optimization perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):76:1–76:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638059.

Xiong:2022:GDL

[XWZ^{+}22] Haoyi Xiong, Ruosi Wan, Jian Zhao, Zeyu Chen, Xingjian Li, Zhanxing Zhu, and Jun Huan. GrOD: Deep learning with gradients orthogonal decomposition for knowledge transfer, distillation, and adversarial training. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):117:1–117:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3530836.

Xu:2024:DLN

[XXG^{+}24] Hui Xu, Liyao Xiang, Xiaoying Gan, Luoyi Fu, Xinbing Wang, and Chenghu Zhou. Distributional learning for network alignment with global constraints. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):88:1–88:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638056.

Xu:2024:OWG

[XXO^{+}24] Hui Xu, Liyao Xiang, Junjie Ou, Yuting Weng, Xinbing Wang, and Chenghu Zhou. Open-world graph active learning for node classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (2):43:1–43:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3607144.

Xue:2019:SAP

[XXZ19] Yu Xue, Bing Xue, and Mengjie Zhang. Self-adaptive particle swarm optimization for large-scale feature selection in classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):50:1–50:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3340848.

Xu:2021:CIN

[XYGC21] En Xu, Zhiwen Yu, Bin Guo, and Helei Cui. Core interest network for click-through rate prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):23:1–23:16, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428079.

Xu:2020:NSR

[XYW^{+}20] Yuanbo Xu, Yongjian Yang, En Wang, Jiayu Han, Fuzhen Zhuang, Zhiwen Yu, and Hui Xiong. Neural serendipity recommendation: Exploring the balance between accuracy and novelty with sparse explicit feedback. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):50:1–50:25, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3396607.

Xiao:2020:MUM

[XYZ^{+}20] Keli Xiao, Zeyang Ye, Lihao Zhang, Wenjun Zhou, Yong Ge, and Yuefan Deng. Multiuser mobile sequential recommendation for route optimization. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):52:1–52:28, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3360048.

Xu:2023:OSF

[XZ23] Wanyue Xu and Zhongzhi Zhang. Optimal scalefree small-world graphs with minimum scaling of cover time. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):93:1– 93:??, August 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3583691.

Xie:2021:UPC

[XZLL21] Hong Xie, Mingze Zhong, Yongkun Li, and John C. S. Lui. Understanding persuasion cascades in online product rating systems: Modeling, analysis, and inference. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):36:1–36:29, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3440887.

Xu:2019:FGA

[XZSY19] Yanan Xu, Yanmin Zhu, Yanyan Shen, and Jiadi Yu. Fine-grained air quality inference with remote sensing data and ubiquitous urban data. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (5):46:1–46:??, October 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3340847.

Xie:2018:SNM

[XZXW18] Wei Xie, Feida Zhu, Jing Xiao, and Jianzong Wang. Social network monitoring for bursty cascade detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(4):40:1–40:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Xu:2012:GME

[XZYL12] Tianbing Xu, Zhongfei Zhang, Philip S. Yu, and Bo Long. Generative models for evolutionary clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 6(2):7:1–7:??, July 2012. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yu:2024:IGC

[YBTL24] Penghang Yu, Bing-Kun Bao, Zhiyi Tan, and Guanming Lu. Improving graph collaborative filtering with directional behavior enhanced contrastive learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):183:1–183:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663574.

Yin:2015:MLB

[YCC^{+}15] Hongzhi Yin, Bin Cui, Ling Chen, Zhiting Hu, and Chengqi Zhang. Modeling location-based user rating profiles for personalized recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3): 19:1–19:??, April 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ye:2008:DSA

[YCJK08] Jieping Ye, Jianhui Chen, Ravi Janardan, and Sudhir Kumar. Developmental stage annotation of Drosophila gene expression pattern images via an entire solution path for LDA. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(1):4:1–4:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yao:2021:SCI

[YCL^{+}21] Liuyi Yao, Zhixuan Chu, Sheng Li, Yaliang Li, Jing Gao, and Aidong Zhang. A survey on causal inference. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(5):74:1– 74:46, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3444944.

Ye:2024:BYC

[YCW^{+}24] Tiandi Ye, Cen Chen, Yinggui Wang, Xiang Li, and Ming Gao. BapFL: You can backdoor personalized federated learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):166:1–166:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649316.

Yildiz:2022:SRR

[YDE^{+}22] Ilkay Yildiz, Jennifer Dy, Deniz Erdogmus, Susan Ostmo, J. Peter Campbell, Michael F. Chiang, and Stratis Ioannidis. Spectral ranking regression. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):120:1–120:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3530693.

Yu:2015:CSF

[YDS^{+}15] Kui Yu, Wei Ding, Dan A. Simovici, Hao Wang, Jian Pei, and Xindong Wu. Classification with streaming features: an emerging-pattern mining approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(4): 30:1–30:??, June 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yu:2023:ITB

[YDW^{+}23] Zhiwen Yu, Minling Dang, Qilong Wu, Liming Chen, Yujin Xie, Yu Wang, and Bin Guo. An information theory based method for quantifying the predictability of human mobility. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):131:1–131:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3597500.

Yang:2022:CCP

[YGL^{+}22] Keyu Yang, Yunjun Gao, Lei Liang, Song Bian, Lu Chen, and Baihua Zheng. CrowdTC: Crowd-powered learning for text classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):15:1–15:23, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457216.

Yu:2012:LLC

[YHCL12] Hsiang-Fu Yu, Cho-Jui Hsieh, Kai-Wei Chang, and Chih-Jen Lin. Large linear classification when data cannot fit in memory. ACM Transactions on Knowledge Discovery from Data (TKDD), 5 (4):23:1–23:??, February 2012. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yu:2015:GGA

[YHL15] Rose Yu, Xinran He, and Yan Liu. GLAD: Group anomaly detection in social media analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 10 (2):18:1–18:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yang:2023:PFL

[YHLC23] Lei Yang, Jiaming Huang, Wanyu Lin, and Jiannong Cao. Personalized federated learning on Non-IID data via group-based meta-learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(4):49:1–49:??, May 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3558005.

Yin:2024:CVK

[YHW^{+}24] Jiao Yin, Wei Hong, Hua Wang, Jinli Cao, Yuan Miao, and Yanchun Zhang. A compact vulnerability knowledge graph for risk assessment. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):194:1–194:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671005.

Yang:2024:MSM

[YHZ^{+}24] Hongwei Yang, Hui He, Weizhe Zhang, Yan Wang, and Lin Jing. Multi-source and multi-modal deep network embedding for crossnetwork node classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):149:1–149:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653304.

Yang:2024:HPL

[YJT^{+}24] Jingfeng Yang, Hongye Jin, Ruixiang Tang, Xiaotian Han, Qizhang Feng, Haoming Jiang, Shaochen Zhong, Bing Yin, and Xia Hu. Harnessing the power of LLMs in practice: a survey on ChatGPT and beyond. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):160:1–160:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649506.

Yao:2024:ALG

[YL24] Kai-Lang Yao and Wu-Jun Li. Asymmetric learning for graph neural network based link prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):106:1–106:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3640347.

Yan:2020:MTI

[YLHY20] Xiaoqiang Yan, Zhengzheng Lou, Shizhe Hu, and Yangdong Ye. Multi-task information bottleneck co-clustering for unsupervised cross-view human action categorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(2):15:1– 15:23, March 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375394.

Yang:2013:EOL

[YLK13] Haiqin Yang, Michael R. Lyu, and Irwin King. Efficient online learning for multitask feature selection. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(2):6:1– 6:??, July 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yang:2019:PCM

[YLL19] Wangdong Yang, Kenli Li, and Keqin Li. A pipeline computing method of SpTV for three-order tensors on CPU and GPU. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):63:1–63:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3363575.

Yan:2021:SAB

[YLL^{+}21a] Ruidong Yan, Yi Li, Deying Li, Yongcai Wang, Yuqing Zhu, and Weili Wu. A stochastic algorithm based on reverse sampling technique to fight against the cyberbullying. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):71:1–71:22, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441455.

Yu:2021:UVC

[YLL21b] Kui Yu, Lin Liu, and Jiuyong Li. A unified view of causal and non-causal feature selection. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):63:1–63:46, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436891.

Yu:2024:FSE

[YLL^{+}24] Kui Yu, Zhaolong Ling, Lin Liu, Peipei Li, Hao Wang, and Jiuyong Li. Feature selection for efficient local-to-global Bayesian network structure learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):37:1–37:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624479.

Yan:2019:RBT

[YLW^{+}19] Ruidong Yan, Yi Li, Weili Wu, Deying Li, and Yongcai Wang. Rumor blocking through online link deletion on social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(2):16:1–16:??, June 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301302.

Yang:2024:DPD

[YMD^{+}24] Lingkai Yang, Sally McClean, Mark Donnelly, Kashaf Khan, and Kevin Burke. Detecting process duration drift using gamma mixture models in a left-truncated and rightcensored environment. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):195:1–195:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3669942.

Yang:2020:CIM

[YMPH20] Yu Yang, Xiangbo Mao, Jian Pei, and Xiaofei He. Continuous influence maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(3):29:1–29:38, May 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3380928.

Yuan:2023:IVD

[YMX^{+}23] Junkun Yuan, Xu Ma, Ruoxuan Xiong, Mingming Gong, Xiangyu Liu, Fei Wu, Lanfen Lin, and Kun Kuang. Instrumental variable-driven domain generalization with unobserved confounders. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(8):118:1–118:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3595380.

Yin:2021:RIR

[YSG^{+}21] Shuai Yin, Yanfeng Sun, Junbin Gao, Yongli Hu, Boyue Wang, and Baocai Yin. Robust image representation via low rank locality preserving projection. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):64:1–64:22, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434768.

Ye:2024:MDF

[YSLL24] Xiaoqing Ye, Yang Sun, Dun Liu, and Tianrui Li. A multisource data fusion-based heterogeneous graph attention network for competitor prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):39:1–39:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3625101.

Yue:2021:EBC

[YSW^{+}21] Lin Yue, Hao Shen, Sen Wang, Robert Boots, Guodong Long, Weitong Chen, and Xiaowei Zhao. Exploring BCI control in smart environments: Intention recognition via EEG representation enhancement learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (5):90:1–90:20, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3450449.

Yang:2022:DTA

[YSY^{+}22] Xu Yang, Chao Song, Mengdi Yu, Jiqing Gu, and Ming Liu. Distributed triangle approximately counting algorithms in simple graph stream. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(4):79:1– 79:43, August 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3494562.

Yang:2018:FFR

[YTH18] Pei Yang, Qi Tan, and Jingrui He. Function-on-function regression with mode-sparsity regularization. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(3): 36:1–36:??, April 2018. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yang:2018:LIC

[YTL18] Yang Yang, Jie Tang, and Juanzi Li. Learning to infer competitive relationships in heterogeneous networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(1):12:1–12:??, February 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yu:2016:STR

[YTW^{+}16] Zhiwen Yu, Miao Tian, Zhu Wang, Bin Guo, and Tao Mei. Shop-type recommendation leveraging the data from social media and locationbased services. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1): 1:1–1:??, August 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yao:2024:TLP

[YW24] Rujing Yao and Ou Wu.

A taxonomy for learning with perturbation and algorithms. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):129:1–129:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3644391.

Yu:2016:FDV

[YWC^{+}16] Zhiwen Yu, Zhitao Wang, Liming Chen, Bin Guo, and Wenjie Li. Featuring, detecting, and visualizing human sentiment in Chinese micro-blog. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4): 48:1–48:??, July 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yang:2024:SGI

[YWC24] Ming-Chuan Yang, Guo-Wei Wong, and Meng Chang Chen. Sparse grid imputation using unpaired imprecise auxiliary data: Theory and application to PM2.5 estimation. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):65:1–65:??, April 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3634751.

Yu:2016:SAO

[YWDP16] Kui Yu, Xindong Wu, Wei Ding, and Jian Pei. Scalable and accurate online feature selection for big data. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2):16:1–16:??, December 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yin:2022:WFM

[YWG^{+}22] Jianfei Yin, Ruili Wang, Yeqing Guo, Yizhe Bai, Shunda Ju, Weili Liu, and Joshua Zhexue Huang. Wealth flow model: Online portfolio selection based on learning wealth flow matrices. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):30:1–30:27, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3464308.

Yu:2015:DIP

[YWH^{+}15] Zhiwen Yu, Zhu Wang, Huilei He, Jilei Tian, Xinjiang Lu, and Bin Guo. Discovering information propagation patterns in microblogging services. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):7:1–7:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yuan:2022:AIC

[YWK^{+}22] Junkun Yuan, Anpeng Wu, Kun Kuang, Bo Li, Runze Wu, Fei Wu, and Lanfen Lin. Auto IV: Counterfactual prediction via automatic instrumental variable decomposition. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):74:1–74:20, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494568.

Yang:2024:MGC

[YWL^{+}24] Xihong Yang, Yiqi Wang, Yue Liu, Yi Wen, Lingyuan Meng, Sihang Zhou, Xinwang Liu, and En Zhu. Mixed graph contrastive network for semi-supervised node classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):162:1–162:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3641549.

Yang:2019:TSV

[YWR^{+}19] Wenmain Yang, Kun Wang, Na Ruan, Wenyuan Gao, Weijia Jia, Wei Zhao, Nan Liu, and Yunyong Zhang. Time-sync video tag extraction using semantic association graph. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (4):37:1–37:??, August 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3332932.

Yang:2023:TCS

[YWR^{+}23] Jie Yang, Zhixiao Wang, Xiaobin Rui, Yahui Chai, Philip S. Yu, and Lichao Sun. Triadic closure sensitive influence maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(6):77:1–77:??, July 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3573011.

Yang:2022:SHS

[YWS^{+}22] Yang Yang, Hongchen Wei, Zhen-Qiang Sun, Guang-Yu Li, Yuanchun Zhou, Hui Xiong, and Jian Yang. S2OSC: a holistic semi-supervised approach for open set classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (2):34:1–34:27, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468675.

Yang:2014:USN

[YWW^{+}14] Zhi Yang, Christo Wilson, Xiao Wang, Tingting Gao, Ben Y. Zhao, and Yafei Dai. Uncovering social network Sybils in the wild. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(1):2:1–2:??, February 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Ying:2021:IKB

[YWW^{+}21] Shi Ying, Bingming Wang, Lu Wang, Qingshan Li, Yishi Zhao, Jianga Shang, Hao Huang, Guoli Cheng, Zhe Yang, and Jiangyi Geng. An improved KNN-based efficient log anomaly detection method with automatically labeled samples. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):34:1–34:22, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441448.

Yang:2024:SBA

[YWZ^{+}24] Yang Yang, Feifei Wang, Enqiang Zhu, Fei Jiang, and Wen Yao. Social behavior analysis in exclusive enterprise social networks by FastHAND. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):134:1–134:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3646552.

Yang:2021:EDR

[YYC^{+}21] Lei Yang, Xi Yu, Jiannong Cao, Xuxun Liu, and Pan Zhou. Exploring deep reinforcement learning for task dispatching in autonomous ondemand services. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):44:1–44:23, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3442343.

Yu:2022:CFS

[YYD22] Kui Yu, Yajing Yang, and Wei Ding. Causal feature selection with missing data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(4):66:1– 66:24, August 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3488055.

Yang:2016:JML

[YYF^{+}16] Pei Yang, Hongxia Yang, Haoda Fu, Dawei Zhou, Jieping Ye, Theodoros Lappas, and Jingrui He. Jointly modeling label and feature heterogeneity in medical informatics. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):39:1– 39:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yan:2024:GMC

[YYS^{+}24] Bo Yan, Cheng Yang, Chuan Shi, Yong Fang, Qi Li, Yanfang Ye, and Junping Du. Graph mining for cybersecurity: a survey. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):47:1–47:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610228.

Ying:2018:FIG

[YZH^{+}18] Josh Jia-Ching Ying, Ji Zhang, Che-Wei Huang, Kuan-Ta Chen, and Vincent S. Tseng. FrauDetector+: an incremental graph-mining approach for efficient fraudulent phone call detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 12 (6):68:1–68:??, October 2018.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Yuan:2022:AMS

[YZL^{+}22] Mu Yuan, Lan Zhang, Xiang-Yang Li, Lin-Zhuo Yang, and Hui Xiong. Adaptive model scheduling for resource-efficient data labeling. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(4):71:1– 71:22, August 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3494559.

Yang:2024:LEE

[YZW^{+}24] Yi Yang, Zhong-Qiu Zhao, Gongqing Wu, Xingrui Zhuo, Qing Liu, Quan Bai, and Weihua Li. A lightweight, effective, and efficient model for label aggregation in crowdsourcing. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4): 81:1–81:??, May 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3630102.

Yang:2022:SWT

[YZZ22] Xing Yang, Chen Zhang, and Baihua Zheng. Segmentwise time-varying dynamic Bayesian network with graph regularization. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):113:1–113:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522589.

Zamzami:2020:PMF

[ZB20] Nuha Zamzami and Nizar Bouguila. Probabilistic modeling for frequency vectors using a flexible shiftedscaled Dirichlet distribution prior. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(6):69:1– 69:35, October 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3406242.

Zugner:2020:AAG

[ZBAG20] Daniel Zügner, Oliver Borchert, Amir Akbarnejad, and Stephan Günnemann. Adversarial attacks on graph neural networks: Perturbations and their patterns. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):57:1–57:31, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3394520.

Zhang:2020:TOO

[ZBL^{+}20] Ping Zhang, Zhifeng Bao, Yuchen Li, Guoliang Li, Yipeng Zhang, and Zhiyong Peng. Towards an optimal outdoor advertising placement: When a budget constraint meets moving trajectories. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (5):51:1–51:32, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3350488.

Zhao:2024:LIT

[ZBX^{+}24] Ziyu Zhao, Yuqi Bai, Ruoxuan Xiong, Qingyu Cao, Chao Ma, Ning Jiang, Fei Wu, and Kun Kuang. Learning individual treatment effects under heterogeneous interference in networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):199:1–199:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673761.

Zhang:2024:DDL

[ZCC^{+}24] Chi Zhang, Linhao Cai, Meng Chen, Xiucheng Li, and Gao Cong. DeepMeshCity: a deep learning model for urban grid prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):148:1–148:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3652859.

Zhang:2017:CDM

[ZCF^{+}17] Tianyang Zhang, Peng Cui, Christos Faloutsos, Yunfei Lu, Hao Ye, Wenwu Zhu, and Shiqiang Yang. comeNgo: a dynamic model for social group evolution. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (4):41:1–41:??, August 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhou:2019:PMM

[ZCL19] Xiren Zhou, Huanhuan Chen, and Jinlong Li. Probabilistic mixture model for mapping the underground pipes. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(5):47:1– 47:??, October 2019. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3344721.

Zhu:2022:PPN

[ZCL^{+}22] Nengjun Zhu, Jian Cao, Xinjiang Lu, Chuanren Liu, Hao Liu, Yanyan Li, Xiangfeng Luo, and Hui Xiong. Predicting a person’s next activity region with a dynamic region-relation-aware graph neural network. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(6):116:1–116:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3529091.

Zhuang:2024:TRR

[ZCL24] Wen-Ming Zhuang, Chih-Yao Chen, and Cheng-Te Li. Towards robust rumor detection with graph contrastive and curriculum learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):175:1–175:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653023.

Zaki:2010:VVO

[ZCS10] Mohammed J. Zaki, Christopher D. Carothers, and Boleslaw K. Szymanski. VOGUE: a variable order hidden Markov model with duration based on frequent sequence mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 4(1):5:1–5:??, January 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhu:2019:AAL

[ZCZQ19] Peisong Zhu, Zhuang Chen, Haojie Zheng, and Tieyun Qian. Aspect aware learning for aspect category sentiment analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):55:1–55:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3350487.

Zhang:2022:SUM

[ZDY^{+}22] Chunkai Zhang, Zilin Du, Yuting Yang, Wensheng Gan, and Philip S. Yu. On-shelf utility mining of sequence data. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):21:1–21:31, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457570.

Zhang:2024:FFF

[ZELS24] Wentai Zhang, HaiHong E., Haoran Luo, and Mingzhi Sun. FulBM: Fast fully batch maintenance for landmarkbased 3-hop cover labeling. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):159:1–159:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3650035.

Zhang:2023:ESM

[ZFH23] Yihong Zhang, Xiu Susie Fang, and Takahiro Hara. Evolving social media background representation with frequency weights and cooccurrence graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):95:1–95:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3585389.

Zhang:2024:TFL

[ZFL^{+}24] Lei Zhang, Lele Fu, Chen Liu, Zhao Yang, Jinghua Yang, Zibin Zheng, and Chuan Chen. Toward few-label vertical federated learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):176:1–176:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3656344.

Zhang:2022:BSF

[ZFW22] Min-Ling Zhang, Jun-Peng Fang, and Yi-Bo Wang. BiLabel-specific features for multi-label classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(1):18:1–18:23, February 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3458283.

Zhao:2018:JRL

[ZFWC18] Wayne Xin Zhao, Feifan Fan, Ji-Rong Wen, and Edward Y. Chang. Joint representation learning for locationbased social networks with multi-grained sequential contexts. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(2):22:1– 22:??, March 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhong:2014:UBL

[ZFY14] Erheng Zhong, Wei Fan, and Qiang Yang. User behavior learning and transfer in composite social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(1):6:1–6:??, February 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2018:SRI

[ZGC18] Ziqi Zhang, Jie Gao, and Fabio Ciravegna. SemRe-Rank: Improving automatic term extraction by incorporating semantic relatedness with personalised PageRank. ACM Transactions on Knowledge Discovery from Data (TKDD), 12(5):57:1–57:??, July 2018. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2021:CHI

[ZGHM21] Benhui Zhang, Maoguo Gong, Jianbin Huang, and Xiaoke Ma. Clustering heterogeneous information network by joint graph embedding and nonnegative matrix factorization. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (4):73:1–73:25, June 2021.

CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441449.

Zhu:2024:STR

[ZGQ^{+}24] Ronghang Zhu, Dongliang Guo, Daiqing Qi, Zhixuan Chu, Xiang Yu, and Sheng Li. A survey of trustworthy representation learning across domains. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):173:1–173:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3657301.

Zhao:2022:STE

[ZGY^{+}22] Liang Zhao, Yuyang Gao, Jieping Ye, Feng Chen, Yanfang Ye, Chang-Tien Lu, and Naren Ramakrishnan. Spatiotemporal event forecasting using incremental multi-source feature learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):40:1–40:28, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3464976.

Zhou:2024:CED

[ZGY^{+}24] Peng Zhou, Yufeng Guo, Haoran Yu, Yuanting Yan, Yanping Zhang, and Xindong Wu. Concept evolution detecting over feature streams. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(8):209:1–209:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3678012.

Zhu:2016:CCS

[ZH16] Yada Zhu and Jingrui He.

Co-clustering structural temporal data with applications to semiconductor manufacturing. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(4):43:1– 43:??, July 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhou:2021:UMF

[ZH21] Zhengze Zhou and Giles Hooker. Unbiased measurement of feature importance in tree-based methods. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):26:1–26:21, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429445.

Zhang:2023:EIS

[ZH23] Yihong Zhang and Takahiro Hara. Explainable integration of social media background in a dynamic neural recommender. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):40:1–40:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3550279.

Zhang:2021:EHQ

[ZHL^{+}21] Haida Zhang, Zengfeng Huang, Xuemin Lin, Zhe Lin, Wenjie Zhang, and Ying Zhang. Efficient and high-quality seeded graph matching: Employing higher-order structural information. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):51:1–51:31, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3442340.

Zhang:2024:BSB

[ZHLZ24] Acong Zhang, Jincheng Huang, Ping Li, and Kai Zhang. Building shortcuts between distant nodes with biaffine mapping for graph convolutional networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):139:1–139:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3650113.

Zhou:2019:RMS

[ZHS^{+}19] Yu Zhou, Jianbin Huang, Heli Sun, Yizhou Sun, Shaojie Qiao, and Stephen Wambura.

Recurrent meta-structure for robust similarity measure in heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):64:1–64:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3364226.

Zhu:2023:NGI

[ZHSL23] Xuliang Zhu, Xin Huang, Longxu Sun, and Jiming Liu. A novel graph indexing approach for uncovering potential COVID-19 transmission clusters. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(2):24:1–24:??, February 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3538492.

Zhang:2020:TWS

[ZHT20] Chen Zhang, Steven C. H.

Hoi, and Fugee Tsung. Timewarped sparse non-negative factorization for functional data analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 14 (6):72:1–72:23, October 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408313.

Zheng:2024:PPA

[ZJ24] Yimei Zheng and Caiyan Jia. ProtoMGAE: Prototypeaware masked graph autoencoder for graph representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):137:1–137:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3649143.

Zhu:2024:NTS

[ZJJ^{+}24] Yichen Zhu, Bo Jiang, Haiming Jin, Mengtian Zhang, Feng Gao, Jianqiang Huang, Tao Lin, and Xinbing Wang. Networked time-series prediction with incomplete data via generative adversarial network. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(5):115:1– 115:??, June 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3643822.

Zhang:2014:DGP

[ZJL^{+}14] Gensheng Zhang, Xiao Jiang, Ping Luo, Min Wang, and Chengkai Li. Discovering general prominent streaks in sequence data. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(2):9:1– 9:??, June 2014. CODEN ????

ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2007:MPP

[ZKCY07] Minghua Zhang, Ben Kao, David W. Cheung, and Kevin Y. Yip. Mining periodic patterns with gap requirement from sequences. ACM Transactions on Knowledge Discovery from Data (TKDD), 1(2):7:1–7:??, August 2007. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zaki:2008:ISI

[ZKYW08] Mohammed J. Zaki, George Karypis, Jiong Yang, and Wei Wang. Introduction to special issue on bioinformatics. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(1):1:1–1:??, March 2008. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhou:2015:SIB

[ZL15] Yang Zhou and Ling Liu. Social influence based clustering and optimization over heterogeneous information networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):2:1–2:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2023:RNR

[ZL23] Liang Zhang and Cheng Long.

Road network representation learning: a dual graph-based approach. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (9):121:1–121:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3592859.

Zhu:2020:SCM

[ZLC20] Tianyu Zhu, Guannan Liu, and Guoqing Chen. Social collaborative mutual learning for item recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):43:1–43:19, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3387162.

Zheng:2014:FHE

[ZLD14] Li Zheng, Tao Li, and Chris Ding. A framework for hierarchical ensemble clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(2):9:1–9:??, September 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2023:SGU

[ZLD^{+}23] Zhijie Zhang, Wenzhong Li, Wangxiang Ding, Linming Zhang, Qingning Lu, Peng Hu, Tong Gui, and Sanglu Lu. STAD-GAN: Unsupervised anomaly detection on multivariate time series with self-training generative adversarial networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):71:1–71:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572780.

Zhou:2023:SPA

[ZLDL23] Peng Zhou, Xinwang Liu, Liang Du, and Xuejun Li. Self-paced adaptive bipartite graph learning for consensus clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(5):62:1–62:??, June 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564701.

Zhang:2024:TOS

[ZLGY24] Chunkai Zhang, Maohua Lyu, Wensheng Gan, and Philip S. Yu. Totally-ordered sequential rules for utility maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):80:1–80:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3628450.

Zhao:2024:HCN

[ZLHZ24] Hong Zhao, Zhengyu Li, Wenwei He, and Yan Zhao. Hierarchical convolutional neural network with knowledge complementation for long-tailed classification. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):154:1–154:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653717.

Zhang:2023:IGL

[ZLLW23] Zan Zhang, Lin Liu, Jiuyong Li, and Xindong Wu. Integrating global and local feature selection for multilabel learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(1):4:1–4:??, January 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532190.

Zhou:2024:NEP

[ZLSC24] Houquan Zhou, Shenghua Liu, Huawei Shen, and Xueqi Cheng. Node embedding preserving graph summarization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(6):145:1–145:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3649505.

Zhang:2009:DGA

[ZLT09] Zhenjie Zhang, Laks V. S. Lakshmanan, and Anthony K. H. Tung. On domination game analysis for microeconomic data mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 2(4):18:1–18:??, January 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2015:OBF

[ZLT^{+}15] Lei Zhang, Ping Luo, Linpeng Tang, Enhong Chen, Qi Liu, Min Wang, and Hui Xiong. Occupancybased frequent pattern mining. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):14:1– 14:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2023:DDQ

[ZLT^{+}23] Sensen Zhang, Xun Liang, Hui Tang, Xiangping Zheng, Alex X. Zhang, and Yuefeng Ma. DuCape: Dual quaternion and capsule network-based temporal knowledge graph embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):104:1–104:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589644.

Zhu:2021:ARD

[ZLW^{+}21a] Xiaoyan Zhu, Yingbin Li, Jiayin Wang, Tian Zheng, and Jingwen Fu. Automatic recommendation of a distance measure for clustering algorithms. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(1):7:1–7:22, January 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418228.

Zhu:2021:SCS

[ZLW21b] Yi Zhu, Lei Li, and Xindong Wu. Stacked convolutional sparse auto-encoders for representation learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(2):31:1–31:21, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434767.

Zhou:2020:EAK

[ZLY^{+}20a] Xu Zhou, Kenli Li, Zhibang Yang, Yunjun Gao, and Keqin Li. Efficient approaches to k representative G-Skyline queries. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(5):58:1– 58:27, August 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3397503.

Zhu:2020:PMG

[ZLY^{+}20b] Hongyuan Zhu, Qi Liu, Nicholas Jing Yuan, Kun Zhang, Guang Zhou, and Enhong Chen. Pop music generation: From melody to multi-style arrangement. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(5):54:1–54:31, August 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3374915.

Zhang:2019:RRH

[ZLZ^{+}19] Xuchao Zhang, Shuo Lei, Liang Zhao, Arnold P. Boedihardjo, and Chang-Tien Lu. Robust regression via heuristic corruption thresholding and its adaptive estimation variation. ACM Transactions on Knowledge Discovery from Data (TKDD), 13 (3):28:1–28:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314105.

Zhang:2024:PAT

[ZLZ^{+}24] Lei Zhang, Yong Liu, Zhiwei Zeng, Yiming Cao, Xingyu Wu, Yonghui Xu, Zhiqi Shen, and Lizhen Cui. Package arrival time prediction via knowledge distillation graph neural network. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (5):108:1–108:??, June 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3643033.

Zhang:2024:FPP

[ZMC^{+}24] Taolin Zhang, Chengyuan Mai, Yaomin Chang, Chuan Chen, Lin Shu, and Zibin Zheng. FedEgo: Privacypreserving personalized federated graph learning with ego-graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):40:1–40:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3624017.

Zhao:2024:DER

[ZMW^{+}24] Chen Zhao, Feng Mi, Xintao Wu, Kai Jiang, Latifur Khan, and Feng Chen. Dynamic environment responsive online meta-learning with fairness awareness. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (6):153:1–153:??, July 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648684.

Zhuang:2023:CLB

[ZMZS23] Jiabo Zhuang, Shunmei Meng, Jing Zhang, and Victor S. Sheng. Contrastive learning based graph convolution network for social recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):120:1–120:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587268.

Zhou:2009:LST

[ZP09] Bin Zhou and Jian Pei. Link spam target detection using page farms. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(3): 13:1–13:??, July 2009. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2015:DAV

[ZP15] Yao Zhang and B. Aditya Prakash. Data-aware vaccine allocation over large networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):20:1– 20:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2023:RCM

[ZP23] Mimi Zhang and Andrew Parnell. Review of clustering methods for functional data. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):91:1– 91:??, August 2023. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3581789.

Zhu:2016:EVM

[ZPC^{+}16] Wen-Yuan Zhu, Wen-Chih Peng, Ling-Jyh Chen, Kai Zheng, and Xiaofang Zhou. Exploiting viral marketing for location promotion in location-based social networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(2):25:1– 25:??, December 2016. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhong:2010:ATS

[ZPSYY10] Ning Zhong, Gregory Piatetsky-Shapiro, Yiyu Yao, and Philip S. Yu. ACM TKDD special issue on knowledge discovery for Web intelligence. ACM Transactions on Knowledge Discovery from Data (TKDD), 5(1):1:1–1:??, December 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhou:2022:UAN

[ZRJ^{+}22] Yang Zhou, Jiaxiang Ren, Ruoming Jin, Zijie Zhang, Jingyi Zheng, Zhe Jiang, Da Yan, and Dejing Dou. Unsupervised adversarial network alignment with reinforcement learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(3):50:1–50:29, June 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477050.

Zareie:2024:MDE

[ZS24] Ahmad Zareie and Rizos Sakellariou. Maximizing the diversity of exposure in online social networks by identifying users with increased susceptibility to persuasion. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(2):42:1–42:??, February 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3625826.

Zhang:2023:MVE

[ZSC^{+}23] Xuanqi Zhang, Qiangqiang Shen, Yongyong Chen, Guokai Zhang, Zhongyun Hua, and Jingyong Su. Multi-view ensemble clustering via low-rank and sparse decomposition: From matrix to tensor. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(7):103:1–103:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589768.

Zhong:2022:CFC

[ZSM22] Sheng Zhong, Vinicius M. A.

Souza, and Abdullah Mueen. Combining filtering and crosscorrelation efficiently for streaming time series. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(5):100:1–100:24, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3502738.

Zafarani:2015:UIA

[ZTL15a] Reza Zafarani, Lei Tang, and Huan Liu. User identification across social media. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(2):16:1– 16:??, October 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2015:WIY

[ZTL^{+}15b] Jing Zhang, Jie Tang, Juanzi Li, Yang Liu, and Chunxiao Xing. Who influenced you? Predicting retweet via social influence locality. ACM Transactions on Knowledge Discovery from Data (TKDD), 9(3): 25:1–25:??, April 2015. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2020:INA

[ZTT^{+}20] Si Zhang, Hanghang Tong, Jie Tang, Jiejun Xu, and Wei Fan. Incomplete network alignment: Problem definitions and fast solutions. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):38:1–38:26, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (elec-

                                        tronic). URL https://dl.

acm.org/doi/abs/10.1145/
3384203.

Zhang:2022:DEL

[ZWB22] Min-Ling Zhang, Jing-Han Wu, and Wei-Xuan Bao. Disambiguation enabled linear discriminant analysis for partial label dimensionality reduction. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (4):72:1–72:18, August 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494565.

Zhang:2022:HVA

[ZWC22] Hangbin Zhang, Raymond K.

Wong, and Victor W. Chu. Hybrid variational autoencoder for recommender systems. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):37:1– 37:37, April 2022. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3470659.

Zhao:2023:CDC

[ZWC^{+}23] Boxiang Zhao, Shuliang Wang, Lianhua Chi, Qi Li, Xiaojia Liu, and Jing Geng. Causal discovery via causal star graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (7):98:1–98:??, August 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3586997.

Zhang:2019:ICS

[ZWG^{+}19] Mingyue Zhang, Xuan Wei, Xunhua Guo, Guoqing Chen, and Qiang Wei. Identifying complements and substitutes of products: a neural network framework based on product embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):34:1–34:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3320277.

Zhao:2016:MPA

[ZWH^{+}16] Wayne Xin Zhao, Jinpeng Wang, Yulan He, Ji-Rong Wen, Edward Y. Chang, and Xiaoming Li. Mining product adopter information from online reviews for improving product recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(3):29:1– 29:??, February 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2022:EHI

[ZWLS22] Yiding Zhang, Xiao Wang, Nian Liu, and Chuan Shi. Embedding heterogeneous information network in hyperbolic spaces. ACM Transactions on Knowledge Discovery from Data (TKDD), 16(2):35:1–35:23, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468674.

Zhu:2024:ELD

[ZWSD24] Weiyao Zhu, Ou Wu, Fengguang Su, and Yingjun Deng. Exploring the learning difficulty of data: Theory and measure. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4): 84:1–84:??, May 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3636512.

Zhao:2020:UFS

[ZWWH20] Peilin Zhao, Dayong Wang, Pengcheng Wu, and Steven C. H. Hoi. A unified framework for sparse online learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(5):59:1– 59:20, August 2020. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3361559.

Zhang:2024:DSA

[ZWYW24] Yuhong Zhang, Jianqing Wu, Kui Yu, and Xindong Wu. Diverse structure-aware relation representation in cross-lingual entity alignment. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):92:1–92:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638778.

Zhang:2024:FTT

[ZWYZ24] Jiaxin Zhang, Yiqi Wang, Xihong Yang, and En Zhu. A fully test-time training framework for semi-supervised node classification on out-of-distribution graphs. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(7):172:1–172:??, August 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3649507.

Zhang:2022:IBL

[ZXL22] Xiaoying Zhang, Hong Xie, and John C. S. Lui. Improving bandit learning via heterogeneous information networks: Algorithms and applications. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (6):111:1–111:??, December 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3522590.

Zhang:2023:CIT

[ZXZ^{+}23] Hao Zhang, Yewei Xia, Kun Zhang, Shuigeng Zhou, and Jihong Guan. Conditional independence test based on residual similarity. ACM Transactions on Knowledge Discovery from Data (TKDD), 17 (8):117:1–117:??, September 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3593810.

Zhang:2013:MRL

[ZY13] Yu Zhang and Dit-Yan Yeung. Multilabel relationship learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 7(2):7:1–7:??, July 2013. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2014:RAL

[ZY14] Yu Zhang and Dit-Yan Yeung.

A regularization approach to learning task relationships in multitask learning. ACM Transactions on Knowledge Discovery from Data (TKDD), 8(3):12:1–12:??, June 2014. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2024:ADA

[ZYC^{+}24] Yang Zhang, Ting Yu, Shengqiang Chi, Zhen Wang, Yue Gao, Ji Zhang, and Tianshu Zhou. Attribute diversity aware community detection on attributed graphs using threeview graph attention neural networks. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (8):196:1–196:??, September 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672081.

Zhang:2024:HSF

[ZYD^{+}24a] Chunkai Zhang, Yuting Yang, Zilin Du, Wensheng Gan, and Philip S. Yu. HUSP-SP: Faster utility mining on sequence data. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):5:1–5:??, January 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597935.

Zhao:2024:PAD

[ZYD24b] Han Zhao, Xu Yang, and Cheng Deng. Parameteragnostic deep graph clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(3):66:1– 66:??, April 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3633783.

Zhou:2019:MTC

[ZYH19] Yao Zhou, Lei Ying, and Jingrui He. Multi-task crowdsourcing via an optimization framework. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(3):27:1–27:??, July 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3310227.

Zhao:2021:SIF

[ZYS^{+}21] Huan Zhao, Quanming Yao, Yangqiu Song, James T. Kwok, and Dik Lun Lee. Side information fusion for recommender systems over heterogeneous information network. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):60:1– 60:32, June 2021. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3441446.

Zhang:2021:SEB

[ZYZZ21] Daokun Zhang, Jie Yin, Xingquan Zhu, and Chengqi Zhang. Search efficient binary network embedding. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(4):61:1–61:27, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436892.

Zhang:2010:MDR

[ZZ10] Yin Zhang and Zhi-Hua Zhou.

Multilabel dimensionality reduction via dependence maximization. ACM Transactions on Knowledge Discovery from Data (TKDD), 4 (3):14:1–14:??, October 2010. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhai:2023:LGR

[ZZ23] Penglong Zhai and Shihua Zhang. Learnable graphregularization for matrix decomposition. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(3):32:1–32:??, April 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3544781.

Zhang:2024:DDG

[ZZF24] Ying Zhang, Zhiqiang Zhao, and Zhuo Feng. diGRASS: Directed graph spectral sparsification via spectrum-preserving symmetrization. ACM Transactions on Knowledge Discovery from Data (TKDD), 18 (4):102:1–102:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3639568.

Zeng:2021:FRD

[ZZG^{+}21] Shaoning Zeng, Bob Zhang, Jianping Gou, Yong Xu, and Wei Huang. Fast and robust dictionary-based classification for image data. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(6):97:1–97:22, July 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3449360.

Zhang:2015:SMB

[ZZL15] Xianchao Zhang, Xiaotong Zhang, and Han Liu. Smart multitask Bregman clustering and multitask kernel clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 10(1):8:1–8:??, July 2015. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2021:POP

[ZZLZ21] Zheng Zhang, Xiaofeng Zhu, Guangming Lu, and Yudong Zhang. Probability ordinalpreserving semantic hashing for large-scale image retrieval. ACM Transactions on Knowledge Discovery from Data (TKDD), 15(3):37:1–37:22, May 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442204.

Zhang:2009:EAG

[ZZW09] Xiang Zhang, Fei Zou, and Wei Wang. Efficient algorithms for genome-wide association study. ACM Transactions on Knowledge Discovery from Data (TKDD), 3(4): 19:1–19:??, November 2009. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zang:2021:JMS

[ZZXY21] Tianzi Zang, Yanmin Zhu, Yanan Xu, and Jiadi Yu. Jointly modeling spatio-temporal dependencies and daily flow correlations for crowd flow prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (4):58:1–58:20, June 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3439346.

Zhou:2021:HOS

[ZZY^{+}21] Dawei Zhou, Si Zhang, Mehmet Yigit Yildirim, Scott Alcorn, Hanghang Tong, Hasan Davulcu, and Jingrui He. High-order structure exploration on massive graphs: a local graph clustering perspective. ACM Transactions on Knowledge Discovery from Data (TKDD), 15 (2):18:1–18:26, April 2021. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3425637.

Zhang:2023:MLF

[ZZY^{+}23] Zan Zhang, Zhe Zhang, Jialu Yao, Lin Liu, Jiuyong Li, Gongqing Wu, and Xindong Wu. Multi-label feature selection via adaptive label correlation estimation. ACM Transactions on Knowledge Discovery from Data (TKDD), 17(9):134:1–134:??, November 2023. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604560.

Zhang:2024:BGL

[ZZY24a] Guixian Zhang, Shichao Zhang, and Guan Yuan. Bayesian graph local extrema convolution with long-tail strategy for misinformation detection. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(4):89:1–89:??, May 2024. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639408.

Zhang:2024:DAD

[ZZY^{+}24b] Qiuyue Zhang, Yunfeng Zhang, Xunxiang Yao, Shilong Li, Caiming Zhang, and Peide Liu. A dynamic attributesdriven graph attention network modeling on behavioral finance for stock prediction. ACM Transactions on Knowledge Discovery from Data (TKDD), 18(1):16:1– 16:??, January 2024. CO-DEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3611311.

Zhou:2022:OSS

[ZZYW22] Peng Zhou, Shu Zhao, Yuanting Yan, and Xindong Wu. Online scalable streaming feature selection via dynamic decision. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (5):87:1–87:20, October 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3502737.

Zhang:2016:SNE

[ZZYY16] Xianchao Zhang, Linlin Zong, Quanzeng You, and Xing Yong. Sampling for Nyström extension-based spectral clustering: Incremental perspective and novel analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 11(1):7:1–7:??, August 2016. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Zhang:2022:KKR

[ZZYY22] Zhenyu Zhang, Lei Zhang, Dingqi Yang, and Liu Yang. KRAN: Knowledge refining attention network for recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD), 16 (2):39:1–39:20, April 2022. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3470783.

Zhu:2020:SGC

[ZZZ^{+}20a] Xiaofeng Zhu, Shichao Zhang, Jilian Zhang, Yonggang Li, Guangquan Lu, and Yang Yang. Sparse graph connectivity for image segmentation. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):46:1–46:19, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3397188.

Zhu:2020:SWM

[ZZZ^{+}20b] Xiaofeng Zhu, Shichao Zhang, Yonghua Zhu, Wei Zheng, and Yang Yang. Self-weighted multi-view fuzzy clustering. ACM Transactions on Knowledge Discovery from Data (TKDD), 14(4):48:1–48:17, July 2020. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3396238.